A Complete Bibliography of Publications in SoftwareX

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

29 May 2024 Version 1.14

Title word cross-reference

1 [BRF^{+}23, KL22, PO19]. 1.5 [SAA18]. 2 [ACPdP22, BBF^{+}19, BBN^{+}22, Can22b, DEV^{+}21, EHGPT19, Elk22, KDHG20, KPN^{+}18, MASB18, MBI20, Ozb17, RDO^{+}19, THT20, ZRK19]. 3 [AT22, Arc20, BB20a, BSV^{+}23, CC20, GRF23, HS20, IIS18, KWZ22, KWB22, LCMD^{+}18, LMS^{+}16, MBC22, MLTF^{+}18, NS22, Ozb17, Ras20, RDO^{+}19, SB21, THT20, Zek17]. 4 [BFSJP^{+}21, KPN^{+}18, Zek17]. < [HVH24]. > [HVH24]. ^{++} [SLBC^{+}20, SH23]. ^{2} [MCC20]. ^{3} [CMDDARG23]. ^{x} [Afr22]. β [WBS21]. F [JS19]. K [BHC^{+}24]. l_{1} [YC19]. l_{2} [YC19]. μ [CBHLG21, OAF20]. q [BHHÖ22]. UnBlocks^{gen} [Ras20].

-amino [WBS21]. -d [CC20, PO19, Zek17].

-dimensional [Elk22]. -peptide [WBS21]. -signature [JS19]. -space [BHHÖ22]. -strategy [Arc20, IIS18].

.NET [Fra22a]. .npy [Rei23]. .npy [?]Reininghaus:2023:CCL. .npz [Rei23]. .STL [AISM21]. .STL-based [AISM21].

1 [ddSNX22]. 1.0 [Ars21, MP20]. 1.1.0-pyfao56 [BTDT23]. 11p [CPUARC20]. 15 [RNK22]. 17 [Rei23, Sho24]. 18 [GCdJAURO21]. 19 [CK23, EBGCT21, GPK^{+}21, IVSV23]. 1D [PBS^{+}24]. 1DCSEMQWE [KL22].

2 [CCTCGCOB23, CCO21, DG21, FNX22, GPW^{+}20, GNPP23, vDPI^{+}18]. 2-D [BSB^{+}23, MBDS20, GNPP23]. 2-Dimensional [CCTCGCOB23]. 2.0 [AGL^{+}23, BB20a, JWJ^{+}22, JHAMF19, Jur23a, LGP^{+}22, RS22]. 2.0-MOOSE [LGP^{+}22]. 2.0-pyFIRI [ZRK22]. 2/low [YLM^{+}24]. 2019 [Ano19d]. 2020 [Ano20d, Ano20c]. 2021 [Ano21a, Ano21f, Ano21h, Ano21g]. 2022 [Ano22a, Ano22f, Ano22h, Ano22g]. 2023 [Ano23a, Ano23f, Ano23g, Ano23h]. 2024 [Ano24c, Ano24d]. 23 [EWT^{+}24]. 24 [JP24]. 2BarG [GLR22]. 2D [ANA16, AANA19, Azi23, Can22a, Can24, Elk22, MFCV20, MBDS20, ddSNX22]. 2D-VSR-Sim [MBDS20]. 2D-xtDCR-LTBEM [Azi23]. 2D/3D [MFCV20]. 2PJIT [AT22].

3 [BÇ23, Kar21, KF22, MCGK19, MSV23, SCD^{+}24]. 3-D [BÇ23, Kar21]. 3-Tau [KF22]. 32/64 [MG22]. 32/64-bit [MG22]. 3D [FNX22, MFCV20]. 3DEM [GRF23]. 3DRSP [BC22]. 3s [IIS18]. 3TM [KF22].

56 [BTDT23, Tho22, TBD23]. 5G [NNDS23]. 5G/6G [NNDS23].

64-bit [MG22]. 6G [NNDS23].

7.1 [BO19].

802.11p [CPUARC20].

9 [KZ18]. 90 [Taq16].

Ab-initio [XSS20, XSC^{+}21, ZJKS23]. ABAQUS [FH23, GTC21, LH23, SMD23]. ABAQUS-machine [FH23]. Abaqus/CAE [LH23]. ABAQUSTM [WMZH23]. ability [PZKK21]. ABMU [Che21]. Aboria [RB17]. absolute [MTS^{+}18]. Absorption [EFTS23]. academic [Kwa19]. accelerated [ACPdP22, CWGM24, CWLG^{+}21, DS20, LAA23, Nev17, MCBM24]. accelerator [LCGA22]. accelerators [RLK18]. access [CPF23, MWJ15, VEM^{+}18]. accessible [CPB^{+}23, FAM^{+}20, BGCS19]. accident [AMA19]. accumulated [Alb19]. accurate [NOÖÇ19, PO19]. accurately [AAN17]. Acid [CGKD23, BSS^{+}23, WBS21]. acidified [BSS^{+}23]. Acme [SB21]. ACORNS [DSJ^{+}22]. acoustic [DDCSH23, LEFSO20, RLJS23, SWK19, SA17]. ACPYPE [BFRK19]. acquire [MGMAG22]. Acquisition [CWM^{+}21a, BHB^{+}21, PS19a, Sur20, UdL20]. acquisitions [ADSG^{+}20]. ACROPOLIS [SCSB24]. Acrossim [SV23]. actin [MBM22]. active [VMS^{+}23, SE23, TBO24]. activity [AGADAG^{+}23, ACNF22, Has18]. Actual [BS19]. actuator [CCBC^{+}21]. Acuity [TMGW23]. Adaptable [GPW^{+}20]. adaptative [BBB^{+}20]. Adapting [FWB^{+}21]. Adaptive [BBJ^{+}18, CUSRCP^{+}22, NMT23, RDO^{+}19]. adaptivity [CBPC23, KBB19]. Add [OCK^{+}23]. Add-in [OCK^{+}23]. additively [OJ22]. ADIOS [GPW^{+}20]. ADITU [GEG22]. Adjoint [Bla21, Bla23]. Adjoint-Based [Bla21, Bla23]. adjoints [RJN^{+}23]. ADM1jl [AMAE24]. advanced [DM19, GAM^{+}24, SLM23, AAA^{+}21, BHB^{+}21, KCP^{+}24]. Advancing [AAF^{+}24a]. adversarial [VDCL20]. advligorts [BHB^{+}21]. Aedes [IDE^{+}21]. aerial [BGAGJGC23]. aerodynamics [BGGL20]. aerosol [DAC^{+}24]. AFDeter [YHC^{+}22]. AFF3CT [CHL^{+}19]. affective [JCFM24, KW23]. AffecTube [KW23]. ag5Tools [BdSvE23]. against [BGM^{+}24, CSZM22]. AGAT [RKDP21]. AGDATABOX [CMdS^{+}23, BJM^{+}19, MBU^{+}19]. AgDataBox-Map [MBU^{+}19]. AGDATABOX-RS [CMdS^{+}23]. Agent [Giu19, COGP19, PBM22, PMBM^{+}23, ZR19, Che21, FRNW^{+}24, GG21]. agent-based [COGP19, PBM22, ZR19, Che21, GG21]. agents [WWB^{+}23]. AgERA5 [BdSvE23]. Aggregate [CVAP22]. Aggregated [KRA23]. aggregation [AAB23, VEM^{+}18]. agile [ASRI22, BBS^{+}22, GZA^{+}22]. agnostic [HR21, SVM^{+}23]. agricultural [PMS23, VFFS23, VEM^{+}18, ZTC^{+}21, dSBSvE23]. agriculture [BJM^{+}19, MBU^{+}19, REK^{+}23]. agro [AMC17]. agro-hYdrological [AMC17]. agroecosystem [Sha23]. agrometeorological [BdSvE23]. AI [AAF^{+}24a, CCO21, Geo17, GBW^{+}24, GSS^{+}23, NNDS23, NSO21, REK^{+}23]. AI-based [REK^{+}23]. AI-enhanced [AAF^{+}24a]. AICodeReview [AAF^{+}24a]. aid [KHM^{+}22]. aided [ZRBCI20]. airflow [YHC^{+}22]. airflow-related [YHC^{+}22]. AKFruitData [MGMAG22]. AKFruitYield [MAGM^{+}23]. ALBOMAURICE [IDE^{+}21]. albopictus [IDE^{+}21]. AlfaMCV [dQRRBdSM20]. Algebra [AMWHD23, Kri22, RV20, LK21]. algorithm [AWO^{+}21, COG19, KHM^{+}22, KPN^{+}18, LHCK18, LSPG22, LSMG19, PG18, SEL^{+}16, TC20, ZY23, ZNS17, dSM22, BGAGJGC23, dSM22]. Algorithms [CKC22, AGC^{+}24, CCFSB20, Gag22, GCC20, HH21, HX22, KPTW19, KFA^{+}24, KG23, KBB19, NGOT24, PMMF19, REFB17, SFF^{+}19, SUM21, Sil20, SBT22, TS21, VMRVF^{+}24, WC23]. algorithms-based [AGC^{+}24]. alignment [BCA19, CC23, MKB^{+}18]. alignments [TBCG21]. ALIIAS [ESE^{+}23]. all-atom [HA23]. allele [HHN20]. AllInOne [NHR23]. allocation [ACNF22, MTJ^{+}22]. ALPHA [WLW23]. alphaBetaLab [MVBF19]. alteration [HHN20]. AMBER [BFRK19, SHvW20]. AMED [KRA23]. America [BGCS19]. amino [WBS21]. AMLBID [GABH22]. AMLCS [NRC23]. AMLCS-DA [NRC23]. amorphous [HA23]. AMPAO [KAR^{+}19]. amplitude [JAC20, LBL^{+}21]. Anaerobic [AS18, AMAE24]. ANALYSE [WWB^{+}23]. analyser [KG24, SBCK17]. analyses [BSdG^{+}22, EFTS23, MCC20, SCK^{+}23].

Analysing [ZTC^{+}21, PBM22]. Analysis [AJR^{+}23, EL20, FST^{+}21, KAR^{+}19, KSFG18, ŁMF^{+}23, PCC^{+}19, UVPB^{+}22, AM19, AMA19, AK24, AS17, AHR^{+}22, ASRI22, ATCA20, BMK23, BÇ23, BBD^{+}22, BHC^{+}24, BT23, BAEBAS19, BJM23b, BlD^{+}20, CRV22, Car21, CC20, CPB^{+}23, CCG^{+}23, CMMF19, DDCSH23, DM19, DKL^{+}21, DTDd19, DPD22, Elk22, EWT^{+}23, EWF^{+}23, EWT^{+}24, FS23, FF24, GZS23, GZA^{+}22, GMF20, GEH19, HT18, HOB23, HRC20, ITL^{+}23, IHT23, Jur22, KSS20, KS21, KXK^{+}23, KHJ^{+}23, KGK^{+}20, KAO^{+}21, LKK24, LKK23, LSMG19, LCMD^{+}18, LRL23, LYZ^{+}22, LVCF24, LS23, LLM^{+}20, LBL^{+}21, MSDP23, MPMC21, MAGM^{+}23, MCBM24, MASB18, Mül23, NK20, OJ22, OLRLB21, Ozb17, PL23, PLC^{+}24, PRSS19, PDM22, PKA^{+}22, Por23, PNC23, RLK18, Ras20, RIHVP23, RXZ^{+}23, RGP^{+}22, RPV^{+}20, RMMG21, RSMW20, Rub16, RL19]. analysis [SAD22, SGB^{+}22, SF16, SK22, SBCK17, SGA^{+}16, SA20, SCG^{+}20, SMKL23, SRML17, SL20b, SAS^{+}22, TŚG^{+}24, TJS18, TJS19, TCCK23, VMRVF^{+}24, VCLS21, WGB16, Web17, Web21, WM21a, WPM^{+}20, Wet20, WKS22, WS24b, XZL24, YFVL24, ZWCQ22, ZS24, dSBSvE23, tHLMN19, LMN18, Alb15, FPBM18, GAE23, LFT21, MP20, PARS22, Vel23, dSBSD17, ABC^{+}20]. analyte [AHR^{+}22]. analytic [NSLD16, TF23, MRMD19]. analytical [FS19, KPN^{+}18, MC18]. Analytics [EPF^{+}23, LFK^{+}24, ÇA23, DVSD22, LFGVGDBD24, MML20]. Analyze [dMOH21, Arc20, ABC^{+}20, Fru21, GMNG^{+}18, IIS18, LMB^{+}19, OBWM23, ZAPS20]. analyzer [KALS24, MTPHH18, VFFS23, CCV22]. analyzing [BHKK23, CCV22, GB24, GPEM23, MK16, PDH16, SMZSW24, WLW23]. ANDez [KK24]. Android [GLS^{+}22, AGVM20, BT23, GJS^{+}22, JRP23, ML20, PRC^{+}23, She19, YLS^{+}18]. Android-based [GJS^{+}22]. ANDURIL [tHLMN19, LMN18, RtHLMN20]. ANDURYL [tHLMN19, RtHLMN20]. angiography [MTS^{+}24]. angle [HDA21, HHB^{+}22b, MTPHH18, OLRLB21, RGP^{+}22, SL20b]. angle-resolved [SL20b]. Angular [Fra22a]. AniLength [Jun21]. AnIMAGE [VRVAC23]. animal [HM23, PLC^{+}24]. animated [IAW^{+}15]. Animation [CVS19]. anisotropic [Alb19, Azi23, RCT20]. anisotropic-DCR [Azi23]. annotating [MP17, SSW^{+}24]. Annotation [MDG22, STC23, CLM^{+}20, GZA^{+}22, KZ18, Ozk23, RKA^{+}23, TT17, WSRMB23]. annotations [KW23]. anomalies [MTS^{+}18]. Anomaly [ASSS24, BSB20, BJM23a]. Anonymization [ESE^{+}23]. Anonymization/Pseudonymization [ESE^{+}23]. ANOXic [KP23]. Ansible [BB22a, DDT20]. AnsibleMetrics [DDT20]. answering [HCD24]. ant [GCC20]. antenna [SDP^{+}20]. antennas [VRVAC23]. anyFish [IAW^{+}15]. AnyMOD.jl [Gök21]. AP [EWT^{+}24, EWT^{+}23]. API [BJM^{+}19, BKM22, KJKK23, MKS^{+}23]. APIs [BBS^{+}22]. app [AGADAG^{+}23, MKMK23, NFASC21, SNLFGVRB24, She19, SNLFGVRB24]. apple [MAGM^{+}23]. Application [MASB18, O’D21, AL16, BSV^{+}23, CS21, Cin24, CRF21, CMdS^{+}23, CBS^{+}16, DZH^{+}23, DPD22, DER^{+}23, EWT^{+}23, EWT^{+}24, FK23, GJS^{+}22, HBS16, IKYY22, JRP23, KK22, KXK^{+}23, LHB24, MBM22, MSLNSMOG24, ML20, MKT^{+}23, MKMK23, MGMAG22, MAAOZMAM22, MAM^{+}23, OE21, PRC^{+}23, PMS23, PP19, Por23, PNC23, RSLL^{+}20, SFF^{+}19, SLL23, YLS^{+}18, ZCCGGV^{+}22, HHH22]. application-oriented [AL16]. application-specific [LHB24]. applications [ABM^{+}21, CPF23, CPV^{+}20, DEV^{+}21, DZZ^{+}22, GBW^{+}24, HVCS^{+}23, LMM22, MGCS23, NNDS23, NSO21, PRC^{+}23, REK^{+}23, RWJ^{+}21, SDL21, SCG^{+}20, SMD^{+}24, Fra22a]. apply [HRC20]. Applying [BHHÖ22, HH21]. approach [AMVB19, ACO21, AGG23, BBHB22, DEV^{+}21, GAE23, HJRBM23, KZ18, LIZ^{+}20, MDL^{+}18, NMT23, SNLFGVRB24, SB22, SCPC18, TK23, XYC22, dRB^{+}20, MSDP23]. approaches [BBG^{+}19, TBO24]. approximant [LF15]. approximation [GWGB22]. apps [AGVM20, BT23]. APRCOIE [LHL^{+}24]. AQUADA [SSC23]. aquatic [KP23]. AQUINAS [FS23]. arbitrary [TSGOC^{+}22]. Arc [AGG23]. ArcGIS [OCK^{+}23]. arches [MMCKK21, STC^{+}18, YSM22]. Architect [SLPR20]. architectural [AI21]. architecture [AOGC^{+}20, CK22b, GWK16, JFJM22, MKS^{+}23, SFG21]. archival [Ver22]. ARCP [AGG23]. area [GZ21a, VJA^{+}18]. ARGAEL [STC23]. ARGument [STC23]. arguments [Bat19]. arithmetic [Rou19]. army [FVA^{+}20, SC17]. arrays [JM20]. art [MSV23, TS22]. ART data analyzer [TLDM19]. arterial [MML23]. artery [MCK^{+}23]. Article [AB24, Ric19]. artifact [ADSG^{+}20]. Artificial [dlTAJCÁBG24, DSB^{+}23, LCGA22, MMS20]. ASAS [dSBSD17]. ASE [KGF^{+}24]. ASKI [SF16]. ASS4HR [MMS20]. assess [HH21, MRGAGP23, dlSVBLdA^{+}17]. assessing [OCK^{+}23, RBQ^{+}24, SCPC18, VDCL20]. Assessment [AS18, NSS^{+}19, SO23b, ASVV23, CFA22, FJ22, GPK^{+}21, LSB^{+}20, PCPC24, RHR^{+}23, SUM21, SO23a]. assessments [VEM^{+}18]. assets [TTHZ23]. assignments [CFA22]. assimilation [NRC23]. assistant [MZSH21]. assisted [SM19, HKM^{+}19]. association [HMK23]. associations [HTWT23]. astronomical [SHvW20]. astrophysical [Duc23]. astrophysics [MAC^{+}21]. ASVtorch [LVK21]. asymptotically [ZKJ^{+}24]. ATM [GMNG^{+}18]. Atmospheric [vdOJP^{+}20, NRC23]. atom [HA23, OT16]. Atomic [AGHK21, dSBSD17, KGF^{+}24]. atomization [HS20]. attack [WWB^{+}23]. attacks [MSS^{+}23]. attenuation [PDTG17]. AUC [GZ21a]. AudExpCreator [NK18]. audio [BCA19, SA17]. auditory [NK18]. Audur [VJA^{+}18]. AUGMECON [FPN^{+}22]. AUGMECON-Py [FPN^{+}22]. augmentation [APLI24]. augmented [YYL^{+}22]. auroral [ZRK21, ZRK22]. Authentication [ESE^{+}23]. author [KK24]. auto [GABH22, DRM20]. auto-explained [GABH22]. Auto-generation [DRM20]. AutoCNet [LRPD18]. autocorrelation [Byk19]. AutoCT [BEPB24]. Automated [AJR^{+}23, BEPB24, CMMF19, EL17, vKMH^{+}20, AK24, ADSG^{+}20, CdSLCC20, CLM^{+}20, GMF20, HGG20, HDA21, MBA^{+}22, REK^{+}23, SSSH16, SSC23, dHGMCGL^{+}24, GABH22, LEK22, RLJS23, SC23]. Automatic [BNC24, KAR^{+}19, SDL21, dHGMCGL^{+}24, ASVV23, BPL^{+}19, CFA22, CCG^{+}23, HKF^{+}20, JGGF23, Jun21, NBCC19, PDM22, RMM18, ZCCGGV^{+}22, MVBF19]. automatically [CRP21]. Automating [JS19, NGK^{+}20, TT17, SPMGA^{+}23, TLDM19]. automation [BLM^{+}22, EBL24, GM24]. autonomous [AA23, FWJ^{+}23, PMS23, PCPC24]. Auxiliary [DK19]. available [AFGH22]. average [Ras21]. averaging [RLM^{+}24]. aware [TGS^{+}19, VC18, OPPZ22]. aXBRL [She19]. axioms [BKM22]. axis [OYW^{+}21, Web23]. axisymmetric [FS23, MFQ^{+}21]. AYNEXT [SAA^{+}23b]. azTotMD [Ras19a, RS22]. Azure [MGMAG22, MAGM^{+}23].

BabelFish [AVB17]. BabiaXR [MLGBR23]. Back [O’D21]. Back-Gated [O’D21]. Backdoor [MSS^{+}23]. bacterium [QTL23]. badcrossbar [JM20]. Badlands [Sal16]. balance [ZNS17]. balanced [CLK21, RDO^{+}19]. balancing [MB23]. balls [YC19]. BAM [DM19]. Bank [Kur23]. BANSHEE [KMLP^{+}23, PMNWR20]. bar [GLR22]. barriers [OCK^{+}23]. Base [YDN21, YD20]. Based [Bla21, Bla23, Giu19, AFGH22, AGC^{+}24, AGADAG^{+}23, AISM21, Agu23, AMA19, AANA19, AK24, AG21, AKC23, ACO21, ATT21, AHR^{+}22, ASVV23, AAF^{+}24b, Bac21, BSdG^{+}22, BÇ23, BBHB22, BGM^{+}24, BOP^{+}23, BBHP^{+}23, BC22, BWMS22, BSS^{+}23, BCF20, BSV^{+}23, CRV22, CCBC^{+}21, CS21, COGP19, CBPC23, CJA23, CLDdM22, CPD^{+}20, CCAC22, CCFSB20, CK22a, CK23, CCO21, CDOBA22, CWM^{+}21b, CKC23, Dav21, Deg20, DZH^{+}23, Ek16, FH23, FWJ^{+}23, FNX22, FK23, GLS^{+}22, GZ21b, GZS23, GZR^{+}19, GACS^{+}23, GP22, GZW^{+}22, GCBDDBF23, GG21, GYWG23, GMNG^{+}18, GJS^{+}22, HKZ24, HJRBM23, HOM^{+}18, HLR^{+}21, HMC^{+}24, HRC20, IZBT21, IIH^{+}23, JLP23, JZGW22, JGGF23, Jun21, KG24, Kar21, KWB22, KG23, KALS24, Kos22, KXK^{+}23, Krä20, KPJ24, Kur21, LK21, LSPG22, LCMD^{+}18, LLP^{+}24, LPL^{+}24, LF15, LSSK16, LYX^{+}18]. based [LYZ^{+}22, LHZT23, LDAL20, LVCF24, LQM22, MKWS22, MSDP23, MKB^{+}18, ML20, MBDS20, MFJ23, MB21, MCK^{+}23, MCPG24, MBI20, Nai17, NMLM18, NGK^{+}20, NK18, NL23, OBWM23, PMS23, PBSB22, PBM22, PLK^{+}23, PS18a, PS18b, PS18c, PS19b, Por23, RBQ^{+}24, REK^{+}23, RJH^{+}20, RB17, SF16, SUM21, SSW^{+}24, SKM19, SGA^{+}16, SA20, SABEh20, SMR22, SADU23, SS17, SLM23, TK23, TOPC23, TTT19, UdL20, VDCL20, VCA^{+}24, VJA^{+}18, VMRVF^{+}24, Ver22, VRVAC23, VEM^{+}18, VCLS21, WBS22a, WBS22b, WBNS23, WRV23, WSAC^{+}23, XZL24, XTD24, YKKD19, YHC^{+}22, YB23, ZCCGGV^{+}22, ZY23, ZNS17, ZLLT23, ZR19, ddSNX22, dPPD^{+}23, dlSVBLdA^{+}17, Che21, FLR22]. Baseliner [OHO16]. BASEveg [CVV23]. basic [LKK23]. basin [Sal16]. Bath [WW17]. Batteries [CGKD23]. Battery [HBA^{+}20, NMT23, TSNRM24]. BayesESS [SMKL23]. Bayesian [KMLP^{+}23, GdVSdSCM20, Hel22, IKYY22, LVT^{+}22, PMNWR20, PLGHG^{+}23, STH^{+}21, SWCP20, SMKL23, SR17, XYC22]. bayesvl [LVT^{+}22]. BBR [ASVV23]. BccFccRaycaster [LMS^{+}16]. BCI [LHZT23]. BDEsize [CLK21]. beaches [BFDD^{+}22]. beam [Dan22, SLM23, THG^{+}23]. beams [DM19]. beamWeldFoam [FPI^{+}22]. bearing [WLW23]. bed [GMNG^{+}18]. Bedforms [GMNG^{+}18]. Bedforms-ATM [GMNG^{+}18]. BEEP [HBA^{+}20]. behaved [LXK23]. behavior [IAW^{+}15, KH19a, SIC24, WMM18, dFB22, PCPC24]. Behavioral [Yaa23, ITL^{+}23]. behavioural [SDCA19]. being [KFA^{+}24]. Bembel [DHK^{+}20]. benchmark [CC21]. benchmarking [AES^{+}22, Gag22, GFRMVL24, JLP23, MBF20, MAGM^{+}23, PMMF19, REFB17]. benchmarks [KPN^{+}18, SNW^{+}21]. bends [RNG^{+}21]. Benopt [JMT22]. Benopt-Heat [JMT22]. BENs [WKR^{+}20]. Bessel [Tak22]. Best [Nis20]. Beta [FK23]. between [BDPZ19, BGAGJGC23, DER^{+}23, EAMSR^{+}23, HTB19, LS16, MP18, Nai17, TBO24]. bias [SLL23]. BiasAdjustCXX [SLL23]. bibliographic [Vel23]. bibliometric [Mül23]. bibliometrics [RK19]. BIcenter [ACO21]. Big [GABH22, HTWT23, HL22, Krä20, SKD22, WLP16]. binary [RKDP21, ZGZvB19, vDPI^{+}18, BCD22]. binding [LEK22]. bindings [AS23]. binomial [SR17]. bio [GSP^{+}17]. bio-economic [GSP^{+}17]. biocontrol [CH23]. BioEmoDetector [ASGC^{+}24]. bioenergy [JMT22]. biofeedback [CK22a, CK23]. biogas [HKC^{+}18]. bioimage [GMF20]. bioinformatics [Cin24, KPSM17, LFGVGDBD24]. BioIntertidal [HJRBM23]. biomass [MTJ^{+}22]. Biomedical [HLW^{+}16]. biometric [SHK^{+}23]. BiometricBlender [SHK^{+}23]. biomolecular [dRB^{+}20]. BioNES [CK22a, CK23]. biophysical [ASVV23, KEwMO24]. bioScience [LFGVGDBD24]. BioSPPy [BSC^{+}24]. bipartite [KKAUA21]. Bird [BGGL20]. bit [MG22]. black [AGVM20, AGdSC20, KCP^{+}24, PRC^{+}23]. black-box [AGVM20, AGdSC20, KCP^{+}24, PRC^{+}23]. blast [dFB22]. BLASTn [TBCG21]. Blaze [GWK16]. Blaze-DEMGPU [GWK16]. blender [TTHZ23]. Blink [SDG^{+}24]. BLNN [SWCP20]. block [MASB18, Por23]. block-in-matrix [MASB18]. blockchain [TWL22]. BlockTrainHK [TWL22]. blood [GMGG19, MML23]. BLUE [Nis20]. blueprints [DDT20]. bluffs [PL23]. bluffs/cliffs [PL23]. BMIL [HLW^{+}16]. board [GBW^{+}24, PS19a, Ano15a, Ano15b, Ano16, Ano17, Ano18, Ano19c, Ano19a, Ano19b, Ano20a, Ano20b, Ano21b, Ano21c, Ano21d, Ano21e, Ano22b, Ano22c, Ano22d, Ano22e, Ano23b, Ano23c, Ano23d, Ano23e, Ano24a, Ano24b]. Bodies [YMCB23, MP18]. body [CCH^{+}19, PDTG17, WBH^{+}23]. body-wave [PDTG17]. Boltzmann [HOM^{+}18]. bond [BBP^{+}18]. bone [Kar21]. boosted [SIC24]. Boosting [NSO21, GAE23]. boostingDEA [GAE23]. Botrytis [ASVV23]. boundary [BÁP^{+}22, DHK^{+}20, KAK21]. bounded [CK21, GEG22]. box [AGVM20, AGdSC20, KCP^{+}24, LHB24, PRC^{+}23, JJT^{+}23]. BOXVIA [IKYY22]. BPMN [OPPZ22]. BPpy [Yaa23]. Bragg [PMM16]. Braid [MC18]. Braided [MC18]. brain [LHZT23, KG23]. brain-computer [LHZT23]. branch [GL15]. break [LQM22]. breast [KAO^{+}21]. bridge [Can24, Can22a]. bridges [HLG22]. broadening [EWT^{+}23, EWT^{+}24]. broader [HL22]. broken [MWM20]. browser [AAF^{+}24b, Nai17]. browser-based [AAF^{+}24b, Nai17]. browsing [APLI24]. brushed [NCGLRR24]. BSuccinct [Bel24]. BuckinghamPy [KS21]. buckling [Ozb17]. buffer [BSS^{+}23]. BufferCapacity3 [Bre23]. buffering [Bre23]. bugs [dHGMCGL^{+}24]. Buhos [NMLM18]. build [FVD20, PMZ21]. Builder [Dav21, HESH19, HR21, SDL21, CUSRCP^{+}22]. Building [Fra22a, RKDP21, BS19, CUSRCP^{+}22, GZ21a, GZ21b, MBI20, PMMF19, RLM^{+}24, WBS21, ZRBCI20]. buildings [Elk22, SK22]. bursts [DTDd19]. business [AHtH20, GFRMVL24, HTWT23].

C [BC20, CM19, Dan20, DHK^{+}20, GMF18, HVB16, KJKK23, Kis20, Kor21, MCC20, MB21, MP18, MMH19, OT16, OE21, Rei23, SLBC^{+}20, SMD23, Sho24, WLW23]. C# [MP18]. C-language [OT16]. C/C [BC20]. C37.118.2 [AVB17, BFV18]. C37.118.2-compliant [AVB17]. CACP [CKC22]. CACPPAF [MBM22]. CAD [CC20]. CAE [LH23]. CAinterprTools [Alb15]. Calc [PKKQ20]. calculate [MWM20]. calculating [Alb19, AKC23, CBS^{+}16, KLY21, Ohn21, Ras21, TKLG19]. calculation [DSI^{+}20, JS19, MGK^{+}20, PDTG17, Tak22, WW17]. Calculations [XSS20, BBP^{+}18, Dan22, KP19, LLM^{+}20, RRSK18, TSCH20, XSC^{+}21, ZJKS23]. calculator [KSP19, SSM22, STC^{+}18, SSM22]. calculus [BJM23b, ZS22]. calibration [CWC^{+}23, CAW^{+}20, KH19b, SMD^{+}24]. call [LPS^{+}24]. called [RLM^{+}24]. caloric [SAA18]. CAM [MC18]. camera [MLTF^{+}18, MKR^{+}21]. cameras [CWC^{+}23, CAW^{+}20, MGMAG22, MAGM^{+}23]. CameraTransform [GRW^{+}19]. can [KP23]. cancer [HHN20]. candidate [HMK23]. canonical [SRK23]. capacity [BSS^{+}23, TSNRM24, Zie19]. captured [CBHLG21]. capturing [RXZ^{+}23]. Caratheodory [BV19]. carbon [WTAG^{+}22]. cardiac [NGK^{+}20, RHR^{+}23, WDZ^{+}20]. cardiomyocyte [FG20]. cardiovascular [RSLL^{+}20]. cards [HMCA15]. care [WSRMB23, CRP21]. Carlo [NBS^{+}21, BCA19, CCE21, DAC^{+}24, IIH^{+}23, KP19, KXK^{+}23, LNS15, NBM^{+}19, vKH20]. Cartesian [BDF^{+}20]. CAS [Rag17]. case [AMA19, BGGL20, CNST20]. cases [MVG^{+}24]. cashocs [Bla21, Bla23]. cat2cat [NG23]. CaTchDes [BV19]. categorical [MLTF^{+}18, NG23]. CATS [WSRMB23]. Causal [ZL23]. causality [RMMG21]. cause [KCP^{+}24]. cause-effect [KCP^{+}24]. cavitating [LXJ^{+}23]. cavities [Deg20]. cavity [dRB^{+}20]. CBE [TSCH20]. CCMORPH [ŁMF^{+}23, TŚG^{+}24]. CCPi [KPTW19]. CCPi-Regularisation [KPTW19]. CDB [EWT^{+}24, EWT^{+}23]. CDB-AP [EWT^{+}24, EWT^{+}23]. cDNA [KSP19]. CEG [KCP^{+}24]. cell [CLM^{+}20, GMGG19, LKSS20, TGS^{+}19, WTZ^{+}21]. cells [CMMF19, PNC23]. cement [TCQOGM23]. CemrgApp [RSLL^{+}20]. centrality [FF24]. centric [SKD22]. CEREI [ICS^{+}24]. CESSIPy [CR22]. CFAR [NOÖÇ19]. CFD [BBB^{+}20, SC23, SV23]. chain [AGL^{+}23, CCE21]. chains [PO19]. Change [STH^{+}21, NR16, XYC22, YBVC^{+}22, SC17]. changes [DL16]. channel [DTDd19, TCCK23]. channels [AM23, CSZM22, RNG^{+}21]. chaotic [Ben15]. characteristic [LYZ^{+}22]. characteristics [BO19, Car21]. characterization [HDA21, LBH^{+}20, OKJ23, RNG^{+}21, SADU23, THT20, UVPB^{+}22, WW17]. characterize [MBM22, RAL^{+}20]. ChargEVal [PBM22]. charging [PBM22]. ChatbotSQL [PMBM^{+}23]. checking [Pos22]. chemical [GV20]. chemicals [TBO24]. chemistry [KP19]. chess [RL19]. ChessY [RL19]. Chinese [LHL^{+}24]. chlorophyll [HLR^{+}21]. chromatography [MMD24]. Chrome [KW23]. chromosomal [HHN20]. chronic [MTS^{+}24]. chronostratigraphic [AS17]. Chronoweb [GB24]. Chunks [RRKA22]. CI [RLM^{+}24]. CIM [GVAO19]. CIM-2-mod [GVAO19]. cinerea [ASVV23]. CIOA [dSM22]. Circle [dSM22]. Circle-Inspired [dSM22]. Circuit [OV24, NMT23]. circuits [AGC^{+}24]. Circular [STC^{+}18, Nas20]. Circulation [NRC23]. CIRN [BB20b]. cities [DV23]. city [SUM21, dMOH21]. CityZones [PCdJSdFR^{+}23]. civil [CR22]. ck [HTWT23]. ck-FARM [HTWT23]. class [HVB16, LYZ^{+}22, MGGHG^{+}24, SHK^{+}23]. Classic [AAN17]. classical [MDL^{+}18, YBVC^{+}22]. classifiability [ZWCQ22]. classification [ÁG19, ATT21, AGH20, BHC^{+}24, CCTCGCOB23, COG19, DZH^{+}23, GZ21a, GZ21b, KMT23, MAM^{+}23, RLF^{+}21, SCSB24, SO23a, SEdCCCRG23, TTT19, YG19, ZGZvB19, CKC22]. classifier [CKC23, Kle24]. ClassNoise [MGGHG^{+}24]. Classwms [BHC^{+}24]. Clava [BC20]. Cliffs [ŁMF^{+}23, PL23, TŚG^{+}24]. climate [AFGH22, ABM^{+}21, VEM^{+}18]. climatic [SLL23]. clinical [MKMK23, TCCK23]. closed [Ver22, XGW23, ZKJ^{+}24]. closed-form [ZKJ^{+}24]. closed-loop [Ver22, XGW23]. clot [MML23]. clotFoam [MML23]. CLoTH [CVD21]. clothing [HGG20]. Cloud [WSRMB23, AGADAG^{+}23, BB22a, KPSM17, Krä20, MGK^{+}20, SSW^{+}24, ZCCGGV^{+}22, TSNRM24]. cloud-based [AGADAG^{+}23, Krä20, ZCCGGV^{+}22]. Cloud-native [WSRMB23]. Cloud2FEM [CLDdM22]. CloudEdgeAssetOptimizer [TSNRM24]. clouds [CLDdM22, HTB19, MIHS21]. CLUMP [ANOU21]. CLUSplus [PLK^{+}23]. Cluster [BLE21, CRV22, ÁG19, HMY24, JGGF23]. clustering [BHC^{+}24, Gag21, Gag22, JGGF23, KKAUA21, MSLNSMOG24, RMV^{+}24, SKD22, TOPC23, TS21]. clusters [KMT24, LHCK18, MHTZ22, vDPI^{+}18]. cmaRs [YÖYB23]. CMIP6 [AFGH22]. CMIP6-D&A [AFGH22]. CNAplot [HHN20]. Cnerator [OE21]. Cnet [RLJS23]. CNN [SEZLMS^{+}23]. CNNs [SH19]. cnpy [Rei23]. CNV [AJK^{+}23]. CNV-Z [AJK^{+}23]. Co [TT17, GCP22, PMMF19, VMS^{+}23, SNLFGVRB24]. co-expression [SNLFGVRB24]. Co-regulation [TT17]. co-simulation [PMMF19]. co-simulations [GCP22]. co-teacher [VMS^{+}23]. Coastal [BB20b, KP20, CAW^{+}20, MAS^{+}22, PL23, SP22, TŚG^{+}24, ŁMF^{+}23]. CoastalImageLib [MAS^{+}22]. coda [PDTG17]. coda-normalization [PDTG17]. CodaNorm [PDTG17]. Code [KSFG18, ACPdP22, AAF^{+}24a, ACC^{+}23, ABC^{+}20, Ars21, Azi23, BC22, Byk19, CA18, CBLI22, DSJ^{+}22, Ek16, GMF18, GM24, HMCA15, HR21, IZBT21, IIH^{+}23, KH19a, KDHG20, KWZ22, KRB^{+}20, LFR^{+}20, MVG^{+}24, MPMC21, MB21, NSLD16, OE21, VMS^{+}23, Rag17, RMM18, SEL^{+}16, SS17, TACH17, WLW23, WMM18, YG19, dSBSD17, ANOU21, DDT20, RMM18]. coded [NG23]. codegen [MB21]. codes [ABC^{+}20, BV19, DK19, KBB19, SF16, VPVNP22]. coding [BFNP24, KSFG18, TDG19]. coefficient [JAC21]. coefficients [Azi23, Hel22]. COFFEE [DFSW19]. CoFly [REK^{+}23]. Cognitive [KXK^{+}23]. coherent [DKL^{+}21]. cohesion [KALS24]. cohesive [ZRK19]. coils [MFQ^{+}21]. coincidence [EWT^{+}23, EWT^{+}24]. CoinTossX [JSG22]. cold [MTPHH18]. cold-neutron [MTPHH18]. Collaboration [HTB19]. collaborative [ACO21, GM24, AGG23]. collapse [STC^{+}18]. Collecting [AGG23, SPMGA^{+}23, YLS^{+}18]. collection [Bat19]. collective [Arc20]. colony [GCC20]. colored [DI22]. Colour [BGAGJGC23]. combine [PZKK21]. Combining [Nis20, HCD24]. combustion [SBJ^{+}23]. Comfort [TSCH20, BDQ^{+}22, TS20]. Command [CC20, SLL23, CWM^{+}21a]. command-line [SLL23]. commercial [CT24]. Commlib [PTS22]. common [AFGH22, MAS^{+}22]. Communication [TDG19, FRNW^{+}24, GPK^{+}21, ML20, PTT20, PTS22, SU23]. Communities [GPEM23]. community [KHM^{+}22]. COMP [BCD^{+}15]. compact [DJCL23]. companion [SOS19]. Comparison [CKC22, O’D21, SA17]. compatible [MBY22, ROMH22]. compensation [ZY23]. competitive [CK22b]. competitiveness [Kur23]. compilation [BC20, PI17]. compiled [GMF18]. compiler [CHP23, CCAC22, Kur23, NBCC19]. compiler-based [CCAC22]. compilers [NBCC19]. complete [ABB^{+}19, UdL20]. complex [Afr22, AZ17, BFNP24, Dan22, NBM^{+}19, NBS^{+}21, NSO21, SAC^{+}21, TC20, VRVAC23, WM21b]. compliant [AVB17, BFV18, Ek16, VBA^{+}16]. components [DW21]. composite [MC18, RLM^{+}24, SSC23, WMZH23]. composites [TCQOGM23, ZRK19]. composition [BWMS22]. composition-structure [BWMS22]. COMPOSTI [NL23]. Compound [dlSVBLdA^{+}17]. comprehensive [BBP^{+}18, FF24, GM24, LSOM18, MBC22, NHR23, SR17, ZS24]. compressible [GP22, Par22]. compression [GCW^{+}23, SGA^{+}16, WBH^{+}23, ASSS24]. compressively [SOS19]. CompSim [AZ17]. compuGUT [ME17]. computable [IIS18]. Computation [RFP^{+}24, Bat19, Gir21, Kul20, OC20, SHTE23]. Computational [BBD^{+}18, Bla21, Bla23, LLP^{+}24, SCG^{+}20, dSBSD17, AGHK21, CMdS^{+}23, LLK^{+}20, MBI20, RLK18, RSMW20, SH23]. Computationally [SL21]. computations [SR17, TLDM19]. compute [ABB^{+}19, BO19, EAMSR^{+}23, HLP^{+}19, HWST22, KH19a, KDHG20, NGOT24, PCdJSdFR^{+}23, CC21, MPX^{+}24]. computed [KPTW19, PFC^{+}18]. computer [Dan22, LHZT23, MTS^{+}18, NIY16, PBS^{+}24, RSLL^{+}20, SV23, SSC23, ZRBCI20, MML20, RLN21]. computer-aided [ZRBCI20]. computers [JLP23, PBL^{+}21]. Computing [MWJ15, HKM^{+}19, JM20, LFGVGDBD24, MAG24, MB23, MDL^{+}18, Nas20, PS19c, Sho24, SAA18, TSNRM24]. COMSOL [MBM22]. concentrated [YSM22]. concentrating [PNC23]. concepts [TWL22]. concrete [JRP23, dQRRBdSM20]. concurrent [KHJ^{+}23]. conduct [GSP^{+}17]. Conducting [AALW23]. ConfigILM [HCD24]. configurable [HCD24]. configuration [VC18]. configurational [RFP^{+}24]. configurations [YFVL24]. confocal [DN17]. ConfocalGN [DN17]. ConForce [RFP^{+}24]. conformal [Nas20]. conformer [LEK22]. conic [COG19]. connected [LHCK18, Nas20]. Connecting [BSVP20, ED22]. Connectivity [AGG23, MCC20, OCK^{+}23]. connector [PPBZ21]. Conservation [BGCS19]. constant [Ras19a]. constrained [GRD^{+}23, PMY^{+}24]. constraint [TSMT19]. constraints [BHHÖ22]. constructing [Cin24, FM24a, Ver22]. construction [KKN^{+}22, LPL^{+}24]. constructivist [Geo17]. consumption [DVSD22]. contact [OLRLB21]. container [KHH^{+}22, MFCSÁM20]. containerized [KPSM17]. containers [CCZ23]. contaminant [FHA17]. content [Agu23]. context [BCFJ24, Jan24]. ContHeart [FG20]. Contiki [ODE^{+}22]. Contiki-NG [ODE^{+}22]. continuation [NSLD16]. continuous [AMH23]. continuum [JWJ^{+}22, RAS19b]. Contour [RSMW20, HLP^{+}19, HWST22]. contract [Gór24]. contrast [TGS^{+}19]. Control [Bla21, Bla23, AISM21, BB22a, BST^{+}17, BE20, BHB^{+}21, CJGA23, CRF21, GRD^{+}23, GCBDDBF23, KKN^{+}22, NCGLRR24, PMS23, PMMF19, SE23, VJA^{+}18]. control-oriented [CJGA23]. Controlled [KL22, GCW^{+}23, OE21]. controls [KHJ^{+}23]. convection [AG21]. Conversational [PMBM^{+}23]. conversations [FSMCFP23]. Conversion [BFRK19, DBJ19, HKZ24, Nai17]. convert [BRF^{+}23, GWBF23]. converter [SMD23]. Converting [CCTCGCOB23]. Convex [CCBC^{+}21, CK21, MP18]. Convolution [PM23]. Convolutional [CCTCGCOB23, DA24]. cooling [LKK23]. coordinates [AT22, LH22]. coordination [Amo21]. Copatrec [KF23]. coppeCosenzaR [TCPC22]. CoppeliaSim [BSVP20]. copula [Cob21]. copy [AJK^{+}23, HHN20]. core [HOM^{+}18]. coronary [MCK^{+}23]. corpora [CJA23]. corporate [LPR21]. corpus [Kwa19]. correct [Liu23]. corrected [THSW24]. Correction [CHL^{+}19, SY20]. corrections [GRW^{+}19, SLL23]. correlated [Nis20]. Correlation [DEV^{+}21, FNX22, CPB^{+}23, GCdJAURO21, KF23, LZN21, OAF20, PR19, PMLM23, RCT23, TACH17, UVPB^{+}22, CWLG^{+}21, ddSNX22]. correspondence [LRPD18, Alb15]. Corrigendum [CK23, EWT^{+}24, JP24, RNK22]. Cortado [SvZvdA23]. cosima [FRNW^{+}24]. cosima-mango [FRNW^{+}24]. Cost [ICS^{+}24, Alb19, DEV^{+}21, MTJ^{+}22, REFB17, SP22, TSNRM24, WCW^{+}23]. cost- [MTJ^{+}22]. Cost-Effective [ICS^{+}24]. cost-surfaces [Alb19]. costly [AL16]. costs [HWT^{+}23]. CoTranslate [CJA23]. Couette [LFR^{+}20]. country [PP19]. coupled [CVV23, PRSS19, Roa23]. coupler [VRG19]. Coupling [RDH^{+}21, HS20, PBS^{+}24]. covariate [ACNF22]. cover [WRV23]. coverage [KSP19]. covert [CSZM22]. COVID [EBGCT21, GPK^{+}21, IVSV23]. COVID-19 [EBGCT21, GPK^{+}21, IVSV23]. Covira [GPK^{+}21]. CoZ [AKM20]. cp [Zie19]. cp-tools [Zie19]. CPL1.0 [DK19]. CPR [BGAGJGC23]. CPU [ACPdP22]. CPU-GPU [ACPdP22]. crack [DRPS21, DBPS22, SS17, ZWM21]. CrackDect [DRPS21, DBPS22]. CRAPPY [CWM^{+}21a, WCW^{+}23]. crawling [KF17]. create [LPG^{+}23, TSGOC^{+}22]. creating [BFNP24, Gök21, LHB24, MAS^{+}22, NK18, PAK22]. creation [GBW^{+}24, LCMD^{+}18, Mül23]. creativity [AOGC^{+}20]. creativity-stimulating [AOGC^{+}20]. Creator [BS19]. crisp [WS24a]. CRISPR [KSP19]. Crispyn [BW22, BWS23]. CRISSPAC [MCK^{+}23]. CristalX [CWM^{+}21b]. criteria [AMA19, BW22, BWS23, KSS23, MSDP23, PZKK21, RJDKJ24, SKS23]. crop [AMC17, ASVV23, CPB^{+}23, VFFS23]. Cross [AZ17, CPV^{+}20, Ohn21, PR19, SV23, Zhu15, ZLLT23]. cross-correlation [PR19]. cross-platform [CPV^{+}20, SV23, Zhu15]. cross-project [ZLLT23]. crossbar [JM20]. crosslinked [HA23]. crossplatform [CC21]. crowd [AKM20]. crowd-powered [AKM20]. crowdsourcing [CJA23]. cryogenic [MSK24]. CryptoAnalytics [DFS24]. Cryptocoins [DFS24]. crystalline [Zie19]. CSM [dARPH^{+}19]. CSSIAR [dlSVBLdA^{+}17]. CSTNU [Pos22]. CT [BEPB24, KPN^{+}18]. CU [CCE21, WKR^{+}20]. CU-BENs [WKR^{+}20]. CU-MSDSp [CCE21]. Cube [ÇA23].

CUDA [GÓ19, MCBM24, RS22, TCH^{+}22]. CUDA-accelerated [MCBM24]. CulSim [Ull16]. cultural [Ull16]. curating [WTZ^{+}21]. Current [BBJ^{+}18, NCGLRR24]. currents [JM20]. curvature [DM19]. curve [GZ21a, MKB^{+}18, PI17]. curves [MM16, PDH16]. curvilinear [LH22]. cuSten [GÓ19]. CUSTOMHyS [CDAOB^{+}20]. customising [CDAOB^{+}20]. Cutting [TTHZ23]. CVIK [JGGF23]. cyber [WWB^{+}23, PTS22]. cyber-physical [WWB^{+}23, PTS22]. CyberGIS [MPX^{+}24, WLP16]. CyberGIS-Compute [MPX^{+}24]. cybersecurity [LOM^{+}23]. CycFlowDec [BS21]. cycle [SEL^{+}16]. cycles [BS21]. CyEnGNet [SNLFGVRB24]. CyEnGNet-App [SNLFGVRB24]. cylindrical [AT22]. cytometry [AAL^{+}22]. Cytoscape [SNLFGVRB24].

D [BB20a, BBF^{+}19, DEV^{+}21, KPN^{+}18, CC20, ACPdP22, AT22, BÇ23, BFSJP^{+}21, BSB^{+}23, BBN^{+}22, BRF^{+}23, BSV^{+}23, Can22b, EHGPT19, GRF23, GNPP23, HS20, Kar21, KDHG20, KWZ22, KPN^{+}18, KWB22, KL22, LCMD^{+}18, LMS^{+}16, MBC22, MLTF^{+}18, MBDS20, MASB18, MBI20, NS22, Ozb17, PO19, Ras20, RDO^{+}19, SAA18, SB21, THT20, ZRK19, Zek17]. D&A [AFGH22]. D-coupling [HS20]. D-O [VMRVF^{+}24]. D-O-II [VMRVF^{+}24]. Daily [BCFJ24, HM23]. Dam [SIC24, LQM22]. damage [MD19]. damages [SSC23]. DARS [AB24]. dashboard [BFNP24, CPCCL24]. dashboards [BFNP24, RLN21]. DASP [BBJ^{+}18]. Data [DZZ^{+}22, EPF^{+}23, PCL22, SB22, WFL^{+}23, AAA^{+}21, AFGH22, AJK^{+}23, ASAA20, AMA19, AVB17, ÁG19, ATT21, AAF^{+}24b, BFV18, BJM^{+}19, Bel24, BHE^{+}19, BGM^{+}24, BB23, BS19, BST^{+}17, BBG^{+}19, BAEBAS19, BE20, BHB^{+}21, BBF^{+}19, BBN^{+}22, BRF^{+}23, BdSvE23, CSK19, CSZM22, CdSLCC20, CPCCL24, CLM22, CMdS^{+}23, CBS^{+}16, CPB^{+}23, DRM20, DVSD22, DAB^{+}19, DRFB^{+}23, DKL^{+}21, EWF^{+}23, EFTS23, FRdN21, FSL16, GWBF23, GPW^{+}20, GM24, GCW^{+}23, GEH19, GJS^{+}22, HHH22, Hac23, HHB^{+}22a, HH21, HDA21, HHB^{+}22b, HT18, HMY24, HTWT23, HL22, ITL^{+}23, JGGF23, KWB22, KPM^{+}22, KBB19, Krä20, KGK^{+}20, Kul20, Kur23, LMB^{+}19, LEFSO20, LBT^{+}21, LRPD18, LIZ^{+}20, LF15, LMS^{+}16, LBH^{+}20, LVCF24, MGGHG^{+}24, MLTF^{+}18, MSK24, MTPHH18, MMG19, MGMAG22, MLGBR23, NKHZ21, NM20, NRC23, OHO16, OLRLB21, PI17, PLC^{+}24]. data [QTL23, RGP^{+}22, Rub16, RKA^{+}23, STH^{+}21, SGP23, SKD22, SMZSW24, SKM19, SGA^{+}16, SA20, SDCA19, SV23, SPMGA^{+}23, SC19, SH19, SHK^{+}23, Sur20, ST21, Tag23, TBCG21, TTT19, UdL20, Ver22, VCLS21, WSRMB23, WLP16, WGB16, Web17, Web21, Wet20, XZL24, YLS^{+}18, YÖYB23, ZEADR23, ZPM23, dSBSvE23, ASAA20, CCTCGCOB23, CCG^{+}23, GABH22, GAE23, HHH22, RIHVP23, TT17]. data-centric [SKD22]. data-driven [BE20, STH^{+}21]. database [AMVB19, BSS^{+}23, BdSvE23, FHB^{+}21, HESH19, PMBM^{+}23, SMD23, TT17]. databases [FWJ^{+}23]. DataCollector [SPMGA^{+}23]. dataflow [SCG^{+}20]. Dataset [DVSD22, ACRSGJ23, GACS^{+}23, MKMK23, NG23, ÖA23, PMP16, SDL21, SMR22, ZWM21, dHGMCGL^{+}24]. datasets [Cin24, FAM^{+}20, GZ21a, GZ21b, MGK^{+}20, NNDS23, NSO21, PAK22, SEdCCCRG23, Vel23, WRV23]. datatypes [FCMTV24]. DATTES [RIHVP23]. Davos [FM24a]. DBH [MIHS21]. DCEM [SKD22]. dcor [RCT23]. DCR [Azi23]. DDRS4PALS [PS19a]. deal.II [SGBH18]. deal2lkit [SGBH18]. debugger [SDG^{+}24]. DEC [BJM23b]. December [Ano19d, Ano20d, Ano21a, Ano22a, Ano23a].

Decentralized [AB24]. Decidim.viz [CPCCL24]. decision [AMA19, DW20, DER^{+}23, HWT^{+}23, KSS23, MSDP23, MFCSÁM20, MMS21, MKZ24, PLK^{+}23, PZKK21, RJDKJ24, SKS23, TCPC22, WKS22, WKS23, WS24a, WKCS24, WS24b]. decision-making [KSS23, SKS23, WKS23, WS24a, WKCS24]. DecisionMaster [PZKK21]. Decisions [tHLMN19, LMN18]. decoding [ZVP22]. DECOLib [DW21]. Decomp2d [OCK22]. decomposing [BS21]. decomposition [COK24, DSI^{+}20, OCK22, VMRVF^{+}24, DW21, NM20]. decomposition-based [VMRVF^{+}24]. Deconvolution [YC19]. Decremental [LHCK18]. dedicated [SvZvdA23]. Deep [ALRM21, FVA^{+}20, SFF^{+}19, CJGA23, GFRMVL24, Jun21, KG24, LVK21, LLP^{+}24, LPL^{+}24, MP17, MHTZ22, NGOT24, TK23, Tem23, WMA24, ZPM23, dPPD^{+}23, vKMH^{+}20, APK^{+}23, CHP23, CCI^{+}22, FLR22, MAM^{+}23]. deep-sea [MP17]. deep-temporal [ZPM23]. Deep-water [FVA^{+}20]. deepfake [EAFJ^{+}22]. Deepness [APK^{+}23]. DeepSR [Tem23]. defect [PAK22, ZLLT23]. defenses [MSS^{+}23]. defined [AGEABD^{+}24]. DeFinetti [Arc20]. deformable [HGS17]. deformation [BÇ23]. deformations [MD19]. DEIP [Tru18]. delay [KKN^{+}22, PR19, TSNRM24]. delayed [RMMG21]. delineation [PNL^{+}21]. delivery [KPSM17, LCGA22]. demand [MW23]. DEMGPU [GWK16]. democracy [CPCCL24]. democratizing [MPX^{+}24]. denoising [GRF23]. DENSER [ALRM21]. densities [DRPS21, DBPS22]. densitometric [PFC^{+}18]. Density [WFD23, FJ22, FS19, FPI^{+}22, LSOM18, LF15, LEK22, SBCK17, SLWS^{+}17]. density-functional [SLWS^{+}17]. Dental [FSAD23a, FSAD23b]. dependencies [MGCS23]. dependent [Alb19, RS22, WLW23]. deployable [AMH23]. deploying [LHB24]. deployment [GGI23, PBM22]. depth [HHN20, MBY22, RPV^{+}20, CCG^{+}23]. derivation [MMD24]. derivatives [RJN^{+}23, Tak22]. derived [ZAPS20]. desalination [ND22]. describe [BKM22]. deserialization [DRM20]. Design [BMR17, KP20, AL16, CLK21, DBJ19, EL17, EBL24, Fru21, HA19, HX22, KXK^{+}23, WS24a, YBVC^{+}22]. Designer [YDN21, YD20]. designing [NK18]. Designs [BV19]. Desktop [MSLNSMOG24, OCK^{+}23, KCP^{+}24]. Destination [HMY24]. detailed [DDCSH23, LS23, SEL^{+}16]. Detecting [DRPS21, AJK^{+}23, ASGC^{+}24, MIHS21, PP19, DBPS22]. Detection [Fra22b, KCNJ24, BSB20, BJM23a, ÇA23, EAFJ^{+}22, HMY24, Jur23a, LZN21, LIZ^{+}20, MTS^{+}24, NOÖÇ19, RMM23, RMMG21, SHvW20, SG24, SNHS20, XYC22, ZWM21, dRB^{+}20, STH^{+}21]. detector [PS18a, PS18b, PS18c, PS19b, YLM^{+}24]. detector-output [PS18a, PS18b, PS18c, PS19b]. detectors [NIY16, RAL^{+}20]. deterioration [BB23, HLG22]. determination [CLK21, YHC^{+}22]. determine [MAAOZMAM22]. determining [AAN17, BW22, BWS23, DSI^{+}20, MMCKK21]. deterministic [MKU22, AM23, MMS21]. developing [CJGA23]. Development [JRP23, dQRRBdSM20, SFF^{+}19, WMZH23, ASRI22, Bjö19, Ek16, GM24, HVCS^{+}23, KHH^{+}22, KFA^{+}24, LHB24, SUM21]. developments [LSOM18]. device [BLM^{+}22]. devices [ED22, HSMF22, ODE^{+}22, Sur20]. DEVSimPy [CS21]. DEVSimPy-mob [CS21]. DfAnalyzer [SCG^{+}20]. DFT [EAFJ^{+}22]. DFT-MF [EAFJ^{+}22]. diagnosis [CPD^{+}20, DER^{+}23, ZCCGGV^{+}22, CPD^{+}20]. diagonalization [Kri22]. diagram [MKZ24]. Diamond [KWB22]. Diary [RMM18]. DIC [DEV^{+}21, OAF20]. DICOM [FM24b, FM24b, JFJM22]. dielectric [MPAK19]. DIETERpy [GMKRS21]. difference [GÓ19]. different [BDPZ19, MD22, MSS^{+}23, PZKK21]. differentiable [SACT23]. Differential [PLGHG^{+}23, DFSW19, DER^{+}23, TTT19, VPVNP22]. differentiation [FSMCFP23]. diffraction [Nev17]. diffractive [DBJ19]. diffusion [AG21, GB24]. Digestion [AS18, AMAE24]. Digital [CPF23, DEV^{+}21, AAL^{+}22, AI21, BHB^{+}21, CPCCL24, LS23, OAF20, OYW^{+}21, PMZ21, TACH17, ZRBCI20, CWLG^{+}21, FNX22, MBA^{+}22, ddSNX22]. digital-twin-building [ZRBCI20]. DIGNEA [MSLH23]. dimensional [AI21, BÁP^{+}22, Elk22, KS21, Kur23, LKSS20, MP18, NL23, NR23, PFC^{+}18, RRS18, SA20, SL21, SHK^{+}23, YMCB23, ZAPS20, ZNS17, CCTCGCOB23]. dimensionality [GMNG^{+}18, MSV23]. dimensions [BC22, MKMK23]. direct [BE20, LFR^{+}20, NCGLRR24]. directed [BMBR21]. directional [TKLG19]. direpack [MSV23]. disambiguation [KDRT23, KK24]. discontinuous [CBLI22, SWK19, Tru18, VPVNP22]. discover [HTWT23]. Discovering [MVG^{+}24, SAC^{+}21]. Discovery [CKM21, CCC^{+}21, KRA23, LBH^{+}20, SvZvdA23, ZS24]. Discrete [AKC23, HGS17, BJM23b, CS21, CNST20, DW21, KDHG20, LS16]. discrete-event [CS21]. discriminative [PLGHG^{+}23]. discriminatory [MSLH23]. disease [DER^{+}23, MCK^{+}23]. diseases [MTS^{+}24, MAM^{+}23, SB22, TBO24]. Dispatch [GMKRS21]. dispatching [Bac21]. displacements [YSM22]. display [TCCK23]. displaying [Tag23]. disposal [GBJ23]. dissipation [OHO16]. distance [MP18, RCT23]. distances [LS16, MCBM24]. distributed [AM23, GNPP23, HT18, MHTZ22, MGK^{+}20, RGdlTQ^{+}23, TK23, MWJ15]. distributing [BFNP24]. distribution [Byk19, EL17, LS16]. Distributions [BO19, Fru21, MMD24]. districts [SUM21]. disturbances [HX24]. diverse [MSLH23]. diversity [Ull16]. DLAE [SFF^{+}19]. DLDiagnosis [MAM^{+}23]. DLTPulseGenerator [PS18a, PS18b, PS18c, PS19b]. DM [AMA19]. DM-MCDA [AMA19]. DNA [HM22, KSFG18]. DNN [FLR22]. DNN-Tuner [FLR22]. Docker [CCZ23]. DockerChannel [CCZ23]. docking [RJH^{+}20]. document [IIS18, AI21]. documentation [RMM18]. documents [AI21, She19]. Dodona [VMS^{+}23]. Dolos [MVG^{+}24]. Domain [FdCBV23, MT19, BCR^{+}18, BÁP^{+}22, HSMF22, Kri22, VRBM16, dCWL^{+}23]. domain-specific [Kri22]. domains [GEG22, MSLH23, MSS^{+}23, Nas20]. Dominican [BGCS19]. Doppler [EWT^{+}24, EWT^{+}23]. dotCall64 [GMF18]. double [JSB20]. download [HM22]. Downloader [PMP16]. downloading [BdSvE23, PMP16]. DPClusSBO [KKAUA21]. DPIVSoft [ACPdP22]. DPIVSoft-OpenCL [ACPdP22]. dplbnDE [PLGHG^{+}23]. DPsim [MVRM19]. DQSEGDB [FHB^{+}21]. draft [GBW^{+}24]. drafting [HGG20]. dream [EDZ^{+}23]. Dreamento [EDZ^{+}23]. drift [BSG20, SG24]. driven [BE20, EWF^{+}23, NR16, STH^{+}21, SL20a, SB22]. driving [AA23, AGADAG^{+}23, PCPC24]. DRM [NIP^{+}23]. DRMSimulator [NIP^{+}23]. drone [TK23]. drones [SAS^{+}22]. DroneTimeline [SAS^{+}22]. Drop [CFA22]. DropPy [OLRLB21]. DRS4 [PS19a]. DRT [DZZ^{+}22]. drtsans [HHB^{+}22b]. drug [LIZ^{+}20]. Dry [NIP^{+}23]. DSAIL [KEwMO24]. DSAIL-TreeVision [KEwMO24]. DSL [CBPC23, CVAP22]. DSL-based [CBPC23]. DSN [Ver22]. DSSAT [dARPH^{+}19]. DSSAT-CSM [dARPH^{+}19]. DssPyLib [GNPP23]. DTT [AKC23]. dtwParallel [EAMSR^{+}23]. dual [MGMAG22]. due [PRSS19]. Duenna [HBS16]. duo [Els23]. Durham [BBJ^{+}18]. during [Dan20, JAC21, MFJ23]. Dutch [vdOJP^{+}20]. Dymium [SR20]. DYNAblast [dFB22]. DynABlock 2D [Por23]. Dynameta [SCK^{+}23]. Dynamic [Fra22a, MMS21, AES^{+}22, AS17, CA18, Dra23, EAMSR^{+}23, GZR^{+}19, LHCK18, MKU22, MVRM19, SAA18, SCK^{+}23]. dynamical [Dan22, KHG21, PKA^{+}22]. dynamics [AMC17, Arc20, Ben15, BCTMA^{+}23, CVV23, GP22, IHT23, IIS18, KMK^{+}21, iNKN^{+}20, OT16, Por23, PBL^{+}21, Ras19a, RS22, RWJ^{+}21, Sal16, SFK^{+}19, THG^{+}23, VRG19]. DynaProg [MMS21]. dynmix [Dra23]. Dyssol [SDH20].

E-ARCP [AGG23]. e-learning [MSLNSMOG24]. E-SCORE [HMC^{+}24]. eadf [SDP^{+}20]. EaRL [EL20]. Early [HBA^{+}20]. Earth [GKK^{+}23, WFL^{+}23, ZRK21, ZRK22]. Earthquake [EL20, AANA19, ÁG19]. Eastman [DK19, RE22]. easy [CA18, DSJ^{+}22, EPF^{+}22, PTS22, PLC^{+}24, PMP16, TS22]. easy-to-use [CA18, DSJ^{+}22, PTS22]. EasyEIS [BDSC22]. EASYMORE [GKK^{+}23]. EBSDPolygonizer [Liu24]. EC [SHTE23]. EC-KitY [SHTE23]. ECG [Ozk23, PNL^{+}21]. ECGdeli [PNL^{+}21]. Eclipse [BBD^{+}18]. EcoDynElec [LPG^{+}23]. ecological [Sha23, SCK^{+}23]. EcoMind [VCA^{+}24]. Economic [AS18, GSP^{+}17, JMT22]. Economies [Giu19]. Ecopath [SSSH16, SCPC18]. Ecosampler [SCPC18]. Ecosim [SSSH16, SCPC18]. ecosystem [CCO21, KHH^{+}22, KP23, SO23a]. ECPC [SMD^{+}24]. Eddy [vdOJP^{+}20]. Eddylicious [ML18]. edgar [LPR21]. edge [MHTZ22, BSB^{+}23, MB23, TSNRM24]. EdgeCloud [KMT24]. editor [OPPZ22, CMDDARG23]. Editorial [SWK15, Ano15a, Ano15b, Ano16, Ano17, Ano18, Ano19c, Ano19a, Ano19b, Ano20a, Ano20b, Ano21b, Ano21c, Ano21d, Ano21e, Ano22b, Ano22c, Ano22d, Ano22e, Ano23b, Ano23c, Ano23d, Ano23e, Ano24a, Ano24b]. education [WLP16]. educational [GP22, MVG^{+}24, SVM^{+}23, SDG^{+}24, MDR^{+}22]. EEG [DZZ^{+}22, CGHGRB21, CRV22, DZZ^{+}22, EDZ^{+}23, ZVP22]. EEG/MEG [ZVP22]. eeglib [CGHGRB21]. effect [HLG22, JAC21, KCP^{+}24, O’D21, WW17, O’D21]. Effective [ICS^{+}24, BDPZ19, KDHG20]. effects [Kul20, SAA18]. efficiency [LKK24]. Efficient [BO19, GKM20, Hel22, KBB19, AHP23, Ars21, CMACMÁT23, CLK21, GMF18, GM24, KWZ22, KGF^{+}24, LIZ^{+}20, LYZ^{+}22, SL18, SLL23, SPG^{+}23, SL21, TKLG19, ZKJ^{+}24, ZS22, WSK22]. Efficiently [WTZ^{+}21, EAMSR^{+}23]. effortless [LPL^{+}24]. Effusion [CRF21]. Eigenpairs [AGH20]. EKF [PG18]. EKF-AUS-NL [PG18]. elastic [RCT20]. elasticity [SBL19]. ElasticMatrix [RCT20]. elastohydrodynamic [YMCB23]. elastomers [MD19]. ELC [ND22]. electric [DHK^{+}20, PMS23]. electrical [XGW23, LYX^{+}18]. Electricity [PZ22, GWBF23, LPG^{+}23, WM21a]. electrocardiogram [ZCCGGV^{+}22]. Electrocardiograms [KCNJ24]. electrochemical [BDSC22, ND22, VCLS21]. Electrode [CGKD23]. electroglottogram [TJS18, TJS19]. Electromagnetic [KL22]. electromagnetics [ILP^{+}24, VRG19]. electromechanical [TCQOGM23]. Electron [SC19, GRF23, MSK24, SBCK17, vKH20]. electron-matter [vKH20]. electronic [MB21, RRSK18]. electronics [JZGW22]. electrophysiology [AAF^{+}24b]. Element [MP20, Afr22, BÁP^{+}22, CLDdM22, DHK^{+}20, FS23, HGS17, KSA23, MMH19, ND22, PMM16, RNK21, RNK22, Tru18, WKR^{+}20, ZRK19, ZMS18]. element-boundary [BÁP^{+}22]. elements [DBJ19, GTC21, MMH19]. Elephant [Jur23a]. Elfun18 [Bat19]. ELICIPY [dMVBTN24]. elicitation [dMVBTN24]. ellipsometry [RGP^{+}22]. elliptic [Bat19]. Elmer [VRG19]. Embedded [NCGLRR24, OT16, WCW^{+}23, ZRK19]. embedding [MLD22]. emergence [Ull16]. Emipy [OBWM23]. EmiR [PS22]. emissions [OBWM23]. emotions [ASGC^{+}24]. emphasis [VEM^{+}18]. EMPIRE [BSdG^{+}22]. empirical [BBP^{+}18]. employing [DET23, RSMW20]. emulated [ED22]. EMulatOr [VDP^{+}19]. enable [CAW^{+}20]. enabled [PMS23, TGS^{+}19]. Enabling [GY23, Liu24, PFC^{+}18, LGP^{+}22, PGA^{+}20]. Encryption [CHP23]. End [Bed24a, DVSD22, WRV23]. end-to-end [WRV23]. endemism [GL15]. Endogenous [GMKRS21]. Energy [ICS^{+}24, LFT21, BSdG^{+}22, BBP^{+}18, BS19, CK22b, FS19, FPI^{+}22, Gök21, GCP22, MWM20, MTPHH18, MTJ^{+}22, RCT23, RIHVP23, RERES23, SUM21, SO21, WM21b, WWB^{+}23, YLM^{+}24, ZNS17, GPEM23, HSMF22]. energysim [GCP22]. Engine [HKM^{+}19, GLS^{+}22, GVAO19, JSG22, SFF^{+}19, dlTAJCÁBG24, MSK24]. Engineering [BFF^{+}21, SCG^{+}20, AANA19, CR22, EDZ^{+}23, HMC^{+}24, NSS^{+}19, ZRBCI20, dSM22]. engineers [AGM^{+}23]. Enhanced [EAFJ^{+}22, AAF^{+}24a, JWJ^{+}22]. Enhancing [APLI24]. EnMAP [JJT^{+}23]. EnMAP-Box [JJT^{+}23]. ennemi [LZN21]. ensemble [KOC21, Kle24, SNLFGVRB24]. ENTIRETY [HSMF22]. entities [KDRT23, ST21]. entropy [BDPZ19, GZ21b, KHG21, SR19]. entropy-based [GZ21b]. Envelopment [GAE23]. Environment [AGHK21, BFF^{+}21, Fra22a, KGF^{+}24, LMB^{+}19, LKSS20, SDCA19, WKS23, WKCS24, YBVC^{+}22, BBD^{+}18, MYKT22]. environmental [CBHLG21, HLP^{+}19, HWST22, LPG^{+}23, SO23a]. environments [Cha17, CT24, KPSM17, MGMAG22, SE23, WS24a]. EODIE [WFL^{+}23]. EOF [VRG19]. EOF-Library [VRG19]. EPCIS [AGL^{+}23]. epilepsy [GAM^{+}24]. epitaxial [Dan20]. EPT [KOC21]. EpyNN [MDR^{+}22]. EQP [MFCV20]. equation [Azi23, DHK^{+}20, GEG22, GNPP23, KSS20, Kor21, MD21, SWK19, TTT19, LH22, YMCB23]. Equations [Can22b, AK24, AG21, DFSW19, KPOD16, SL18, SAA^{+}23a, VPVNP22]. equilibria [NBM^{+}19, NBS^{+}21]. equilibrium [SP19]. equivalent [NMT23]. era [WLP16]. erasure [Ars21]. eRDF [SBCK17]. Ergo [RRSK18]. ergodicity [MWM20]. Error [CHL^{+}19, GCW^{+}23, PG18]. error-controlled [GCW^{+}23]. errors [Liu23]. ESP3 [LEFSO20]. EspyInsideFunction.jl [Mai22]. essential [DER^{+}23]. Establishing [MKS^{+}23]. estate [HESH19]. estimate [CCFSB20, HWT^{+}23, Nis20]. estimates [EB24, Nis20]. estimating [BSS^{+}23, HLG22, KSP19, MIHS21, PMLM23]. Estimation [HR21, ACEAV24, CAG^{+}24, Dra23, FJ22, FRdN21, JAC21, KHG21, MVBF19, MAGM^{+}23, NMT23, SO23a, VBA^{+}16, ZTC^{+}21, BL16]. estimators [ZKJ^{+}24]. ETCAL [KH19b]. ETF [KCP^{+}24]. ETF-RI-CEG-Advanced [KCP^{+}24]. ethnicity [Xie22]. ETL [ACO21]. Euclidean [Kis20]. Eulerian [LXJ^{+}23]. Europe [OBWM23, PZ22]. EvalNE [MLD22]. evaluating [AES^{+}22, CCZ23, FHA17, RKDP21, MSS^{+}23]. Evaluation [GMKRS21, HBA^{+}20, KNW^{+}23, LMN^{+}20, SIC24, STC23, CGS19, FMFCEG^{+}23, FF24, GSP^{+}17, GTG21, HL22, JC23, MLD22, MBC22, MPAK19, PS19a, SAA^{+}23b, SD23, Zaj20]. evapotranspiration [BTDT23, Tho22, TBD23]. event [CS21, CNST20, EB24, HKZ24, LSSK16, MML20]. event-based [LSSK16]. events [LS16]. EvoDyn [IIS18]. EvoDyn-3s [IIS18]. Evolution [TTT19, DFSW19, Kor21, TLDM19, ATT21, PLGHG^{+}23]. Evolutionary [ALRM21, IIS18, AGC^{+}24, PS22, SHTE23]. Evolver [AMDN23]. evolving [BJM23a]. Evoplex [COGP19]. EVOVAQ [AGC^{+}24]. exact [AGG23, Kar21, Kri22, TF23]. ExactPack [TF23]. Examination [TMGW23]. Exasim [VPVNP22]. excitations [Taq16]. excited [KLY21]. exclusion [MMD24]. Executable [BKM21, IKYY22]. execution [BB22a, BCD22, CHP23, DM20, GGI23, Gór24]. exercises [SVM^{+}23, TTHZ23]. exfiltration [CSZM22]. existing [CLDdM22, CAW^{+}20, LS23]. existing/historical [CLDdM22]. exome [CdSLCC20]. Expansion [RAL22, WTAG^{+}22, BBB21, TC20]. Expectation [SKD22]. experience [APLI24, BCFJ24]. Experiences [CMDDARG23]. experiencing [TWL22]. experiment [ACC^{+}23, BST^{+}17, SSM22, WGB16, Web17, Web21]. experimental [BHE^{+}19, CWM^{+}21a, HBS16, MKR^{+}21, WCW^{+}23]. experimentally [CWM^{+}21b, PKA^{+}22]. experimentation [AMH23, CMACMÁT23, JC23, WMA24, dSBSvE23, dlTAJCÁBG24]. experiments [CLK21, EFTS23, MFJ23, MLGBR23, NK18, dPPD^{+}23]. expert [tHLMN19, LMN18, dMVBTN24, YD20, YDN21]. explainable [DSB^{+}23, PDS^{+}22]. explained [GABH22]. Explanation [CMDDARG23, SEdCCCRG23]. explanations [MKWS22]. explicit [BSG20, ZR19]. Exploiting [NNDS23]. exploration [AI21, GLGG22, KLL23, NR23]. exploratory [KRA23]. explore [Jan24, KSFG18]. Explorer [KOZ^{+}24, BBD^{+}22, CMACMÁT23]. Explorer.py [WM21b]. Exploring [SB22, MVG^{+}24]. exponential [Byk19, DET23]. exposure [BSG20]. expression [Agu23, CPD^{+}20, CKC23, SNLFGVRB24]. extendable [KH19b, PPBZ21]. extended [LEK22, SAA^{+}23a, TMGW23]. extending [IA17]. Extensible [Giu19, KPOD16, BFV18, BJM23a, GJS^{+}22, JP23, JP24, Jan24, PRC^{+}23, RERES23, VBA^{+}16, JZGW22]. Extension [KNW^{+}23, LFK^{+}24, AGADAG^{+}23, KW23, WBS21]. extensive [CMACMÁT23, YMA23]. extent [BNC24]. extents [GL15]. exterior [BJM23b]. external [JAC21, SF16]. externalities [PS19c]. extract [ABC^{+}20, MGMAG22]. extracting [BdSvE23, KEwMO24, Mai22]. Extraction [BFF^{+}20, HHH22, BO19, CGHGRB21, VVDV22, HKF^{+}20, LHL^{+}24, PRSS19, SKM19, SA20, WRV23, YG19]. extractor [Agu23, WFL^{+}23]. Extrapolation [KL22]. Extreme [AMH23]. extremely [CKC23]. extremes [HLP^{+}19, HWST22]. ExWave [SWK19]. eye [AALW23, KH19b]. eye-tracking [AALW23]. EyeMind [AALW23]. EZFF [KMK^{+}21].

FaaS [MB23]. Face [ZY23]. FacetModeller [LCMD^{+}18]. Facial [FSAD23a, SB22]. facies [JAMP23, JAMP23]. facilitate [EFTS23, HGG20, TBCG21]. Facilitating [AAF^{+}24b, CWM^{+}21b, MLGBR23]. factor [ESE^{+}23]. factorisation [LB22]. factorizations [Roa23]. factors [BFRK19, CBS^{+}16, SO23a]. Failure [NSS^{+}19]. FAIMS [BSRSC18]. fake [RMM23]. falling [RRS18]. FAO [BTDT23, Tho22, TBD23]. FAO-56 [BTDT23, Tho22, TBD23]. far [SDP^{+}20]. far-field [SDP^{+}20]. FARM [HTWT23]. fascicles [KH19a]. fashion [RBQ^{+}24]. Fast [ALRM21, BFDD^{+}22, CHL^{+}19, Gag21, NOÖÇ19, SM19, Tak22, AHR^{+}22, DHK^{+}20, KWB22, Kle24, MSK24, PO19, RHR^{+}23, Rag17, SLL23, WMA24, ALRM21]. Fast-DENSER [ALRM21]. fastmat [WSK22]. FastRealBoostBins [Kle24]. FastSPM [BRF^{+}23]. fatigue [SSC23]. fault [DK19, YB23]. FBG SiMul [PMM16]. fcc [OT16]. FD [MT19]. FDT [RMM23]. featsel [REFB17]. Feature [BFF^{+}20, YG19, BOP^{+}23, CGHGRB21, CMACMÁT23, Cha17, VVDV22, GWGB22, KPJ24, Kur21, MGK^{+}20, REFB17, SA20, SHK^{+}23, ZY23]. feature-rich [Cha17]. features [Ars21, KMLP^{+}23, LSB^{+}20, NGOT24, PMFÁM22, PDM22, SA17, Tag23]. featuring [AGADAG^{+}23, KHH^{+}22]. February [Ano23f, Ano24c]. Federated [KNW^{+}23, DSB^{+}23]. federations [HTB19]. Feedback [BE20]. FEFLOW [PRSS19]. FEM [RNK21, RNK22, GNPP23, KBB19, RFP^{+}24, VRG19]. FEniCS [DJCL23, RDH^{+}21, RAS19b]. FEniCS-preCICE [RDH^{+}21]. fermentation [ME17]. ferrofluid [JAC21]. ferrofluids [JAC20]. fetal [BlD^{+}20]. FeView [RNK21, RNK22]. FFA [MTS^{+}24]. FFA-Lens [MTS^{+}24]. FH [THSW24]. fiber [DRPS21, DBPS22, ZRK19]. fiber-reinforced [DRPS21, DBPS22]. Fibre [PMM16]. fibrous [YHC^{+}22]. fiducial [LKSS20]. FieLd [O’D21, CVS19, BSRSC18, BFRK19, BTMB21, HLR^{+}21, JAC21, KAS23, ML20, MFQ^{+}21, NHR23, OC20, Ras19a, RS22, RLN21, SDP^{+}20, BHCT16]. field-based [HLR^{+}21]. fields [JAC20, LD24, NR23]. FieldView [NR23]. filaments [MBM22]. File [BS19, SS19, DJCL23]. files [DBJ19, GZS23, Krä20, Liu23, Rei23, SC17, Ver22, VC18]. filings [LPR21]. film [LLP^{+}24]. films [Dan20, RRS18]. filtered [YC19]. Find plasma membrane [AHP23]. Finding [CKM21]. fine [GBJ23]. Finite [RNK21, RNK22, Afr22, AG21, BÁP^{+}22, CLDdM22, Dra23, FS23, GÓ19, KSA23, MMS21, MMH19, ND22, PMM16, Tru18, WKR^{+}20, ZRK19, ZMS18, MP20]. finite-element [ND22]. finite-volume [AG21]. FIPEX [OCK^{+}23]. fire [WSAC^{+}23]. first [AAA^{+}21]. fish [IAW^{+}15, OCK^{+}23]. fisheries [GSP^{+}17, NFASC21]. fit [MM16, YB23]. Fitting [SSSH16, DLH18, ENCS20, HDA21]. Fixed [MT19, LS16]. fixedTimeEvents [LS16]. FLAGR [AAB23]. flame [Car21, Car21]. Flash [DWO^{+}22]. Flash-X [DWO^{+}22]. FLBEIA [GSP^{+}17]. FleCSIble [LLK^{+}20]. FleCSPH [LLK^{+}20]. FleetRL [CT24]. fleets [CT24, IVSV23]. FleXbox [KPC^{+}20]. Flexible [BSRSC18, Afr22, AAB23, ASGC^{+}24, CCE21, Jan24, KPM^{+}22, Kos22, MSK24, NFASC21, SO21, SAA18, Sur20, VDP^{+}19, VVDV22]. Flickr [FAM^{+}20]. FLIP [HLR^{+}21]. flooding [BGM^{+}24, BNC24]. Floodsar [BNC24]. Flow [MBF20, NK20, BDPZ19, BS21, HKF^{+}20, JWJ^{+}22, KWZ22, KRB^{+}20, LH22, LD24, LXJ^{+}23, LFR^{+}20, MW23, MML23, NR16, Par22, SHH23, SE23, SH23, Jur23a, MBC22, Jur22]. Flow-models [Jur23a, Jur22]. FlowDenoising [GRF23]. Flower [KNW^{+}23]. flows [BBB^{+}20, Jur22, Jur23a, SL21, SRK23]. flowsheet [SDH20]. FLT [FSAD23a]. FLUBIO [AG21]. Fluid [HOM^{+}18, HS20, GP22, LDM20, SC23, VRG19]. Fluid-Structure [HOM^{+}18, SC23]. fluids [NBM^{+}19, NBS^{+}21]. fluorescein [MTS^{+}24]. fluorescence [HLR^{+}21, IIH^{+}23, LBG20, HLR^{+}21]. flux [OHO16]. FM [Elk22, LKSS20]. FM-2D [Elk22]. FM-Track [LKSS20]. FMC [ÁG19]. FMC-Earthquake [ÁG19]. FMI [VBA^{+}16]. fMRI [ADSG^{+}20, MCC20]. FNML [AGEABD^{+}24]. focal [ÁG19]. focus [SH23]. focused [ASSS24, Bel24]. Fog [MB23]. follow [EB24]. follow-up [EB24]. FonaDyn [TJS18, TJS19]. food [Bre23, BSS^{+}23]. forbidden [VV19]. force [BFSJP^{+}21, BFRK19, Ras19a, RS22]. forcefields [KMK^{+}21]. forces [RFP^{+}24]. forecasting [GTOM23, DFS24]. Forensic [SAS^{+}22]. forest [HWT^{+}23, MIHS21, ZR19, BGCS19]. ForestSim [ZR19]. form [GZR^{+}19, ZKJ^{+}24]. form-based [GZR^{+}19]. format [AFGH22, DJCL23, NSO21]. formation [MML23]. formatting [Liu23]. forms [GMNG^{+}18]. formulation [GEG22]. formulations [BSS^{+}23]. Fortran [Azi23, GMF18, Nai17, Taq16]. Forward [CHL^{+}19, SF16]. Founsure [Ars21]. Fourier [MT19, YC19]. FPGA [MB21, UdL20]. FPGA-based [MB21]. fracture [KAK21, KAS23]. frame [CRF21, JSB20]. Framework [Che21, ACR22, ABM^{+}21, AES^{+}22, AMVB19, AFA22, ASRI22, BCD^{+}15, BBD^{+}22, BDF^{+}20, BPL^{+}19, BK19, BT23, BBS^{+}22, CCZ23, CCC^{+}21, Can24, CBPC23, CM19, Cha17, CPV^{+}20, CPB^{+}23, CDOBA22, DAC^{+}24, Dan20, DAB^{+}19, DS20, FVA^{+}20, FPN^{+}22, GMKRS21, Gag22, GCC20, GQCP^{+}18, GPW^{+}20, GCW^{+}23, GWK16, GV20, HLG22, HGG20, HT18, HVCS^{+}23, HX22, HYR^{+}23, IHT23, JWJ^{+}22, JC23, Jur22, LSMG19, LYX^{+}18, LVCF24, MLD22, MB23, MGCS23, MCPG24, MML23, MD19, NR16, NHR23, NBCC19, PTT20, PBM22, PLK^{+}23, PPBZ21, RHR^{+}23, REFB17, RSMW20, SUM21, SAA18, SR20, SDH20, SL20b, SVM^{+}23, SMD^{+}24, TCH^{+}22, VCLS21, WHG^{+}23, XTD24, YLM^{+}24, ZWM21, ZMS18, dPPD^{+}23, KGK^{+}20]. frameworks [PPBZ21]. fraudulent [She19]. free [BBP^{+}18, GEG22, OC20, Rub16, TGS^{+}19, ZRK21, ZRK22]. frequencies [HHN20]. Frequency [BO19, BAEBAS19, MT19, JAC20, MRMD19, Ozb17, SOS19]. friction [KAK21]. Friendly [MCC20, HMK23, LYZ^{+}22, Ozk23, XTD24]. fringe [OYW^{+}21]. Frouros [SG24]. fruit [MGMAG22, MAGM^{+}23]. FS [CMACMÁT23]. FS-Studio [CMACMÁT23]. FTIR [CBHLG21]. full [DPD22, MGFO23, SF16, ZNS17]. full-spectrum [DPD22]. full-waveform [MGFO23]. Fully [PRSS19, BB20a]. function [Byk19, MFCV20, RHR^{+}23, SBCK17, Tak22, Zek17]. Functional [MCC20, WFD23, LSOM18, LEK22, SLWS^{+}17]. functionality [DRM20]. functionally [MP20]. functionals [BT21, LSOM18]. functions [AGEABD^{+}24, ABB^{+}19, Bat19, BT21, COG19, DW21, Mai22, REFB17]. Fundamental [TS21]. fusion [FPI^{+}22]. future [VEM^{+}18]. FuzConvSteganalysis [DA24]. fuzzy [DA24, CKC23, DGK^{+}24, KXK^{+}23, RJDKJ24, WS24a, WKCS24, WKS23].

Gabriel [Jur23b]. GABRIELE [Giu19]. GAGI [dlTAJCÁBG24]. galaxy [Duc23]. Galclaim [Duc23]. Galerkin [CBLI22, SWK19, VPVNP22]. galkin [PI17]. game [EHGPT19, FVVL23, KFA^{+}24, TWL22, HKM^{+}19, dlTAJCÁBG24]. Game-Engine-Assisted [HKM^{+}19]. games [CK22a, CK23, IIS18]. GAN [GLS^{+}22]. GANG [GLS^{+}22]. GANG-MAM [GLS^{+}22]. gas [EWF^{+}23, MFQ^{+}21, MTJ^{+}22]. gas-driven [EWF^{+}23]. gases [SLBC^{+}20]. gate [JLP23]. gate-based [JLP23]. Gated [O’D21]. Gaussian [DI22, VRP23]. GCPU OpticalFlow [CWGM24]. GDOESII [DBJ19]. GDSII [DBJ19]. GEARS [HKM^{+}19]. GenClass [ATT21]. GenConstraint [TSMT19]. gene [CKC23, HMK23, TT17]. General [NRC23, dlTAJCÁBG24, CGS19, CSK19, Cer19, HCD24, HA23, PMZ21, Giu19]. general-purpose [Cer19]. generalised [KAS23, LB22]. generalized [Hel22, MD21, MD22]. generate [ANOU21, GTC21, IAW^{+}15, JCFM24, KPN^{+}18, MSLH23, SSP21, SO21, TTHZ23, ZRK19, ZAPS20]. Generating [ST21, Bac21, MGGHG^{+}24, MBF20, MFQ^{+}21, NNDS23, VRVAC23, VPVNP22]. Generation [RAL22, WTAG^{+}22, AJK^{+}23, AS17, BBS^{+}22, Byk19, Can22b, CA18, DRM20, DRFB^{+}23, GZR^{+}19, GACS^{+}23, Gór24, HOB23, LH23, LLK^{+}20, MB21, ML18, NGK^{+}20, ODE^{+}22, OE21, PS19c, Rag17, Ras20, RL19, SMR22, Tru18, UVPB^{+}22, dHGMCGL^{+}24]. generative [GBW^{+}24]. generator [BB23, CLDdM22, DSJ^{+}22, DI22, DN17, Ek16, HR20, HA23, LS23, ÖA23, SHK^{+}23, LSF^{+}23, SMR22, GBW^{+}24, MBF20]. generic [BSB^{+}23]. Genetic [TCH^{+}22, CM19, HX22, KFA^{+}24, KSFG18, KSFG18]. genieclust [Gag21]. genomic [APO^{+}20, GQCP^{+}18]. genotype [SADU23]. Geo [ÇA20]. Geo-MST [ÇA20]. geocoding [MZSH21]. geocomputation [MPX^{+}24]. geodetic [BÇ23]. geographic [CLM22]. geographical [ÇA20]. geographically [LHY^{+}23, DMFG22]. geomagnetic [HX24]. geometric [BSB^{+}23, BJM23b, CM19, BOP^{+}23, TCH^{+}22]. geometrical [AZ17, LS23]. geometries [SRK23, TSGOC^{+}22]. Geometry [Ohn21, JAC20, Kis20, RNG^{+}21, SB21]. geomorphic [PL23]. Geophysics [KL22]. georeferenced [GL15]. GeoReVi [LBH^{+}20]. Georges [THG^{+}23]. GeoRocket [Krä20]. geosciences [Zek17]. Geospatial [LFK^{+}24, Krä20, LMB^{+}19, PCdJSdFR^{+}23, WLP16]. geovisualization [CLM22]. GeoWeightedModel [DMFG22]. German [JMT22]. GetSensorData [GJS^{+}22]. Gideon [LMM22]. GIN [Cin24]. GIS [CC20, SDCA19, VEF^{+}23]. Git [SS19]. GitHub [ÖA23]. given [Byk19, TSNRM24]. glass [JP23, JP24, WMM18]. glasses [BWMS22]. Glauber [LNS15]. global [GZS23, GLCBHTHO23, HMY24, HRC20, HX24, Ozb17, CC23, HX24]. Global-GMDs [HX24]. glow [PDH16]. GLYCAM06 [BFRK19]. glycoprotein [KZ18]. GMDs [HX24]. GMP [AS23]. GNLStools.py [MD22]. GNNkeras [PBSB22]. GNNs [BOP^{+}23]. GNSS [GZS23]. GO [TCQOGM23, DRM20]. goal [KBB19]. goal-oriented [KBB19]. goes [WCW^{+}23]. Golang [BB22a]. GOMC [NBM^{+}19, NBS^{+}21]. goniometry [OLRLB21]. GOOSE [HYR^{+}23]. GOPY [MBI20]. gosset [dSBSvE23]. governance [Kur23]. govErning [HSMF22]. government [CPF23]. GPFniCS [KAS23]. GPR [CKC23]. GPRlab [XZL24]. GPS [HMY24, YLS^{+}18]. GPU [NBS^{+}21, ACPdP22, CC21, CWGM24, CWLG^{+}21, DS20, GWK16, LAA23, NBM^{+}19, Nev17, NSLD16, WSAC^{+}23]. GPU-accelerated [DS20, Nev17]. GPU-based [WSAC^{+}23]. GPUCorrel [CWLG^{+}21]. GPUs [PO19]. Grace [Zhu15]. grade [PMFÁM22]. graded [MP20]. gradient [TFCZ24]. gradients [DSJ^{+}22]. grain [CWM^{+}21b]. grain-based [CWM^{+}21b]. Grammatical [TTT19, ATT21]. Graph [PCC^{+}19, PM23, COK24, FVD20, Jan24, JZGW22, Jur23b, KPKS22, LPS^{+}24, PBSB22, RJH^{+}20, BOP^{+}23, KPK23, PBSB22]. graph-based [JZGW22]. graphene [MBI20, O’D21, TCQOGM23]. graphene-based [MBI20]. grapher [KPS23]. graphic [JCFM24]. Graphical [ADSG^{+}20, ANA16, BHKK23, CPV^{+}20, FM24b, KCP^{+}24, MTS^{+}18, RtHLMN20, SMZSW24, SCSB24, SPMGA^{+}23, SSC23, WS24a, MM16]. graphics [HMCA15, MML20, SV23, VPVNP22, Zhu15]. graphs [BJM23a, DL16, FVD20, GLGG22, KKAUA21, KCP^{+}24, RL19]. GraSPI [JZGW22]. grating [PMM16]. Gravitational [CKM21, CCC^{+}21, DKL^{+}21, FHB^{+}21, MAC^{+}21, RAL^{+}20, Wet20, XTD24]. gravitational-wave [DKL^{+}21, MAC^{+}21, RAL^{+}20, Wet20]. gravity [MTS^{+}18]. greenhouse [MTJ^{+}22]. grid [BFV18, GWBF23, ROMH22, FWB^{+}21]. grids [BB20a, PKA^{+}22]. GridTool [GWBF23]. GridTools [ABM^{+}21]. GROMACS [AMS^{+}15, BFRK19]. ground [ACR22, XZL24]. ground-motions [ACR22]. groundwater [FHA17, JAMP23, PRSS19, SHH23]. group [FSMCFP23, LSF^{+}23, ZC20, Kur23]. groups [GACS^{+}23]. growth [CCG^{+}23, Dan20, CCG^{+}23]. GSD [COK24]. GSGP [CM19, TCH^{+}22]. GSGP-C [CM19]. GSGP-CUDA [TCH^{+}22]. GSimPy [ZC20]. GSITK [ASRI22]. Gsolve [MTS^{+}18]. GstLAL [CCC^{+}21]. GTO [APO^{+}20]. GTOPX [SNW^{+}21]. GUI [AFGH22, AANA19, BHC^{+}24, BBHP^{+}23, Bre23, BSS^{+}23, EWF^{+}23, FK23, GZR^{+}19, GCBDDBF23, Jun21, NK18, Ozk23, SHH23, SBCK17, SSC23, SLM23]. GUI-based [BBHP^{+}23, BSS^{+}23, Jun21, NK18, SLM23]. guide [ACNF22]. guided [CCBC^{+}21]. guided-wave [CCBC^{+}21]. GURU [MM16]. GVIDI [GACS^{+}23]. GWGPM [LSF^{+}23]. GWmodelS [LHY^{+}23]. GWpy [MAC^{+}21]. Gxsview [Ohn21]. Gym [SE23]. Gym-preCICE [SE23].

habitat [HJRBM23]. habitats [BSG20]. Half [BSB^{+}23]. Half-Edge [BSB^{+}23]. Han [MM16]. hand [Bed24b, KM21]. hand-held [Bed24b]. handle [VC18]. Handling [NG23, WKS23, WKCS24]. hardware [Bed24b, MDL^{+}18, WCW^{+}23]. harmonic [GEH19]. harmony [LSPG22]. harness [MDG22]. harnessing [FVVL23]. Harvester [TT17]. harvesting [FAM^{+}20]. Hausdorff [MCBM24]. HC [PTT20]. HC-4-PM [PTT20]. hdData360r [Kur23]. health [ASGC^{+}24, BB23, CCBC^{+}21, EHGPT19, KFA^{+}24, Ozk23]. hearing [KHM^{+}22]. heart [BlD^{+}20]. heat [JMT22, KRB^{+}20, MFJ23, SAA18, Zie19, JMT22]. Heatrapy [SAA18]. HEDSATS [FS19]. HEFactory [CHP23]. held [Bed24b]. Helium [Ras21]. Helmholtz [DHK^{+}20, GEG22]. help [Alb15]. HeNDS [Ras21]. Hermes [KPSM17]. Hessians [DSJ^{+}22]. heterarchical [PTT20]. heterogeneity [MDG22]. heterogeneous [MWJ15, PBSB22, Sur20]. HETOOL [BSB^{+}23]. heuristic [CDAOB^{+}20, CDOBA22]. HexagDLy [SH19]. hexagonally [SH19]. Hierarchical [PCC^{+}19, Gag21, TCPC22]. hierarchies [BKM22, GMNG^{+}18]. hierarchy [LPS^{+}24]. High [AMS^{+}15, DTD17, EPF^{+}23, KWB22, Kur21, OLRLB21, Afr22, AAB23, BBB^{+}20, CJA23, CZ22, DPD22, EWF^{+}23, FPI^{+}22, GQCP^{+}18, GPW^{+}20, GCW^{+}23, GWK16, JWJ^{+}22, JSG22, JAC20, JHAMF19, KH19a, Kur23, LFGVGDBD24, LFR^{+}20, MB21, MMH19, Par22, PZ22, SGBH18, SFG21, SWK19, Sho24, SA20, SHK^{+}23, FS19]. high-dimensional [Kur23]. High-level [DTD17]. high-order [BBB^{+}20, MMH19]. high-performance [AAB23, CZ22, GPW^{+}20, GCW^{+}23, JWJ^{+}22, LFGVGDBD24, LFR^{+}20, Par22, SFG21, Sho24]. high-quality [CJA23]. high-renewable [JHAMF19]. High-throughput [OLRLB21, DPD22, EWF^{+}23, GQCP^{+}18, JSG22]. Higher [GWGB22]. highly [KGF^{+}24, VDP^{+}19]. highRES [PZ22]. highRES-Europe [PZ22]. Higra [PCC^{+}19]. hikers [GACS^{+}23]. Hilbert [JS19]. histolab [MBA^{+}22]. Historical [MZSH21, CLDdM22, LPG^{+}23]. History [CKM21]. HiTIME [LIZ^{+}20]. HiVecMap [CLM22]. hms [SH23]. HMSE [PBS^{+}24]. HOFS [GWGB22]. hole [Taq16]. holographic [OYW^{+}21]. Home2 [ACRSGJ23]. homogeneous [MASB18, PBSB22]. Homomorphic [CHP23]. Hopkinson [GLR22]. Hopping [KGF^{+}24]. horizon [BSdG^{+}22, MMS21]. hospital [CNST20]. host [Duc23]. hot [ÇA23]. HPexome [CdSLCC20]. HS [LSPG22]. HS-Solver [LSPG22]. HSIToolbox [DZH^{+}23]. HTC [KPSM17]. HTML [BKM21, Nai17]. HTPolyNet [HA23]. huge [Tag23]. Human [RLN21]. Human-Computer [RLN21]. humanoid [MMS20]. humans [FVA^{+}20]. Hybrid [NMT23, DK19, KPSM17]. hydrate [WLW23]. hydrate-bearing [WLW23]. hydrate-saturation-dependent [WLW23]. hydraulic [EL17]. hydro [LEFSO20, PRSS19]. hydro-acoustic [LEFSO20]. hydro-mechanical [PRSS19]. hydrochemical [JAMP23]. Hydrodynamics [LQM22, LLK^{+}20]. hydrogen [EWF^{+}23, EFTS23, MTJ^{+}22, THSW24, EFTS23]. hydrographic [LSB^{+}20]. hYdrological [AMC17]. HYDROSCAPE [FHA17]. HYDRUS [PBS^{+}24]. HYDRUS-1D [PBS^{+}24]. HyPAT [EFTS23, EWF^{+}23]. hyper [CDAOB^{+}20, CDOBA22]. hyper-heuristic [CDAOB^{+}20, CDOBA22]. hyperbox [KG23, KG23]. hyperbox-based [KG23]. hyperbox-brain [KG23]. hypergraph [BMK23]. Hyperomet [SK22]. hyperparameter [HCS^{+}20]. hyperparameters [FLR22]. hyperspectral [CBHLG21, DZH^{+}23].

I-STL2MOOSE [SGP23]. iamxt [SRML17]. IASM [Has18]. IbIPP [IZBT21]. iBluff [PL23]. ICF [COG19]. iCorrVision [FNX22, ddSNX22]. iCorrVision-2D [ddSNX22]. iCorrVision-3D [FNX22]. ICU [STH^{+}21]. IDARC [ANA16, AANA19]. IDARC-2D [ANA16, AANA19]. ideal [GP22]. IdentiCyte [GMGG19]. identification [AHP23, AGdSC20, CR22, GMGG19, KS23, LRPD18, LDM20, PR19, RLJS23]. identify [Duc23, JMT22, Liu23, TBO24]. identities [OTdMP22]. IDOL [KK22]. IEC [HYR^{+}23, ROMH22]. IEEE [AVB17, BFV18, CPUARC20]. IFMCDM [RJDKJ24]. IGAOD [HX22]. II [ESE^{+}23, VMRVF^{+}24]. II-factor [ESE^{+}23]. IM [YB23]. IM-FIT [YB23]. IMage [LK21, ACPdP22, CWLG^{+}21, DEV^{+}21, FNX22, IZBT21, LD24, SEZLMS^{+}23, SO23b, ddSNX22, AVCP^{+}21, BGGL20, Car21, CCTCGCOB23, DN17, GRW^{+}19, GTG21, HCD24, JFJM22, Jun21, KPN^{+}18, KPTW19, KAO^{+}21, LRPD18, MAS^{+}22, MBF20, MCBM24, NGK^{+}20, NGOT24, OCK22, OAF20, PDM22, PPBZ21, RMM23, RKDP21, RSLL^{+}20, RJN^{+}23, SRML17, SS17, Tem23, ZY23]. Image-based [IZBT21, NGK^{+}20]. image-related [PPBZ21]. Image2DEM [LS23]. ImageJ [AHP23, THT20]. imagery [BNC24, LHZT23, SDL21]. Images [KAR^{+}19, BGAGJGC23, CBHLG21, CRP21, DZH^{+}23, DRPS21, DBPS22, HLR^{+}21, HKF^{+}20, KEwMO24, KAO^{+}21, LHB24, MTS^{+}24, NOÖÇ19, THT20, TKLG19, SO23b]. Imaging [AVCP^{+}21, BHCT16, BB20b, HLR^{+}21, KP20, PPBZ21, AAN17, BBD^{+}22, BHHÖ22, CSK19, CCG^{+}23, GZW^{+}22, HA19, LBG20, RSLL^{+}20, SFF^{+}19, SBL19, JJT^{+}23]. imbalance [RLF^{+}21]. imbalanced [GZ21a, GZ21b, ZGZvB19]. ImbTreeAUC [GZ21a]. ImbTreeEntropy [GZ21b]. imitate [SHK^{+}23]. immersive [SD23, SMD^{+}24, HVCS^{+}23]. immunoassays [VCLS21]. impact [CPF23, SMKL23]. impacts [LPG^{+}23, OCK^{+}23]. Impedance [LYX^{+}18, BDSC22, KG24]. implement [AFA22]. Implementation [MT19, AWO^{+}21, AMAE24, BHHÖ22, CKC23, KMLP^{+}23, LB22, MKU22, SA20, TKLG19, VJA^{+}18, WC23]. Implemented [Giu19, Kle24, PG18]. implementing [AI21, LVT^{+}22, MD22, RNR17]. improve [DW20, GACS^{+}23, MKS^{+}23, MCPG24, ZAPS20]. Improved [LNS15, MCBM24, APLI24, ACRSGJ23]. Improvement [SCD^{+}24, FM24b]. improvements [RtHLMN20]. impulse [FRdN21]. impulseest [FRdN21]. ImWIP [RJN^{+}23]. In-Cognitive [KXK^{+}23]. In-depth [RPV^{+}20]. in-lab [MHTZ22]. in-plane [FNX22, WMZH23, ddSNX22]. INACITY [dMOH21]. inclined [BFDD^{+}22]. Including [WCW^{+}23]. incompressible [BBB^{+}20]. inconsistently [NG23]. incremental [MAAOZMAM22]. Indel [DG21]. Indel-sensitive [DG21]. independence [KHH^{+}22]. independent [MDL^{+}18]. index [FWJ^{+}23, HKF^{+}20, JGGF23]. Indicator [BLE21]. indicators [Kur23, RLM^{+}24]. indices [BDQ^{+}22]. individual [NS22]. Individualized [Has18]. Indonesian [Kwa19]. induced [FRPC23, FRE^{+}24, Kul20]. inducing [FPI^{+}22]. Industrial [GABH22, OBWM23, SBT22, ZR19]. industries [NGOT24]. industry [PMFÁM22, RBQ^{+}24]. industry-grade [PMFÁM22]. inertial [CMMF19]. Infection [KJW21, SB22]. infectious [SB22]. inference [DGK^{+}24, SWCP20, SR17, ZL23, ZKJ^{+}24, AWO^{+}21]. inferences [MHTZ22]. inferential [Hac23]. inferring [Amo21]. inflow [ML18]. influence [PP19]. INFOBAR [ADSG^{+}20]. Informatics [HLW^{+}16]. Information [NK20, WFL^{+}23, ABC^{+}20, BDPZ19, CCZ23, CPF23, GGI23, GWGB22, GB24, HKZ24, Kur21, LZN21, LHL^{+}24, OTdMP22]. information-based [Kur21]. informative [MGMAG22]. informed [RWJ^{+}21]. infrastructure [LLK^{+}20, LS23, ZWM21, MWJ15, DDT20]. Infrastructure-as-Code [DDT20]. ingredient [Bre23]. IngredientDB [BSS^{+}23]. inherent [HHH22]. inhomogeneous [AZ17]. Initialization [IZBT21]. initio [XSS20, XSC^{+}21, ZJKS23]. injection [YB23]. Input [GPW^{+}20, AL16, FRdN21, MD22]. insect [VFFS23]. insects [CH23]. insertion [Tru18]. insertions [WBS21]. Insight [DPD22]. insights [HL22]. INSPECT [ANA16, AANA19]. INSPECT-PBEE [AANA19]. inspection [Bed24a, HHN20]. Inspired [dSM22]. instability [AT22]. Installer [MYKT22]. Instance [BCR^{+}18, FdCBV23, TGS^{+}19, MSLH23, She19]. Instance-aware [TGS^{+}19]. Instantaneous [SAD22]. institutional [GBW^{+}24]. instrument [CRF21, DWO^{+}22]. integer [KK22, AKC23]. integral [KAK21, SF16]. integrals [Bat19]. integrate [TSGOC^{+}22]. Integrated [BBD^{+}18, DRM20, FNX22, FRNW^{+}24, KKAUA21, LFT21, SSW^{+}24, SGP23, SR20, SO23a, ZWCQ22, ddSNX22]. Integrating [BJM23b]. Integration [BJM^{+}19, BFF^{+}21, BLM^{+}22, Els23, ESE^{+}23, EPF^{+}22, MFCV20, MKS^{+}23, PO19, dARPH^{+}19, SFF^{+}19, SHTE23, SV23]. integrators [DET23]. intelligence [DSB^{+}23, HTWT23, LCGA22, dlTAJCÁBG24]. intelligent [HKF^{+}20, PMFÁM22, PMS23, WWB^{+}23]. intensity [Dan20, Dan22]. intensive [WSRMB23]. Interaction [HOM^{+}18, Can22a, Can24, DL16, GGI23, YLS^{+}18, vKH20, RLN21]. interactions [Arc20, LMM22]. Interactive [YSM22, AALW23, AS17, AAF^{+}24b, BFNP24, Bre23, Cha17, DAB^{+}19, FSL16, GGI23, GP22, HX22, Kul20, MKT^{+}23, MKMK23, MM16, NR23, OHO16, RSLL^{+}20, RE22, SvZvdA23, SBCK17, DW20, GLGG22]. interatomic [KMK^{+}21]. interconnected [Cin24]. interface [ANA16, ADSG^{+}20, BHKK23, CPV^{+}20, DAC^{+}24, DLH18, EK23, FM24b, GdCF16, GMF18, GMF20, GTC21, JP23, JP24, JCFM24, LHZT23, MP17, MTS^{+}18, MM16, RtHLMN20, RK19, SK22, SCSB24, SPMGA^{+}23, SMR22, SSC23, WHG^{+}23, vdOJP^{+}20]. interfaces [Ek16, GZR^{+}19, Wet20]. interfacial [Tru18]. interfacing [SBT22]. interferometers [BJR^{+}20]. interferometry [OYW^{+}21]. intermediate [Mai22]. Internet [KJKK23]. interoperability [MKS^{+}23]. interoperable [BCD^{+}15, PPBZ21]. interpolation [BGAGJGC23, CZ22, DET23]. Interpretability [ZEADR23]. Interpretation [SIC24, JAMP23, ZVP22]. interpreting [Alb15]. interThermalPhaseChangeFoam [NR16]. intertidal [HJRBM23]. interval [DGK^{+}24, FHB^{+}21, MKB^{+}18, SABEh20]. interval-based [MKB^{+}18]. interval-valued [DGK^{+}24, SABEh20]. intervention [STH^{+}21]. interventions [HLG22]. intestinal [ME17]. introduce [BMBR21]. Introducing [Han20, Ric19]. Introduction [SLL23]. Intrusion [CCI^{+}22]. intuitionistic [RJDKJ24, WKCS24, WKS23]. invariant [YYL^{+}22]. invasive [CH23]. inversion [BBF^{+}19, BBN^{+}22, MGFO23, NL23, SF16]. INvestigate [dMOH21]. Investigating [FRNW^{+}24, SH23]. investigation [KPS23, LMN^{+}20]. Investment [GMKRS21]. Investments [ICS^{+}24]. invocation [CA18]. involving [SAA18]. ionic [CCFSB20]. ionized [KLY21, SLBC^{+}20]. ionosphere [ZRK21, ZRK22]. IOS [AGADAG^{+}23, PRC^{+}23]. ioT [HSMF22, ASSS24, ED22, ODE^{+}22, PMS23, PMZ21].

IoT-focused [ASSS24]. IP [Jur22]. iPSL [BCR^{+}18, VRBM16, dCWL^{+}23]. IPv6 [CSZM22]. IPv6CC [CSZM22]. ir [WBH^{+}23]. IRB [GBW^{+}24]. IRB-draft-generator [GBW^{+}24]. IREX [SEdCCCRG23]. IRIC [ZGZvB19]. IRRigation [AMC17]. irs [CCI^{+}22]. irs-partition [CCI^{+}22]. ISA.jl [SAD22]. iScore [RJH^{+}20]. iSeeE [CMDDARG23]. Ising [JP24, JP23]. Isobenefit [DV23]. Isobenefit-cities [DV23]. isogeometric [DHK^{+}20]. isolated [FG20]. isosurface [MZL^{+}23]. isotope [WW17]. Isotopes [dlSVBLdA^{+}17]. IssyOsborne [YMCB23]. Isula [GCC20]. iterative [KWB22, Sil20, SEdCCCRG23]. iTesla [BCR^{+}18, VRBM16, dCWL^{+}23]. Itinerum [PFJM19]. IV [SD23]. IV-PSNR [SD23]. ixpeobssim [BBD^{+}22].

J [KGK^{+}20, SACT23]. J-PET [KGK^{+}20]. j-Wave [SACT23]. Jacobian [Bat19]. January [Ano20c, Ano21f, Ano22f]. JAVA [WFD23, Agu23, AS23, CBPC23, CK21, CPV^{+}20, GCC20, Gór24, KPK23, LKK24, Pos22, PPBZ21]. JavaScript [MSDP23, dARPH^{+}19]. JAX [TFCZ24]. JaxSGMC [TFCZ24]. JBlockCreator [HGG20]. JDFTx [SLWS^{+}17]. jDSSAT [dARPH^{+}19]. jet [AT22]. JGMP [AS23]. JGNN [KPK23]. JHeaps [Mic21]. JIMWLK [Kor21]. JMetaBFOP [GLCBHTHO23]. joint [SLWS^{+}17]. Joseph [KOZ^{+}24]. js [OPPZ22]. Json [GZR^{+}19]. Json-GUI [GZR^{+}19]. judgments [tHLMN19, LMN18]. Julia [AMAE24, Gök21, LQM22, Rou19, SAD22]. Julia-based [LQM22]. July [Ano19d, Ano20d, Ano21g, Ano22g, Ano23g]. jump [CCE21]. June [Ano20c, Ano21h, Ano22h]. juSPH [LQM22]. JuWaPfl [HWT^{+}23]. juxtaposed [HHN20].

K3s [KMT24]. Kaplan [EB24]. Keras [PBSB22]. Keras-based [PBSB22]. kernel [RJH^{+}20, ZRBCI20, PF23]. kernels [AKC23]. KGC [AGEABD^{+}24]. kind [Tak22]. Kinect [MGMAG22, MAGM^{+}23]. Kinematic [STC^{+}18]. Kinetic [DSI^{+}20, HRC20]. KineticPy [MWM20]. kinetics [MWM20]. KipTool [CSK19]. kit [Bjö19, SHTE23]. Kit4DL [WMA24]. KitY [SHTE23]. KLFromRecordingDays [SA17]. KMstability [EB24]. knife [FVA^{+}20, SC17]. KNIME [LFK^{+}24]. knobs [DTD17]. Knowledge [YDN21, FVD20, GLGG22, LBH^{+}20, RBQ^{+}24, YD20]. knowledge-based [RBQ^{+}24]. Knudsen [CRF21]. Kochina [SHH23]. Kolsky [GLR22]. Korean [KALS24]. kPWorkbench [KMIG20]. KrakN [ZWM21]. Krypton [LH22]. Kt [CCO21]. Kufast [AA23]. Kunai [BT23]. Kunz [JS19].

Lab [HLW^{+}16, MHTZ22]. Labeling [NSO21, DRFB^{+}23, VCA^{+}24]. laboratories [EPF^{+}22, MCGK19]. Laboratory [Giu19, WHP24, PFC^{+}18, HHB^{+}22b]. labs [SPMGA^{+}23]. LabVIEW [CRF21, GCdJAURO21, KG24]. LAF [BS19]. Lagrangian [HS20, LXJ^{+}23, YYL^{+}22]. LALSuite [Wet20]. lambda [MFJ23]. LAMMPS [THSW24]. land [PRSS19, WRV23]. landmark [FSAD23a]. landscape [BSG20, DL16, Sal16, YFVL24]. landscapes [MWM20, WM21b, YFVL24]. lane [PMS23]. language [FdCBV23, HCD24, HBS16, Kri22, Kwa19, MDL^{+}18, OT16, OGPSQ22, PMBM^{+}23, SVM^{+}23]. language-agnostic [SVM^{+}23]. Laplace [DHK^{+}20]. laptops [AMS^{+}15]. LARA [BC20]. Large [PMP16, vdOJP^{+}20, AALW23, AMVB19, Bed24b, BD22, COG19, DZZ^{+}22, FAM^{+}20, KWB22, LBT^{+}21, LVCF24, LBL^{+}21, RNG^{+}21, SNLFGVRB24]. Large-Eddy [vdOJP^{+}20]. large-scale [AMVB19, BD22, DZZ^{+}22, LVCF24]. Laser [FRE^{+}24, SLM23, FRPC23]. LaserbeamFoam [FRE^{+}24, FRPC23]. Laserchicken [MGK^{+}20]. latency [JSG22]. Latin [BGCS19]. Lattice [HOM^{+}18]. Lattice-Boltzmann [HOM^{+}18]. LatticeLibrary [LMS^{+}16]. LatticeMech [KDHG20]. lattices [LMS^{+}16, OJ22]. law [KPJ24]. law-based [KPJ24]. layer [CPUARC20, MWJ15, MKS^{+}23]. layered [RCT20]. layers [BSB20, LXK23]. LBM [SGB^{+}22]. LBSA [SLM23]. LC [LIZ^{+}20]. LC-MS [LIZ^{+}20]. Lead [CGKD23, KCNJ24]. Lead-Acid [CGKD23]. leadership [Amo21]. leakages [CCZ23]. Learn [VMS^{+}23]. Learning [CHP23, GABH22, KNW^{+}23, LMN18, ZWM21, ACC^{+}23, AGM^{+}23, CCH^{+}19, CT24, DSB^{+}23, EAFJ^{+}22, FH23, FWJ^{+}23, FVA^{+}20, GFRMVL24, GY23, HH21, HCS^{+}20, Jur23a, KG23, KK24, KLL23, LLP^{+}24, LPL^{+}24, MKWS22, MSLNSMOG24, MHTZ22, MCPG24, NGOT24, PMBM^{+}23, VMS^{+}23, PDS^{+}22, PLGHG^{+}23, RSLL^{+}20, DFS24, SFF^{+}19, SDL21, SE23, SCSB24, SHTE23, SG24, TK23, Tem23, TS22, TGS^{+}19, TWL22, WMA24, WSRC24, ZL23, ZLLT23, dPPD^{+}23, vKMH^{+}20, tHLMN19, MAM^{+}23, Roa23, WWB^{+}23]. learning-based [FWJ^{+}23]. least [Alb19, AL16]. least-cost [Alb19]. least-costly [AL16]. led [MKS^{+}23]. Lefever [MD21]. legacy [DDCSH23]. LEGO [WTAG^{+}22]. LEI2JSON [HKZ24]. Leja [DET23]. length [Jun21]. Lens [MTS^{+}24]. Lesion [MTS^{+}24]. Lethe [BBB^{+}20]. LETICIA [GGI23]. level [AMS^{+}15, DTD17, KG24, LFT21, dRB^{+}20]. Leveraging [KFA^{+}24]. LeXInt [DET23]. lib [WSRC24]. libcommute [Kri22]. libcommute/pycommute [Kri22]. libraries [AI21, Bel24, Kis20, KSP19, Wet20].

Library [ANOU21, BCR^{+}18, VRBM16, dCWL^{+}23, AM19, Afr22, AAB23, AS23, ABC^{+}20, Ars21, AGdSC20, BW22, BWS23, BFV18, BJM23a, BBG^{+}19, BB22b, BDQ^{+}22, BCF20, CAG^{+}24, CA18, CK21, DM20, DW21, DHK^{+}20, DGK^{+}24, EPF^{+}23, FCMTV24, FSL16, FF24, GFRMVL24, GKM20, GÓ19, HCD24, HOB23, HBA^{+}20, IA17, JLP23, KH19b, KSS23, KMT23, KM21, KMK^{+}21, LKK24, LMM22, LRPD18, LSOM18, LB22, LFGVGDBD24, MBA^{+}22, MFCV20, MSK24, Mic21, MMH19, MKZ24, MK16, DDT20, PTS22, PBSB22, PS18a, PS18b, PS18c, PS19b, PMZ21, Pos22, Ras20, RV20, Rei23, RMV^{+}24, SGB^{+}22, SU23, SGBH18, SKS23, SL20a, SG24, SBJ^{+}23, SB21, THG^{+}23, TOPC23, TF23, WDZ^{+}20, WKS22, WS24b, WKR^{+}20, YMA23, ZMS18, ZGZvB19, Zie19, ZS22, ZPM23, BFF^{+}20]. Library [BCR^{+}18, BSB^{+}23, FdCBV23, FBA^{+}22, RLF^{+}21, RRKA22, VRG19, BLE21]. Libtorch [GY23]. libVersioningCompiler [CA18]. libxc [LSOM18]. LiDAR [MGK^{+}20]. life [Afr22]. Lifecycle [EL20]. lifetime [PS18a, PS18b, PS18c, PS19a, PS19b]. Light [KWB22, Agu23, SLM23, BHCT16]. Light-Field [BHCT16]. Lightning [CVD21]. lightweight [CBLI22, FM24a, KMT24, SDCA19, BCTMA^{+}23]. LIGO [AAA^{+}21, BHB^{+}21, FWB^{+}21]. like [AG21, TSGOC^{+}22]. limb [HVCS^{+}23]. LimeSurvey [ESE^{+}23]. limit [Por23]. Line [CC20, MDG22, SLL23, YG19]. lineage [CLM^{+}20]. Linear [AMWHD23, LK21, Nis20, BÁP^{+}22, FPN^{+}22, Hel22, KK22, KM21, KPJ24, LZN21, LCGA22, LSB^{+}20, PG18, RLK18, RV20, RRSK18, SRK23, WSK22]. linear-operator [RV20]. linear-scaling [RRSK18]. link [ACEAV24, SAA^{+}23b]. Linux [LHB24]. LiPTool [FWJ^{+}23]. liquid [RRS18, SP19]. liquids [CCFSB20, vDPI^{+}18]. literature [HSD^{+}22, NMLM18]. lithography [DBJ19]. litstudy [HSD^{+}22]. Little [Geo17]. LIVE [MYKT22]. lived [LXK23]. LiveDewStream [MHTZ22]. livestock [HKZ24]. LLT [KPJ24]. Load [RDO^{+}19, GTOM23, MB23, STC^{+}18]. loading [MFCSÁM20, MAAOZMAM22, SSC23]. loads [JCFM24, MMCKK21, dFB22]. local [FH23, HMY24, TACH17, LFT21]. localization [IHT23]. Localized [BS19, RPV^{+}20]. locally [BS19]. logging [GGI23]. logic [CCO21, DA24, CKC23]. logic-based [CCO21]. long [BCFJ24, GMF18, MWM20, MKMK23]. long-term [BCFJ24]. long-time [MWM20]. longitudinal [NG23, OCK^{+}23]. Loop [FSAD23a, FSAD23b, Ver22, XGW23]. Loss [EL20, ACEAV24]. low [ACC^{+}23, BFNP24, DEV^{+}21, JSG22, SP22, WCW^{+}23, YLM^{+}24, ZAPS20, WTAG^{+}22]. Low-carbon [WTAG^{+}22]. low-code [ACC^{+}23]. low-cost [DEV^{+}21, SP22, WCW^{+}23]. low-dimensional [ZAPS20]. low-latency [JSG22]. lower [ZRK21, ZRK22]. LTBEM [Azi23]. Lugiato [MD21]. LUMA [HOM^{+}18]. Lung [AJR^{+}23]. Lyapunov [ABB^{+}19]. LyapXool [ABB^{+}19, AGH20]. lying [MLTF^{+}18]. Lyo [Ek16].

M [GBJ23, NCGLRR24, XSS20, ZJKS23]. M-SPARC [XSS20, ZJKS23]. Machine [GABH22, ACC^{+}23, AGM^{+}23, CCH^{+}19, FH23, FVA^{+}20, FVD20, GY23, HH21, HCS^{+}20, Jur23a, KG23, KK24, KLL23, LCGA22, MKWS22, MCPG24, PDS^{+}22, RSLL^{+}20, DFS24, SDL21, SCSB24, SHTE23, SG24, TS22, TGS^{+}19, WMA24, ZL23]. machines [RJH^{+}20]. machining [UdL20]. macros [AHP23]. MaCySTe [LOM^{+}23]. made [DDCSH23, HGG20, PMP16, TS22]. made-to-measure [HGG20]. magcoilcalc [MFQ^{+}21]. magnet [MFQ^{+}21]. magnetic [BBN^{+}22, JAC20, JAC21, MFQ^{+}21, OC20, WDZ^{+}20]. Magpylib [OC20]. maintainability [ABC^{+}20]. maintenance [HWT^{+}23]. MakeDecision [WS24a]. Making [vDPI^{+}18, KSS23, RJDKJ24, SKS23, WKS23, WS24a, WKCS24]. Malaccha [WRV23]. MALTA [KSP19]. Malware [GLS^{+}22]. MAM [GLS^{+}22]. MaMiCo [JWJ^{+}22]. Mammut [DTD17]. manage [LMB^{+}19]. management [AHtH20, AMC17, ÁG19, ADSG^{+}20, CMdS^{+}23, DTD17, GSP^{+}17, GPW^{+}20, KPM^{+}22, LBT^{+}21, LBH^{+}20, MBU^{+}19, NMLM18, PTT20, Rub16]. manager [BFNP24, BLM^{+}22]. managing [MGCS23]. mango [FRNW^{+}24]. manifolds [ZAPS20]. manipulation [BSV^{+}23, HKF^{+}20, LCMD^{+}18, NGK^{+}20]. manipulations [Kis20]. Mann [CRV22]. manual [LCMD^{+}18]. manufacture [HGG20]. manufactured [OJ22]. manufacturing [AISM21, BOP^{+}23]. ManufacturingNet [AGM^{+}23]. many [CCH^{+}19, HOM^{+}18, VMRVF^{+}24, WBH^{+}23]. many-body [CCH^{+}19, WBH^{+}23]. many-core [HOM^{+}18]. MAO [SGDC22]. map [HX24, KXK^{+}23, OYW^{+}21, MBU^{+}19]. Mapper [HJRBM23]. Mapping [GL15, WDZ^{+}20, Bed24b, BNC24, BD22, CUSRCP^{+}22, DG21, GRW^{+}19, GVAO19, HJRBM23, IDE^{+}21, LD24, MZL^{+}23, YSM22, WM21b]. mappings [Nas20]. maps [BGAGJGC23]. MapX [LMB^{+}19]. MaranStable [SRK23]. marine [BBHP^{+}23, SO23a]. MarineEpi [BBHP^{+}23]. maritime [LOM^{+}23]. marker [LKSS20]. Market [WM21a]. markets [CK22b, WM21a]. Markov [CCE21, WRV23]. MartMi [LHZT23]. MartMi-BCI [LHZT23]. mask [DBJ19]. MaskDensity14 [LF15]. Masked [OTdMP22]. Masked-Piper [OTdMP22]. Masking [OTdMP22]. MaSMaker [TSGOC^{+}22]. masonry [LS23, Por23, SK22, YSM22]. Mass [CRF21, PMP16, CMMF19, Ras20]. masses [BHC^{+}24]. massive [CLM22, MGK^{+}20, DRFB^{+}23]. massively [LGP^{+}22, Mos20, PGA^{+}20]. Master [KHM^{+}22]. MATBOX [UVPB^{+}22]. Matching [MSK24, JSG22]. MatCoupLy [Roa23]. material [BHE^{+}19, MTJ^{+}22, QTL23, WHG^{+}23]. MaterialModeler [BHE^{+}19]. MaterialModeler-From [BHE^{+}19]. materials [DM19, JAC20, MP20, MASB18, MYKT22, SSP21, SDH20, TLDM19, WM21b, YHC^{+}22, ZS24]. MateriApps [MYKT22]. Mathematica [Arc20, Byk19, IIS18, Ozb17, RL19]. mathematical [AK24]. MatHH [CDOBA22]. MATLAB [CK23, AI21, BFSJP^{+}21, Bat19, BHC^{+}24, BGGL20, BAEBAS19, BV19, BlD^{+}20, CPUARC20, CH23, CK22a, Cob21, Deg20, EK23, FJ22, FS23, FM24b, tHLMN19, HKF^{+}20, HL22, IZBT21, JSB20, JAC20, KMLP^{+}23, LMN18, Liu23, MMD24, MKU22, Nas20, Ozk23, PMNWR20, PNL^{+}21, Por23, RCT20, SY20, SBL19, VRVAC23, XZL24, YHC^{+}22, Zek17, ZLLT23, BBHP^{+}23, BC22, BO19, BBF^{+}19, BBN^{+}22, BSS^{+}23, BD22, Can22a, Can22b, Can24, CDOBA22, JGGF23, KAR^{+}19, LHZT23, MP18, NK18, SLM23, XSS20, ZJKS23]. MATLAB-based [CK23, CK22a, VRVAC23, YHC^{+}22, BC22, CDOBA22, JGGF23, LHZT23]. Matlab-Simulation [ZJKS23, XSS20]. MatlabHTM [BSB20]. MatNexus [ZS24]. Matrix [Kos22, RXZ^{+}23, RRKA22, LB22, MASB18, Roa23, WRV23]. Matrix-based [Kos22]. matter [vKH20]. MATVines [Cob21]. Mauritius [IDE^{+}21]. MAWA [APLI24]. Max [SRML17]. Max-tree [SRML17]. Maximization [SKD22]. maximum [SR19]. May [Ano23h, Ano24d]. maze [TSGOC^{+}22, AGHK21]. maze-like [TSGOC^{+}22]. MC [Cer19, DM19]. MC-BAM [DM19]. MCDA [AMA19, WBS22a, WBS22b, WBNS23]. McDAPS [TCCK23]. Mcfly [vKMH^{+}20]. MCMC [TFCZ24]. Meander [RNG^{+}21]. meandering [RNG^{+}21]. MEAS [AK24]. MeaSSUre [OKJ23]. measure [BDSC22, Car21, HGG20]. measure plasma membrane [AHP23]. measured [HHH22]. Measurement [NK20, CWGM24, HKC^{+}18, Jun21, Liu23]. measurements [AHP23, CMMF19, FNX22, MTS^{+}18, MKR^{+}21, O’D21, PFC^{+}18, SADU23, TCQOGM23, ddSNX22]. measures [AMVB19, EB24]. measuring [DDT20, RLM^{+}24, ZC20]. mechanical [BÁP^{+}22, KH19a, PRSS19]. mechanics [Bjö19, BWMS22, BJM23b, KDHG20, LKSS20, LDM20, RAS19b]. mechanisms [ÁG19]. mechanochemical [DyH21]. media [KRB^{+}20, MPAK19, RCT20, SGB^{+}22, WLW23]. mediation [AVB17, BFV18]. medical [ML20, MCBM24, RLK18, RSLL^{+}20, SFF^{+}19]. Medicare [MKS^{+}23]. MedVoxelHD [MCBM24]. meets [MML20]. MEG [ZVP22]. Meier [EB24]. melting [CCFSB20]. membrane [AHP23, KMIG20]. memoization [BPL^{+}19]. memory [BSB20, IA17, SPG^{+}23]. memory-efficient [SPG^{+}23]. Mental [TTHZ23]. MergeBathy [ZHP^{+}18]. mesh [BDF^{+}20, CLDdM22, GEG22, HR20, NGK^{+}20, RDO^{+}19, Tru18]. mesh-free [GEG22]. meshes [SGP23]. meshing [SC23, UVPB^{+}22]. meshless [RB17]. Meshtool [NGK^{+}20]. meso [KSA23]. meso-scale [KSA23]. MesoGen [JRP23]. Mesoscopic [JRP23]. message [SU23]. Meta [AMDN23, SCK^{+}23]. meta-analyses [SCK^{+}23]. Meta-optimizing [AMDN23]. metabolites [LIZ^{+}20]. MetaCluster [TOPC23]. metadata [FHB^{+}21]. metagraphs [RNR17]. metaheuristic [TOPC23]. metaheuristic-based [TOPC23]. metaheuristics [AMDN23, CDAOB^{+}20]. metamaterial [KDHG20]. metamodeling [EBL24]. Meteorological [BS19]. meter [OKJ23]. methane [HKC^{+}18, NIP^{+}23]. Method [HOM^{+}18, KL22, RSMW20, BE20, CWC^{+}23, CCE21, DSI^{+}20, FAM^{+}20, HLP^{+}19, HWST22, KAK21, KXK^{+}23, KAS23, LB22, Nis20, NSLD16, OT16, PMM16, PPRE17, PDTG17]. methodology [BGAGJGC23]. methods [AES^{+}22, AGH20, BW22, BWS23, BO19, BCF20, JC23, MSV23, RB17, RJDKJ24, VCLS21, WKS22, WS24b]. metric [KHG21, ÖA23]. MetricHunter [ÖA23]. metrics [MRGAGP23, PCPC24]. MF [EAFJ^{+}22]. mFLICA [Amo21]. MFPP [Kos22]. MGARD [GCW^{+}23]. mgpr [VRP23]. MGtoolkit [RNR17]. MICA [MKB^{+}18]. MICR [KP23]. micro [PFC^{+}18]. micro-computed [PFC^{+}18]. microbes [SSW^{+}24]. Microbi [SSW^{+}24]. MicroFEA [MP20]. MicroFract [SS17]. micromagnetic [Zhu15]. microscopes [CBHLG21, HL22]. microscopy [BFSJP^{+}21, GRF23, MSK24, QTL23, Rub16, TGS^{+}19, AVCP^{+}21]. microsimulation [SR20]. microstrip [VRVAC23]. MicroStructPy [HR20]. microstructural [Liu24, SS17]. Microstructure [FPBM18, FST^{+}21, UVPB^{+}22, HR20]. microstructures [CWM^{+}21b, SSP21]. MicroVIP [AVCP^{+}21]. microwave [KAO^{+}21]. Microxanox [KP23]. mics [BLE21]. mics-library [BLE21]. MIDAS [HT18, ZRK19]. MIDAS-VT-Pre [ZRK19]. Middleware [MPX^{+}24, SBT22, YDMC15]. Mie [THSW24]. Mie-FH [THSW24]. MIEP [GTG21]. Milky [PI17]. minima [LEK22, KGF^{+}24]. minimal [Kar21]. Minimalist [AMC17]. minimalistic [DN17, Rag17]. minimization [PS22]. minimize [IVSV23]. minimum [ÇA20, MMCKK21]. minimum-thickness [MMCKK21]. mining [AMA19, MKT^{+}23, SvZvdA23, YÖYB23, ZS24]. Minkowski [BT21]. MiRNA [CPD^{+}20]. MiRNA-QC-and-Diagnosis [CPD^{+}20]. MIROR [SBT22]. MISA [GMF20]. mission [SNW^{+}21]. MISTIQS [PBL^{+}21]. MitC [KKN^{+}22]. mitigate [EBGCT21]. mitigation [BGM^{+}24, KKN^{+}22]. mixed [BÁP^{+}22, KK22]. mixed-dimensional [BÁP^{+}22]. mixes [LPG^{+}23]. mixture [BL16]. mixtures [Dra23, THSW24, vDPI^{+}18]. MJOLNIR [LJML20]. mk [LHB24]. ML [MAG24, VDCL20, ZL23]. ML-based [VDCL20]. MLEce [ZKJ^{+}24]. MLOps [GM24]. mlr3shiny [TS22]. mlr3shiny-State-of-the-art [TS22]. MMS [ROMH22]. MNEflow [ZVP22]. mob [CS21]. Mobile [KCNJ24, APLI24, Bed24a, Bed24b, BCTMA^{+}23, CS21, ML20, MAM^{+}23, SBT22, BSRSC18, CPF23]. MobiQ [YLS^{+}18]. MobiScout [AGADAG^{+}23]. MOBOpt [GdVSdSCM20]. mod [GVAO19]. Modal [PARS22]. Model [AMAE24, AMC17, BSdG^{+}22, BSB20, BHE^{+}19, CPUARC20, CH23, CVV23, Dan22, DK19, Ek16, GSP^{+}17, GVAO19, GBJ23, IDE^{+}21, JMT22, KWB22, KAS23, LCGA22, LIZ^{+}20, LPL^{+}24, MKWS22, MC18, MM16, MTJ^{+}22, NMT23, PG18, PZ22, RNK21, RNK22, Sal16, SO21, SH23, SBJ^{+}23, TCPC22, WTAG^{+}22, ZRK19, ZNS17, WLW23, HGWM18, KF22]. model-2-model [GVAO19]. model-based [Ek16, KWB22]. model-selection [LIZ^{+}20]. modeler [Has18]. Modelica [BCR^{+}18, DRFB^{+}23, FdCBV23, GVAO19, SU23, VRBM16, VBA^{+}16, dCWL^{+}23]. ModelicaGridData [DRFB^{+}23]. modeling [AZ17, BK19, BSB^{+}23, COGP19, Elk22, FdCBV23, FVD20, GG21, HGS17, HRC20, JRP23, Jur22, Jur23a, KAK21, LBT^{+}21, MGGHG^{+}24, MFQ^{+}21, MMT24, NGK^{+}20, ND22, dQRRBdSM20, SP19, SL21, Tru18, WMM18, WKR^{+}20, YMA23, ZR19, AMWHD23]. Modelling [CH23, Che21, Bre23, BJR^{+}20, Liu24, LFT21, MVBF19, PBM22, SR20, WLW23]. models [AALW23, BSS^{+}23, DRM20, DSB^{+}23, FVA^{+}20, GKK^{+}23, Gök21, HCD24, HVB16, Hel22, IAW^{+}15, LCMD^{+}18, LHY^{+}23, LMN^{+}20, MAG24, MD22, MBI20, NNDS23, PMM16, RJH^{+}20, SSSH16, SO21, SEdCCCRG23, VBA^{+}16, WS24a, DMFG22, NRC23, Jur22, Jur23a]. moderate [JAC20]. Modern [FS23, DDCSH23, HMCA15, JHAMF19, NBCC19]. MODFLOW [PBS^{+}24]. modification [SKD22]. modified [KSS20, Tak22]. Modifying [GLS^{+}22]. Modular [GWK16, MAGM^{+}23, TFCZ24, BFV18, BST^{+}17, BKM22, CJGA23, CRF21, KHH^{+}22, NBCC19, SFG21, SF16, SR20, THG^{+}23, VBA^{+}16, XTD24, YLS^{+}18, Zaj20]. modulators [SLM23]. Module [PMY^{+}24, AGEABD^{+}24, BS21, CGHGRB21, CR22, CWC^{+}23, CWM^{+}21a, DyH21, GZR^{+}19, ILP^{+}24, LGG^{+}23, MCGK19, MD22, PARS22, dARPH^{+}19, SPG^{+}23, TBO24]. MODULO [NM20]. MoebInv [Kis20]. moisture [AMC17, MPMC21]. MoLE [VDCL20]. molecular [AMS^{+}15, DFC18, DPF^{+}21, FC17, HA23, JWJ^{+}22, KMK^{+}21, LSF^{+}23, MMD24, iNKN^{+}20, OT16, Ras19a, RWJ^{+}21, SFK^{+}19, SGA^{+}16, ST21, TBO24, RS22]. molecular-continuum [JWJ^{+}22]. molecule [IHT23]. molecules [KLY21]. Moment [DM19, BCF20, SR19]. Moment-curvature [DM19]. Mon [KMT24]. MONAI [LFP^{+}23]. MoniThor [UdL20]. monitor [GCBDDBF23, MLTF^{+}18, RLN21]. monitoring [AGADAG^{+}23, CCBC^{+}21, CAW^{+}20, FG20, GFRMVL24, GACS^{+}23, KMT24, MGFO23, MLTF^{+}18, RLJS23, STH^{+}21, SP22, UdL20, VFFS23, YKC^{+}19]. monocular [MBY22]. Monte [BCA19, LNS15, NBM^{+}19, NBS^{+}21, CCE21, DAC^{+}24, IIH^{+}23, KP19, KXK^{+}23, vKH20]. Monte-Carlo [IIH^{+}23]. Moodle [MSLNSMOG24]. MOOSE [GY23, ILP^{+}24, LGP^{+}22, LGG^{+}23, PGA^{+}20, PMY^{+}24, SGP23, SPG^{+}23]. MOOSE2 [AL16]. Morph [AGEABD^{+}24]. Morph-KGC [AGEABD^{+}24]. Morpheme [KALS24]. Morpheme-based [KALS24]. morphodynamics [CVV23]. Morphogenesis [DV23]. morphological [BlD^{+}20, MCBM24]. Morphology [ŁMF^{+}23, JZGW22, OJ22, TŚG^{+}24]. MOSA [VMRVF^{+}24]. MOSA/D [VMRVF^{+}24]. MOSA/D-O [VMRVF^{+}24]. MOSAD [VMRVF^{+}24]. MOSAD/D [VMRVF^{+}24]. MOSAD/D-O-II [VMRVF^{+}24]. mosquito [IDE^{+}21]. Motion [Web23, ADSG^{+}20, Can22b, EPF^{+}22, FK23, FK23]. MotionPerfection [GZA^{+}22]. motions [ACR22]. Motor [FdCBV23, GZA^{+}22, LHZT23]. motors [NCGLRR24]. Mouse [KOZ^{+}24, KOZ^{+}24]. mouth [EAFJ^{+}22]. movecost [Alb19]. movement [EAFJ^{+}22, JAC21]. MoveR [PLC^{+}24]. movies [BRF^{+}23]. MPI [DFSW19, RJH^{+}20]. MPI-parallelized [DFSW19]. MPLBM [SGB^{+}22]. MPLBM-T [SGB^{+}22]. MPSLIB [HVB16]. Mr.CAS [Rag17]. MS [LIZ^{+}20]. MSAT [RXZ^{+}23]. MSCPDPLab [ZLLT23]. MSDSp [CCE21]. MSGLib [SU23]. MST [ÇA20]. MStaT [RNG^{+}21]. Multi [AGG23, ANOU21, FdCBV23, AMS^{+}15, AA23, AMDN23, BSdG^{+}22, BÁP^{+}22, FPN^{+}22, GdVSdSCM20, GCP22, GJS^{+}22, KSS23, KMK^{+}21, LRPD18, LYZ^{+}22, MSK24, MTPHH18, MMS21, PBM22, PCdJSdFR^{+}23, PZKK21, RJDKJ24, SUM21, SKS23, SHK^{+}23, TCCK23, ZLLT23, FRNW^{+}24]. Multi-Agent [FRNW^{+}24]. multi-channel [TCCK23]. multi-class [LYZ^{+}22, SHK^{+}23]. multi-criteria [KSS23, PZKK21, RJDKJ24, SKS23]. Multi-Domain [FdCBV23, BÁP^{+}22]. multi-energy [GCP22, SUM21]. multi-horizon [BSdG^{+}22]. multi-image [LRPD18]. multi-level [AMS^{+}15]. multi-objective [AMDN23, FPN^{+}22, GdVSdSCM20, KMK^{+}21]. multi-purpose [MSK24]. multi-sensor [GJS^{+}22]. multi-source [ZLLT23]. Multi-sphere [ANOU21]. multi-stage [MMS21]. multi-tenancy [AA23]. multi-tier [PCdJSdFR^{+}23]. multi-user [PBM22]. Multi-vehicle [AGG23]. multiagents [BCTMA^{+}23]. multicamera [SMD^{+}24]. multichannel [BAEBAS19]. multicomponent [SLBC^{+}20]. multicore [ACPdP22]. multidimensional [ZEADR23]. Multifaceted [KAR^{+}19]. MultiFEBE [BÁP^{+}22]. MultiFLEXX [MTPHH18]. multiflexxlib [MTPHH18]. MULTIFRAC [THT20]. MultiGlobAl [CC23]. multigrid [GCW^{+}23, ZNS17]. multilayer [CC23]. multimedia [KRA23]. multimodal [AES^{+}22, CK22a, CK23, OTdMP22, SL20b]. multimode [SL20a]. multiobjective [KK22]. Multiphase [SGB^{+}22, Kar21, SRK23]. multiphysics [DWO^{+}22, LGP^{+}22, PGA^{+}20]. multiplatform [BB20a]. Multiple [BL16, MKB^{+}18, O’D21, AMA19, BCA19, CA18, HVB16, KM21, LBG20, MSDP23, MKMK23, BLE21]. multiple-point [HVB16]. multiplexing [LJML20]. multiplicity [JS19]. multiplied [LF15]. multiply [Nas20]. Multiresolution [BCA19]. multiscale [LXJ^{+}23, THT20, AGHK21, NM20]. Multisensor [BAEBAS19]. MultispeQ [SADU23]. multitemporal [BNC24]. multivariate [ASAA20, FMFCEG^{+}23, HLP^{+}19, HWST22, PR19, SA20, VRP23]. MultiVehicle [BCTMA^{+}23]. MUPen2DTool [BBN^{+}22]. MUSEN [DS20]. music [HOB23]. Musicaiz [HOB23]. Mutagenesis [BMBR21]. mutation [YB23, SLBC^{+}20]. mutation-based [YB23]. mutual [GWGB22, LZN21]. mutually [LHCK18]. MVSim [BCTMA^{+}23]. MVTS [ASAA20]. MVTS-Data [ASAA20]. MWSegEval [KAO^{+}21]. MY [AMC17]. myDLTS [KG24]. myoelectric [SEZLMS^{+}23].

name [KK24]. named [KDRT23]. NAMELIST [Nai17]. names [Xie22]. Nancy [ZS22]. NanoDroplet [Ras21]. Nanoindentation [CGS19]. NANomaterials [RPV^{+}20]. nanoparticles [Nev17]. NANOPTICS [RPV^{+}20]. narratives [ASGC^{+}24]. National [HHB^{+}22b]. native [KPK23, WSRMB23]. natural [GMNG^{+}18, LMB^{+}19]. Navier [AG21, LGG^{+}23, SAA^{+}23a, Wil17]. navigation [Bed24a]. Navigator [STH^{+}21]. NCBI [HM22, PMP16]. NCC [KHJ^{+}23]. NDE [SFG21]. NDECOAX [MPAK19]. NDefA [BÇ23]. NDVI [HJRBM23]. NDVI-based [HJRBM23]. Near [BV19, ML20]. near-field [ML20]. Near-Optimal [BV19]. Nearest [BHC^{+}24]. neat [vDPI^{+}18]. Nebula [vKH20]. NEFBDAA [Fra22a]. negative [LB22, CGKD23]. Neighbors [BHC^{+}24]. NEMO [VDP^{+}19]. neocortical [BSB20]. NEOLAB [CGKD23]. NeoViewer [AAF^{+}24b]. Neper2CAE [GTC21]. NES [CK23, CK22a]. NESIM [RGdlTQ^{+}23]. NESIM-RT [RGdlTQ^{+}23]. NET [MD19]. net chan [AM23]. NetCausality [RMMG21]. NetCDF [ZNS17]. NetCenLib [FF24]. netDFT [WFD23]. NetKet [CCH^{+}19]. NetLogo [GG21]. NetOnZeroDXC [PR19]. Network [ALRM21, CVD21, GV20, KJW21, AFGH22, AM23, BMK23, CC20, DA24, FF24, GZS23, HTV22, Jun21, Jur22, MW23, MLD22, NNDS23, RWJ^{+}21, RGdlTQ^{+}23, RMMG21, SCD^{+}24, TSNRM24, VDP^{+}19, YKC^{+}19, ZS22]. networking [Jur23b]. Networks [PMNWR20, PLGHG^{+}23, BÇ23, BS21, BB22b, BD22, COGP19, CSZM22, CJGA23, CC23, Cin24, EL17, FVD20, LVK21, Mül23, PR19, Pos22, SNLFGVRB24, SWCP20, TC20, WC23, ZVP22, BOP^{+}23, CCI^{+}22, CCTCGCOB23, FLR22, KMLP^{+}23, KPK23, MDR^{+}22, PBSB22]. NetworkX [BOP^{+}23]. Neural [BOP^{+}23, CCTCGCOB23, FLR22, KPK23, MDR^{+}22, PBSB22, APK^{+}23, BB22b, CJGA23, DA24, FVD20, Jun21, LVK21, RWJ^{+}21, RGdlTQ^{+}23, RMMG21, SWCP20, WC23, ZVP22]. neural-network [RWJ^{+}21]. neuroevolutionary [CCFSB20]. NeuroFramework [CCFSB20]. neuroimaging [GAM^{+}24, MCC20]. NeuroSpeech [OAVCVB^{+}19]. NeuTomPy [MMG19]. neutron [CSK19, HHB^{+}22b, LJML20, MTPHH18]. neutrosophic [SABEh20]. news [PP19]. NewsCompare [PP19]. next [AJK^{+}23, LLK^{+}20, ODE^{+}22]. NG [ODE^{+}22]. NiceProp [GP22]. Niget [CGS19]. Niimpy [ITL^{+}23]. Ninja [BCD22]. NL [PG18]. NL4Py [GG21]. nlchains [PO19]. NLP [PF22]. NMR [BBF^{+}19]. NNC [TTT19]. no [BFNP24]. noble [MFQ^{+}21]. node [KPKS22]. noise [DI22, LF15, MGGHG^{+}24, MD22, RAL^{+}20]. noisy [Ben15]. Non [LZN21, PMNWR20, BAEBAS19, BSG20, CRV22, DI22, FRdN21, GP22, GEH19, Kis20, KHJ^{+}23, LB22, Ras19a, ZRK21, ZRK22, ZR19, KMLP^{+}23]. non-auroral [ZRK21, ZRK22]. non-concurrent [KHJ^{+}23]. non-constant [Ras19a]. non-Euclidean [Kis20]. non-Gaussian [DI22]. non-ideal [GP22]. non-industrial [ZR19]. Non-linear [LZN21]. non-negative [LB22]. Non-Parametric [PMNWR20, CRV22, FRdN21, KMLP^{+}23]. non-stationary [BAEBAS19]. non-target-species [BSG20]. non-uniformly [GEH19]. nonadiabatic [SFK^{+}19]. nondestructive [MPAK19]. Nonio [NBCC19]. nonlinear [AGdSC20, GRD^{+}23, KF23, MD22, PO19, Por23, SK22, LH22]. nonparametric [FJ22]. NonParRolCor [PMLM23]. nonstationary [SOS19]. nonuniform [BFRK19]. normalization [BBHB22, PDTG17, SABEh20]. notebooks [FM24a]. novel [KSFG18, PP19]. nozzle [FH23]. NS [SCD^{+}24]. NS-3 [SCD^{+}24]. nsCouette [LFR^{+}20]. NSDPY [HM22]. nuclear [Bed24a, BBN^{+}22]. nuclei [Taq16]. numba.cuda [Kle24]. numba.jit [Kle24]. number [AJK^{+}23, BL16, HHN20]. numbering [Ric19]. Numerical [FPI^{+}22, GRD^{+}23, Kor21, Afr22, Azi23, Byk19, DFSW19, Elk22, LFR^{+}20, Mai22, dQRRBdSM20, VPVNP22]. NURBS [BK19]. NURBS-Python [BK19].

O [VMRVF^{+}24, KPS23, VMRVF^{+}24]. Oak [HHB^{+}22b]. Obfuscapk [AGVM20]. obfuscation [AGVM20]. obial [KP23]. ObiWan [SSW^{+}24]. ObiWan-Microbi [SSW^{+}24]. Object [BSB^{+}23, AM19, BK19, CCV22, HVH24, MGCS23, NS22, KJW21]. Object-Oriented [BSB^{+}23, AM19, BK19, CCV22, MGCS23, NS22, KJW21]. objective [AMVB19, AMDN23, BW22, BWS23, FPN^{+}22, GdVSdSCM20, KMK^{+}21, SA17, SD23, VMRVF^{+}24]. Objects [WBS22b, WBNS23, BKM22, WBS22a]. observables [Nis20, Ver22]. Observation [WFL^{+}23]. observations [SMZSW24]. observatory [FHB^{+}21]. observed [BS19]. observing [AAA^{+}21]. Obstacl [MVBF19]. obtain [dFB22]. obtained [CWM^{+}21b]. OCCI [GdCF16]. occupy [KP23]. ocean [MVBF19, DPD22]. Ocelet [DL16]. OCT [HKF^{+}20]. OCT-A [HKF^{+}20]. Octave [Wet20]. Ocular [KAR^{+}19, MTS^{+}24]. ODAMNet [TBO24]. ODMeans [HMY24]. ODT [SL21]. oedometer [MAAOZMAM22]. off [OYW^{+}21, SCSB24]. off-axis [OYW^{+}21]. off-stream [SCSB24]. Oliot [AGL^{+}23]. OMERO [SSW^{+}24]. OMERO-based [SSW^{+}24]. omicR [TBCG21]. Omicron [RAL^{+}20]. Omnisolver [JP24, JP23]. One [SL21, YMCB23, Dan22]. one-beam [Dan22]. One-dimensional [SL21, YMCB23]. Online [TSCH20, WS24a, BÇ23, HKS^{+}22, HT18, HX22, RKA^{+}23, TWL22, dMVBTN24, XYC22, STH^{+}21]. onlineBcp [XYC22]. onto [Nas20]. ontologies [DRM20]. ooi [GdCF16]. OONIS [KJW21]. OOP [BKM22]. op [HVH24]. OPA [MSDP23]. Open [AAA^{+}21, BÇ23, BCR^{+}18, Bed24b, BTMB21, DAB^{+}19, HT18, KKN^{+}22, KHM^{+}22, LFP^{+}23, OKJ23, RJN^{+}23, SH23, VRG19, WLP16, AHtH20, ACPdP22, AGL^{+}23, AGVM20, ASVV23, BSdG^{+}22, BSRSC18, BDF^{+}20, BGGL20, BK19, BBB^{+}20, CR22, Cer19, CBLI22, CBS^{+}16, DEV^{+}21, DVSD22, DS20, Elk22, EPF^{+}22, EDZ^{+}23, FNX22, FK23, GZS23, GWBF23, GB24, GNPP23, GV20, GMNG^{+}18, HGG20, HYR^{+}23, ICS^{+}24, IAW^{+}15, JWJ^{+}22, JSG22, KH19a, KRB^{+}20, KK24, KPM^{+}22, KSA23, LMB^{+}19, LEFSO20, LSB^{+}20, LHL^{+}24, LQM22, MP17, ML20, MAS^{+}22, MMCKK21, Mic21, MKR^{+}21, MRGAGP23, MML23, MAAOZMAM22, MBY22, NFASC21, ODE^{+}22, OHO16, OAF20, OBWM23, PL23, PCPC24, PFJM19, PM23, PNL^{+}21, PKA^{+}22, PBL^{+}21, RHR^{+}23, REK^{+}23, RERES23, ROMH22, RSMW20, Rub16, RRSK18, SFK^{+}19, SG24, SDH20, SMR22, SC19, SACT23, TSGOC^{+}22]. open [TOPC23, WGB16, Web17, Web21, WDZ^{+}20, WTAG^{+}22, ZRK21, ZRK22, ddSNX22, LK21, BHKK23, FdCBV23, FWB^{+}21, KPS23, LPG^{+}23, RAL22, UVPB^{+}22]. Open-Instance [BCR^{+}18]. Open-source [BÇ23, HT18, KKN^{+}22, LFP^{+}23, RJN^{+}23, SH23, VRG19, ACPdP22, AGL^{+}23, ASVV23, BSdG^{+}22, BSRSC18, BDF^{+}20, BGGL20, BK19, BBB^{+}20, CR22, Cer19, CBLI22, CBS^{+}16, DEV^{+}21, DS20, Elk22, EDZ^{+}23, FNX22, FK23, GZS23, GWBF23, GB24, GNPP23, GV20, HYR^{+}23, ICS^{+}24, IAW^{+}15, JWJ^{+}22, JSG22, KRB^{+}20, KK24, KPM^{+}22, KSA23, LEFSO20, LSB^{+}20, LQM22, MP17, MAS^{+}22, MMCKK21, Mic21, MKR^{+}21, MML23, MAAOZMAM22, MBY22, NFASC21, OHO16, OAF20, OBWM23, PL23, PCPC24, PM23, PKA^{+}22, PBL^{+}21, RHR^{+}23, REK^{+}23, RERES23, RSMW20, Rub16, RRSK18, SFK^{+}19, SG24, SDH20, SMR22, SACT23, TSGOC^{+}22, TOPC23, WGB16, Web17, Web21, WTAG^{+}22, ddSNX22, AGVM20, RAL22, UVPB^{+}22, BHKK23]. OpenArch [MMCKK21]. OpenCL [ACPdP22]. OpenCLC [LSB^{+}20]. OpenFL [DSB^{+}23]. OpenFL-XAI [DSB^{+}23]. OpenFOAM [GKM20, HS20, LXJ^{+}23, RDO^{+}19, SAA^{+}23a, SNHS20, VRG19]. OpenGJK [MP18]. OpenIMDML [FdCBV23]. OpenImpala [LK21]. OpenIPDM [HLG22]. OpenIPSL [BCR^{+}18, dCWL^{+}23]. OpenMechanochem [DyH21]. openMHA [KHM^{+}22]. OPENMMF [SL20a]. OpenMoist [MPMC21]. Openpipeflow [Wil17]. OpenPIV [BGGL20]. OpenPIV-MATLAB [BGGL20]. OpenPLC61850 [ROMH22]. OpenQSEI [SBL19]. OpenSe [ZMS18]. OpenSees [RNK22, RNK21, SK22]. OpenSeesPy [GYWG23, ZMS18]. OpenSeesPyView [GYWG23]. OpenStack [GdCF16, HTB19]. OpenTEPES [RAL22]. Operating [LD24, LYZ^{+}22, ODE^{+}22]. Operational [PARS22]. operations [BBHB22, MCPG24, SABEh20]. operator [Kri22, RV20]. Opportunistic [TDG19]. OptFROG [MRMD19]. Opti [BBJ^{+}18]. Optical [MBF20, OV24, AAN17, BJR^{+}20, Deg20, DBJ19, HA19, SPMGA^{+}23, RPV^{+}20].

OptiGUI [SPMGA^{+}23]. Optimal

[Bla21, Bla23, BV19, GRD^{+}23, LMS^{+}16, MTJ^{+}22, PPRE17, WBH^{+}23]. Optimel [PPRE17]. OptiMic [SSP21]. optimisation [CDAOB^{+}20, FPN^{+}22, MSLH23, SUM21]. optimisation-based [SUM21]. Optimization [Bla21, Bla23, PMY^{+}24, dSM22, AES^{+}22, CCBC^{+}21, GdVSdSCM20, GLCBHTHO23, HFS^{+}21, HCS^{+}20, IZBT21, IKYY22, JMT22, KK22, KHG21, LSPG22, LXK23, MBDS20, MFQ^{+}21, Por23, RV20, Sil20, SBT22, THG^{+}23, TSMT19, ZY23, TK23, WTAG^{+}22]. optimization-based [Por23]. optimize [KFA^{+}24, TSNRM24]. Optimized [NBM^{+}19, NBS^{+}21, MRMD19, iNKN^{+}20, SSP21]. optimizer [HVH24]. optimizing [AMDN23, BB22b, FLR22]. OptiSens [CCBC^{+}21]. OPTool [Sil20]. orchard [MGMAG22]. orchards [MAGM^{+}23]. orchestration [BB22a]. order [BBB^{+}20, FSMCFP23, GWGB22, MMH19]. ordered [LYZ^{+}22, RLM^{+}24]. ordering [NBCC19]. ordinal [BBHB22, ZEADR23, MSDP23]. Ordinalysis [ZEADR23]. oreo [LBL^{+}21]. organic [JZGW22]. Oriented [BSB^{+}23, AM19, AL16, BK19, CCV22, CJGA23, KBB19, MGCS23, NS22, KJW21]. Origin [HMY24]. Origin-Destination [HMY24]. Orthogonal [NM20, AKC23]. ORTiS [MB21]. OSCAR [Deg20]. OScillating [YMCB23]. oscillations [Dan20, Dan22]. oscillatory [LBL^{+}21]. OSLC [Ek16]. OSLC-compliant [Ek16]. OSM [AAN17]. OSM-Classic [AAN17]. other [RDH^{+}21]. OTSunWebApp [PNC23]. out-of-plane [FNX22]. outcrops [AS17]. Output [GPW^{+}20, FRdN21, PS18a, PS18b, PS18c, PS19b, SMD23]. outputs [PLK^{+}23]. ov [LSMG19]. ov-SSA [LSMG19]. overlap [TBO24, LSMG19]. Overlap-SSA [LSMG19]. Oversampling [RLF^{+}21]. OWA [RLM^{+}24]. OWL [BKM22, DRM20]. OWL2Go [DRM20]. OWLOOP [BKM22]. owner [ZR19]. OXI [RKA^{+}23]. OXic [KP23]. oxide [BWMS22, TCQOGM23]. oxide-cement [TCQOGM23].

P [CCO21]. P-Kt [CCO21]. P2PFaaS [MB23]. Package [AJR^{+}23, KAR^{+}19, WBS22b, WBNS23, XSS20, XSC^{+}21, ZJKS23, ASAA20, ANA16, Alb15, Alb19, AKC23, ASSS24, Amo21, ATCA20, BMK23, BLE21, BO19, BAEBAS19, BE20, BRF^{+}23, BJR^{+}20, BdSvE23, BSG20, CVV23, CPD^{+}20, CCFSB20, CJGA23, CZ22, COK24, CLK21, Cob21, CBHLG21, CPB^{+}23, DMFG22, Deg20, DFC18, DPF^{+}21, DFSW19, Dra23, DTDd19, EBL24, EAMSR^{+}23, FSMCFP23, FRdN21, FM24a, FC17, FAM^{+}20, FVVL23, GZ21a, GZ21b, GMF18, GRW^{+}19, GKK^{+}23, Gir21, Gök21, Gór24, Hac23, HFS^{+}21, HHB^{+}22a, Han20, HM22, HSD^{+}22, Hel22, HMY24, HTWT23, HKF^{+}20, HRC20, JSB20, KPN^{+}18, KF17, KF23, KOC21, Kor21, KPKS22, KHJ^{+}23, KP23, KPJ24, LVT^{+}22, LBT^{+}21, LJML20, LPG^{+}23, LD24, LS16, LF15, LPR21, LLM^{+}20, LBL^{+}21, LQM22, MKWS22, MAC^{+}21, MMS20, MAG24, MGFO23, MGGHG^{+}24, MPAK19]. package [MAS^{+}22, MMCKK21, MP20, MBF20, MTPHH18, MFQ^{+}21, MSV23, MMG19, Mos20, ML18, NKHZ21, NRC23, OT16, OC20, PL23, PDH16, PR19, PLC^{+}24, PF22, PLGHG^{+}23, PKA^{+}22, PMLM23, PDTG17, QTL23, RNR17, RE22, RGP^{+}22, RAS19b, RJDKJ24, SWCP20, SKD22, SBL19, SO23a, SADU23, SR17, Tag23, TS20, TBO24, TTHZ23, THSW24, VEF^{+}23, VRP23, VC18, VV19, WBS22a, WKS23, WKCS24, Xie22, XYC22, YFVL24, YÖYB23, ZTC^{+}21, ZC20, ZL23, ZRK21, ZRK22, dSBSvE23, DET23]. PackageCargo [MFCSÁM20]. packages [CCV22]. packing [BC22, PMFÁM22].

PADC [NIY16]. PADRES [PCL22]. page [HR21]. pair [CJA23, THSW24]. PaleoGeomorphology [HGWM18]. Pan [SO23b]. Pan-Sharpened [SO23b]. Panda [Els23]. panel [Kur23]. PanNDE [SFG21]. PAPARA [MP17]. Parabolized [LH22]. paradigm [AI21]. ParaFEM [BJM23b]. paraffinic [SP19]. ParaGEMS [BJM23b]. parallel [AG21, ATT21, BBB^{+}20, CLM22, HHB^{+}22a, KWZ22, KF17, LGP^{+}22, LLK^{+}20, LQM22, PGA^{+}20, Sal16, SPG^{+}23, Tak22, ZS22, dRB^{+}20, AMWHD23, Par22]. parallelism [AMS^{+}15, HBS^{+}20]. parallelizable [GG21]. Parallelized [CWM^{+}21a, CCE21, DFSW19, TLDM19]. PArallisable [LK21]. parameter [Bac21, NMT23, PLGHG^{+}23, SCPC18, VBA^{+}16, HR21]. parameterizable [BPL^{+}19]. parameterization [KMK^{+}21]. parameters [ASVV23, DSI^{+}20, HFS^{+}21, KEwMO24, LCGA22, YHC^{+}22]. Parametric [PMNWR20, CRV22, FRdN21, PG18, SMKL23, WDZ^{+}20, KMLP^{+}23]. parasweep [Bac21]. ParaView [SMD23]. Parkinson’s [DER^{+}23]. ParKVFinder [dRB^{+}20]. ParSD [Fru21]. parser [RMM18]. ParShift [FSMCFP23]. parsing [LPR21]. Part [FNX22, ddSNX22]. Part2Track [JSB20]. partial [VPVNP22]. Participatory [VEF^{+}23]. Particle [LQM22, RB17, SNHS20, BGGL20, DAC^{+}24, Fru21, HFS^{+}21, HS20, LXK23, LLK^{+}20, Taq16, YYL^{+}22, ZRK19, ZY23, ACPdP22, JSB20, LD24]. Particle-based [RB17]. particle-resolved [DAC^{+}24]. particle/fiber [ZRK19]. particles [HGS17, ANOU21]. particulate [SDH20]. partition [RKDP21, CCI^{+}22]. partitions [CCI^{+}22]. passage [OCK^{+}23]. passing [SU23]. passive [JM20, RLJS23]. passport [ML20]. patch [KOC21]. path [AISM21, ACEAV24, GV20, SS17, VV19]. pathogen [BBHP^{+}23]. Pathology [MBA^{+}22]. paths [Alb19, VV19]. Pattern [BGAGJGC23, FSAD23b, HGG20, KF23]. patterns [LSF^{+}23, Nev17]. PBEE [AANA19]. PC [YBVC^{+}22]. PC-TAS [YBVC^{+}22]. PCA [SGA^{+}16, ZAPS20]. PCA-based [SGA^{+}16]. PCA-derived [ZAPS20]. PCAfold [ZAPS20]. Peacemaker [vDPI^{+}18]. peaks [KCNJ24]. PeakSwift [KCNJ24]. PECT [PS19c]. PÉEK [ZCCGGV^{+}22]. peer [MB23]. peer-to-peer [MB23]. PENalty [BBF^{+}19]. penetrating [XZL24]. peptide [WBS21]. PeptideMind [HH21]. perception [YFVL24]. perceptual [SO23b]. perform [BDSC22]. Performance [EPF^{+}23, KNW^{+}23, VMRVF^{+}24, AMS^{+}15, Afr22, AAB23, AANA19, CZ22, GPW^{+}20, GCW^{+}23, GWK16, JWJ^{+}22, KH19a, KWB22, Kur21, LFGVGDBD24, LFR^{+}20, MB21, Par22, SGBH18, SFG21, SWK19, Sho24, SADU23, SH23, Zaj20, RtHLMN20]. performance-based [AANA19]. performance-ranking [SADU23]. performed [MGMAG22]. Performing [KGF^{+}24, PBL^{+}21, SPG^{+}23]. Peridigm [WHG^{+}23]. Peridynamic [WHP24, WHG^{+}23]. PeriLab [WHP24]. periodic [Kar21, KWZ22]. periodicity [WMZH23]. permanent [MD19]. permeability [WLW23]. permeable [SGB^{+}22]. Permeation [EFTS23, EWF^{+}23]. permutation [CRV22]. permutations [Han20]. Personal [YD20, YDN21, OTdMP22]. personalised [RLN21]. perspective [GRW^{+}19]. pesticide [BSG20]. PESummary [HR21]. PET [KGK^{+}20]. pH [Bre23, BSS^{+}23]. phantoms [KPN^{+}18]. Phase [OYW^{+}21, AT22, DBJ19, KAS23, NR16, NBM^{+}19, NBS^{+}21, NBCC19, TGS^{+}19, YBVC^{+}22, YC19]. phase-change [YBVC^{+}22]. PhaseWare [OYW^{+}21]. phasor [BCR^{+}18, MVRM19, VRBM16, dCWL^{+}23]. phenomena [BBB21]. phenotyping [NHR23]. PHOBOS [LNS15]. phonetic [FSAD23b]. Photoelectron [SC19]. photoemission [SL20b]. photographic [YLS^{+}18]. photographs [MP17]. photon [GCdJAURO21]. photonics [SPMGA^{+}23]. PhotonSTR [GCdJAURO21]. PhotonSTR-18 [GCdJAURO21]. photosearcher [FAM^{+}20]. photovoltaic [PNC23]. phyloDB [LVCF24]. phylogenetic [GL15, LVCF24]. physical [CPUARC20, EHGPT19, IA17, NBM^{+}19, NBS^{+}21, WWB^{+}23, PTS22]. Physics [PMY^{+}24, RWJ^{+}21, MCGK19, Nis20, YMA23]. Physics-constrained [PMY^{+}24]. Physics-informed [RWJ^{+}21]. PhySim [CPUARC20]. PhySim-11p [CPUARC20]. physiological [BBG^{+}19, BSC^{+}24, TCCK23]. physiologically [HRC20]. physiology [BBG^{+}19]. PiCO [FSL16]. Picoastal [SP22]. piezoelectric [TCQOGM23]. PII [DPF^{+}21, FST^{+}21, Web17, Web21, YDN21]. pipeline [DKL^{+}21, MG22, SHvW20, CKC22, HLR^{+}21]. pipelines [APO^{+}20, KPC^{+}20, MBC22]. Piper [OTdMP22]. PIV [LD24, MBF20]. piv-image-generator [MBF20]. PKgui [SHH23]. pksensi [HRC20]. placement [CCBC^{+}21, FWJ^{+}23]. plagiarism [MVG^{+}24]. planar [MBF20]. plane [BFDD^{+}22, FNX22, WMZH23, ddSNX22]. planetary [LRPD18]. planner [Boe23]. planning [AISM21, BSV^{+}23, CNST20, JHAMF19, Kos22, LFT21, WGB16, Web17, Web21, Web23, RAL22]. planning-level [LFT21]. plant [Bed24a, NHR23]. plasma [AHP23]. plasmon [AHR^{+}22, RPV^{+}20]. plates [WMZH23, dFB22]. Platform [BBJ^{+}18, AGADAG^{+}23, AMA19, ASGC^{+}24, ACC^{+}23, BÇ23, BO19, COGP19, CPCCL24, CPV^{+}20, EHGPT19, Elk22, GB24, GZW^{+}22, HKS^{+}22, HKM^{+}19, IAW^{+}15, JHAMF19, JFJM22, KS23, KHH^{+}22, KHM^{+}22, KHJ^{+}23, KGK^{+}20, Kwa19, LMB^{+}19, LHZT23, MHTZ22, MKMK23, MCK^{+}23, ME17, PMMF19, PFJM19, REK^{+}23, SEZLMS^{+}23, SKM19, SCK^{+}23, SV23, SADU23, VJA^{+}18, Zhu15, AMWHD23, AVCP^{+}21]. PlatformCommander [EPF^{+}22]. platforms [EPF^{+}22, Kur23]. play [CK22a, CK23]. Playing [Geo17]. PLC [ROMH22]. plenoptic [HA19]. PlenoptiSign [HA19]. PlgCirMap [Nas20]. plotting [Cer19, JM20]. PlotXY [Cer19]. Plug [PRSS19, CK22a, CK23, GZW^{+}22, LH23, WMZH23]. plug-and-play [CK22a, CK23]. Plug-in [PRSS19, GZW^{+}22, LH23, WMZH23]. Pluggable [KJKK23]. plugin [APK^{+}23, BGAGJGC23, BCD22, ÇA20, CCV22, DJCL23, THT20]. PM [PTT20]. PmlBeta [WBS21]. PND [RWJ^{+}21]. PNW [RLJS23]. PNW-Cnet [RLJS23]. PocketFinderGNN [BOP^{+}23]. Point [STH^{+}21, CLDdM22, HVB16, MGK^{+}20, MIHS21, XYC22, CRP21]. Point-Of-Care [CRP21]. Poisson [GNPP23]. Polar [YLM^{+}24]. Polar-2 [YLM^{+}24]. Polar-2/low [YLM^{+}24]. polarimetry [BBD^{+}22]. polarization [YLM^{+}24]. polarizers [MFQ^{+}21]. policies [ZR19]. Policy [TK23]. Polubarinova [SHH23]. PolyBench [CC21]. PolyBench/GPU [CC21]. polycrystalline [SSP21]. polycrystals [GTC21]. polyelectrolyte [MBM22]. polygonal [Nas20]. polymers [DRPS21, DBPS22, HA23]. polynomials [AKC23, MFCV20]. Polyrun [CK21]. polytopes [CK21]. POMATO [WM21a]. Pony [MSS^{+}23]. pool [Sho24]. populations [IDE^{+}21]. porosity [Kar21]. Porous [FST^{+}21, KRB^{+}20, WLW23, FPBM18]. port [CC21]. portable [ABM^{+}21, TSGOC^{+}22]. position [MLTF^{+}18]. positional [RL19]. positivity [BHHÖ22]. Post [IZBT21, Bac21, Cer19, GYWG23, LDM20, RNK21, RNK22]. Post-Processing [IZBT21, Bac21, Cer19, GYWG23, LDM20, RNK21, RNK22]. posture [GCBDDBF23]. potential [THSW24]. Power [BCR^{+}18, VRBM16, dCWL^{+}23, ATCA20, Bed24a, DRFB^{+}23, GTOM23, JHAMF19, LBT^{+}21, MB21, MVRM19, PS19c, PKA^{+}22, SUM21, VJA^{+}18, YG19, CK22b, WM21a]. PowerCS [RERES23]. PowerDynamics.jl [PKA^{+}22]. powered [AKM20, LEK22]. powerful [Els23, YÖYB23]. PowerSystems.jl [LBT^{+}21]. PowTrAn [ATCA20]. PPGISr [VEF^{+}23]. PPRPA [Taq16]. PPWSim [HTV22]. Practical [NSS^{+}19, Ben15, SA20]. Practically [Rou19]. practice [GZA^{+}22]. Praznik [Kur21]. pre [CAW^{+}20, SO21, ZCCGGV^{+}22, NHR23, ZRK19]. pre-diagnosis [ZCCGGV^{+}22]. pre-existing [CAW^{+}20]. pre-processing [SO21, NHR23]. preCICE [RDH^{+}21, SE23]. Precise [BSV^{+}23]. precision [BJM^{+}19, BJR^{+}20, CCAC22, EB24, MBU^{+}19, REK^{+}23]. preconsolidation [MAAOZMAM22]. predation [VFFS23]. predict [LCGA22]. predicting [HM23, MCK^{+}23, PLK^{+}23, SB22, WMZH23, Xie22, Zie19]. Prediction [HBA^{+}20, BWMS22, FH23, HKC^{+}18, MW23, SAA^{+}23b, SS17, ZLLT23]. predictions [Kle24, KGF^{+}24, PAK22]. predictive [IDE^{+}21, YÖYB23]. predictors [SM19]. PrepAnnECG [Ozk23]. preparation [SEZLMS^{+}23]. prepare [SEZLMS^{+}23]. preprocessing [ASAA20, HHB^{+}22a, LRL23, MBA^{+}22, NHR23, NKHZ21, Ozk23, PF22]. preprocessor [AISM21]. presence [PG18]. preserving [CHP23, GRF23, HTV22, OTdMP22]. Pressure [FH23, MAAOZMAM22]. pressurized [EL17]. PRETUS [GZW^{+}22]. prevalence [CAG^{+}24]. price [DFS24]. Primer [BMBR21]. principles [BGM^{+}24]. printed [Kar21]. prioritization [BGCS19].

Priority [MSDP23, BBHB22, Mic21]. priors [SMKL23]. prisms [LDAL20]. Privacy [HTV22, CHP23, GSS^{+}23, PCL22]. privacy-preserving [CHP23]. Prize [AGG23]. probabilistic [HLG22, GTOM23]. probability [FJ22, HLW^{+}16]. probe [Rub16]. probes [OHO16]. problem [MFCSÁM20, VV19, AGG23]. problems [Afr22, BDF^{+}20, BÁP^{+}22, GRD^{+}23, GLCBHTHO23, KSS23, LSPG22, LQM22, MMS21, SKS23, TOPC23, TSMT19, VMRVF^{+}24]. ProbLog [FLR22]. ProbLog-based [FLR22]. procedure [CNST20, SSSH16]. process [AALW23, AHtH20, Byk19, CBHLG21, DK19, DSI^{+}20, GFRMVL24, GLR22, LHB24, MKT^{+}23, NIP^{+}23, NGOT24, RE22, SvZvdA23, SPMGA^{+}23, VRP23]. processes [AISM21, DL16, FPI^{+}22, HWT^{+}23, OPPZ22, SAA18]. Processing [BAEBAS19, IZBT21, SH19, AFGH22, AHR^{+}22, Bac21, BST^{+}17, BBG^{+}19, BSC^{+}24, CSK19, CdSLCC20, Cer19, VVDV22, GYWG23, Jun21, LEFSO20, LDM20, LMS^{+}16, MCC20, MHTZ22, MMG19, NGOT24, OHO16, PBSB22, PLC^{+}24, QTL23, RNK21, RNK22, RSLL^{+}20, SOS19, SO21, SRML17, SC19, TSNRM24, TCQOGM23, XZL24, Zhu15, NHR23]. processors [VPVNP22]. Product [MDG22, NGOT24]. production [HKC^{+}18, ZTC^{+}21]. products [MAS^{+}22]. professionals [Ozk23]. profiles [LPG^{+}23]. profilometry [OYW^{+}21]. PrognosEase [BB23]. prognosis [BB23, DK19]. program [AS17, Arc20, ABB^{+}19, Bre23, BSS^{+}23, Car21, FSL16, FHA17, GLR22, GTG21, Liu23, MTS^{+}18, NIY16, NL23, Ras21, RRSK18, SFK^{+}19, Taq16, dHGMCGL^{+}24, Tru18]. Programming [CVAP22, BSdG^{+}22, BCD^{+}15, CM19, CFA22, CKC23, Els23, GYWG23, LPL^{+}24, MKU22, OGPSQ22, SGBH18, SVM^{+}23, TSMT19, UdL20, Yaa23, MMS21, TCH^{+}22]. programming-based [GYWG23]. programs [Bel24, CNST20, PBS^{+}24]. project [GBJ23, KKN^{+}22, Kos22, PTT20, SC19, ZLLT23, CFA22]. projection [BBB21, OYW^{+}21, SA20]. projection-based [SA20]. projections [VEM^{+}18]. projects [KPS23]. ProLoaF [GTOM23]. promotion [EHGPT19]. Propagation [Fra22b, EBGCT21, RCT20]. propagators [WBH^{+}23]. Proper [NM20]. properties [KDHG20, KSFG18, MBM22, NBM^{+}19, NBS^{+}21, SLBC^{+}20, YHC^{+}22]. proportion [SR17, SR17]. ProQualAI [NGOT24]. Proscene [Cha17]. ProtectLeaf [VFFS23]. protein [RJH^{+}20]. protein-protein [RJH^{+}20]. proteomic [APO^{+}20, HH21]. protocols [RLK18]. proton [NIY16]. prototype [CPCCL24]. Prototyping [KPC^{+}20, CBLI22, Rag17, WMA24, WSRC24, YD20, YDN21]. PROVER [GBJ23]. PROVER-M [GBJ23]. provides [HDA21]. providing [GMF18]. pRovisioning [HSMF22]. proximal [KPTW19, TK23]. pruned [BB22b]. Pseudonymization [ESE^{+}23]. PSI [SO23b]. PSNR [SD23]. PSPSO [HFS^{+}21]. Psychophysi [NK18]. PTRAIL [HHB^{+}22a]. public [CPF23, IVSV23, ÖA23, VEF^{+}23]. publication [BRF^{+}23, RMM18]. publishing [HYR^{+}23]. publishing/subscribing [HYR^{+}23]. pulse [MD22, SHvW20]. pulses [PS18a, PS18b, PS18c, PS19b]. PuMA [FPBM18, FST^{+}21]. pure [Rag17]. purpose [CSK19, Cer19, HCD24, MSK24, PMZ21]. purposes [AG21]. Py [FPN^{+}22]. PyARPES [SL20b]. PyBanshee [KMLP^{+}23]. pyBaram [Par22]. pyBIMstab [MASB18]. pybliometrics [RK19]. PyblioNet [Mül23]. PYCASO [CWC^{+}23]. PyCiGen [GTC21]. PyCinemetrics [LLP^{+}24]. pycity scheduling [SUM21]. pyCLAMs [ZWCQ22]. pyCLARA [SO21]. pyCM [RSMW20]. pycommute [Kri22]. PycWB [XTD24]. PyDDRBG [AES^{+}22]. PyECOM [GPEM23]. pyEIA [VCLS21]. pyEIT [LYX^{+}18]. pyfao56 [BTDT23, Tho22, TBD23]. PyfastSPM [BRF^{+}23]. pyFDM [WKS22, WS24b]. pyFIRI [ZRK21, ZRK22]. PyFWI [MGFO23]. pyGLLE [MD21]. pyGNMF [LB22]. Pygpc [WPM^{+}20]. PyGran [AM19]. pygrank [KPKS22]. pyGRETA [SO21]. PyIFDM [WKS23, WKCS24]. pyIMD [CMMF19]. PyIVNS [SABEh20]. Pykat [BJR^{+}20]. PyLESA [LFT21]. PyLops [RV20]. PyMaxEnt [SR19]. pymcdm [KSS23, SKS23]. pyMeshFOAM [SC23]. pymetamodels [EBL24]. pymodconn [CJGA23]. PyModPDE [KSS20]. PyMOL [WBS21]. PyOMA [PARS22]. PyOMA GUI [PARS22]. PyOPAsolver [BBHB22]. PyOphidia [EPF^{+}23]. PyPartMC [DAC^{+}24]. pyPcazip [SGA^{+}16]. pyphysio [BBG^{+}19]. Pypiezo [TCQOGM23]. Pypiezo-GO [TCQOGM23]. pyPOCQuant [CRP21]. pyPRIMA [SO21]. PyQBench [JLP23]. PyRamanGUI [BHKK23]. pyrepo [WBS22a, WBS22b, WBNS23]. pyrepo-mcda [WBS22a, WBS22b, WBNS23]. PySide2 [LLP^{+}24]. pySigmaP [MAAOZMAM22]. pysimm [DFC18, DPF^{+}21, FC17]. PySpike [MK16]. PySTPrism [LDAL20]. pySuStaIn [AWO^{+}21]. pyTEP [RE22]. PyTES [XGW23]. pythermalcomfort [TS20]. Python [DFC18, DPF^{+}21, MDR^{+}22, ASAA20, AES^{+}22, AWO^{+}21, ASSS24, AGEABD^{+}24, BW22, BWS23, BB22a, BBHB22, BS21, BJM23a, BST^{+}17, BK19, BLE21, BWMS22, BE20, BT21, Boe23, BSC^{+}24, BB22b, BTDT23, BRF^{+}23, BDQ^{+}22, BJR^{+}20, CGHGRB21, ÇA23, CVV23, CR22, CWC^{+}23, CAG^{+}24, CWGM24, CJGA23, CZ22, CBLI22, CC20, CWM^{+}21a, CWLG^{+}21, CKC23, DSB^{+}23, Dav21, DyH21, DM20, DRFB^{+}23, DFSW19, DGK^{+}24, EPF^{+}23, Els23, EBL24, EAMSR^{+}23, FCMTV24, FSMCFP23, FNX22, FRdN21, FM24a, FPN^{+}22, FC17, FLR22, FF24, GMKRS21, GFRMVL24, GRW^{+}19, GKK^{+}23, Gir21, GP22, GQCP^{+}18, GPEM23, GCBDDBF23, GNPP23, GEH19, GG21, GYWG23, Hac23, HBS^{+}20, HHB^{+}22a, HR20, HM22, HSD^{+}22, HOB23, HBA^{+}20, IHT23, JLP23, JM20, KSS20, KS21, KF23, KG23, Kle24, KMLP^{+}23, KXK^{+}23, KPKS22]. Python [KMK^{+}21, LRPD18, LPG^{+}23, LB22, LPS^{+}24, LYX^{+}18, LYZ^{+}22, LFGVGDBD24, LFT21, MZL^{+}23, MAC^{+}21, MAG24, MBA^{+}22, MGFO23, MCGK19, MSK24, MAS^{+}22, MTS^{+}18, MD21, MD22, MPMC21, MTPHH18, MFQ^{+}21, MSV23, MMG19, MCPG24, Mos20, ML18, MK16, NS22, NRC23, NL23, OE21, OC20, OBWM23, DDT20, Par22, PARS22, PF22, PDS^{+}22, RMM23, RCT23, RNR17, Ras20, RV20, RE22, RMV^{+}24, Roa23, RGP^{+}22, RK19, SR19, SUM21, SDP^{+}20, SMZSW24, SMD23, SAA18, SHTE23, SG24, SABEh20, TS20, TBO24, THG^{+}23, TOPC23, Tho22, TBD23, TF23, TTHZ23, TSG^{+}23, Vel23, Ver22, VC18, VCLS21, dMVBTN24, WSK22, WPM^{+}20, Wet20, WKS22, WS24b, WMM18, XGW23, XTD24, Yaa23, YKKD19, YMA23, ZTC^{+}21, ZAPS20, ZC20, ZWCQ22, ZL23, ZMS18, Zie19, ZRK21, ZRK22, ddSNX22, dPPD^{+}23, vdOJP^{+}20]. Python-3 [MCGK19]. Python-based [Dav21, FNX22, GP22, GCBDDBF23, MCPG24, NL23, OBWM23, Ver22, VCLS21, XTD24, ddSNX22]. Pythonic [DAC^{+}24, ENCS20]. PyTME [MSK24]. PyTOPS [YKKD19]. PyTorch [BOP^{+}23, SH19]. pytrax [TKLG19]. PyVecContour [MZL^{+}23]. PyVisVue3D3 [LPS^{+}24]. pyvrft [BE20]. PyVT [LRL23].

Q [CCI^{+}22, JAMP23]. Q-Facies [JAMP23]. Q-Networks [CCI^{+}22]. Q6 [BBP^{+}18]. QALMA [RLK18]. QBMMlib [BCF20]. QC [CPD^{+}20]. QCOBJ [VC18]. QExpy [MCGK19]. QGIS [APK^{+}23, BGAGJGC23, ÇA20, ÇA23, JJT^{+}23]. QKDNetSim [SCD^{+}24]. QL [FSL16]. QMC [KP19]. QMC-SW [KP19]. qNoise [DI22]. qo [HVH24]. QoE [BCFJ24]. QPSO [LXK23]. QTI [BHHÖ22]. Quadrature [KL22, BCF20]. quadrature-based [BCF20]. Quail [CBLI22]. qualitative [GZS23, NGOT24]. Quality [SO23b, AISM21, AMVB19, AAF^{+}24a, BRF^{+}23, BCFJ24, CJA23, HH21, MRGAGP23, RLK18, RLM^{+}24, SD23]. Quantification [SY20, BEPB24, CAG^{+}24, JZGW22, KMK^{+}21, SSC23, TSG^{+}23]. QuantificationLib [CAG^{+}24]. quantify [CRP21]. quantifying [SMKL23, BDPZ19]. QuantImPy [BT21]. Quantitative [BB20b, CAW^{+}20, JAMP23, LEFSO20, NGOT24]. quantity [VC18]. quantity-aware [VC18]. Quantum [LXK23, OV24, ACC^{+}23, Bjö19, CCH^{+}19, HVH24, JLP23, KP19, Kri22, MDL^{+}18, MD22, PBL^{+}21, SFK^{+}19, SL20a, SCD^{+}24, THSW24, AGC^{+}24]. Quantum-behaved [LXK23]. QuantuMoonLight [ACC^{+}23]. quasi [MD19, SBL19]. quasi-static [MD19]. QUBO [JP24, JP23]. Qudi [BST^{+}17]. queries [FSL16, MP18]. query [PMBM^{+}23]. question [HCD24]. queues [Mic21]. quickly [MKMK23]. QuITO [GRD^{+}23]. QXMD [SFK^{+}19].

R [KCNJ24, AFGH22, Alb15, Alb19, Amo21, ATCA20, AGdSC20, BdSvE23, BSG20, CPD^{+}20, CCV22, COK24, CLK21, CBHLG21, CPB^{+}23, DMFG22, Dra23, DTDd19, EK23, EB24, FAM^{+}20, FVVL23, GZ21a, GZ21b, GWGB22, GMF18, GL15, GAE23, Han20, Hel22, HMY24, HTWT23, HRC20, HYR^{+}23, KF17, KOC21, KHJ^{+}23, KP23, KPJ24, LVT^{+}22, LS16, LF15, LPR21, LLM^{+}20, LBL^{+}21, MGGHG^{+}24, MKMK23, NKHZ21, OCK22, PS22, PL23, PDH16, PLC^{+}24, PLGHG^{+}23, PMLM23, RJDKJ24, SWCP20, SKD22, SA20, SCK^{+}23, SO23a, SADU23, SR17, Tag23, VEF^{+}23, VRP23, VV19, WRV23, Xie22, XYC22, YFVL24, YÖYB23, ZKJ^{+}24, ZGZvB19, dSBSvE23]. R-based [AFGH22, WRV23]. R-GOOSE [HYR^{+}23]. R-package [KHJ^{+}23]. R-peaks [KCNJ24]. R-Shiny [DMFG22, MKMK23]. R-SV [HYR^{+}23]. Rad [LKK23]. radar [NOÖÇ19, XZL24]. radiation [Ohn21, RLK18]. radiative [LKK23, RS22]. Radio [ACEAV24]. radiologists [HKS^{+}22]. Raman [BHKK23, SSM22]. Rambrain [IA17]. random [BC22, Byk19, RJH^{+}20, SA20, TBO24, TKLG19]. RandPro [SA20]. Ranflood [BGM^{+}24]. range [TJS18, TJS19]. rank [AAB23]. RankEval [LMN^{+}20]. ranking [KPKS22, LMN^{+}20, RJH^{+}20, SADU23, dSBSvE23]. RankspeQ [SADU23]. ransomware [BGM^{+}24]. Rapid [WSRC24, BLM^{+}22, YHC^{+}22, VBA^{+}16]. rare [TBO24]. raster [BGAGJGC23]. rate [BlD^{+}20]. ratios [HHN20]. raw [BHE^{+}19]. ray [BBD^{+}22, FRPC23, FRE^{+}24, GTG21, IIH^{+}23, KLY21, KPC^{+}20, LBG20, PFC^{+}18, PKKQ20, PNC23, YLM^{+}24, PKKQ20]. ray-tracing [FRPC23, FRE^{+}24, IIH^{+}23]. raytracing [Gir21]. Rcall [EK23]. RCE [BFF^{+}21]. RCrawler [KF17]. reaction [GV20, GV20]. reactive [MGCS23, iNKN^{+}20]. read [DG21, HHN20]. reading [Rei23]. Real [CWM^{+}21a, XSS20, XSC^{+}21, ZJKS23, AKC23, AVB17, AM23, Bat19, BHB^{+}21, CLM22, GACS^{+}23, GZW^{+}22, Jur23b, LHZT23, MB21, MVRM19, RGdlTQ^{+}23, SHvW20, TJS18, TJS19, YKC^{+}19, YDMC15, ZRBCI20, HESH19]. Real-estate [HESH19]. Real-space [XSS20, XSC^{+}21, ZJKS23]. Real-Time [CWM^{+}21a, AVB17, AM23, BHB^{+}21, CLM22, GACS^{+}23, GZW^{+}22, LHZT23, MB21, MVRM19, RGdlTQ^{+}23, SHvW20, TJS18, TJS19, YKC^{+}19, YDMC15, ZRBCI20]. real-world [Jur23b]. Realistic [CT24, Deg20, Liu24]. Reality [HKM^{+}19, SDCA19, TMGW23, BSVP20, HVCS^{+}23]. RealModel [MMT24]. RealPi2dDIC [DEV^{+}21]. receiver [LYZ^{+}22]. recognition [BOP^{+}23, FSAD23b, MML20, SB22, YG19, ZY23]. recognizer [KF23]. reconstruction [KPN^{+}18, KPTW19, KWB22, KPC^{+}20, KGK^{+}20, LFP^{+}23, MBC22, MMG19, MBY22, SR19, SNLFGVRB24]. record [GZA^{+}22]. recorded [BL16]. recorder [LSSK16]. recording [SA17]. recordings [OTdMP22]. red [GMGG19]. redistribution [dlSVBLdA^{+}17]. Reduced [AMWHD23, SBCK17, TCQOGM23]. reduction [HHB^{+}22b, MTPHH18, MSV23]. refactoring [GCW^{+}23]. refellips [RGP^{+}22]. Reference [BE20, WBS22a, WBS22b, WBNS23, Jur23b]. Refinement [SEdCCCRG23, RDO^{+}19]. reflective [ACO21]. reflectivity [DLH18, PKKQ20]. Reforming [NIP^{+}23]. REFRAME [MGCS23]. reg [PM23]. reg-sgc [PM23]. Region [Kar21]. regional [LPG^{+}23]. regions [SO21]. RegionTPMS [Kar21]. registration [BEPB24, GJS^{+}22]. Regression [BV19, VRP23, BGAGJGC23, SIC24]. regular [Agu23, PMLM23]. Regularisation [KPTW19]. regularized [PM23]. REgulation [PCL22, TT17]. Rehab [HVCS^{+}23]. Rehab-Immersive [HVCS^{+}23]. rehabilitation [HVCS^{+}23]. reinforced [DRPS21, DBPS22, dQRRBdSM20, WSRC24]. Reinforced-lib [WSRC24]. reinforcement [CT24, TK23, WSRC24, dPPD^{+}23, SE23]. related [BBP^{+}18, BO19, HWT^{+}23, PS19c, PPBZ21, YHC^{+}22]. relational [PMBM^{+}23]. relations [KF23]. relationships [TBO24]. relative [MTS^{+}18]. relaxation [BBF^{+}19, BBN^{+}22, WMM18]. RelaxPy [WMM18]. Relays [TDG19]. release [ACRSGJ23]. relevant [SAC^{+}21]. reliability [NSS^{+}19]. Reliable [MP18]. remapping [GKK^{+}23]. remote [APK^{+}23, CAW^{+}20, CMdS^{+}23, ZPM23]. REMP [BMBR21]. Renamer [FM24b]. renaming [FM24b]. rendering [HMCA15, JFJM22]. Renewable [ICS^{+}24, JHAMF19]. Renewables [GMKRS21]. Renku [SGDC22]. ReNView [GV20]. repair [Ars21, dHGMCGL^{+}24]. Repast [Giu19]. repeated [YLS^{+}18]. replay [LMM22, LMM22]. replicate [HH21]. replication [ASRI22]. replicator [Arc20]. RepoFS [SS19]. report [EB24]. reporting [HKS^{+}22, NFASC21]. repositories [ÖA23, SS19]. repository [DAB^{+}19, Jur23b]. Representation [ALRM21, SDP^{+}20, KMT23]. reproducibility [HDA21]. Reproducible [AMVB19, BLE21, CPB^{+}23, FM24a, FAM^{+}20, MBA^{+}22, SGDC22]. reprogramming [IVSV23]. Republic [BGCS19]. requirements [CNST20, HMC^{+}24, TSNRM24]. Research [HKM^{+}19, O’D21, AMVB19, AG21, APO^{+}20, BSRSC18, BCTMA^{+}23, BCFJ24, ESE^{+}23, EPF^{+}22, FVVL23, GZW^{+}22, HKC^{+}18, Jan24, Jur23b, KHM^{+}22, KLL23, Par22, RSLL^{+}20, ROMH22, SLL23, TS20, WLP16]. researchers [TCCK23]. resection [BSV^{+}23]. ResEntSG [KHG21]. reservoir [MGFO23]. reservoirs [SCSB24]. residual [RSMW20]. resilience [BD22, KPS23, Ull16, ZTC^{+}21]. ResiPy [ZTC^{+}21]. ResMapper [BD22]. resolution [MRMD19, Tem23, PZ22]. resolved [DAC^{+}24, GTG21, JSB20, SL20b]. resonance [AHR^{+}22, BBB21, BBN^{+}22, RPV^{+}20, WDZ^{+}20]. resources [LMB^{+}19, TSNRM24]. Response [CCI^{+}22, FRdN21]. responses [SDP^{+}20]. ReSS [SAC^{+}21]. RESTful [BBS^{+}22].

Restoration [KHG21]. REstriction [BMBR21]. results [Alb15, GLR22, Mai22, RFP^{+}24]. rethnicity [Xie22]. Reticula [BMK23]. Retrieval [AB24, GGI23, LPR21, OYW^{+}21]. RETSManager [HESH19]. reusable [MKU22]. reuse [AAF^{+}24b]. Reversible [CCE21]. review [GBW^{+}24, NMLM18]. reviews [AAF^{+}24a, FVVL23, HSD^{+}22]. Rewriting [HMCA15]. Reynolds [YMCB23]. RGB [CCG^{+}23]. RGB-Depth [CCG^{+}23]. rheed [Dan20, Dan22]. rheological [HHH22, HHH22]. Rheological-Data [HHH22]. rheometer [HHH22, MM16]. rheometer-inherent [HHH22]. Rhizomer [GLGG22]. RI [KCP^{+}24]. Ricgraph [Jan24]. rich [Cha17]. Ridership [IVSV23]. ridesharing [MW23]. Ridge [HHB^{+}22b]. Riemannian [KHG21]. Riesz [BBB21]. right [KM21]. right-hand [KM21]. RINEX [GZS23]. RINEXAV [GZS23]. ring [Liu23]. riparian [CVV23]. RISC [MG22]. RISC-V [MG22]. Risk [EL20, GPK^{+}21, PCdJSdFR^{+}23, SCSB24]. River [SLPR20, BNC24, CVV23]. rivers [LD24, OCK^{+}23, PDM22]. RML [AGEABD^{+}24]. RML-FNML [AGEABD^{+}24]. RmSAT [NOÖÇ19]. RmSAT-CFAR [NOÖÇ19]. Road [Can24, AMA19, BD22]. Roborueda [GCBDDBF23]. robot [AISM21, AOGC^{+}20, Bed24a, Geo17, MMS20, YDMC15, ACRSGJ23, LD24]. robotics [AKM20, BCTMA^{+}23, BSVP20, Els23]. robots [MBDS20, SBT22]. Robust [HCS^{+}20, Gag21, JMT22, LKK24, NKHZ21]. robustDEA [LKK24]. robustness [FRNW^{+}24]. ROC [GZ21a]. rock [Ras20, YMA23]. rocking [Por23]. rockphypy [YMA23]. rocks [MMT24]. rolling [PMLM23]. ROOT [Nis20]. ROS [LD24, MMS20, MBY22]. rotating [JAC20, JAC21]. rotation [PI17, YYL^{+}22]. rough [LH23]. roughness [OJ22]. Routing [AGG23, SBT22]. RPaSDT [Fra22b]. RPExpand [BBB21]. RS [CMdS^{+}23]. rsdtlib [ZPM23]. rsppfp [VV19]. RT [RGdlTQ^{+}23]. RTransferEntropy [BDPZ19]. rubber [MM16]. Ruby [Rag17]. RufGen [LH23]. rule [MKWS22, SKM19]. rule-based [MKWS22, SKM19]. Rules [SSM22, PZKK21]. RuleXAI [MKWS22]. Rumor [Fra22b]. run [FWB^{+}21]. Runner [EHGPT19]. running [MHTZ22]. runs [AAA^{+}21]. Runtime [SCG^{+}20, CBPC23, FSL16]. RuralSpeedSafetyX [DW20]. Rust [Bel24, ABC^{+}20]. rust-code-analysis [ABC^{+}20]. rwl read [Liu23]. RXMD [iNKN^{+}20]. RxpsG [SC19].

S [RLM^{+}24]. S-CI-OWA [RLM^{+}24]. S2352711016300152 [Web17, Web21]. S2352711016300395 [DPF^{+}21]. S2352711018300281 [FST^{+}21]. S2352711019303334 [YDN21]. SA [AK24]. SA-MEAS [AK24]. SADEAT [AS18]. safety [DW20, GACS^{+}23, SIC24]. SAGA [MWJ15]. SAML [HTB19]. sample [CLK21, Tag23]. sampled [GEH19, SH19]. samplers [BCA19]. samples [CBHLG21, LBG20]. sampling [CK21, LMS^{+}16, NSLD16, WBH^{+}23, WC23]. sap [OHO16]. SAR [BNC24]. SAS [RMM18, HDA21]. Sas-temper [HDA21]. satellite [HJRBM23, RKA^{+}23, SDL21]. saturation [WLW23]. saving [HHH22]. Savu [KWB22]. SBTAnalyzer [QTL23]. Scaffolding [RLN21]. scaffolds [Kar21]. ScaFi [CVAP22]. Scala [CVAP22, RLF^{+}21]. Scalable [ENCS20, AGADAG^{+}23, KBB19, Krä20, MPX^{+}24, Mos20, iNKN^{+}20, RV20, VDP^{+}19, HGWM18, TDG19]. Scale [BO19, AMVB19, Bed24b, BFRK19, BD22, COG19, DZZ^{+}22, EPF^{+}23, GMNG^{+}18, KSA23, LBT^{+}21, LVCF24, YYL^{+}22]. scale-based [GMNG^{+}18]. scales [BSG20]. scaling [MYKT22, RRSK18]. scan [MFJ23]. scanning [Rub16]. scattering [HDA21, HHB^{+}22b, SF16]. scattering-integral-based [SF16]. scbursts [DTDd19]. scenario [KPM^{+}22]. scenarios [PBM22]. scene [YG19]. scheduling [MB23, SUM21, dPPD^{+}23]. Schema [HKZ24]. Schema-based [HKZ24]. scheme [GNPP23]. schemes [RXZ^{+}23]. schlably [dPPD^{+}23]. Schrödinger [MD22]. Science [BFF^{+}21, FWB^{+}21, SCG^{+}20, AHtH20, BBG^{+}19, GM24, LFGVGDBD24, MYKT22]. Scientific [HKM^{+}19, DAB^{+}19, FVVL23, GY23, Sho24, ESE^{+}23]. Scilab [CGKD23]. Scopus [RK19]. SCORE [HMC^{+}24, MCK^{+}23]. scoring [MLTF^{+}18]. SCOUT [SY20]. SCRAMBLE [O’D21]. scraping [KF17]. Scratch [SDG^{+}24]. screening [BMBR21]. script [HL22]. Scriptable [RK19]. Scripts [GTC21]. SCSV IPS [JAC21]. sDNA [CC20]. SDPTool [PAK22]. sea [MP17]. Seamless [KPSM17, Els23, JFJM22, SHTE23, THG^{+}23]. Search [FBA^{+}22, She19, CDAOB^{+}20, GGI23, KRA23, LSPG22, XTD24]. searching [LEK22]. SEBASTiAn [PRC^{+}23]. SEC [LPR21]. SEC2MWD [MMD24]. secml [PDS^{+}22]. second [AAA^{+}21, Tak22]. section [Ohn21]. sectional [AZ17]. sections [AS17, dQRRBdSM20]. Secure [PDS^{+}22]. security [HMC^{+}24, KJKK23, PRC^{+}23, PCL22]. SED [Dav21]. SEDBYS [Dav21]. sedimentary [MMT24]. sediments [GBJ23]. seedling [CCG^{+}23]. SEG [SC17]. SEG-Y [SC17]. segmentation [BEPB24, JC23, RKDP21, UVPB^{+}22]. Segy [SC17]. Segy-change [SC17]. seismic [ACR22, AS17, BD22, Elk22, KSA23, MMCKK21, NL23, SF16]. Seismo [KSA23]. Seismo-VLAB [KSA23]. selecting [PPRE17]. Selection [SSM22, CMACMÁT23, GWGB22, GZS23, HFS^{+}21, Kur21, LIZ^{+}20, NBCC19, PDM22, REFB17, YG19]. self [NFASC21]. self-reporting [NFASC21]. sEmanNTIc [HSMF22]. Semantic [TCH^{+}22, CM19, FVD20, GLGG22]. semi [CLM^{+}20, FS19, FVD20]. semi-analytical [FS19]. semi-automated [CLM^{+}20]. SENinja [BCD22]. sensed [SOS19]. sensing [APK^{+}23, CMdS^{+}23, ZPM23, RPV^{+}20]. sensitive [DG21]. sensitivity [HRC20, SF16, WPM^{+}20]. sensor [CCBC^{+}21, GJS^{+}22, HTV22]. sensors [CCG^{+}23, DTD17, Sur20]. sentence [CJA23]. sentiment [ASRI22]. SentinAir [Sur20]. separately [AMH23]. SepINav [STH^{+}21]. sepsis [STH^{+}21, STH^{+}21]. Sequence [PMP16, AS17, BSB20, CJGA23, LVCF24, TBCG21]. sequence-to-sequence [CJGA23]. sequences [BCA19, HM22, KSFG18, WBS21]. sequencing [AJK^{+}23, CdSLCC20, HHN20]. sequential [HVB16, BCA19]. serialization [DRM20]. SerialTrack [YYL^{+}22]. Series [BFF^{+}20, FBA^{+}22, ASAA20, Amo21, BDPZ19, CRV22, VVDV22, EAMSR^{+}23, FMFCEG^{+}23, NKHZ21, PR19, PMLM23, RKA^{+}23, SM19, WSRMB23, vKMH^{+}20]. services [MW23]. set [SEZLMS^{+}23]. sets [SAC^{+}21]. setups [CWM^{+}21a, WCW^{+}23]. severity [MCK^{+}23, SB22]. SfM [MBC22]. SFRAT [NSS^{+}19]. sgc [PM23]. ShakerMaker [ACR22]. shallow [BFDD^{+}22, SH23]. Shape [Bla21, Bla23]. shapes [VRVAC23]. shaping [SLM23]. share [IAW^{+}15]. Sharpened [SO23b]. shear [JAC21]. shells [FS23]. SHEMAT [KRB^{+}20]. SHEMAT-Suite [KRB^{+}20]. Sherpa [HCS^{+}20]. Shiny [DMFG22, MKMK23, SCK^{+}23]. Shiny4SelfReport [NFASC21]. ShinyGAStool [HMK23]. shinyHugePlot [Tag23]. Shk [KZ18]. Shk-9 [KZ18]. shock [RXZ^{+}23]. shock-capturing [RXZ^{+}23]. short [DG21, LXK23, MT19]. short-lived [LXK23]. Short-Time [MT19]. shortCardiac [RHR^{+}23]. shortening [FG20]. shortest [VV19]. shotgun [HH21]. shRNA [KSP19]. SIAR [dlSVBLdA^{+}17]. SIC [PMFÁM22]. side [KM21]. SIFM [NK20]. Signal [BAEBAS19, AHR^{+}22, BBG^{+}19, BO19, BSC^{+}24, COK24, LSSK16, MRMD19, PMM16, SEZLMS^{+}23, SY20]. signals [SOS19, TCCK23]. signature [JS19]. significance [BW22, BWS23]. SiLA [BLM^{+}22, LHB24]. Silence [BBS^{+}22]. silico [ME17]. Sim [MBDS20, RERES23]. Sim-PowerCS [RERES23]. SimApi [PMMF19]. SimCES [KHH^{+}22]. similarity [FMFCEG^{+}23, LSB^{+}20, ZC20]. SimilarityTS [FMFCEG^{+}23]. Simple [GMGG19, BS21, CKC23, GBJ23, KKAUA21, KP19, TKLG19, YHC^{+}22, PM23]. simplified [RERES23, SSM22]. simplifies [ACR22]. simplify [GQCP^{+}18, BB22b]. Simplifying [AGHK21, GCP22]. SimPrily [GQCP^{+}18]. Simu5G [NNDS23]. simulate [BBHP^{+}23, BSG20, CGKD23, Deg20, EBGCT21, MML23, MD19]. Simulating [DL16, GPEM23, KRB^{+}20, KP23, ME17, PBM22]. Simulation [AGHK21, BBJ^{+}18, CPUARC20, KGF^{+}24, XSS20, dSBSD17, ACR22, AM19, AVCP^{+}21, BBD^{+}22, CH23, Can22a, Can24, CS21, CK22b, CNST20, DAC^{+}24, Dan20, DFC18, DPF^{+}21, DRFB^{+}23, DWO^{+}22, FPI^{+}22, FRPC23, FRE^{+}24, FC17, FRNW^{+}24, GSP^{+}17, GSF21, HVB16, HS20, IIH^{+}23, JWJ^{+}22, KHH^{+}22, KXK^{+}23, KHJ^{+}23, LGP^{+}22, MG22, MBDS20, MFJ23, MB21, NIP^{+}23, NBM^{+}19, NBS^{+}21, PMMF19, PKKQ20, PMM16, PGA^{+}20, PS18a, PS18b, PS18c, PS19a, PS19b, RDH^{+}21, RRS18, SFG21, SGA^{+}16, SDH20, SV23, SL21, SBJ^{+}23, THSW24, WSAC^{+}23, YLM^{+}24, YKC^{+}19, XSC^{+}21, ZJKS23, vdOJP^{+}20]. simulations [AMS^{+}15, BCR^{+}18, BS19, CWM^{+}21b, DyH21, DS20, GQCP^{+}18, GCP22, HA23, KWZ22, KSA23, LFR^{+}20, MKMK23, MYKT22, NR16, OT16, PBL^{+}21, RE22, RAS19b, SC23, SSP21, SPG^{+}23, SB22, THG^{+}23, VRBM16, dCWL^{+}23, SNHS20]. Simulator [BCTMA^{+}23, CVD21, OV24, BSVP20, DV23, HTV22, MVRM19, NS22, Nev17, iNKN^{+}20, RERES23, RGdlTQ^{+}23, SCD^{+}24, SACT23, Ull16, Zhu15, vKH20, KJW21]. simulators [AFA22]. Single [KCNJ24, AFA22, CLM^{+}20, CMMF19, DTDd19, IHT23, QTL23, SHvW20, SR17]. single-bacterium [QTL23]. single-cell [CLM^{+}20]. single-channel [DTDd19]. single-molecule [IHT23]. single-tree [AFA22]. singular [LSMG19, LAA23]. SIP [YKC^{+}19]. SIR [SL18]. SIRR [AMC17]. Sismic [DM20]. site [BMBR21]. site-directed [BMBR21]. SiTree [AFA22]. situ [DEV^{+}21, SPG^{+}23]. SIVR [JFJM22]. Size [MT19, CLK21, Fru21, MMD24, MAGM^{+}23, Tag23]. Sizes [Ras21]. skin [YDMC15]. Skinware [YDMC15]. skull [LFP^{+}23]. Sleep [MLTF^{+}18, EDZ^{+}23]. slender [AZ17]. Slitflow [IHT23]. slope [Alb19, MASB18]. slope-dependent [Alb19]. SLSL [LXK23]. SLSL-QPSO [LXK23]. slurry [HM23]. small [HDA21, HHB^{+}22b, SB21]. small-angle [HDA21, HHB^{+}22b]. smart [CK22b, Gór24, HSMF22, LFT21, MKS^{+}23, ROMH22, JCFM24]. SMART-TeleLoad [JCFM24]. smartgrid [PMMF19]. smartphone [MHTZ22, PFJM19]. SmarTS [Gór24]. SMEs [RBQ^{+}24]. smoothed [LLK^{+}20, LQM22]. SMS [CUSRCP^{+}22]. SMS-Builder [CUSRCP^{+}22]. smuggler [FM24a]. Smyrida [MKT^{+}23]. Snap4City [BFNP24]. SnapperML [MCPG24]. SNEToolkit [KDRT23]. SNNTool [WC23].

SnP [FSAD23b]. So-called [RLM^{+}24]. soccer [MML20, MML20]. social [AKM20, Sha23, YLS^{+}18]. social-ecological [Sha23]. soft [MBDS20]. SoftCtrlDC [NCGLRR24]. SoftCtrlDC-M [NCGLRR24]. Software [AOGC^{+}20, BBB21, Bla21, Bla23, CKM21, DBJ19, DSI^{+}20, EL20, HKC^{+}18, HDA21, KLY21, LCMD^{+}18, LMS^{+}16, MDG22, MBU^{+}19, Mül23, PPRE17, RLM^{+}24, SA17, SD23, SLWS^{+}17, WBS22b, WBNS23, YDN21, dlSVBLdA^{+}17, AFGH22, ANA16, ACO21, AAL^{+}22, AMH23, AI21, ACNF22, BMK23, BSRSC18, BMBR21, BFV18, BJM^{+}19, BGGL20, BOP^{+}23, Bjö19, BO19, BWMS22, BDSC22, CUSRCP^{+}22, CCC^{+}21, CWGM24, CWLG^{+}21, DDCSH23, DWO^{+}22, DPD22, EPF^{+}22, FH23, FPBM18, FST^{+}21, FSL16, FHA17, GZS23, GRD^{+}23, GMGG19, GNPP23, GMNG^{+}18, HHH22, HM23, HJRBM23, HLP^{+}19, HWST22, IAW^{+}15, JAC21, JZGW22, Jun21, KG24, KKN^{+}22, KSS20, KS21, KKAUA21, KPN^{+}18, KWB22, KRA23, KPOD16, KEwMO24, KMIG20, KPS23, KSA23, LEFSO20, LHB24, LJML20, LKSS20, LSB^{+}20, LBG20, LYZ^{+}22, LHY^{+}23, MSDP23, MP17, MPAK19]. software [MP20, MBF20, MFJ23, MGCS23, MGMAG22, MAGM^{+}23, MRGAGP23, MASB18, MAAOZMAM22, NSS^{+}19, NMLM18, NK20, NCGLRR24, NGOT24, NM20, NR23, OKJ23, ÖA23, PAK22, PMMF19, PARS22, PCdJSdFR^{+}23, PKKQ20, PS19a, PM23, PBL^{+}21, PDTG17, QTL23, RWJ^{+}21, dQRRBdSM20, RJH^{+}20, RDH^{+}21, SR19, STH^{+}21, SOS19, SHH23, SP19, SMR22, SGDC22, SDG^{+}24, Sur20, SBT22, SLM23, TSGOC^{+}22, TCQOGM23, WLP16, WBS22a, WGB16, Web17, Web21, WW17, WC23, XZL24, YB23, YSM22, ZAPS20, ZRK21, ZRK22, ddSNX22, dFB22, AJR^{+}23, CK22b, CLM^{+}20, FG20, FNX22, HHN20, KF22, KPS23, KGK^{+}20, NSS^{+}19, NIP^{+}23, ND22, Ras19a, SKM19, YD20, ZRK19]. SoftwareX [CK23, EWT^{+}24, JP24, RNK22, Ric19]. Soil [AMC17, CPB^{+}23, dlSVBLdA^{+}17]. soils [MAAOZMAM22]. soiltestcorr [CPB^{+}23]. solar [PNC23]. solar-thermal [PNC23]. SOLDIER [SIC24]. Solid [GSF21, SP19, BJM23b, WFD23]. solids [MFJ23]. Soloff’s [CWC^{+}23]. solution [Afr22, ACO21, AGG23, CCG^{+}23, GEG22, KJKK23, iNKN^{+}20, OCK22, SIC24]. solutions [Azi23, FS19, MP18, SHH23, TF23, VPVNP22, WSRC24]. solve [GNPP23]. solver [AK24, AG21, BBB^{+}20, BÁP^{+}22, BFDD^{+}22, HOM^{+}18, LH22, LK21, LSPG22, LXJ^{+}23, MB21, MMS21, Par22, SL18, SWK19, SAA^{+}23a, SRK23, Wil17, LSPG22, MSDP23]. solvers [JP23, JP24]. solveTruss [Ozb17]. sOlving [YMCB23, BDF^{+}20, GLCBHTHO23, KSS23, Kor21, KM21, MD21, SKS23, TTT19]. somatosensory [MMS20]. SonarQube [CCV22, dHGMCGL^{+}24]. soot [SBJ^{+}23]. SootLib [SBJ^{+}23]. SOQCS [OV24]. sorting [DTDd19, DPD22]. SOUL [RLF^{+}21]. Sound [KOZ^{+}24, BMR17]. Source [Fra22b, KL22, KPS23, MVBF19, AHtH20, ACPdP22, AGL^{+}23, ABC^{+}20, ASVV23, BSdG^{+}22, BÇ23, BSRSC18, BDF^{+}20, Bed24b, BGGL20, BK19, BC20, BBB^{+}20, BL16, BHKK23, BDSC22, BTMB21, CR22, Cer19, CBLI22, CBS^{+}16, DEV^{+}21, DAB^{+}19, DS20, Elk22, EPF^{+}22, EDZ^{+}23, FNX22, FK23, GZS23, GWBF23, GB24, GNPP23, GV20, GMNG^{+}18, HGG20, HT18, HYR^{+}23, ICS^{+}24, IAW^{+}15, JWJ^{+}22, JSG22, KKN^{+}22, KH19a, KRB^{+}20, KK24, KPM^{+}22, KSA23, LEFSO20, LK21, LSB^{+}20, LFP^{+}23, LQM22, MVG^{+}24, MP17, ML20, MAS^{+}22, MMCKK21, Mic21, MKR^{+}21, MRGAGP23, MML23, MAAOZMAM22, MBY22, NFASC21, OKJ23, ODE^{+}22, OHO16, OAF20, OE21, OBWM23, PL23, PCPC24, PFC^{+}18, PM23, PNL^{+}21, PKA^{+}22, PBL^{+}21, RHR^{+}23, REK^{+}23, RJN^{+}23, RMM18, RERES23, ROMH22, RSMW20, Rub16, RRSK18, SFK^{+}19, SG24, SDH20, SMR22]. source [SACT23, SH23, TSGOC^{+}22, TOPC23, VRG19, WGB16, Web17, Web21, WDZ^{+}20, WTAG^{+}22, ZRK21, ZRK22, ZLLT23, ddSNX22, AGVM20, KWB22, KPS23, RAL22, UVPB^{+}22]. source-meter [OKJ23]. Source-o-grapher [KPS23]. source-to-source [BC20]. SourceMonitor [ÖA23]. sources [Duc23]. SP [SP19]. SP-Wax [SP19]. SPAALUV [AJR^{+}23]. space [BHHÖ22, CNST20, Has18, KBB19, KPJ24, LDAL20, MP18, SNW^{+}21, SHK^{+}23, XSS20, XSC^{+}21, ZJKS23, ÇA23]. space-time [KBB19, LDAL20]. spanning [ÇA20]. SPARC [XSS20, XSC^{+}21, ZJKS23]. sparse [LRPD18, SOS19, WBH^{+}23, WBH^{+}23]. sparse-ir [WBH^{+}23]. spatial [BGCS19, CZ22, CC20, GEH19, PS19c, PZ22, SLM23, YFVL24, KDRT23]. SpatialCluster [RMV^{+}24]. spatially [BSG20, ZR19]. spatio [Azi23, LRL23]. spatio-temporal [Azi23, LRL23]. Speaker [LVK21]. specialization [NBCC19]. species [BSG20, KRB^{+}20, RLJS23]. Specific [dlSVBLdA^{+}17, HHH22, Kri22, LHB24]. spectra [BHKK23, IIH^{+}23, PS18a, PS18b, PS18c, PS19a, PS19b, dSBSD17]. SpectraFox [Rub16]. spectral [KAK21, KWZ22, SAD22]. spectral-boundary-integral [KAK21]. SpectraSorter [DPD22]. spectrograms [MRMD19]. spectrometer [DPD22]. spectrometers [LJML20, Web23]. Spectrometry [CRF21]. spectroscopic [RGP^{+}22]. spectroscopies [SL20b]. Spectroscopy [SC19, BDSC22, EWT^{+}23, EWT^{+}24, GCdJAURO21, JJT^{+}23, KG24, Rub16, SSM22]. spectrum [DPD22, LSMG19]. Speech [FSAD23b, BL16, DDCSH23]. SPGM [HGWM18]. SPH [LQM22]. SpharaPy [GEH19]. sphere [ANOU21, BC22, Hac23]. sphstat [Hac23]. spike [MK16]. spiking [RGdlTQ^{+}23]. spin [JP23, JP24, JAC20, JAC21, MFQ^{+}21]. spin-glass [JP23, JP24]. spin-polarizers [MFQ^{+}21]. spin-up [JAC20]. Spine [KPM^{+}22]. SpinUpFlowDescriptor [JAC20]. split [GLR22]. splitting [KPTW19]. sports [MML20]. spot [ÇA23, DAB^{+}19]. spray [BSG20, HS20]. spray-drift [BSG20]. spreadsheet [SKM19, LSPG22]. Squeeze [YMCB23]. SRC Num TDOA [BL16]. SSA [LSMG19, LSMG19]. SST [KJKK23]. StabilisatION [PDM22]. Stability [LH22, EB24, MFCSÁM20, MASB18, RXZ^{+}23, SRK23]. Stable [dlSVBLdA^{+}17]. stack [Bed24a, KMT24, THT20]. StaDyn [OGPSQ22]. Stage [AWO^{+}21, MMS21]. stain [TGS^{+}19]. stain-free [TGS^{+}19]. standalone [OT16]. standard [ASVV23, OE21, SEL^{+}16, DZZ^{+}22]. Standardized [GAM^{+}24, GMF20, MWJ15, RHR^{+}23]. Star [YLM^{+}24]. Star-XP [YLM^{+}24]. Stars [Dav21]. starting [MYKT22, TT17]. Starvation [YMCB23]. State [TS22, FRPC23, FRE^{+}24, MSV23]. state-of-the-art [MSV23]. statechart [DM20]. states [KLY21, KP23, LEK22, PS19c]. Static [Ozb17, AES^{+}22, BT23, KDHG20, MD19, PRC^{+}23, Por23, SBL19]. Station [KP20, PBM22]. stationary [BAEBAS19]. Statistical [ZKJ^{+}24, BWMS22, CRV22, HVB16, HR20, MSV23]. Statistics [RNG^{+}21, Hac23, LVT^{+}22, RCT23]. StatMechGlass [BWMS22]. status [BBJ^{+}18]. steady [SHH23]. Steganalysis [DA24]. stegomalware [CSZM22]. stencil [GÓ19]. step [MFCV20]. Stepwise [SSSH16]. stereo [BL16]. stereoscopic [KEwMO24]. STFT [MT19]. STFT-FD [MT19]. stiffness [WMZH23]. stimulating [AOGC^{+}20]. stimulation [XGW23]. StimVis [Kul20]. STL [SGP23]. STL2MOOSE [SGP23]. STL Process [AISM21]. STM [HM23]. Stochastic [OV24, SPG^{+}23, CR22, KPOD16, OE21, TFCZ24, AS18]. stock [LS23]. Stokes [AG21, LGG^{+}23, SAA^{+}23a, Wil17]. storage [RIHVP23]. store [Krä20]. stored [HM23]. storing [FHB^{+}21]. Stplanpy [Boe23]. strain [AAN17, CWGM24]. strategic [MKS^{+}23]. strategies [GSP^{+}17]. strategy [Arc20, IIS18, KFA^{+}24]. stratified [MPAK19]. stratigraphic [AS17]. stream [MHTZ22, SCSB24]. streaming [SM19]. streamline [GKK^{+}23, GBW^{+}24]. streamlining [SAA^{+}23b]. streams [BRF^{+}23, HT18]. street [SMZSW24]. stress [FH23, RSMW20]. STRON [BFV18]. structural [CCBC^{+}21, LS23, WKR^{+}20, YHC^{+}22]. Structure [GRF23, HOM^{+}18, BWMS22, KGF^{+}24, RRSK18, SC23, DZZ^{+}22]. Structure-preserving [GRF23]. Structured [ALRM21, HKS^{+}22, PLK^{+}23, ST21]. structures [AZ17, Bel24, CLDdM22, KDHG20, KBB19, OT16, Por23, SSC23]. strut [OJ22]. StrutSurf [OJ22]. student [MSLNSMOG24]. studies [AALW23, BLE21, CUSRCP^{+}22, EB24, GGI23, HMK23, LLP^{+}24, SGDC22]. STUDIO [SMZSW24, CMACMÁT23]. study [AMA19, FSMCFP23, IAW^{+}15, TLDM19]. studying [LKSS20, NIY16]. SUB [PRSS19]. subjected [dFB22]. subscale [MVBF19]. subscribing [HYR^{+}23]. Subsequence [FBA^{+}22]. subsidence [PRSS19]. substances [Zie19]. subsurface [LBH^{+}20]. Subtype [AWO^{+}21]. succinct [Bel24]. suit [KMIG20]. suite [BST^{+}17, CC21, Gór24, SO21, TS21, WW17, ZS24, KRB^{+}20]. suites [MSLH23, SVM^{+}23]. SULISO [WW17].

Summary [HR21]. super [Tem23].

supercomputers [AMS^{+}15]. Superscalar [BCD^{+}15]. supply [AGL^{+}23]. support [DW20, DER^{+}23, HVCS^{+}23, HWT^{+}23, MCGK19, MFCSÁM20, PMBM^{+}23, PZKK21, RJH^{+}20, SO23a]. supported [RJH^{+}20]. Supporting [HBS^{+}20, GMF18]. supports [VMS^{+}23]. surface [AHR^{+}22, LCMD^{+}18, LH23, OJ22, RPV^{+}20]. surface-based [LCMD^{+}18]. surfaces [Alb19, Kar21, QTL23, TSGOC^{+}22]. SurfRCaT [CAW^{+}20]. surgical [BSV^{+}23]. surreal [Rou19, Rou19]. survey [MTS^{+}18, PFJM19, YLS^{+}18, BLE21]. surveys [Bed24b]. sustainability [RBQ^{+}24]. Sustainable [AMC17, Boe23, CK22b, KFA^{+}24]. SustainMeter [RBQ^{+}24]. SV [HYR^{+}23]. SW [KP19]. SWANe [GAM^{+}24]. swarm [HFS^{+}21, LXK23, ZY23]. Sweep [O’D21]. sweeps [Bac21]. Swift [KNW^{+}23]. SWIGLAL [Wet20]. Swiss [FVA^{+}20, SC17]. Switch [JHAMF19]. swSim [GSF21]. Symbolic [FLR22, BCD22, CHP23, CCO21, HOB23, NK20, NK20]. Sympy [AK24]. Sympy-based [AK24]. synchronizer [HESH19]. synchrony [MK16]. synchrophasor [AVB17, BFV18, VJA^{+}18]. synchrophasor-based [VJA^{+}18]. SYNTAX [MCK^{+}23]. synthesis [dSBSvE23]. synthetic [Cin24, SMR22, SHK^{+}23]. System [AB24, BCR^{+}18, LD24, PZ22, AKM20, AHtH20, AGL^{+}23, AHR^{+}22, AGdSC20, Bed24b, BHB^{+}21, CS21, CCI^{+}22, CR22, DTD17, DRFB^{+}23, FLR22, GACS^{+}23, Gök21, GCP22, HA23, KPM^{+}22, LBT^{+}21, LHL^{+}24, MMS20, MMT24, MKR^{+}21, ODE^{+}22, PMFÁM22, PMS23, PZKK21, RBQ^{+}24, SS19, SO21, SP22, Sur20, SMD^{+}24, TJS18, TJS19, TCCK23, VJA^{+}18, WS24a, YKC^{+}19, CCI^{+}22, FRNW^{+}24, GKK^{+}23, GPW^{+}20]. systematic [CUSRCP^{+}22, NMLM18].

Systems [BCR^{+}18, LFT21, VRBM16, dCWL^{+}23, AM23, CCH^{+}19, DFC18, DPF^{+}21, FC17, GV20, GTOM23, Jan24, JHAMF19, KHG21, KMIG20, KM21, MB21, MVRM19, SAC^{+}21, SL18, SUM21, SP19, SAA18, SL20a, SG24, WWB^{+}23, YBVC^{+}22, YD20, YDN21, ZTC^{+}21, PTS22].

T [IVSV23, SGB^{+}22]. T-Ridership [IVSV23]. TabbyXL [SKM19]. tabular [SEdCCCRG23]. TAC [ASSS24, CK22b]. TAFFO [CCAC22]. Tailored [YB23]. Takin [Web17, Web21, WGB16]. takos [LLM^{+}20]. talk [vDPI^{+}18]. Taming [Els23]. tangent [PG18]. TARDIS [LCGA22]. target [BSG20, NOÖÇ19]. Targeting [BGCS19]. TAS [YBVC^{+}22]. task [GGI23, HBS^{+}20]. Tasks [RRKA22, NGK^{+}20, PCPC24, PF22]. Tau [KF22]. TauFactor [CBS^{+}16]. Taylor [LFR^{+}20]. TBTK [Bjö19]. Tchakaloff [BV19]. Tchebichef [AKC23]. TDF [DJCL23]. TDOA [BL16]. teacher [VMS^{+}23]. teaching [AG21, CBLI22, HBS16, Par22]. team [KFA^{+}24]. TeamWellX [KFA^{+}24]. TechMiner [Vel23]. technique [AAN17]. techniques [DFS24, SAA^{+}23b]. technologies [JMT22]. TeDDy [MKZ24]. tegdet [BJM23a]. TeleLoad [JCFM24]. teleoperation [JCFM24]. temper [HDA21]. temperature [CCFSB20, HM23, RS22]. temperature-dependent [RS22]. Template [MSK24, Bac21, GM24]. template-based [Bac21]. Templated [MKZ24]. temporal [Azi23, BMK23, GB24, LRL23, OPPZ22, PS19c, Pos22, PZ22, ZPM23]. tenancy [AA23]. tendon [KH19a]. TendonMech [KH19a]. Tennessee [DK19, RE22]. term [BCFJ24, MVBF19]. terminal [MFJ23]. terminal-based [MFJ23]. test [ACNF22, BGGL20, Ben15, CPB^{+}23, GLR22, Gór24, SVM^{+}23, TTHZ23]. testbed [LOM^{+}23]. testbeds [AA23]. TESTed [SVM^{+}23]. testing [CSZM22, DM20, KCP^{+}24, MBA^{+}22, MAAOZMAM22, NNDS23, PRC^{+}23, SVM^{+}23]. tests [CRV22, MGMAG22, RIHVP23, CRP21]. text [KALS24, ZS24]. TextCL [PF22]. TFInterpy [CZ22]. TFMLAB [BFSJP^{+}21]. TFSAP [BO19]. tgcd [PDH16]. their [BL16, CPF23, CAW^{+}20, MIHS21]. theory [LSOM18, LEK22, SLWS^{+}17, WFD23]. therapy [RLK18]. Thermal [TSCH20, BDQ^{+}22, DSI^{+}20, FS19, LLM^{+}20, OHO16, PNC23, TS20]. thermally [FRPC23, FRE^{+}24, NR16]. thermoacoustic [YBVC^{+}22]. thermodynamic [SLBC^{+}20, SP19]. Thermofeel [BDQ^{+}22]. thermography [SSC23]. thermoluminescence [PDH16]. thermostat [RS22]. thickness [MMCKK21]. thin [Dan20, ZWM21]. Things [KJKK23]. Think [LHB24]. thread [Sho24, dRB^{+}20]. thread-level [dRB^{+}20]. three [AI21, BC22, LKSS20, MKS^{+}23, MP18, NR23, PFC^{+}18, RRS18]. three-dimensional [AI21, LKSS20, MP18, NR23, PFC^{+}18, RRS18]. thresholding [LAA23]. throughput [DPD22, EWF^{+}23, GQCP^{+}18, JSG22, OLRLB21]. Tidy [CCTCGCOB23, MKMK23]. tier [PCdJSdFR^{+}23]. tight [LEK22]. Time [BO19, BAEBAS19, ÇA23, CWM^{+}21a, ASAA20, AVB17, Amo21, AM23, BCR^{+}18, BDPZ19, BJM23a, BHB^{+}21, CLM22, VVDV22, EB24, EAMSR^{+}23, FMFCEG^{+}23, FHB^{+}21, GACS^{+}23, GZW^{+}22, GTG21, HHH22, Hel22, JSB20, KBB19, LS16, LHZT23, LDAL20, MWM20, MRMD19, MB21, MVRM19, NKHZ21, iNKN^{+}20, PR19, PO19, PMLM23, RGdlTQ^{+}23, RMMG21, RKA^{+}23, SHvW20, SOS19, SM19, TJS18, TJS19, VRBM16, WSRMB23, YKC^{+}19, YDMC15, ZRBCI20, dCWL^{+}23, vKMH^{+}20, BFF^{+}20, FBA^{+}22, MT19, OPPZ22].

Time-Aware [OPPZ22]. time-delayed [RMMG21]. time-domain [BCR^{+}18, VRBM16, dCWL^{+}23]. Time-Frequency [BO19, BAEBAS19, MRMD19, SOS19]. time-interval [FHB^{+}21]. time-resolved [GTG21]. time-saving [HHH22]. Time-Scale [BO19]. time-series [WSRMB23]. time-to-event [EB24]. time-to-solution [iNKN^{+}20]. time-varying [Hel22]. TimeAwareBPMN [OPPZ22]. TimeAwareBPMN-js [OPPZ22]. timeline [SAS^{+}22]. times [HWT^{+}23, MFCV20]. TinkercadNetConnector [ED22]. TINTO [CCTCGCOB23]. Tiny [ASSS24]. TIQA [SO23b]. TIQA-PSI [SO23b]. TMS [Kul20]. TMS-induced [Kul20]. TOMATO [WDZ^{+}20]. tomographic [CBS^{+}16, KPTW19, KPC^{+}20, MMG19]. tomography [KGK^{+}20, PFC^{+}18, LYX^{+}18]. TomoPhantom [KPN^{+}18]. TomoWarp2 [TACH17]. Tool [AS18, AMWHD23, EFTS23, GMKRS21, NSS^{+}19, TSNRM24, WM21a, AA23, AJK^{+}23, AT22, ATT21, AGVM20, ACNF22, BFNP24, BBHB22, BGM^{+}24, BS19, BAEBAS19, BB20a, BBF^{+}19, BBN^{+}22, BDSC22, BCFJ24, BD22, CUSRCP^{+}22, Can22a, Can22b, Can24, CSK19, CJA23, CdSLCC20, Cer19, CMACMÁT23, CK22a, CK23, CLM22, CFA22, CAW^{+}20, CGKD23, CRP21, DW20, DRFB^{+}23, DSI^{+}20, Duc23, Ek16, FJ22, FWJ^{+}23, Fru21, GZA^{+}22, GLCBHTHO23, GABH22, GWBF23, GPK^{+}21, GP22, GBW^{+}24, GPEM23, GYWG23, HHH22, HM23, HA19, HMC^{+}24, HMK23, HWT^{+}23, ICS^{+}24, IVSV23, JAMP23, JM20, KZ18, KK24, KEwMO24, KSFG18, KPS23, KCP^{+}24, Kul20, LHB24, LSB^{+}20, LLP^{+}24, LPL^{+}24, LDM20, LBH^{+}20, LYZ^{+}22, LFT21, MTS^{+}24, MW23, MG22, MSLH23, MFCSÁM20, MLTF^{+}18, MWM20, MBDS20, MGK^{+}20, MIHS21, MBI20, MBY22, NS22, Nai17]. tool [NK18, NR23, OPPZ22, OHO16, OJ22, OBWM23, PAK22, PRC^{+}23, PCPC24, PBS^{+}24, PCdJSdFR^{+}23, PMM16, PS19c, RNK21, RNK22, RXZ^{+}23, RAL^{+}20, RRS18, RMMG21, Rub16, RKA^{+}23, STH^{+}21, SAC^{+}21, SvZvdA23, SLL23, SSP21, SMZSW24, SCG^{+}20, SHTE23, SABEh20, SMKL23, Sur20, SBT22, TBCG21, TSCH20, Tem23, TCQOGM23, TSMT19, TTT19, UdL20, VDCL20, VCA^{+}24, Ver22, VRVAC23, dMVBTN24, WSRMB23, WC23, WSAC^{+}23, YKKD19, YHC^{+}22, YB23, dlSVBLdA^{+}17, BGCS19, PCL22, Pos22, TSCH20, STC23, CGS19]. Toolbox [EBGCT21, KP20, RNG^{+}21, WDZ^{+}20, AGC^{+}24, AL16, AGM^{+}23, BFSJP^{+}21, BBHP^{+}23, BSC^{+}24, BlD^{+}20, DZZ^{+}22, EDZ^{+}23, FS23, GGI23, GCdJAURO21, GEH19, tHLMN19, HL22, ITL^{+}23, JAC20, JGGF23, KG23, KMLP^{+}23, KAO^{+}21, LMN18, MMD24, MCC20, MMG19, NMT23, Nas20, PMNWR20, PNL^{+}21, RCT20, RJN^{+}23, RL19, SF16, Sil20, SRML17, TŚG^{+}24, VBA^{+}16, WPM^{+}20, XGW23, ZLLT23, BB20b, CHL^{+}19, KPM^{+}22, ŁMF^{+}23, NK18, SY20, SO23b, UVPB^{+}22]. Toolkit [CVAP22, HMCA15, APO^{+}20, BBP^{+}18, CCH^{+}19, FRPC23, FRE^{+}24, GSS^{+}23, HHB^{+}22b, KDRT23, KHH^{+}22, KPTW19, Kri22, LVK21, LRL23, MZL^{+}23, MD21, MRGAGP23, OAF20, RLK18, RMM23, SGBH18, SGA^{+}16, SV23, Zaj20, ZWCQ22, ASAA20, BMR17, BHCT16, FMFCEG^{+}23, Fra22b, KSFG18, PF23]. toolkits [RSLL^{+}20]. Tools [AVB17, HKC^{+}18, ACRSGJ23, CAW^{+}20, DVSD22, EB24, LKK23, MB21, RIHVP23, SAA^{+}23b, SDCA19, DVSD22, LDAL20, SPG^{+}23, Zie19]. toolset [MBC22]. topic [KS23]. TopicTracker [KS23]. topographical [Zek17]. TopoHub [Jur23b]. Topological [BSB^{+}23]. topologies [Jur23b]. topology [BTMB21, IZBT21, TC20]. TopoZeko [Zek17]. TOPSIS [YKKD19]. torcpy [HBS^{+}20]. tortuosity [CBS^{+}16, TKLG19]. Total [WDZ^{+}20]. TOUGH [BB20a]. TOUGH2Viewer [BB20a]. Towers [AI21]. TPMS [Kar21]. trace [ATCA20, LHCK18]. tracing [FRPC23, FRE^{+}24, IIH^{+}23, PNC23]. Track [Can22a, SDCA19, LKSS20]. TRACK P [NIY16]. Tracker [FK23, KH19b, Car21]. Tracking [JSB20, AALW23, CLM^{+}20, FSAD23a, FK23, HS20, LKSS20, PLC^{+}24, QTL23, TK23, TGS^{+}19, Ver22, YYL^{+}22, Zaj20]. TrackPad [CLM^{+}20]. tracks [NIY16, WTZ^{+}21]. TraClets [KMT23]. traction [BFSJP^{+}21]. tractography [Kul20]. TraCurate [WTZ^{+}21]. trade [Kur23]. traffic [LSSK16, MW23]. train [MK16, SEZLMS^{+}23, Can22a]. training [NNDS23, SWCP20]. trajectories [LRL23]. trajectory [BHHÖ22, HHB^{+}22a, KS23, KMT23]. trans [NL23]. trans-dimensional [NL23]. transcranial [XGW23]. transcriptome [TT17]. Transfer [ZWM21, BDPZ19, EAFJ^{+}22, MPMC21, MFJ23, SAA18, ZLLT23]. Transform [MT19, KOC21, MTS^{+}18, SM19, AKC23, YC19]. transformation [GVAO19, KPJ24, SKM19]. transforms [WSK22]. transient [Duc23, KG24, MPMC21, RAL^{+}20, YMCB23]. transients [SHvW20]. transistor [OKJ23]. transistors [O’D21]. transition [BSdG^{+}22, FRPC23, FRE^{+}24, JMT22, LEK22, WRV23]. Transmission [RAL22, BBHP^{+}23, IVSV23]. transonic [LH22]. transparencies [MVBF19]. transparency [AGL^{+}23]. transparent [MFJ23]. transport [FHA17, KRB^{+}20, Ohn21, SLBC^{+}20, FAPANG22]. transportation [Boe23, IVSV23, MW23]. TransRNM [MW23]. TraSER [LSSK16]. travel [PFJM19]. traveltime [Gir21]. treatment [LCGA22]. trec [SO23a]. tree [AFA22, ÇA20, KEwMO24, LPS^{+}24, Liu23, Mos20, NS22, PLK^{+}23, SRML17]. tree-based [PLK^{+}23]. tree-ring [Liu23]. trees [GZ21a, GZ21b, MIHS21, RKDP21, SIC24]. TreeSim [NS22]. TreeSwift [Mos20]. TreeTool [MIHS21]. TreeVision [KEwMO24]. tremor [DER^{+}23]. TremorSoft [DER^{+}23]. trend [SO23a]. trial [MKMK23, RLN21]. trials [KHJ^{+}23, MKMK23, RLN21]. triggered [DPD22]. triple [Web23]. triple-axis [Web23]. triply [Kar21]. TROVA [FAPANG22]. trusses [Ozb17]. TrustML [MAG24]. trustworthiness [MAG24]. TSFEL [BFF^{+}20]. tsflex [VVDV22]. tsrobprep [NKHZ21]. TSSEARCH [FBA^{+}22]. TTB [Can22a]. TTB-2D [Can22a]. ttcrpy [Gir21]. TumorDecon [AAL^{+}22]. tumour [BSV^{+}23]. tuner [CCAC22, FLR22]. Tuning [PF23, BE20]. turbulence [BDF^{+}20, SL21]. turbulent [LFR^{+}20, ML18, SL21]. tutorial [ND22]. twin [LS23, ZRBCI20]. TWINKLE [ZRBCI20]. twins [PMZ21]. two [NR16, PMLM23, Taq16, ZNS17, AT22]. two-dimensional [ZNS17]. two-hole [Taq16]. two-particle [Taq16]. two-phase [NR16, AT22]. tx2 fcnn node [MBY22]. typing [LVCF24].

U.S. [LPR21]. UAV [REK^{+}23]. UAVradio [ACEAV24]. UAVs [ACEAV24]. UBUMonitor [MSLNSMOG24]. UF [MCC20]. uFTIR [CBHLG21]. UGUCA [KAK21]. Ultra [SHK^{+}23]. Ultra-high [SHK^{+}23]. Ultrasonic [KOZ^{+}24, CCBC^{+}21]. Ultrasound [AJR^{+}23, FM24b, GZW^{+}22, FM24b]. Unbiased [Nis20]. unbounded [GEG22]. uncertain [FCMTV24]. Uncertainty [SY20, TSG^{+}23, FPN^{+}22, KXK^{+}23, KMK^{+}21, SCPC18, WPM^{+}20, WKS22, WS24b, tHLMN19, LMN18]. unconfined [SHH23]. undergraduate [MCGK19]. Undersampling [RLF^{+}21]. Undulate [AMH23]. uniform [MFQ^{+}21, BBF^{+}19]. uniformly [GEH19]. unify [APO^{+}20]. unit [BDSC22, Hac23]. United [PS19c]. units [OKJ23, Zhu15]. Unity3D [Che21]. univariate [LF15]. Universal [ANOU21, KSS23, SKS23, SMD23]. unknown [LIZ^{+}20]. unmodeled [DKL^{+}21]. unmodelled [XTD24]. unreinforced [SK22]. Unresolved [MVBF19]. unsteady [Azi23]. unstructured [AG21]. unsynchronized [BCA19]. Update [AGH20, BCR^{+}18, Dan22, DFC18, DPF^{+}21, FST^{+}21, tHLMN19, HWST22, NBS^{+}21, PS18b, PS18c, PS19b, RtHLMN20, TJS19, Web17, Web21, YDN21, Ars21, BFRK19, EFTS23]. updated [KMLP^{+}23, LCGA22]. Upen2DTool [BBF^{+}19]. upon [ÖA23]. upper [HVCS^{+}23]. UPStream [EL17]. UQpy [TSG^{+}23]. urban [PCdJSdFR^{+}23, RMV^{+}24, SR20]. urbanism [DV23]. usability [ACRSGJ23]. USE [DVSD22, CK22a, CK23, CA18, CAW^{+}20, DSJ^{+}22, PTS22]. User [MCC20, ANA16, APLI24, AGEABD^{+}24, BHKK23, FM24b, GGI23, GCBDDBF23, HMK23, JCFM24, LYZ^{+}22, MTS^{+}18, Ozk23, PBM22, RtHLMN20, SMZSW24, SCSB24, SPMGA^{+}23, SMR22, SSC23, WHG^{+}23, XTD24, MM16]. user-defined [AGEABD^{+}24]. User-Friendly [MCC20, HMK23, LYZ^{+}22, XTD24]. Usiigaci [TGS^{+}19]. Using [KL22, AALW23, AMVB19, BCA19, BDPZ19, BHC^{+}24, BS21, BJM23a, BOP^{+}23, BS19, BC20, BWMS22, BDSC22, CH23, DL16, DK19, DRFB^{+}23, DSI^{+}20, EAFJ^{+}22, GZ21a, GY23, GNPP23, GL15, GCP22, HFS^{+}21, HHN20, HLP^{+}19, HWST22, Jun21, KG24, KK24, KGF^{+}24, KCP^{+}24, MSDP23, MLTF^{+}18, MKS^{+}23, MIHS21, MAM^{+}23, NK20, NGK^{+}20, NGOT24, Nis20, OKJ23, OLRLB21, PFC^{+}18, PBM22, PS19a, PDM22, RLM^{+}24, RK19, STH^{+}21, SNLFGVRB24, SWCP20, SCSB24, SDCA19, SSC23, TK23, TBO24, TTHZ23, Vel23, XYC22, YC19, dlSVBLdA^{+}17]. UT [HLW^{+}16]. utility [Bac21]. utilized [LHB24]. utilizing [CCI^{+}22, FdCBV23, ÖA23, SF16]. UTypes [FCMTV24]. UV [DPD22]. UV-VIS [DPD22].

V [MG22, OKJ23]. v.1.00 [dlSVBLdA^{+}17]. v1.0 [Ozb17, PMM16, MD19]. v1.0.0 [KJKK23]. v1.1 [PS18b]. v1.2 [PS18c]. v1.3 [PS19b]. v10.4 [OCK^{+}23]. v2.0 [EFTS23, MM16, TŚG^{+}24]. v4 [RLJS23]. v4.1 [TSG^{+}23]. valence [BBP^{+}18]. validated [PKA^{+}22]. validating [CNST20, MGGHG^{+}24]. validation [HKZ24]. validity [JGGF23]. value [LAA23]. valued [DGK^{+}24, SABEh20]. values [Kar21, MTS^{+}18]. VApor [FAPANG22]. vapourisation [FPI^{+}22]. vapourisation-inducing [FPI^{+}22]. variable [NG23, RGP^{+}22]. variables [GKK^{+}23]. variant [HHN20]. variation [AJK^{+}23, PS19c, YC19]. VAriational [AGC^{+}24]. various [Jan24, LSPG22]. varying [Hel22]. VBI [Can24]. VBI-2D [Can24]. vector [BTMB21, CLM22, NR23, RJH^{+}20]. vectorial [Hac23]. vectorized [MZL^{+}23]. vectors [GMF18, KM21, TDG19]. vegetation [CVV23]. vehicle [AGG23, Can24, CT24, PMS23]. vehicle-bridge [Can24]. vehicles [Can22b]. VELCRO [LPL^{+}24]. Velocimetry [LD24, BGGL20, PDM22, ACPdP22, JSB20]. VEqMon2D [Can22b]. verification [Can24, LVK21, OPPZ22]. VERONA [GFRMVL24]. versatile [FHA17, KH19b, SMD^{+}24]. Version [BWS23, Bla23, BTDT23, DBPS22, FRE^{+}24, SKS23, TBD23, WBS22b, WBNS23, WKCS24, WS24b, ZJKS23, ZRK22, dCWL^{+}23, KMLP^{+}23, LNS15, RS22, ZNS17]. versions [CA18]. vessel [FH23, HKF^{+}20, LRL23]. vFirelib [WSAC^{+}23]. vGameReviews [FVVL23]. via [CCG^{+}23, DA24, CDAOB^{+}20, KHG21, Kle24, LD24, PR19, SKD22]. Viability [Sha23]. ViBA [LKK23]. vibrational [WW17]. Video [BCFJ24, MRGAGP23, AMVB19, FVVL23, KW23, MAGM^{+}23, PLC^{+}24, RMM23, SMZSW24, SD23, SP22, SMD^{+}24, TDG19, PDM22, TDG19]. video-tracking [PLC^{+}24]. Videos [AJR^{+}23]. view [SS19]. Viewer [GV20, Ohn21]. viewing [LMS^{+}16]. Viewscape [YFVL24]. viewsheds [YFVL24]. vine [Cob21]. vineyards [ASVV23]. Virgo [AAA^{+}21]. ViroCon [HLP^{+}19, HWST22]. Virtual [AVCP^{+}21, HKM^{+}19, SDCA19, BSVP20, HVCS^{+}23, LOM^{+}23, VMS^{+}23, BE20]. virtually [IA17]. VIS [DPD22]. ViSA [Sha23]. viscosity} [JAC21]. VisExpA [TC20]. Visibility [TC20]. vision [RSLL^{+}20, SSC23, PDM22]. viskillz [TTHZ23]. viskillz-blender [TTHZ23]. visual [AI21, HCD24, HHN20, LPL^{+}24, MSLNSMOG24, OTdMP22, Taq16, Zaj20]. visual-based [LPL^{+}24]. visualisation [GZA^{+}22, KS23, WGB16, Web17, Web21]. visualizable [IKYY22]. Visualization [LKK23, AAF^{+}24b, BSV^{+}23, DAB^{+}19, GPK^{+}21, GYWG23, GV20, HOB23, LSF^{+}23, LPS^{+}24, MTPHH18, MKT^{+}23, MLGBR23, Mül23, NS22, RNK21, RNK22, RL19, RKA^{+}23, TŚG^{+}24, UVPB^{+}22, VEM^{+}18, WSAC^{+}23, Zek17, HMCA15]. visualizations [TSCH20]. visualize [LMB^{+}19, SDCA19]. Visualizing [SMD23, LVT^{+}22, MMT24, MKMK23]. vkpolybench [CC21]. VLAB [KSA23]. vocal [SA17]. VOF [SNHS20]. VOF-simulations [SNHS20]. voice [TJS18, TJS19]. voltages [JM20]. Volume [HS20, AG21, JFJM22, TACH17]. Volume-of-Fluid [HS20]. volumetrical [SGP23]. Voronoi [BB20a]. vortex [LDM20]. VortexFitting [LDM20]. voxel [IIH^{+}23, LDAL20, MBDS20]. voxel-based [IIH^{+}23, LDAL20, MBDS20]. voxTrace [IIH^{+}23]. VSR [MBDS20]. VT [ZRK19]. VTK [HMCA15]. vulcanization [MM16]. Vulkan [CC21]. vulnerabilities [dHGMCGL^{+}24]. Vyazovkin [DSI^{+}20].

WACline [MDG22]. WAF [VDCL20]. WAF-A-MoLE [VDCL20]. WAFs [VDCL20]. walk [RJH^{+}20, TBO24, TKLG19]. walker [Hel22]. wallpaper [LSF^{+}23]. warping [EAMSR^{+}23, RJN^{+}23]. WaspSim [CH23]. waste [VCA^{+}24]. WatchOS [AGADAG^{+}23]. Water [SEL^{+}16, BHC^{+}24, BFDD^{+}22, DVSD22, EL17, FVA^{+}20, FAPANG22, KLL23, SH23, DVSD22]. water-wave [BFDD^{+}22]. watex [KLL23]. Wave [CKM21, BFDD^{+}22, CCC^{+}21, CCBC^{+}21, DHK^{+}20, DKL^{+}21, FHB^{+}21, GSF21, MAC^{+}21, MVBF19, PDTG17, RCT20, RAL^{+}20, RRS18, SWK19, SACT23, Wet20, XTD24, SACT23]. WaveBurst [DKL^{+}21]. waveform [MGFO23, SF16]. wavelet [SM19]. WaveMaker [RRS18]. Wax [SP19]. Way [PI17]. wearables [EDZ^{+}23]. weather [ABM^{+}21, BS19]. web [Agu23, AMA19, APLI24, BÇ23, BBS^{+}22, BSV^{+}23, CS21, CJA23, Cin24, DZH^{+}23, GZR^{+}19, GB24, HMC^{+}24, IVSV23, JFJM22, KK22, KF17, KXK^{+}23, LMM22, MSDP23, MKT^{+}23, MCK^{+}23, MAM^{+}23, NMLM18, PNC23, SMR22, ACO21, AHR^{+}22, BÇ23, DLH18, MDG22, RLN21, SMR22, VCA^{+}24, VEM^{+}18]. web-application [Cin24]. web-applications [LMM22]. web-based [AMA19, BÇ23, BSV^{+}23, CS21, CJA23, DZH^{+}23, HMC^{+}24, KXK^{+}23, MSDP23, MCK^{+}23, NMLM18, SMR22, AHR^{+}22, VCA^{+}24, VEM^{+}18]. Web-NDefA [BÇ23].

WebCollectives [Agu23]. WebExtension [JC23]. webpage [JC23]. WebRISC [MG22]. WebRISC-V [MG22]. weeks [SA17]. weight [MMD24]. Weighted [DMFG22, LHY^{+}23, RLM^{+}24]. weighting [BW22, BWS23]. Weka [CMACMÁT23]. well [AS17, EB24, KFA^{+}24]. well-being [KFA^{+}24]. WENO [GKM20]. WEUSEDTO [DVSD22]. wheel [HLW^{+}16]. wheelchair [GCBDDBF23]. WheelerLab [AS17]. while [OTdMP22]. Whisper [DG21]. Whitney [CRV22]. whole [CdSLCC20]. whole-exome [CdSLCC20]. wide [MTPHH18, VJA^{+}18]. wide-area [VJA^{+}18]. Wilcoxon [CRV22]. Window [MT19]. wireless [HTV22]. WITec [HL22]. within [ADSG^{+}20, Nis20]. without [PG18]. WITio [HL22]. WLDT [PMZ21]. Workflow [SDCA19, ASVV23, BLM^{+}22, FM24b, GAM^{+}24, KPM^{+}22, KP19, LEK22, SSW^{+}24]. workflows [AGHK21, FWB^{+}21, KPSM17, NGK^{+}20]. working [FVA^{+}20]. workspaces [GG21]. world [ED22, Jur23b, Kur23]. worm [Jun21]. wrappers [AS23]. writing [Rei23]. Written [CKM21, CWLG^{+}21, IZBT21].

X [BBD^{+}22, DWO^{+}22, GTG21, IIH^{+}23, KLY21, KPC^{+}20, LBG20, PFC^{+}18, PKKQ20, YLM^{+}24, AKC23]. x-DTT [AKC23]. X-Ray [PKKQ20, BBD^{+}22, GTG21, IIH^{+}23, KLY21, KPC^{+}20, LBG20, PFC^{+}18, PKKQ20, YLM^{+}24]. XAI [DSB^{+}23]. XAMG [KM21]. XaNSoNS [Nev17]. XBRL [She19]. Xcompact3D [BDF^{+}20]. XDrift [BSG20]. XISMuS [LBG20]. xkl [DDCSH23]. XMHFL [KLY21]. XML [Nai17]. XML/HTML [Nai17]. XP [YLM^{+}24]. XR [MLGBR23]. XRAE [TMGW23]. XReport [HKS^{+}22]. xSPDE [KPOD16]. XTBDFT [LEK22]. xtDCR [Azi23].

YADE [HGS17]. YADPF [MKU22].

YAWL [AHtH20]. Year [BS19]. yield

[CPB^{+}23, MGMAG22, MAGM^{+}23]. YOLO [AOGC^{+}20]. Young [Dav21, HWT^{+}23]. young-forest [HWT^{+}23]. YouTube [KW23].

Z [AJK^{+}23, MFJ23]. Z-lambda [MFJ23]. Z-scan [MFJ23]. Zeolite [AGHK21]. zero [PR19]. zero-delay [PR19]. zfit [ENCS20]. ZigBee [PMS23]. zonal [WM21a]. zones [PCdJSdFR^{+}23, ZRK19]. ZWT [CPV^{+}20]. ZZ [MP17].

References

Acar:2023:KTM

[AA23] Benjamin Acar and Marc Guerreiro Augusto. Kufast: a tool for multi-tenancy in autonomous driving testbeds. SoftwareX, 24 (??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

                                                        www.sciencedirect.com/
                                                        science/article/pii/
                                                        S2352711023002649.

Abbott:2021:ODF

[AAA^{+}21] Rich Abbott, Thomas D. Abbott, Sheelu Abraham, Fausto Acernese, and 1243 others. Open data from the first and second observing runs of Advanced LIGO and Advanced Virgo. SoftwareX, 13(??): Article 100658, January 2021. CODEN ???? ISSN 2352-7110. URL http://

                                                        www.sciencedirect.com/
                                                        science/article/pii/
                                                        S2352711021000030.

Akritidis:2023:FFH

[AAB23] Leonidas Akritidis, Miltiadis Alamaniotis, and Panayiotis Bozanis. FLAGR: a flexible high-performance library for rank aggregation. SoftwareX, 21 (??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023000158.

Almeida:2024:AAC

[AAF^{+}24a] Yonatha Almeida, Danyllo Albuquerque, Emanuel Dantas Filho, Felipe Muniz, Katyusco de Farias Santos, Mirko Perkusich, Hyggo Almeida, and Angelo

Perkusich. AICodeReview:

Advancing code quality with AI-enhanced reviews. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024000487.

Ates:2024:NFR

[AAF^{+}24b] Onur Ates, Shailesh Appukuttan, Hélissande Fragnaud, Corentin Fragnaud, and Andrew P. Davison. NeoViewer: Facilitating reuse of electrophysiology data through browserbased interactive visualization. SoftwareX, 26 (??):??, May 2024. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711024000815.

Aronow:2022:TDC

[AAL^{+}22] Rachel A. Aronow, Shaya Akbarinejad, Trang Le, Sumeyye Su, and Leili Shahriyari. TumorDecon: a digital cytometry software. SoftwareX, 18(??):??, June 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022000528.

Abbad-Andaloussi:2023:CET

[AALW23] Amine Abbad-Andaloussi, Daniel Lübke, and Barbara Weber. Conducting eye-tracking studies on large and interactive process models using Eye-Mind. SoftwareX, 24 (??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023002601.

Aldrich:2017:OCO

[AAN17] Daniel R. Aldrich, Cagri Ayranci, and David S.

                                        Nobes. OSM-Classic: an

optical imaging technique for accurately determining strain. SoftwareX, 6 (??):217–224, ???? 2017. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/

science/article/pii/
S2352711017300407.

AlHamaydeh:2019:IPP

[AANA19] Mohammad AlHamaydeh, Nader Aly, Mohamad Najib, and Sameer Alawnah. INSPECT-PBEE: a performance-based earthquake engineering GUI for IDARC-2D. SoftwareX, 9(??):132–144, January/ June 2019. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018300323.

Alexandrescu:2024:DDA

[AB24] Adrian Alexandrescu and Cristian Nicolae Butincu. DARS: Decentralized article retrieval system. SoftwareX, 25(??):??, February 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023003205.

Argaez:2019:LPC

[ABB^{+}19] Carlos Argáez, Jean-Claude Berthet, Hjörtur Björnsson, Peter Giesl, and Sigurdur Freyr Hafstein. LyapXool — a program to compute complete Lyapunov functions. SoftwareX, 10(??):Article 100325, July/December 2019. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711019302109.

Ardito:2020:RCA

[ABC^{+}20] Luca Ardito, Luca Barbato, Marco Castelluccio, Riccardo Coppola, Calixte Denizet, Sylvestre Ledru, and Michele Valsesia.

                                        rust-code-analysis: a

Rust library to analyze and extract maintainability information from source codes. SoftwareX, 12(??):Article 100635, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711020303484.

Afanasyev:2021:GFP

[ABM^{+}21] Anton Afanasyev, Mauro Bianco, Lukas Mosimann, Carlos Osuna, Felix Thaler, Hannes Vogt, Oliver Fuhrer, Joost VandeVondele, and Thomas C. Schulthess. GridTools: a framework for portable weather and climate applications. SoftwareX, 15(??):??, July 2021. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021000522.

Amato:2023:QLC

[ACC^{+}23] Francesco Amato, Matteo Cicalese, Luca Contrasto, Giacomo Cubicciotti, Gerardo D’Ambola, Antonio La Marca, Giuseppe Pagano, Fiorentino Tomeo, Gennaro Alessio Robertazzi, Gabriele Vassallo, Giovanni Acampora, Autilia Vitiello, Gemma Catolino, Giammaria Giordano, Stefano Lambiase, Valeria Pontillo, Giulia Sellitto, Filomena Ferrucci, and Fabio Palomba.

QuantuMoonLight: a low-

code platform to experiment with quantum machine learning. SoftwareX, 22(??):??, May 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271102300095X.

Alaez:2024:URL

[ACEAV24] Daniel Aláez, Mikel Celaya-Echarri, Leyre Azpilicueta, and Jesús Villadangos. UAVradio: Radio link path loss estimation for UAVs. SoftwareX, 25 (??):??, February 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023003242.

Aubertine:2022:CST

[ACNF22] Jacob Aubertine, Kenan Chen, Vidhyashree Nagaraju, and Lance Fiondella. A covariate software tool to guide test activity allocation. SoftwareX, 17(??):??, January 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021001588.

Almeida:2021:BCW

[ACO21] João R. Almeida, Leonardo Coelho, and José L. Oliveira. BIcenter: a collaborative Web ETL solution based on a reflective software approach. SoftwareX, 16(??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021001497.

Aguilar-Cabello:2022:DOM

[ACPdP22]  Jorge  Aguilar-Cabello,

Luis Parras, and Carlos del

                                        Pino. DPIVSoft-OpenCL:

a multicore CPU-GPU accelerated open-source code for 2D Particle Image Velocimetry. SoftwareX, 20 (??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022001741.

Abell:2022:SFS

[ACR22] José A. Abell, Jorge G. F. Crempien, and Matías Re-

                                        cabarren. ShakerMaker:

a framework that simplifies the simulation of seismic ground-motions. SoftwareX, 17(??):??, January 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271102100159X.

Ambrosio-Cestero:2023:RDN

[ACRSGJ23] Gregorio Ambrosio-Cestero, Jose-Raul Ruiz-Sarmiento, and Javier Gonzalez-Jimenez. The Robot@Home2 dataset: a new release with improved usability tools. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023001863.

Anand:2020:GIA

[ADSG^{+}20] Manish Anand, Jed A.

Diekfuss, Alexis B. Slutsky-Ganesh, Scott Bonnette, Dustin R. Grooms, and Gregory D. Myer. Graphical interface for automated management of motion artifact within fMRI acquisitions: INFOBAR. SoftwareX, 12(??):Article 100598, July/December 2020. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020303113.

Ahrari:2022:PPF

[AES^{+}22] Ali Ahrari, Saber Elsayed, Ruhul Sarker, Daryl Essam, and Carlos A. Coello Coello. PyDDRBG: a Python framework for benchmarking and evaluating static and dynamic multimodal optimization methods. SoftwareX, 17(??): ??, January 2022. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021001850.

Anton-Fernandez:2022:SFI

[AFA22] Clara Antón-Fernández and Rasmus Astrup. SiTree: a framework to implement single-tree simulators. SoftwareX, 18(??): ??, June 2022. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021001667.

Abdollahi:2022:CRB

[AFGH22] Masoud Abdollahi, Babak Farjad, Anil Gupta, and Quazi K. Hassan. CMIP6-D&A: an R-based software with GUI for processing climate data available in network common data format. SoftwareX, 18(??): ??, June 2022. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022000395.

Africa:2022:LFH

[Afr22] Pasquale Claudio Africa. life ^{x}: a flexible, high performance library for the numerical solution of complex finite element problems. SoftwareX, 20(??): ??, December 2022. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022001704.

Alvarez-Gomez:2019:FEF

[ÁG19] José A. Álvarez-Gómez.

FMC-Earthquake focal mech-

anisms data management, cluster and classification. SoftwareX, 9(??):299–307, January/June 2019. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018302590.

Alinovi:2021:FUP

[AG21] Edoardo Alinovi and Joel Guerrero. FLUBIO — an unstructured, parallel, finite-volume based Navier–Stokes and convection– diffusion like equations solver for teaching and research purposes. SoftwareX, 13(??):Article 100655, January 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271102030368X.

Adu-Gyamfi:2023:MSC

[AGADAG^{+}23] Kojo Konadu Adu-Gyamfi, Karo Ahmadi-Dehrashid, Yaw Okyere Adu-Gyamfi, Pujitha Gunaratne, and Anuj Sharma. MobiScout: a scalable cloud-based driving and activity monitoring platform featuring an IOS app and a WatchOS extension. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023002844.

Acampora:2024:EEA

[AGC^{+}24] Giovanni Acampora, Carlos Cano Gutiérrez, Angela Chiatto, José Manuel Soto Hidalgo, and Autilia Vitiello. EVOVAQ: EVOlutionary algorithms-based toolbox for VAriational Quantum circuits. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024001274.

Ayala:2020:RLN

[AGdSC20] Helon Vicente Hultmann Ayala, Marcos Cesar Gritti, and Leandro dos Santos Coelho. An R library for nonlinear black-box system identification. SoftwareX, 11 (??):Article 100495, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019302663.

Arenas-Guerrero:2024:RFM

[AGEABD^{+}24] Julián Arenas-Guerrero, Paola Espinoza-Arias, José Antonio Bernabé-Diaz, Prashant Deshmukh, José Luis Sánchez-Fernández, and Oscar Corcho. An RML-FNML module for Python user-defined functions in Morph-KGC. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024000803.

Almeida:2023:AES

[AGG23] Luana Souza Almeida, Alireza Ghasemi, and Floris Goerlandt. E-ARCP: an exact solution approach to the Collaborative Multivehicle Prize Collecting Arc Routing for Connectivity Problem. SoftwareX, 23(??):??, July 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023001619.

Argaez:2020:ULE

[AGH20] Carlos Argáez, Peter Giesl, and Sigurdur Freyr Hafstein. Update (2.0) to LyapXool: Eigenpairs and new classification methods. SoftwareX, 12(??):

Article 100616, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711020303290.

Antonio:2021:SCW

[AGHK21] Dexter D. Antonio, Jiawei Guo, Sam J. Holton, and Ambarish R. Kulkarni. Simplifying computational workflows with the Multiscale Atomic Zeolite Simulation Environment (MAZE). SoftwareX, 16(??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021001011.

Ahn:2023:OEO

[AGL^{+}23] Jaehyun Ahn, Haifa Gaza, Juhyeok Lee, Hyeongchan Kim, and Jaewook Byun. Oliot EPCIS: an opensource EPCIS 2.0 system for supply chain transparency. SoftwareX, 23 (??):??, July 2023. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023001735.

Antony:2023:MML

[AGM^{+}23] Akshay Antony, Chakradhar Guntuboina, Rishikesh Magar, Lalit Ghule, Ruchit Doshi, Aman Khalid, Sharan Seshadri, and Amir Barati Farimani.

ManufacturingNet: a ma-

chine learning toolbox for engineers. SoftwareX, 23 (??):??, July 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023001747.

Agun:2023:WLR

[Agu23] Hayri Volkan Agun. WebCollectives:

a light regular expression based web content extractor in Java. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023002650.

Aonzo:2020:OOS

[AGVM20] Simone Aonzo, Gabriel Claudiu Georgiu, Luca Verderame, and Alessio Merlo.

Obfuscapk:  an open-

source black-box obfuscation tool for Android apps. SoftwareX, 11 (??):Article 100403, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019302791.

Adler:2023:FMI

[AHP23] Jeremy Adler, Ainsley Huang, and Ingela Parm-

ryd. Find plasma membrane

and measure plasma membrane:

ImageJ macros for efficient identification of and measurements at and around the plasma membrane. SoftwareX, 24 (??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023002662.

Anshori:2022:WBS

[AHR^{+}22] Isa Anshori, Suksmandhira Harimurti, Mahendra Baleno Rama, Rey Ezra Langelo, Jessika, Lenny Putri Yulianti, Gilang Gumilar, Muhammad Yusuf, Silmina Prastriyanti, Brian Yuliarto, Husna Nugrahapraja, Wyanda Arnafia, and Irvan Faizal. Web-based surface plasmon resonance signal processing system for fast analyte analysis. SoftwareX, 18(??):??, June 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022000462.

Adams:2020:YOS

[AHtH20] Michael Adams, Andreas V. Hense, and Arthur H. M. ter Hofstede. YAWL: an open source business process management system from science for science. SoftwareX, 12 (??):Article 100576, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020302892.

Atanasiu:2021:DTM

[AI21] Vlad Atanasiu and Rolf Ingold. Document Towers: a MATLAB software implementing a threedimensional architectural paradigm for the visual exploration of digital documents and libraries. SoftwareX, 14(??):??, June 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021000297.

Afroz:2021:SSB

[AISM21] Abanti Shama Afroz, Francesco Inglese, Cesare Stefanini, and Mario Mi-

lazzo. STL Process: a

.STL-based preprocessor for robot path planning in manufacturing and quality control processes. SoftwareX, 15(??):??, July 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021000650.

Adolfsson:2023:CZN

[AJK^{+}23] Emma Adolfsson, Jon Jonasson, Aniruddh Kashyap, Anna Nordensköld, and Anna Greén. CNV-Z; a new tool for detecting copy number variation in next generation sequencing data. SoftwareX, 24 (??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023002261.

Anto:2023:SSP

[AJR^{+}23] Anito Anto, Linda Rose Jimson, Tanya Rose, Mohammed Jafrin, and Mahesh Raveendranatha Panicker. SPAALUV: Software Package for Automated Analysis of Lung Ultrasound Videos. SoftwareX, 23(??):??, July 2023. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023001565.

Ali:2024:SMS

[AK24] Amjad Ali and Shah Khusro. SA-MEAS: Sympybased automated mathematical equations analysis and solver. SoftwareX, 25(??):??, February 2024. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023002923.

Aliouat:2023:XDP

[AKC23] Ahcen Aliouat, Nasreddine Kouadria, and Doru Florin Chiper. x-DTT: a package for calculating real and Integer Discrete Tchebichef Transform kernels based on orthogonal polynomials. SoftwareX, 23(??): ??, July 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023001371.

Abbas:2020:CCP

[AKM20] Tahir Abbas, Vassilis-Javed Khan, and Panos Markopoulos. CoZ: a crowd-powered system for social robotics. SoftwareX, 11(??):Article 100421, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019302699.

Annergren:2016:PMT

[AL16] Mariette Annergren and Christian A. Larsson. MOOSE2 — a toolbox for least-costly applicationoriented input design. SoftwareX, 5(??):89–95, ???? 2016. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711016300115.

Alberti:2015:CRP

[Alb15] Gianmarco Alberti. CAinterprTools:

an R package to help interpreting Correspondence Analysis’ results. SoftwareX, 1–2(??):19–25, September 2015. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711015000060.

Alberti:2019:MRP

[Alb19] Gianmarco Alberti. movecost: an R package for calculating accumulated slopedependent anisotropic costsurfaces and least-cost paths. SoftwareX, 10 (??):Article 100331, July/ December 2019. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711019302341.

Assuncao:2021:FDF

[ALRM21] Filipe Assunção, Nuno Lourenço, Bernardete Ribeiro, and Penousal Machado.

                                        Fast-DENSER: Fast deep

evolutionary network structured representation. SoftwareX, 14(??):??, June 2021. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S235271102100039X.

Abi-Mansour:2019:POO

[AM19]  Andrew  Abi-Mansour.

PyGran: an objectoriented library for DEM simulation and analysis. SoftwareX, 9(??):168–174, January/June 2019. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018301080.

Austad:2023:NDN

[AM23] Henrik Austad and Geir Mathisen. net chan: Deterministic network channels for distributed realtime systems. SoftwareX, 23(??):??, July 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023001322.

Ait-Mlouk:2019:DMW

[AMA19] Addi Ait-Mlouk and Tarik Agouti. DM-MCDA: a webbased platform for data mining and multiple criteria decision analysis: a case study on road accident. SoftwareX, 10 (??):Article 100323, July/ December 2019. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019301438.

Allen:2024:AJI

[AMAE24] Courtney Allen, Alexandra Mazanko, Niloofar Abdehagh, and Hermann Eberl. ADM1jl: a Julia implementation of the anaerobic digestion model 1. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711024000530.

Albano:2017:MSM

[AMC17] Raffaele Albano, Salvatore Manfreda, and Giuseppe Celano. MY SIRR: Minimalist agro-hYdrological model for Sustainable IR-Rigation management — soil moisture and crop dynamics. SoftwareX, 6 (??):98–106, ???? 2017. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711017300122.

Aldana-Martin:2023:EMO

[AMDN23] José F. Aldana-Martín, Juan J. Durillo, and Antonio J. Nebro. Evolver: Meta-optimizing multiobjective metaheuristics. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023002479.

Asikainen:2023:UEC

[AMH23] Timo Asikainen, Tomi Männistö, and Eetu Huovila. Undulate: Extreme continuous experimentation of separately deployable software. SoftwareX, 22 (??):??, May 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023000961.

Amornbunchornvej:2021:MRP

[Amo21] Chainarong Amornbunchornvej. mFLICA: an R package for inferring leadership of coordination from time series. SoftwareX, 15(??):??, July 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021000935.

Abraham:2015:GHP

[AMS^{+}15] Mark James Abraham, Teemu Murtola, Roland Schulz, Szilárd Páll, Jeremy C. Smith, Berk Hess, and Erik Lindahl. GROMACS: High performance molecular simulations through multi-level parallelism from laptops to supercomputers. SoftwareX, 1–2(??): 13–18, September 2015. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711015000059.

Aldahdooh:2019:RRF

[AMVB19] Ahmed Aldahdooh, Enrico Masala, Glenn Van Wallendael, and Marcus Barkowsky. Reproducible research framework for objective video quality measures using a large-scale database approach. SoftwareX, 8(??):64–68, ???? 2019. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711017300468.

Arnold-Medabalimi:2023:PPL

[AMWHD23] Nicholas Arnold-Medabalimi, Christopher R. Wentland, Cheng Huang, and Karthik Duraisamy. PLATFORM: Parallel Linear Algebra Tool FOr Reduced Modeling. SoftwareX, 21(??): ??, February 2023. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023000092.

AlHamaydeh:2016:IGU

[ANA16] Mohammad AlHamaydeh, Mohamad Najib, and Sameer Alawnah. INSPECT: a graphical user interface software package for IDARC-2D. SoftwareX, 5(??):234–242, ???? 2016. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711016300334.

Anonymous:2015:EBa

[Ano15a]  Anonymous.  Editorial

Board. SoftwareX, 1– 2(??):1–32, September 2015. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711015000084.

Anonymous:2015:EBb

[Ano15b]  Anonymous.  Editorial

Board. SoftwareX, 3–4 (??):1–44, December 2015. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711015000187.

Anonymous:2016:EB

[Ano16]  Anonymous.  Editorial

Board. SoftwareX, 5(??): 1–260, ???? 2016. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711016300383.

Anonymous:2017:EB

[Ano17]  Anonymous.  Editorial

Board. SoftwareX, 6(??): 1–292, ???? 2017. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711017300663.

Anonymous:2018:EB

[Ano18]  Anonymous.  Editorial

Board. SoftwareX, 7(??):i, January/June 2018. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018302772.

Anonymous:2019:EBa

[Ano19a]  Anonymous.  Editorial

Board. SoftwareX, 8(??): i, ???? 2019. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711018302802.

Anonymous:2019:EBb

[Ano19b]  Anonymous.  Editorial

Board. SoftwareX, 9(??):i, January/June 2019. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711019301608.

Anonymous:2019:EB

[Ano19c]  Anonymous.  Editorial

Board. SoftwareX, 10 (??):Article 100372, July/ December 2019. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711019303735.

Anonymous:2019:JD

[Ano19d] Anonymous. July–December 2019. SoftwareX, 10(??): ??, July/December 2019. CODEN ???? ISSN 2352-7110.

Anonymous:2020:EBa

[Ano20a]  Anonymous.  Editorial

Board. SoftwareX, 11 (??):Article 100529, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020302272.

Anonymous:2020:EBb

[Ano20b]  Anonymous.  Editorial

Board. SoftwareX, 12 (??):Article 100649, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020303629.

Anonymous:2020:JJ

[Ano20c]  Anonymous.  January–

June 2020. SoftwareX, 11(??):??, January/June 2020. CODEN ???? ISSN 2352-7110.

Anonymous:2020:JD

[Ano20d] Anonymous. July–December 2020. SoftwareX, 12(??): ??, July/December 2020. CODEN ???? ISSN 2352-7110.

Anonymous:2021:D

[Ano21a] Anonymous. December 2021. SoftwareX, 16(??): ??, December 2021. CO-DEN ???? ISSN 2352-7110.

Anonymous:2021:EBa

[Ano21b]  Anonymous.  Editorial

Board. SoftwareX, 13(??): Article 100670, January 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021000157.

Anonymous:2021:EBb

[Ano21c]  Anonymous.  Editorial

Board. SoftwareX, 14 (??):??, June 2021. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021000583.

Anonymous:2021:EBc

[Ano21d]  Anonymous.  Editorial

Board. SoftwareX, 15 (??):??, July 2021. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021001047.

Anonymous:2021:EBd

[Ano21e]  Anonymous.  Editorial

Board. SoftwareX, 16 (??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021001746.

Anonymous:2021:Ja

[Ano21f]  Anonymous.  January

2021. SoftwareX, 13(??): ??, January 2021. CODEN ???? ISSN 2352-7110.

Anonymous:2021:Jc

[Ano21g] Anonymous. July 2021. SoftwareX, 15(??):??, July 2021. CODEN ???? ISSN 2352-7110.

Anonymous:2021:Jb

[Ano21h] Anonymous. June 2021.

SoftwareX, 14(??):??, June 2021. CODEN ???? ISSN 2352-7110.

Anonymous:2022:D

[Ano22a] Anonymous. December 2022. SoftwareX, 20(??): ??, December 2022. CO-DEN ???? ISSN 2352-7110. Anonymous:2022:EBa

[Ano22b]  Anonymous.  Editorial

Board. SoftwareX, 17(??): ??, January 2022. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022000309.

Anonymous:2022:EBb

[Ano22c]  Anonymous.  Editorial

Board. SoftwareX, 18 (??):??, June 2022. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022000772.

Anonymous:2022:EBc

[Ano22d]  Anonymous.  Editorial

Board. SoftwareX, 19 (??):??, July 2022. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022001418.

Anonymous:2022:EBd

[Ano22e]  Anonymous.  Editorial

Board. SoftwareX, 20 (??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022002060.

Anonymous:2022:Ja

[Ano22f]  Anonymous.  January

2022. SoftwareX, 17(??): ??, January 2022. CODEN ???? ISSN 2352-7110.

Anonymous:2022:Jc

[Ano22g] Anonymous. July 2022. SoftwareX, 19(??):??, July 2022. CODEN ???? ISSN 2352-7110.

Anonymous:2022:Jb

[Ano22h] Anonymous. June 2022. SoftwareX, 18(??):??, June 2022. CODEN ???? ISSN 2352-7110.

Anonymous:2023:D

[Ano23a] Anonymous. December 2023. SoftwareX, 24(??): ??, December 2023. CO-DEN ???? ISSN 2352-7110. Anonymous:2023:EBa

[Ano23b]  Anonymous.  Editorial

Board. SoftwareX, 21 (??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023000316.

Anonymous:2023:EBb

[Ano23c]  Anonymous.  Editorial

Board. SoftwareX, 22 (??):??, May 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023001139.

Anonymous:2023:EBc

[Ano23d]  Anonymous.  Editorial

Board. SoftwareX, 23 (??):??, July 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023002248.

Anonymous:2023:EBd

[Ano23e]  Anonymous.  Editorial

Board. SoftwareX, 24 (??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023003114.

Anonymous:2023:F

[Ano23f]  Anonymous.  February

2023. SoftwareX, 21(??): ??, February 2023. CO-DEN ???? ISSN 2352-7110.

Anonymous:2023:J

[Ano23g] Anonymous. July 2023.

SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110.

Anonymous:2023:M

[Ano23h] Anonymous. May 2023. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110.

Anonymous:2024:EBa

[Ano24a]  Anonymous.  Editorial

Board. SoftwareX, 25 (??):??, February 2024. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711024000220.

Anonymous:2024:EBb

[Ano24b]  Anonymous.  Editorial

Board. SoftwareX, 26 (??):??, May 2024. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711024001390.

Anonymous:2024:F

[Ano24c]  Anonymous.  February

2024. SoftwareX, 25(??): ??, February 2024. CO-DEN ???? ISSN 2352-7110.

Anonymous:2024:M

[Ano24d] Anonymous. May 2024. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110.

Angelidakis:2021:CCL

[ANOU21] Vasileios Angelidakis, Sadegh Nadimi, Masahide Otsubo, and Stefano Utili. CLUMP: a Code Library to generate Universal Multi-sphere Particles. SoftwareX, 15 (??):??, July 2021. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021000704.

Alves-Oliveira:2020:SAY

[AOGC^{+}20] Patrícia Alves-Oliveira, Samuel Gomes, Ankita Chandak, Patrícia Arriaga, Guy Hoffman, and Ana Paiva. Software architecture for YOLO, a creativity-stimulating robot. SoftwareX, 11 (??):Article 100461, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019302468.

Aszkowski:2023:DDN

[APK^{+}23] Przemysław Aszkowski, Bartosz Ptak, Marek Kraft, Dominik Pieczy’nski, and Paweł Drapikowski. Deepness: Deep neural remote sensing plugin for QGIS. SoftwareX, 23 (??):??, July 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023001917.

Aldalur:2024:MEM

[APLI24] Iñigo Aldalur, Alain Perez, Felix Larrinaga, and Miren Illarramendi. MAWA: Enhancing mobile web browsing through web augmentation for improved user experience. SoftwareX, 25(??):??, February 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024000049.

Almeida:2020:GTU

[APO^{+}20] João R. Almeida, Armando J. Pinho, José L. Oliveira, Olga Fajarda, and Diogo Pratas. GTO: a toolkit to unify pipelines in genomic and proteomic research. SoftwareX, 12 (??):Article 100535, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020301473.

Archetti:2020:DMP

[Arc20]  Marco Archetti.  De-

Finetti: a Mathematica program to analyze the replicator dynamics of 3strategy collective interactions. SoftwareX, 11 (??):Article 100415, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271101930216X.

Arslan:2021:FEC

[Ars21] Suayb S. Arslan. Founsure 1.0: an erasure code library with efficient repair and update features. SoftwareX, 13(??):Article 100662, January 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021000078.

Amosu:2017:WIP

[AS17] Adewale Amosu and Yue-

feng Sun. WheelerLab:

an interactive program for sequence stratigraphic analysis of seismic sections, outcrops and well sections and the generation of chronostratigraphic sections and dynamic chronostratigraphic sections. SoftwareX, 6 (??):13–18, ???? 2017. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711016300401.

Abdul-Salam:2018:SAD

[AS18] Yakubu Abdul-Salam. A Stochastic Anaerobic Digestion Economic Assessment Tool (SADEAT). SoftwareX, 7(??):190–197, January/June 2018. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/

science/article/pii/
S235271101830058X.

Amato:2023:JJB

[AS23] Gianluca Amato and Francesca Scozzari. JGMP: Java bindings and wrappers for the GMP library. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023001243.

Ahmadzadeh:2020:MDT

[ASAA20] Azim Ahmadzadeh, Kankana Sinha, Berkay Aydin, and Rafal A. Angryk.

MVTS-Data Toolkit: a

Python package for preprocessing multivariate time series data. SoftwareX, 12(??):Article 100518, July/December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020300157.

Alshouha:2024:BFP

[ASGC^{+}24] Bashar Alshouha, Jesus Serrano-Guerrero, Francisco Chiclana, Francisco P. Romero, and

Jose A. Olivas. BioEmoDetector:

a flexible platform for detecting emotions from health narratives. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/

science/article/pii/
S2352711024000414.

Araque:2022:GSA

[ASRI22] Oscar Araque, J. Fernando Sánchez-Rada, and Carlos A. Iglesias. GSITK: a sentiment analysis framework for agile replication and development. SoftwareX, 17(??):??, January 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021001643.

Amaral:2024:TPP

[ASSS24] Miguel Amaral, Gabriel Signoretti, Marianne Silva, and Ivanovitch Silva. TAC: a Python package for IoTfocused Tiny Anomaly Compression. SoftwareX, 26(??):??, May 2024. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024001183.

Ariza-Sentis:2023:BOS

[ASVV23] Mar Ariza-Sentís, Sergio Vélez, and João Valente. BBR: an open-source standard workflow based on biophysical crop parameters for automatic Botrytis cinerea assessment in vineyards. SoftwareX, 24 (??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S2352711023002388.

Ananth:2022:TPJ

[AT22] Mohan Ananth and Mario F. Trujillo. 2PJIT: Twophase 3D jet instability tool in cylindrical coordinates. SoftwareX, 18 (??):??, June 2022. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022000206.

Ardito:2020:PRP

[ATCA20] Luca Ardito, Marco Torchiano, Riccardo Coppola, and Giulio Antoniol. PowTrAn: an R package for power trace analysis. SoftwareX, 12(??):Article 100512, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S235271102030025X.

Anastasopoulos:2021:GPT

[ATT21] Nikolaos Anastasopoulos, Ioannis G. Tsoulos, and Alexandros Tzallas. GenClass: a parallel tool for data classification based on Grammatical Evolution. SoftwareX, 16(??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021001199.

Almas:2017:BTI

[AVB17] M. S. Almas, L. Vanfretti, and M. Baudette.

BabelFish — tools for

IEEE C37.118.2-compliant real-time synchrophasor data mediation. SoftwareX, 6(??):203–208, ???? 2017. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711017300328.

Ahmad:2021:MMI

[AVCP^{+}21] Ali Ahmad, Guillaume Vanel, Sorina Camarasu-Pop, Axel Bonnet, Carole Frindel, and David Rousseau. MicroVIP: Microscopy image simulation on the Virtual Imaging Platform. SoftwareX, 16 (??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271102100131X.

Aksman:2021:PPI

[AWO^{+}21] Leon M. Aksman, Peter A.

Wijeratne, Neil P. Oxtoby, Arman Eshaghi, Cameron Shand, Andre Altmann, Daniel C. Alexander, and Alexandra L. Young. pySuStaIn: a Python implementation of the Subtype and Stage Inference algorithm. SoftwareX, 16 (??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021001096.

Ashuri:2017:CCS

[AZ17] Turaj Ashuri and Jie Zhang. CompSim: Cross sectional modeling of geometrical complex and inhomogeneous slender structures. SoftwareX, 6 (??):150–154, ???? 2017. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711017300109.

Azis:2023:XLF

[Azi23] Mohammad Ivan Azis.

                                        2D-xtDCR-LTBEM: a For-

tran code for numerical solutions to the unsteady anisotropic-DCR equation of spatio-temporal coefficients. SoftwareX, 23 (??):??, July 2023. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023002066.

Bach:2021:PTB

[Bac21] Eviatar Bach. parasweep: a template-based utility for generating, dispatching, and post-processing of parameter sweeps. SoftwareX, 13(??):Article 100631, January 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020303447.

Boashash:2019:MTF

[BAEBAS19] Boualem Boashash, Abdeldjalil Aïssa-El-Bey, and Mohammad F. Al-Sa’d. Multisensor time-frequency signal processing MAT-LAB package: an analysis tool for multichannel non-stationary data. SoftwareX, 8(??):53–58, ???? 2019. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711017300687.

Bordon:2022:MMD

[BÁP^{+}22] Jacob D. R. Bordón, Guillermo M. Álamo, Luis A. Padrón, Juan J. Aznárez, and Orlando

Maeso.  MultiFEBE:

a multi-domain finite element-boundary element solver for linear mixeddimensional mechanical problems. SoftwareX, 20 (??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022001832.

Batista:2019:ECM

[Bat19] Milan Batista. Elfun18 — a collection of MATLAB functions for the computation of elliptic integrals and Jacobian elliptic functions of real arguments. SoftwareX, 10(??):Article 100245, July/December 2019. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018302796.

Bondua:2020:TMT

[BB20a] S. Bonduà and V. Bortolotti. TOUGH2Viewer 2.0: a multiplatform tool for fully 3D Voronoi TOUGH grids. SoftwareX, 12(??):Article 100596, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020303095.

Bruder:2020:CQC

[BB20b] Brittany L. Bruder and Katherine L. Brodie. CIRN

Quantitative Coastal
Imaging Toolbox. Soft-

wareX, 12(??):Article 100582, July/December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020302958.

Badalyan:2022:AEC

[BB22a] David Badalyan and Oleg Borisenko. Ansible execution control in Python and Golang for cloud orchestration. SoftwareX, 19 (??):??, July 2022. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022000826.

Bragagnolo:2022:SPL

[BB22b] Andrea Bragagnolo and Carlo Alberto Barbano.

Simplify:  a Python

library for optimizing pruned neural networks. SoftwareX, 17(??):??, January 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021001576.

Berghout:2023:PDG

[BB23] Tarek Berghout and Mohamed Benbouzid. PrognosEase: a data generator for health deterioration prognosis. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023001577.

Blais:2020:LOS

[BBB^{+}20] Bruno Blais, Lucka Barbeau, Valérie Bibeau, Simon Gauvin, Toni El Geitani, Shahab Golshan, Rajeshwari Kamble, Ghazaleh Mirakhori, and Jamal Chaouki. Lethe: an opensource parallel high-order adaptative CFD solver for incompressible flows. SoftwareX, 12(??):Article 100579, July/December 2020. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711020302922.

Betz:2021:RSR

[BBB21] Fridtjof Betz, Felix Binkowski, and Sven Burger. RPExpand: Software for Riesz projection expansion of resonance phenomena. SoftwareX, 15(??):??, July 2021. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021000844.

Billings:2018:EIC

[BBD^{+}18] Jay Jay Billings, Andrew R. Bennett, Jordan Deyton, Kasper Gammeltoft, Jonah Graham, Dasha Gorin, Hari Krishnan, Menghan Li, Alexander J. McCaskey, Taylor Patterson, Robert Smith, Gregory R. Watson, and Anna Wojtowicz. The Eclipse Integrated Computational Environment. SoftwareX, 7(??):234–244, January/June 2018. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S235271101830133X.

Baldini:2022:ISA

[BBD^{+}22] Luca Baldini, Niccolò Bucciantini, Niccolò Di Lalla, Steven Ehlert, Alberto Manfreda, Michela Negro, Nicola Omodei, Melissa Pesce-Rollins, Carmelo Sgrò, and Stefano Silvestri.

ixpeobssim: a simula-

tion and analysis framework for the imaging Xray polarimetry explorer. SoftwareX, 19(??):??, July 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022001169.

Bortolotti:2019:UUP

[BBF^{+}19] V. Bortolotti, L. Brizi, P. Fantazzini, G. Landi, and F. Zama. Upen2DTool: a Uniform PENalty Matlab tool for inversion of 2D NMR relaxation data. SoftwareX, 10(??):Article 100302, July/December 2019. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019302006.

Bizzego:2019:PPS

[BBG^{+}19] Andrea Bizzego, Alessandro Battisti, Giulio Gabrieli, Gianluca Esposito, and Cesare Furlanello. pyphysio: a physiological signal processing library for data science approaches in physiology. SoftwareX, 10 (??):Article 100287, July/ December 2019. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711019301839.

Bendarag:2022:PPB

[BBHB22] Abdesadik Bendarag, Jamal Bakkas, Mohamed Hanine, and Omar Boutkhoum.

                                        PyOPAsolver: a Python

based tool for ordinal priority approach operations and normalization. SoftwareX, 20(??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022001443.

Bidegain:2023:MGB

[BBHP^{+}23] Gorka Bidegain, Tal Ben-Horin, Eric N. Powell, John M. Klinck, and Eileen E. Hofmann. MarineEpi: a GUI-based Matlab toolbox to simulate marine pathogen transmission. SoftwareX, 22(??):??, May 2023. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023000535.

Basden:2018:DAO

[BBJ^{+}18] A. G. Basden, N. A. Bharmal, D. Jenkins, T. J. Morris, J. Osborn, J. Peng, and L. Staykov. The Durham Adaptive Opti’s Simulation Platform (DASP): Current status. SoftwareX, 7(??):63– 69, January/June 2018. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018300232.

Bortolotti:2022:MNM

[BBN^{+}22] Villiam Bortolotti, Leonardo Brizi, Anastasiia Nagmutdinova, Fabiana Zama, and Germana Landi. MUPen2DTool: a new Matlab tool for 2D nuclear magnetic resonance relaxation data inversion. SoftwareX, 20 (??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022001583.

Bauer:2018:QCT

[BBP^{+}18] Paul Bauer, Alexandre Barrozo, Miha Purg, Beat Anton Amrein, Mauricio Esguerra, Philippe Barrie Wilson, Dan Thomas Major, Johan Åqvist, and Shina Caroline Lynn Kamerlin. Q6: a comprehensive toolkit for empirical valence bond and related free energy calculations. SoftwareX, 7(??): 388–395, January/June 2018. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711017300675.

Borrego:2022:SWF

[BBS^{+}22] Agustín Borrego, Miguel Bermudo, Fernando Sola, Daniel Ayala, Inma Hernández, and David Ruiz. Silence — a web framework for an agile generation of RESTful APIs. SoftwareX, 20 (??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022001789.

Bispo:2020:CCC

[BC20] João Bispo and João M. P. Cardoso. Clava: C/C++ source-to-source compilation using LARA. SoftwareX, 12(??):Article 100565, July/December 2020. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711019302122.

Black:2022:MBR

[BC22] Travis J. Black and Alexei F. Cheviakov. 3DRSP: Matlab-based random sphere packing code in three dimensions. SoftwareX, 18(??):??, June 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022000437.

Bak:2023:WNO

[BÇ23] Mehmet Bak and Rahmi Nurhan Çeli̇k. Web-NDefA: Opensource and web-based online platform for 3-D deformation analysis of geodetic networks. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023002194.

Basaran:2019:MAM

[BCA19] Dogac Basaran, Ali Taylan Cemgil, and Emin Anarim. Multiresolution alignment for multiple unsynchronized audio sequences using Sequential Monte Carlo samplers. SoftwareX, 8(??): 33–38, ???? 2019. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271101730064X.

Badia:2015:CSI

[BCD^{+}15] Rosa M. Badia, Javier Conejero, Carlos Diaz, Jorge Ejarque, Daniele Lezzi, Francesc Lordan, Cristian Ramon-Cortes, and Raul Sirvent. COMP Superscalar, an interoperable programming framework. SoftwareX, 3–4(??): 27–31, December 2015. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711015000151.

Borzacchiello:2022:SSE

[BCD22] Luca Borzacchiello, Emilio Coppa, and Camil Demetrescu. SENinja: a symbolic execution plugin for

                                        Binary Ninja. SoftwareX,

20(??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022001376.

Bryngelson:2020:QLQ

[BCF20] Spencer H. Bryngelson, Tim Colonius, and Rodney O. Fox. QBMMlib: a library of quadraturebased moment methods. SoftwareX, 12(??):Article 100615, July/December 2020. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711020303289.

Bulat:2024:DVT

[BCFJ24] Jarosław Bułat, Natalia Cieplińska, Rafał Figlus, and Lucjan Janowski. Daily video: a tool for quality of experience (QoE) in long-term context research. SoftwareX, 25(??):??, February 2024. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711024000086.

Baudette:2018:OOI

[BCR^{+}18] Maxime Baudette, Marcelo Castro, Tin Rabuzin, Jan Lavenius, Tetiana Bogodorova, and Luigi Vanfretti. OpenIPSL: Open-Instance Power System Library — update 1.5 to “iTesla Power Systems Library (iPSL): a Modelica library for phasor timedomain simulations”. SoftwareX, 7(??):34–36, January/June 2018. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018300050.

Blanco-Claraco:2023:MSM

[BCTMA^{+}23] José-Luis Blanco-Claraco, Borys Tymchenko, Francisco José Mañas-Alvarez, Fernando Cañadas-Aránega, Ángel López-Gázquez, and José Carlos Moreno. MultiVehicle Simulator (MVSim): Lightweight dynamics simulator for multiagents and mobile robotics research. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023001395.

Byun:2022:RMT

[BD22] Ji-Eun Byun and Dina

D’Ayala.  ResMapper:

Matlab tool for seismic resilience mapping of large-scale road networks. SoftwareX, 20 (??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022001674.

Bartholomew:2020:XOS

[BDF^{+}20] Paul Bartholomew, Georgios Deskos, Ricardo A. S. Frantz, Felipe N. Schuch, Eric Lamballais, and Sylvain Laizet. Xcompact3D: an open-source framework for solving turbulence problems on a Cartesian mesh. SoftwareX, 12 (??):Article 100550, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019303620.

Behrendt:2019:RQI

[BDPZ19] Simon Behrendt, Thomas Dimpfl, Franziska J. Peter, and David J. Zimmermann. RTransferEntropy — Quantifying information flow between different time series using effective transfer entropy. SoftwareX, 10(??):Article 100265, July/December 2019. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019300779.

Brimicombe:2022:TPT

[BDQ^{+}22] Chloe Brimicombe, Claudia Di Napoli, Tiago Quintino, Florian Pappenberger, Rosalind Cornforth, and Hannah L.

Cloke. Thermofeel: a

Python thermal comfort indices library. SoftwareX, 18(??):??, June 2022. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022000176.

Buchicchio:2022:EST

[BDSC22] Emanuele Buchicchio, Alessio De Angelis, Francesco Santoni, and Paolo Carbone. EasyEIS: a software tool to perform electrochemical impedance spectroscopy using a source measure unit. SoftwareX, 18(??): ??, June 2022. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022000553.

Brown:2023:ARP

[BdSvE23] David Brown, Kauê de Sousa, and Jacob van Etten. ag5Tools: an R package for downloading and extracting agrometeorological data from the AgERA5 database. SoftwareX, 21 (??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022001856.

Boeira:2020:PPP

[BE20] Emerson Boeira and Diego Eckhard. pyvrft: a Python package for the Virtual Reference Feedback Tuning, a direct datadriven control method. SoftwareX, 11(??):Article 100383, January/June 2020. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711019302894.

Bedkowski:2024:EEN

[Bed24a] Janusz Bedkowski. End to end navigation stack for nuclear power plant inspection with mobile robot. SoftwareX, 26 (??):??, May 2024. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711024001213.

Bedkowski:2024:OSO

[Bed24b] Janusz Bedkowski. Open source, open hardware hand-held mobile mapping system for large scale surveys. SoftwareX, 25 (??):??, February 2024. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S235271102300314X.

Beling:2024:BRL

[Bel24] Piotr Beling. BSuccinct:

Rust libraries and programs focused on succinct data structures. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024000529.

BenSaida:2015:PTN

[Ben15] Ahmed BenSaïda. A practical test for noisy chaotic dynamics. SoftwareX, 3– 4(??):i, December 2015. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711015000096.

Bai:2024:AAC

[BEPB24] Zhe Bai, Abdelilah Essiari, Talita Perciano, and Kristofer E. Bouchard. AutoCT: Automated CT registration, segmentation, and quantification. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271102400044X.

Bueler-Faudree:2022:FSW

[BFDD^{+}22] Thomas Bueler-Faudree, Sam Delamere, Denys Dutykh, Alexei Rybkin, and Alexander Suleimani. Fast shallow water-wave solver for plane inclined beaches. SoftwareX, 17 (??):??, January 2022. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022000061.

Barandas:2020:TTS

[BFF^{+}20] Marília Barandas, Duarte Folgado, Letícia Fernandes, Sara Santos, Mariana Abreu, Patrícia Bota, Hui Liu, Tanja Schultz, and Hugo Gamboa. TSFEL: Time Series Feature Extraction Library. SoftwareX, 11 (??):Article 100456, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020300017.

Boden:2021:RIE

[BFF^{+}21] Brigitte Boden, Jan Flink, Niklas Först, Robert Mischke, Kathrin Schaffert, Alexander Weinert, Annika Wohlan, and Andreas Schreiber. RCE: an integration environment for engineering and science. SoftwareX, 15(??):??, July 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021000820.

Bellini:2024:SDM

[BFNP24] Pierfrancesco Bellini, Marco Fanfani, Paolo Nesi, and Gianni Pantaleo. Snap4City dashboard manager: a tool for creating and distributing complex and interactive dashboards with no or low coding. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024001006.

Bernardi:2019:AUN

[BFRK19] Austen Bernardi, Roland Faller, Dirk Reith, and Karl N. Kirschner. ACPYPE update for nonuniform 1– 4 scale factors: Conversion of the GLY-CAM06 force field from AMBER to GROMACS. SoftwareX, 10(??):Article 100241, July/December 2019. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018300736.

Barrasa-Fano:2021:TMT

[BFSJP^{+}21] Jorge Barrasa-Fano, Apeksha Shapeti, Álvaro Jorge-Peñas, Mojtaba Barzegari, José Antonio Sanz-Herrera, and Hans Van Oosterwyck. TFMLAB: a MATLAB toolbox for 4D traction force microscopy. SoftwareX, 15(??):??, July 2021. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021000625.

Baudette:2018:SGL

[BFV18] Maxime Baudette, Seyed Reza Firouzi, and Luigi Vanfretti. The STRON grid library: a modular and extensible software library for IEEE C37.118.2 compliant synchrophasor data mediation. SoftwareX, 7(??): 281–286, January/June 2018. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711018301705.

Blanco-Gomez:2023:CAN

[BGAGJGC23] Pablo Blanco-Gómez, Constancio Amurrio-Garcia, José Luis Jiménez-García, and Jose M. Cecilia. CPR Algorithm — a new interpolation methodology and QGIS plugin for Colour Pattern Regression between aerial images and raster maps. SoftwareX, 22(??):??, May 2023. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023000523.

Blackman:2019:FCT

[BGCS19] Allen Blackman, Leonard Goff, Jessica Chu, and Juha Siikamäki. The Forest Conservation Targeting Tool: Accessible spatial prioritization for Latin America and the Dominican Republic. SoftwareX, 10(??):Article 100293, July/December 2019. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019300226.

Ben-Gida:2020:OMO

[BGGL20] Hadar Ben-Gida, Roi Gurka, and Alex Liberzon. OpenPIV-MATLAB — an open-source software for particle image velocimetry; test case: Birds’ aerodynamics. SoftwareX, 12 (??):Article 100585, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020302983.

Berardi:2024:RMT

[BGM^{+}24] Davide Berardi, Saverio Giallorenzo, Andrea Melis, Simone Melloni, and Marco Prandini. Ranflood: a mitigation tool based on the principles of data flooding against ransomware. SoftwareX, 25(??):??, February 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S2352711023003011.

Bork:2021:AAL

[BHB^{+}21] Rolf Bork, Jonathan Hanks, David Barker, Joseph Betzwieser, Jameson Rollins, Keith Thorne, and Erik von Reis. advligorts: the Advanced LIGO realtime digital control and data acquisition system. SoftwareX, 13(??):Article 100619, January 2021. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711020303320.

Belattmania:2024:CMG

[BHC^{+}24] Ayoub Belattmania, Mounir Hakkou, Taoufiq Chtioui, Abdelhaq Aangri, Assia Abdenour, El mehdi Latni, Abdessalam Benharra, Abdelkrim EL Arrim, Azdine Dahaoui, Ahmed Raissouni, Lamiae Khali Issa, and Lhoussaine Ed-daoudy. Classwms: a MATLAB GUI for water masses classification using clustering analysis and K nearest neighbors. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711024001122.

Bolan:2016:LFI

[BHCT16] Jeffrey Bolan, Elise Hall, Chris Clifford, and Brian Thurow. Light-Field Imaging Toolkit. SoftwareX, 5 (??):96–100, ???? 2016. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711016300127.

Benkert:2019:MER

[BHE^{+}19] T. Benkert, C. Hartmann, M. Eder, F. Speckmaier, and W. Volk. MaterialModeler-From experimental raw data to a material model. SoftwareX, 10(??):Article 100249, July/December 2019. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019300147.

Boito:2022:APC

[BHHÖ22] Deneb Boito, Magnus Herberthson, Tom Dela Haije, and Evren Özarslan. Applying positivity constraints to q-space trajectory imaging: the QTI+ implementation. SoftwareX, 18(??):??, June 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022000322.

Brehm:2023:POS

[BHKK23] Simon Brehm, Cameliu Himcinschi, Jakob Kraus, and Jens Kortus. PyRamanGUI:

Open-source graphical user interface for analyzing Raman spectra. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023001826.

Bazzi:2019:AAI

[BJM^{+}19] Claudio Leones Bazzi, Erminio Pita Jasse, Paulo S. Graziano Magalhães, Gabriela Karoline Michelon, Eduardo Godoy de Souza, Kelyn Schenatto, and Ricardo Sob-

jak. AgDataBox API —

integration of data and software in precision agriculture. SoftwareX, 10 (??):Article 100327, July/ December 2019. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018302838.

Bernardi:2023:TEP

[BJM23a] Simona Bernardi, Raúl Javierre, and José Merseguer. tegdet: an extensible Python library for anomaly detection using time evolving graphs. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023000596.

Boom:2023:PID

[BJM23b] Pieter D. Boom, Andrey P.

Jivkov, and Lee Margetts. ParaGEMS: Integrating discrete exterior calculus (DEC) into ParaFEM for geometric analysis of solid mechanics. SoftwareX, 21(??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022001984.

Bjornson:2019:TQM

[Bjö19] Kristofer Björnson. TBTK:

a quantum mechanics software development kit. SoftwareX, 9(??):205–210, January/June 2019. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271101830164X.

Brown:2020:PPP

[BJR^{+}20] Daniel D. Brown, Philip Jones, Samuel Rowlinson, Sean Leavey, Anna C. Green, Daniel Töyrä, and Andreas Freise. Pykat: Python package for modelling precision optical interferometers. SoftwareX, 12(??):Article 100613, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020303265.

Bingol:2019:NPO

[BK19] Onur Rauf Bingol and Adarsh Krishnamurthy.

                                        NURBS-Python: an open-

source object-oriented NURBS modeling framework in Python. SoftwareX, 9 (??):85–94, January/June 2019. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711018301778.

Batalas:2021:EH

[BKM21] Nikolaos Batalas, Vassilis-Javed Khan, and Panos Markopoulos. Executable HTML. SoftwareX, 14 (??):??, June 2021. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021000364.

Buoncompagni:2022:OMA

[BKM22] Luca Buoncompagni, Syed Yusha Kareem, and Fulvio Mastrogiovanni. OWLOOP: a modular API to describe OWL axioms in OOP objects hierarchies. SoftwareX, 17(??):??, January 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021001801.

Bouafif:2016:PSM

[BL16] Mariem Bouafif and Zied

                                        Lachiri. SRC Num TDOA:

Multiple speech sources’ number and their TDOA Estimation from a stereo recorded mixture. SoftwareX, 5(??):227–233, ???? 2016. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711016300322.

Blauth:2021:CCA

[Bla21] Sebastian Blauth. cashocs:

a computational, adjointbased shape optimization and optimal control software. SoftwareX, 13(??): Article 100646, January 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020303599.

Blauth:2023:VCC

[Bla23] Sebastian Blauth. Version 2.0 — cashocs: Computational, adjointbased shape optimization and optimal control software. SoftwareX, 24(??): ??, December 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271102300273X.

Boudet:2020:FHR

[BlD^{+}20] S. Boudet, A. Houzé l’Aulnoit, R. Demailly, A. Delgranche, L. Peyrodie, R. Beuscart, and D. Houzé de l’Aulnoit. A fetal heart rate morphological analysis toolbox for MATLAB. SoftwareX, 11 (??):Article 100428, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018302498.

Bizzego:2021:MLP

[BLE21] Andrea Bizzego, Mengyu Lim, and Gianluca Espos-

ito. mics-library: a

Python package for reproducible studies on the Multiple Indicator Cluster Survey. SoftwareX, 16 (??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021001187.

Bromig:2022:SMR

[BLM^{+}22] Lukas Bromig, David Leiter, Alexandru-Virgil Mardale, Nikolas von den Eichen, Emmeran Bieringer, and Dirk Weuster-Botz.

The SiLA 2 Manager for

rapid device integration and workflow automation. SoftwareX, 17(??): ??, January 2022. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022000103.

Bangaru:2021:RSI

[BMBR21] Madhavi Latha Yadav Bangaru, Ravi Kumar Medabalimi, Sobhan Babu, and Nidhanapati K. Raghavendra. REMP software to introduce a screening REstriction site in site-directed Mutagenesis Primer. SoftwareX, 16 (??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021001448.

Badie-Modiri:2023:RTN

[BMK23] Arash Badie-Modiri and Mikko Kivelä. Reticula: a temporal network and hypergraph analysis software package. SoftwareX, 21(??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022002199.

Baldan:2017:SDT

[BMR17] Stefano Baldan, Stefano Delle Monache, and Davide Rocchesso. The Sound Design Toolkit. SoftwareX, 6(??):248–254, ???? 2017. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711017300195.

Berezowski:2024:FAM

[BNC24] Tomasz Berezowski, Szymon Niemiec, and Andrzej Chybicki. Floodsar: Automatic mapping of river flooding extent from multitemporal SAR imagery. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711024000888.

Boashash:2019:ESP

[BO19] Boualem Boashash and Samir Ouelha. Efficient software platform TFSAP 7.1 and Matlab package to compute time-frequency distributions and related time-scale methods with extraction of signal characteristics. SoftwareX, 8 (??):48–52, ???? 2019. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711017300353.

Boelens:2023:SST

[Boe23] Arnout M. P. Boelens. Stplanpy: a sustainable transportation planner for Python. SoftwareX, 22 (??):??, May 2023. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023000353.

Betkier:2023:PMF

[BOP^{+}23] Igor Betkier, Mateusz Oszczypała, Janusz Pobozniak, Sergiusz Sobieski, and Przemysław Betkier.

PocketFinderGNN: a man-

ufacturing feature recognition software based on Graph Neural Networks (GNNs) using PyTorch Geometric and NetworkX. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023001620.

Besnard:2019:FAP

[BPL^{+}19] Loïc Besnard, Pedro Pinto, Imane Lasri, João Bispo, Erven Rohou, and João M. P. Cardoso. A framework for automatic and parameterizable memoization. SoftwareX, 10 (??):Article 100322, July/ December 2019. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018301559.

Breidt:2023:BIG

[Bre23] Fred Breidt. BufferCapacity3 an interactive GUI program for modelling food ingredient buffering and pH. SoftwareX, 22(??): ??, May 2023. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S235271102300047X.

Briegel:2023:PPP

[BRF^{+}23] Karl D. Briegel, Felix Riccius, Jakob Filser, Alexander Bourgund, Robert Spitzenpfeil, Mirco Panighel, Carlo Dri, Barbara A. J. Lechner, and Friedrich

                                        Esch.  PyfastSPM: a

Python package to convert 1D FastSPM data streams into publication quality movies. SoftwareX, 21(??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S235271102200187X.

Bianchi:2019:LAM

[BS19] Carlo Bianchi and Amanda D. Smith. Localized actual meteorological year file creator (LAF): a tool for using locally observed weather data in building energy simulations. SoftwareX, 10(??):Article 100299, July/December 2019. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711018300839.

Bernardi:2021:CPM

[BS21] Austen Bernardi and Jessica M. J. Swanson. CycFlowDec: a Python module for decomposing flow networks using simple cycles. SoftwareX, 14(??): ??, June 2021. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021000212.

Bautista:2020:MSM

[BSB20] Ilia Bautista, Sudeep Sarkar, and Sanjukta

Bhanja. MatlabHTM: a

sequence memory model of neocortical layers for anomaly detection. SoftwareX, 11(??):Article 100491, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018302863.

Bomfim:2023:HHE

[BSB^{+}23] Danilo S. Bomfim, Rodrigo L. Soares, Luiz F. Bez, Pedro C. F. Lopes, André M. B. Pereira, and Luiz F. Martha. HETOOL: a Half-Edge Topological Object-Oriented Library for generic 2-D geometric modeling. SoftwareX, 21(??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023000031.

Bota:2024:BPT

[BSC^{+}24] Patrícia Bota, Rafael Silva, Carlos Carreiras, Ana Fred, and Hugo Plácido da Silva. BioSPPy: a Python toolbox for physiological signal processing. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024000839.

Backe:2022:EOS

[BSdG^{+}22] Stian Backe, Christian Skar, Pedro Crespo del Granado, Ozgu Turgut, and Asgeir Tomasgard. EMPIRE: an open-source model based on multihorizon programming for energy transition analyses. SoftwareX, 17(??): ??, January 2022. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021001424.

Bub:2020:XRP

[BSG20] Sascha Bub, Thorsten Schad, and Zhenglei Gao. XDrift — an R package to simulate spatially explicit pesticide spray-drift exposure of non-target-species habitats at landscape scales. SoftwareX, 12 (??):Article 100610, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271102030323X.

Ballsun-Stanton:2018:FMF

[BSRSC18] Brian Ballsun-Stanton, Shawn A. Ross, Adela Sobotkova, and Penny Crook. FAIMS Mobile: Flexible, open-source software for field research. SoftwareX, 7(??):47–52, January/June 2018. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711017300869.

Breidt:2023:IGB

[BSS^{+}23] Fred Breidt, Caitlin R.

Skinner, Mileah Shriner, Mollie Ruinsky, Seo Young Yang, Robert P. Wine, and Lynette Johnston.

IngredientDB: a GUI-

based Matlab database program for estimating the pH of acid or acidified food formulations from buffer capacity models. SoftwareX, 24(??): ??, December 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023002418.

Binder:2017:QMP

[BST^{+}17] Jan M. Binder, Alexander Stark, Nikolas Tomek, Jochen Scheuer, Florian Frank, Kay D. Jahnke, Christoph Müller, Simon Schmitt, Mathias H. Metsch, Thomas Unden, Tobias Gehring, Alexander Huck, Ulrik L. Andersen, and Lachlan J. Rogers Fedor Jelezko. Qudi: a modular Python suite for experiment control and data processing. SoftwareX, 6 (??):81–84, ???? 2017. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711017300055.

Busato:2023:PWB

[BSV^{+}23] Angela Busato, Michaela Servi, Yary Volpe, Rocco Furferi, and Maurizio Scorianz. Precise: a webbased 3D visualization and manipulation application for surgical planning of tumour resection. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023000948.

Bogaerts:2020:CCR

[BSVP20] Boris Bogaerts, Seppe Sels, Steve Vanlanduit, and Rudi Penne. Connecting the CoppeliaSim robotics simulator to virtual reality. SoftwareX, 11 (??):Article 100426, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019302511.

Boelens:2021:QMF

[BT21] Arnout M. P. Boelens and Hamdi A. Tchelepi. QuantImPy: Minkowski functionals and functions with Python. SoftwareX, 16(??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021001151.

Blazquez:2023:KSA

[BT23] Eduardo Blázquez and Juan Tapiador. Kunai: a static analysis framework for Android apps. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023000663.

Brekel:2023:VPF

[BTDT23] Josh Brekel, Kelly R.

Thorp, Kendall C. De-Jonge, and Thomas J. Trout. Version 1.1.0pyfao56: FAO-56 evapotranspiration in Python. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023000328.

Bujack:2021:OSV

[BTMB21] Roxana Bujack, Karen Tsai, Steven K. Morley, and Etienne Bresciani.

Open source vector field topology. SoftwareX, 15 (??):??, July 2021. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021000960.

Bos:2019:CMC

[BV19] Len Bos and Marco Vianello. CaTchDes: MATLAB codes for Caratheodory– Tchakaloff near-optimal regression designs. SoftwareX, 10(??):Article 100349, July/December 2019. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019302274.

Baczkiewicz:2022:CPL

[BW22] Aleksandra Baczkiewicz and Jarosław Watróbski. Crispyn — a Python library for determining criteria significance with objective weighting methods. SoftwareX, 19(??):??, July 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022001029.

Bodker:2022:SPB

[BWMS22] Mikkel S. Bødker, Collin J. Wilkinson, John C. Mauro, and Morten M. Smedskjaer. StatMechGlass: Python based software for composition-structure prediction in oxide glasses using statistical mechanics. SoftwareX, 17(??): ??, January 2022. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021001606.

Baczkiewicz:2023:VCP

[BWS23] Aleksandra Baczkiewicz, Jarosław Watróbski, and Wojciech Sałabun. Version [1.1] — [Crispyn — a Python library for determining criteria significance with objective weighting methods]. SoftwareX, 24 (??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023002376.

Bykhovsky:2019:MCN

[Byk19] D. Bykhovsky. Mathematica code for numerical generation of random process with given distribution and exponential autocorrelation function. SoftwareX, 8(??):18–20, ???? 2019. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271101730033X.

Cherubin:2018:LEU

[CA18] S. Cherubin and G. Agosta.

libVersioningCompiler: an easy-to-use library for dynamic generation and invocation of multiple code versions. SoftwareX, 7(??): 95–100, January/June 2018. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711018300426.

Caliskan:2020:GMG

[ÇA20] Murat Çaliskan and Berk Anbaroglu. Geo-MST: a geographical minimum spanning tree plugin for QGIS. SoftwareX, 12(??):Article 100553, July/December 2020. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711020301771.

Caliskan:2023:STC

[ÇA23] Murat Çaliskan and Berk Anbaroglu. Space Time Cube analytics in QGIS and Python for hot spot detection. SoftwareX, 24 (??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023001942.

Castano:2024:QPL

[CAG^{+}24] Alberto Castaño, Jaime Alonso, Pablo González, Pablo Pérez, and Juan José

                                        del Coz. QuantificationLib:

a Python library for quantification and prevalence estimation. SoftwareX, 26 (??):??, May 2024. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024000992.

Cantero:2022:TTT

[Can22a] Daniel Cantero. TTB-2D:

Train–Track–Bridge interaction simulation tool for Matlab. SoftwareX, 20 (??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022001716.

Cantero:2022:VEM

[Can22b] Daniel Cantero. VEqMon2D — equations of motion generation tool of 2D vehicles with Matlab. SoftwareX, 19(??):??, July 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022000693.

Cantero:2024:VRV

[Can24] Daniel Cantero. VBI-2D — road vehicle-bridge interaction simulation tool and verification framework for Matlab. SoftwareX, 26 (??):??, May 2024. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024000967.

Carmignani:2021:FTI

[Car21] Luca Carmignani. Flame Tracker: an image analysis program to measure flame characteristics. SoftwareX, 15(??):??, July 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021000984.

Conlin:2020:STR

[CAW^{+}20] Matthew P. Conlin, Peter N. Adams, Benjamin Wilkinson, Gregory Dusek, Margaret L. Palmsten, and Jenna A. Brown. SurfRCaT: a tool for remote calibration of preexisting coastal cameras to enable their use as quantitative coastal monitoring tools. SoftwareX, 12 (??):Article 100584, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020302971.

Corradini:2021:URP

[CBHLG21] Fabio Corradini, Nicolas Beriot, Esperanza Huerta-Lwanga, and Violette Geissen. uFTIR: an R package to process hyperspectral images of environmental samples captured with μ FTIR microscopes. SoftwareX, 16 (??):??, December 2021.

CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021001321.

Ching:2022:QLO

[CBLI22] Eric J. Ching, Brett Bornhoft, Ali Lasemi, and Matthias Ihme. Quail: a lightweight open-source discontinuous Galerkin code in Python for teaching and prototyping. SoftwareX, 17(??):??, January 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271102200005X.

Carvalho:2023:DBR

[CBPC23] Tiago Carvalho, João Bispo, Pedro Pinto, and João M. P. Cardoso. A DSL-based runtime adaptivity framework for Java. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023001929.

Cooper:2016:TOS

[CBS^{+}16] S. J. Cooper, A. Bertei, P. R. Shearing, J. A. Kilner, and N. P. Bran-

don.  TauFactor: an

open-source application for calculating tortuosity factors from tomographic data. SoftwareX, 5(??):195–202, ???? 2016.

CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711016300280.

Cooper:2020:SDS

[CC20] Crispin H. V. Cooper and Alain J. F. Chiaradia. sDNA: 3-d spatial network analysis for GIS, CAD, command line and Python. SoftwareX, 12(??):Article 100525, July/December 2020. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711019303401.

Capodieci:2021:VCV

[CC21] Nicola Capodieci and Roberto Cavicchioli. vkpolybench: a crossplatform Vulkan Compute port of the Poly-Bench/GPU benchmark suite. SoftwareX, 15(??): ??, July 2021. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021000996.

Cinaglia:2023:MGA

[CC23] Pietro Cinaglia and Mario

                                        Cannataro. MultiGlobAl:

Global alignment of multilayer networks. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/

science/article/pii/
S2352711023002480.

Cattaneo:2022:TCB

[CCAC22] Daniele Cattaneo, Michele Chiari, Giovanni Agosta, and Stefano Cherubin. TAFFO: the compiler-based precision tuner. SoftwareX, 20(??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271102200156X.

Cantero-Chinchilla:2021:OCO

[CCBC^{+}21] Sergio Cantero-Chinchilla, James L. Beck, Juan Chiachío, Manuel Chiachío, and Dimitrios Chronopoulos. OptiSens — convex optimization of sensor and actuator placement for ultrasonic guided-wave based structural health monitoring. SoftwareX, 13 (??):Article 100643, January 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020303563.

Cannon:2021:GSF

[CCC^{+}21] Kipp Cannon, Sarah Caudill, Chiwai Chan, Bryce Cousins, Jolien D. E. Creighton, Becca Ewing, Heather Fong, Patrick Godwin, Chad Hanna, Shaun Hooper, Rachael Huxford, Ryan Magee, Duncan Meacher, Cody Messick, Soichiro Morisaki, Debnandini Mukherjee, Hiroaki Ohta, Alexander Pace, Stephen Privitera, Iris de Ruiter, Surabhi Sachdev, Leo Singer, Divya Singh, Ron Tapia, Leo Tsukada, Daichi Tsuna, Takuya Tsutsui, Koh Ueno, Aaron Viets, Leslie Wade, and Madeline Wade. GstLAL: a software framework for gravitational wave discovery. SoftwareX, 14(??):??, June 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271102100025X.

Chavis:2021:CMF

[CCE21] John Taylor Chavis, Amy Louise Cochran, and Christopher James Earls. CU-MSDSp: a flexible parallelized Reversible jump Markov chain Monte Carlo method. SoftwareX, 14 (??):??, June 2021. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021000091.

Cerecedo-Cordoba:2020:NPB

[CCFSB20] Jorge Alberto Cerecedo-Cordoba, Juan Frausto-Solís, and Juan Javier González

Barbosa. NeuroFramework:

a package based on neuroevolutionary algorithms to estimate the melting temperature of ionic liquids. SoftwareX, 11 (??):Article 100448, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019302924.

Couasnet:2023:GDA

[CCG^{+}23] Geoffroy Couasnet, Mathis Cordier, Hadhami Garbouge, Félix Mercier, David Pierre, Angelina El Ghaziri, Pejman Rasti, and David Rousseau.

Growth Data — an auto-

matic solution for seedling growth analysis via RGB-Depth imaging sensors. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023002686.

Carleo:2019:NML

[CCH^{+}19] Giuseppe Carleo, Kenny Choo, Damian Hofmann, James E. T. Smith, Tom Westerhout, Fabien Alet, Emily J. Davis, Stavros Efthymiou, Ivan Glasser, Sheng-Hsuan Lin, Marta Mauri, Guglielmo Mazzola, Christian B. Mendl, Evert van Nieuwenburg, Ossian O’Reilly, Hugo Théveniaut, Giacomo Torlai, Filippo Vicentini, and Alexander Wietek. NetKet: a machine learning toolkit for manybody quantum systems. SoftwareX, 10(??):Article 100311, July/December 2019. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711019300974.

Cardellini:2022:IPI

[CCI^{+}22] Valeria Cardellini, Emiliano Casalicchio, Stefano Iannucci, Matteo Lucantonio, Sudip Mittal, Damodar Panigrahi, and

                                        Andrea Silvi. irs-partition:

an Intrusion Response System utilizing Deep Q-Networks and system partitions. SoftwareX, 19 (??):??, July 2022. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022000796.

Ciatto:2021:PKL

[CCO21] Giovanni Ciatto, Roberta Calegari, and Andrea Omicini. 2 P-Kt: a logicbased ecosystem for symbolic AI. SoftwareX, 16 (??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021001126.

Castillo-Cara:2023:TCT

[CCTCGCOB23] Manuel Castillo-Cara, Reewos Talla-Chumpitaz, Raúl García-Castro, and Luis Orozco-Barbosa. TINTO: Converting Tidy Data into image for classification with 2-Dimensional Convolutional Neural Networks. SoftwareX, 22 (??):??, May 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023000870.

Chandramouli:2022:ASP

[CCV22] Pranav Chandramouli, Zadia Codabux, and Melina Vidoni. analyzeR: a SonarQube plugin for analyzing object-oriented R packages. SoftwareX, 19 (??):??, July 2022. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022000747.

Cambiaso:2023:DFE

[CCZ23] Enrico Cambiaso, Luca Caviglione, and Marco

Zuppelli. DockerChannel:

a framework for evaluating information leakages of Docker containers. SoftwareX, 24(??): ??, December 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S2352711023002728.

Cruz-Duarte:2020:CCO

[CDAOB^{+}20] Jorge M. Cruz-Duarte, Ivan Amaya, José C. Ortiz-Bayliss, Hugo Terashima-Marín, and Yong Shi. CUSTOMHyS: customising optimisation metaheuristics via hyper-heuristic search. SoftwareX, 12 (??):Article 100628, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711020303411.

Cruz-Duarte:2022:MMB

[CDOBA22] Jorge M. Cruz-Duarte, José C. Ortiz-Bayliss, and Ivan Amaya. MatHH: a Matlab-based Hyper-Heuristic framework. SoftwareX, 18(??):??, June 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022000413.

Cendes:2020:HAT

[CdSLCC20] Lucas L. Cendes, Welliton de Souza, Iscia Lopes-Cendes, and Benilton S. Carvalho. HPexome: an automated tool for processing whole-exome sequencing data. SoftwareX, 11 (??):Article 100478, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019303279.

Ceraolo:2019:MPG

[Cer19] Massimo Ceraolo. MC’s PlotXY — general-purpose plotting and post-processing open-source tool. SoftwareX, 9(??):282–287, January/June 2019. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019300287.

Cipriano:2022:DPA

[CFA22] Bruno Pereira Cipriano, Nuno Fachada, and Pe-

dro Alves. Drop Project:

an automatic assessment tool for programming assignments. SoftwareX, 18 (??):??, June 2022. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022000577.

Cabanero-Gomez:2021:EPM

[CGHGRB21] Luis Cabañero-Gomez, Ramon Hervas, Ivan Gonzalez, and Luis Rodriguez-Benitez. eeglib: a Python module for EEG feature extraction. SoftwareX, 15 (??):??, July 2021. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/

science/article/pii/
S2352711021000753.

Cugnet:2023:NST

[CGKD23] Mikaël Cugnet, Florian Gallois, Angel Kirchev, and Denys Dutykh. NEOLAB: a Scilab tool to simulate the Negative Electrode of Lead-Acid Batteries. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023000900.

Campbell:2019:NNG

[CGS19] Anna Charvátová Campbell, Petr Grolich, and Radek Slesinger. Niget: Nanoindentation general evaluation tool. SoftwareX, 9(??):248–254, January/June 2019. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018301353.

Cacho:2023:MBI

[CH23] Oscar J. Cacho and Susan M. Hester. Modelling biocontrol of invasive insects using Wasp-Sim: a MATLAB simulation model. SoftwareX, 21(??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023000171.

Charalambos:2017:PFR

[Cha17] Jean Pierre Charalambos.

Proscene: a feature-rich framework for interactive environments. SoftwareX, 6(??):42–47, ???? 2017. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271101730002X.

Cheliotis:2021:AAB

[Che21] Kostas Cheliotis. ABMU:

an Agent-Based Modelling Framework for Unity3D. SoftwareX, 15(??):??, July 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021000881.

Cassagne:2019:AFF

[CHL^{+}19] Adrien Cassagne, Olivier Hartmann, Mathieu Léonardon, Kun He, Camille Leroux, Romain Tajan, Olivier Aumage, Denis Barthou, Thibaud Tonnellier, Vincent Pignoly, Bertrand Le Gal, and Christophe Jégo. AFF3CT: A Fast Forward Error Correction Toolbox! SoftwareX, 10(??):Article 100345, July/December 2019. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019300457.

Cabrero-Holgueras:2023:HSE

[CHP23] José Cabrero-Holgueras and Sergio Pastrana. HEFactory: a symbolic execution compiler for privacy-preserving Deep Learning with Homomorphic Encryption. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023000924.

Cinaglia:2024:GWA

[Cin24] Pietro Cinaglia. GIN: a web-application for constructing synthetic datasets of interconnected networks in bioinformatics. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711024000189.

Carvallo:2023:CWB

[CJA23] Andrés Carvallo, Ignacio Jorquera, and Carlos As-

                                        pillaga. CoTranslate: a

web-based tool for crowdsourcing high-quality sentence pair corpora. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023002042.

Chaudhary:2023:PPP

[CJGA23] Gaurav Chaudhary, Hicham Johra, Laurent Georges, and Bjørn Austbø. pymodconn: a Python package for developing modular sequenceto-sequence control-oriented deep neural networks. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023002959.

Ciomek:2021:PJL

[CK21] Krzysztof Ciomek and Miłosz Kadziński. Polyrun: a Java library for sampling from the bounded convex polytopes. SoftwareX, 13 (??):Article 100659, January 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021000042.

Chand:2022:BPP

[CK22a] Kulbhushan Chand and Arun Khosla. BioNES: a plug-and-play MATLABbased tool to use NES games for multimodal biofeedback. SoftwareX, 19(??):??, July 2022. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/

                                        S235271102200111X. See

corrigendum [CK23].

Collins:2022:PTS

[CK22b] John Collins and Wolfgang Ketter. Power TAC: Software architecture for a competitive simulation of sustainable smart energy markets. SoftwareX, 20(??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022001352.

Chand:2023:CBP

[CK23] Kulbhushan Chand and Arun Khosla. Corrigendum to “BioNES: a plugand-play MATLAB-based tool to use NES games for multimodal biofeedback” [SoftwareX 19 (2022) 101184]. SoftwareX, 21 (??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S235271102200214X. See

[CK22a].

Czmil:2022:CCA

[CKC22] Sylwester Czmil, Jacek Kluska, and Anna Czmil. CACP: Classification Algorithms Comparison Pipeline. SoftwareX, 19(??):??, July 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/

science/article/pii/
S2352711022000863.

Czmil:2023:GPI

[CKC23] Anna Czmil, Jacek Kluska, and Sylwester Czmil. GPR: a Python implementation of an extremely simple classifier based on fuzzy logic and gene expression programming. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023000584.

Couvares:2021:FGW

[CKM21] Peter Couvares, Kate Keahey, and Frédérique Marion. Finding the gravitational wave: a history of discovery written in software. SoftwareX, 15 (??):??, July 2021. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021000601.

Castellazzi:2022:CFE

[CLDdM22] Giovanni Castellazzi, Nicolò Lo Presti, Antonio Maria D’Altri, and Stefano de Miranda. Cloud2FEM: a finite element mesh generator based on point clouds of existing/historical structures. SoftwareX, 18(??): ??, June 2022. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S235271102200067X.

Chung:2021:BRP

[CLK21] Jong-Hee Chung, Yong-Bin Lim, and Donghoh Kim. BDEsize: an R package for efficient determination of sample size in a balanced design of experiments. SoftwareX, 16 (??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021001229.

Cornwell:2020:TSS

[CLM^{+}20] J. A. Cornwell, J. Li, S. Mahadevan, J. S. Draper, G. L. Joun, H. Zoellner, N. S. Asli, R. P. Harvey, and R. E. Nordon. TrackPad: Software for semi-automated single-cell tracking and lineage annotation. SoftwareX, 11(??):Article 100440, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711019302390.

Chen:2022:HPT

[CLM22] Luo Chen, Zebang Liu, and Mengyu Ma. HiVecMap: a parallel tool for real-time geovisualization of massive geographic vector data. SoftwareX, 19(??):??, July 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022000917.

Castelli:2019:GCG

[CM19] Mauro Castelli and Luca Manzoni. GSGP-C++ 2.0: a geometric semantic genetic programming framework. SoftwareX, 10 (??):Article 100313, July/ December 2019. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019301736.

Chacon-Maldonado:2023:FSE

[CMACMÁT23] Andrés Manuel Chacón-Maldonado, Gualberto Asencio-Cortés, Francisco Martínez-Álvarez, and Alicia Troncoso. FS-Studio: an extensive and efficient feature selection experimentation tool for Weka Explorer. SoftwareX, 23 (??):??, July 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023000973.

Caro-Martinez:2023:EEE

[CMDDARG23] Marta Caro-Martinez, Jesus M. Darias, Belen Diaz-Agudo, and Juan A. Recio-Garcia. iSeeE^{3} — The Explanation Experiences Editor. SoftwareX, 21 (??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023000079.

Conti:2023:ARC

[CMdS^{+}23] Giuvane Conti, Erivelto Mercante, Eduardo Godoy de Souza, Ricardo Sobjak, and Claudio Leones Bazzi.

                                        AGDATABOX-RS computa-

tional application: Remote sensing data management. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023001310.

Cuny:2019:PAA

[CMMF19] Andreas P. Cuny, David Martínez-Martín, and Gotthold Fläschner. py-IMD: Automated analysis of inertial mass measurements of single cells. SoftwareX, 10(??):Article 100303, July/December 2019. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711019300871.

Cubukcuoglu:2020:DES

[CNST20] Cemre Cubukcuoglu, Pirouz Nourian, I. Sevil Sariyildiz, and M. Fatih Tasgetiren. A discrete event simulation procedure for validating programs of requirements: the case of hospital space planning. SoftwareX, 12(??):Article 100539, July/December 2020. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019303723.

Coblenz:2021:MVC

[Cob21] Maximilian Coblenz. MATVines:

a vine copula package for MATLAB. SoftwareX, 14 (??):??, June 2021. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021000455.

Cimen:2019:IAL

[COG19] Emre Cimen, Gurkan Ozturk, and Omer Nezih Gerek. ICF: an algorithm for large scale classification with conic functions. SoftwareX, 8(??):59–63, ???? 2019. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711017300699.

Cardinot:2019:EPA

[COGP19] Marcos Cardinot, Colm O’Riordan, Josephine Griffith, and Matjaz Perc. Evoplex: a platform for agent-based modeling on networks. SoftwareX, 9(??):199–204, January/ June 2019. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711018302437.

Cho:2024:GPG

[COK24] Hyeonglae Cho, Hee-Seok Oh, and Donghoh Kim. GSD: an R package for graph signal decomposition. SoftwareX, 26(??): ??, May 2024. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711024001249.

Correndo:2023:SRP

[CPB^{+}23] Adrian A. Correndo, Austin Pearce, Carl H. Bolster, John T. Spargo, Deanna Osmond, and Ignacio A. Ciampitti. The

                                        soiltestcorr R pack-

age: an accessible framework for reproducible correlation analysis of crop yield and soil test data. SoftwareX, 21(??): ??, February 2023. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022001935.

Chamorro-Padial:2024:DVD

[CPCCL24] Jorge Chamorro-Padial, Emanuele Cozzo, and Antonio Calleja-López.

                                        Decidim.viz: a data

dashboard prototype for a digital democracy platform. SoftwareX, 25(??): ??, February 2024. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023002960.

Castelluzzo:2020:MQD

[CPD^{+}20] Michele Castelluzzo, Alessio Perinelli, Simone Detassis, Michela Alessandra Denti, and Leonardo Ricci.

MiRNA-QC-and-Diagnosis:

an R package for diagnosis based on MiRNA expression. SoftwareX, 12 (??):Article 100569, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020301977.

Castilla:2023:DGM

[CPF23] Rodrigo Castilla, Alex Pacheco, and Jorge Franco. Digital government: Mobile applications and their impact on access to public information. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271102300078X.

Cabezas:2020:PSM

[CPUARC20] Xavier Alejandro Flores Cabezas, Martha Cecilia Paredes Paredes, Luis Felipe Urquiza-Aguiar, and Diego Javier Reinoso-Chisaguano. PhySim-11p: Simulation model for IEEE 802.11p physical layer in MATLAB. SoftwareX, 12(??):Article 100580, July/December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020302934.

Cirani:2020:ZNC

[CPV^{+}20] Simone Cirani, Marco Picone, Luca Veltri, Luca Zaccomer, and Francesco Zanichelli. ZWT: a new cross-platform graphical interface framework for Java applications. SoftwareX, 12(??):Article 100599, July/December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020303125.

Carini:2022:CPO

[CR22] Matheus Roman Carini and Marcelo Maia Rocha. CESSIPy: a Python opensource module for stochastic system identification in civil engineering. SoftwareX, 18(??):??, June 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022000632.

Colle:2021:MLA

[CRF21] Jean-Yves Colle, Jouni Rautio, and Daniel Freis. A modular LabVIEW application frame for Knudsen Effusion Mass Spectrometry instrument control. SoftwareX, 16(??): ??, December 2021. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021001412.

Cuny:2021:PTA

[CRP21] Andreas P. Cuny, Fabian Rudolf, and Aaron Ponti.

pyPOCQuant — a tool

to automatically quantify Point-Of-Care Tests from images. SoftwareX, 15 (??):??, July 2021. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021000558.

Candia-Rivera:2022:CPA

[CRV22] Diego Candia-Rivera and Gaetano Valenza. Cluster permutation analysis for EEG series based on non-parametric Wilcoxon– Mann–Whitney statistical tests. SoftwareX, 19 (??):??, July 2022. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022001042.

Capocchi:2021:WBS

[CS21] Laurent Capocchi and Jean-Francois Santucci. A web-based simulation of discrete-event system of system with the mobile application DEVSimPymob. SoftwareX, 13(??): Article 100625, January 2021. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711020303381.

Carminati:2019:KGP

[CSK19] Chiara Carminati, Markus Strobl, and Anders Kaestner. KipTool, a general purpose processing tool for neutron imaging data. SoftwareX, 10(??):Article 100279, July/December 2019. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711019300718.

Caviglione:2022:IIC

[CSZM22] Luca Caviglione, Andreas Schaffhauser, Marco Zuppelli, and Wojciech Mazurczyk. IPv6CC: IPv6 covert channels for testing networks against stegomalware and data exfiltration. SoftwareX, 17(??): ??, January 2022. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/

science/article/pii/
S2352711022000024.

Cording:2024:FRR

[CT24] Enzo Cording and Jagruti Thakur. FleetRL: Realistic reinforcement learning environments for commercial vehicle fleets. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024000426.

Candela-Uribe:2022:SBA

[CUSRCP^{+}22] Christian A. Candela-Uribe, Luis E. Sepúlveda-Rodríguez, Julio C. Chavarro-Porras, John A. Sanabria-Ordoñez, José Luis Garrido, Carlos Rodríguez-Domínguez, and Gabriel Guerrero-Contreras. SMS-Builder: an adaptive software tool for building systematic mapping studies. SoftwareX, 17(??):??, January 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021001710.

Casadei:2022:SSD

[CVAP22] Roberto Casadei, Mirko Viroli, Gianluca Aguzzi, and Danilo Pianini. ScaFi: a Scala DSL and toolkit for aggregate programming. SoftwareX, 20(??): ??, December 2022. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022001662.

Conoscenti:2021:CLN

[CVD21] Marco Conoscenti, Antonio Vetrò, and Juan Carlos De Martin. CLoTH: a lightning network simulator. SoftwareX, 15(??):??, July 2021. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021000613.

Creati:2019:FA

[CVS19] N. Creati, R. Vidmar, and P. Sterzai. Field animation. SoftwareX, 9(??): 211–216, January/June 2019. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S235271101830205X.

Caponi:2023:BPP

[CVV23] Francesco Caponi, David F. Vetsch, and Davide Vanzo. BASEveg: a Python package to model riparian vegetation dynamics coupled with river morphodynamics. SoftwareX, 22 (??):??, May 2023. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023000572.

Caron:2023:PPM

[CWC^{+}23] Eddy Caron, Jean-François Witz, Christophe Cuvier, Arnaud Beaurain, Vincent Magnier, and Ahmed El Bartali. PYCASO: Python module for calibration of cameras by Soloff’s method. SoftwareX, 23 (??):??, July 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271102300136X.

Chabib:2024:GGA

[CWGM24] Ahmed Chabib, Jean-François Witz, Pierre Gosselet, and Vincent Mag-

nier. GCPU OpticalFlow:

a GPU accelerated Python software for strain measurement. SoftwareX, 26 (??):??, May 2024. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024000591.

Couty:2021:GGA

[CWLG^{+}21] Victor Couty, Jean-François Witz, Pauline Lecomte-Grosbras, Julien Berthe, Eric Deletombe, and Mathias Brieu. GPUCorrel: a GPU accelerated Digital Image Correlation software written in Python. SoftwareX, 16(??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021001102.

Couty:2021:CCR

[CWM^{+}21a] Victor Couty, Jean-François Witz, Corentin Martel, François Bari, and Antoine Weisrock. CRAPPY: Command and Real-Time Acquisition in Parallelized Python, a Python module for experimental setups. SoftwareX, 16(??): ??, December 2021. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021001278.

Csati:2021:CFS

[CWM^{+}21b] Zoltan Csati, Jean-François Witz, Vincent Magnier, Ahmed El Bartali, Nathalie Limodin, and Denis Najjar. CristalX: Facilitating simulations for experimentally obtained grain-based microstructures. SoftwareX, 14(??):??, June 2021. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021000145.

Chen:2022:THP

[CZ22] Zhiwen Chen and Baorong

                                        Zhong. TFInterpy: a

high-performance spatial interpolation Python package. SoftwareX, 20(??):

??, December 2022. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022001479.

Croix:2024:FSF

[DA24] Ntivuguruzwa Jean De La Croix and Tohari Ahmad.

FuzConvSteganalysis: Ste-

ganalysis via fuzzy logic and convolutional neural network. SoftwareX, 26 (??):??, May 2024. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024000840.

Diblen:2019:SOS

[DAB^{+}19] Faruk Diblen, Jisk Attema, Rena Bakhshi, Sascha Caron, Luc Hendriks, and Bob Stienen. spot: Open source framework for scientific data repository and interactive visualization. SoftwareX, 9(??):328–331, January/ June 2019. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018302097.

DAquino:2024:PPI

[DAC^{+}24] Zachary D’Aquino, Sylwester Arabas, Jeffrey H. Curtis, Akshunna Vaishnav, Nicole Riemer, and Matthew West. PyPartMC: a Pythonic interface to a particle-resolved, Monte Carlo aerosol simulation framework. SoftwareX, 25(??):??, February 2024. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023003096.

Daniluk:2020:RCF

[Dan20] Andrzej Daniluk. rheed++: a C++ framework to simulation of RHEED intensity oscillations during the growth of thin epitaxial films. SoftwareX, 12 (??):Article 100593, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S235271102030306X.

Daniluk:2022:URC

[Dan22] Andrzej Daniluk. Update 2.0 to rheed++: a complex computer model for dynamical one-beam calculations of RHEED intensity oscillations. SoftwareX, 18(??):??, June 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022000371.

Resenes:2019:JJM

[dARPH^{+}19] Jonas de Abreu Resenes, Willingthon Pavan, Carlos Amaral Hölbig, José Maurício Cunha Fernandes, Vakhtang Shelia, Cheryl Porter, and Gerrit Hoogenboom. jDSSAT: a JavaScript module for DSSAT-CSM integration. SoftwareX, 10(??):Article 100271, July/December 2019. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271101930158X.

Davies:2021:SPB

[Dav21] Claire L. Davies. SEDBYS:

a Python-based SED Builder for Young Stars. SoftwareX, 14(??):??, June 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021000327.

Dharmavarapu:2019:GSD

[DBJ19] Raghu Dharmavarapu, Shanti Bhattacharya, and Saulius Juodkazis. GDOE-SII: Software for design of diffractive optical elements and phase mask conversion to GDSII lithography files. SoftwareX, 9(??): 126–131, January/June 2019. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018302711.

Drvoderic:2022:VCD

[DBPS22] Matthias Drvoderic, Jens J.

Bender, Martin Pletz, and Clara Schuecker. Version

0.2 — CrackDect: De-

tecting crack densities in images of fiber-reinforced polymers. SoftwareX, 19 (??):??, July 2022. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022001182.

deCastro:2023:VOI

[dCWL^{+}23] Marcelo de Castro, Dietmar Winkler, Giuseppe Laera, Luigi Vanfretti, Sergio A. Dorado-Rojas, Tin Rabuzin, Biswarup Mukherjee, and Manuel Navarro. Version [OpenIPSL 2.0.0] — [iTesla Power Systems Library (iPSL): a Modelica library for phasor time-domain simulations]. SoftwareX, 21 (??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022001959.

DeNardis:2023:XLS

[DDCSH23] Luca De Nardis, Maria-Gabriella Di Benedetto, Jeung-Yoon Choi, and Stefanie Shattuck-Hufnagel. xkl: a legacy software for detailed acoustic analysis of speech made modern. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/

science/article/pii/
S2352711023001887.

deDeusFilho:2022:PII

[ddSNX22] João Carlos Andrade de Deus Filho, Luiz Carlos da Silva Nunes, and José Manuel Cardoso Xavier.

iCorrVision-2D: an in-

tegrated Python-based open-source Digital Image Correlation software for in-plane measurements (Part 1). SoftwareX, 19 (??):??, July 2022. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022000851.

Palma:2020:APL

[DDT20] Stefano Dalla Palma, Dario Di Nucci, and Damian A. Tamburri.

AnsibleMetrics: a Python

library for measuring Infrastructure-as-Code blueprints in Ansible. SoftwareX, 12 (??):Article 100633, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020303460.

Degallaix:2020:OMB

[Deg20] Jérôme Degallaix. OSCAR:

a MATLAB based package to simulate realistic optical cavities. SoftwareX, 12 (??):Article 100587, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711020303009.

Duque:2023:TDS

[DER^{+}23] Julián D. Loaiza Duque, Antonio J. Sánchez Egea, Hernán A. González Rojas, Pedro Chaná-Cuevas, Joaquim J. Ferreira, and João Costa. TremorSoft: an decision support application for differential diagnosis between Parkinson’s disease and essential tremor. SoftwareX, 22 (??):??, May 2023. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023000894.

Deka:2023:LPE

[DET23] Pranab J. Deka, Lukas Einkemmer, and Mayya Tokman. LeXInt: Package for exponential integrators employing Leja interpolation. SoftwareX, 21(??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022002205.

Das:2021:RLC

[DEV^{+}21] Partha Pratim Das, Muthu Ram Prabhu Elenchezhian, Vamsee Vadlamudi, Kenneth Reifsnider, and Ras-

sel Raihan. RealPi2dDIC:

a low-cost and open-source approach to in situ 2D Digital Image Correlation (DIC) applications. SoftwareX, 13(??):Article 100645, January 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020303587.

deQueirozRamiroReis:2022:PDS

[dFB22] Ana Waldila de Queiroz Ramiro Reis, Maria Fernanda Figueiredo de Oliveira, and Rodrigo Bird Burgos. DYNAblast — a software to obtain the behavior of plates subjected to blast loads. SoftwareX, 20(??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022001510.

Demidov:2018:UPP

[DFC18] Alexander G. Demidov, Michael E. Fortunato, and Coray M. Colina. Update 0.2 to “pysimm: a Python package for simulation of molecular systems”. SoftwareX, 7 (??):70–73, January/June 2018. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018300141.

Rosa:2024:CCP

[DFS24] Pasquale De Rosa, Pascal Felber, and Valerio Schi-

                                        avoni. CryptoAnalytics:

Cryptocoins price forecasting with machine learning techniques. SoftwareX, 26 (??):??, May 2024. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711024000347.

Doulis:2019:CMP

[DFSW19] Georgios Doulis, Jörg Frauendiener, Chris Stevens, and Ben Whale. COFFEE — an MPI-parallelized Python package for the numerical evolution of differential equations. SoftwareX, 10(??):Article 100283, July/December 2019. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711019300950.

Deorowicz:2021:WIS

[DG21] Sebastian Deorowicz and Adam Gudyś. Whisper 2: Indel-sensitive short read mapping. SoftwareX, 14 (??):??, June 2021. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021000376.

Dyczkowski:2024:PLI

[DGK^{+}24] Krzysztof Dyczkowski, Piotr Grochowalski, Dawid Kosior, Dorota Gil, Wojciech Kozioł, Barbara Pekala, Uzay Kaymak, Caro Fuchs, and Marco S. Nobile. Python library for interval-valued fuzzy inference. SoftwareX, 26 (??):??, May 2024. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024001018.

del-Hoyo-Gabaldon:2024:ADG

[dHGMCGL^{+}24] Jesus-Angel del Hoyo-Gabaldon, Antonio Moreno-Cediel, Eva Garcia-Lopez, Antonio Garcia-Cabot, and David de Fitero-Dominguez. Automatic dataset generation for automated program repair of bugs and vulnerabilities through SonarQube. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024000359.

Dolz:2020:BFI

[DHK^{+}20] J. Dölz, H. Harbrecht, S. Kurz, M. Multerer, S. Schöps, and F. Wolf. Bembel: the fast isogeometric boundary element C++ library for Laplace, Helmholtz, and electric wave equation. SoftwareX, 11(??):Article 100476, January/June 2020. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711019301797.

Deza:2022:QGN

[DI22] Juan Ignacio Deza and Hisham Ihshaish. qNoise: a generator of non-Gaussian colored noise. SoftwareX, 18(??):??, June 2022. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022000346.

Dowling:2023:TCF

[DJCL23] Anthony Dowling, Lin Jiang, Ming-Cheng Cheng, and Yu Liu. TDF: a compact file format plugin for FEniCS. SoftwareX, 22 (??):??, May 2023. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023000250.

Don:2019:ACF

[DK19] Mihiran Galagedarage Don and Faisal Khan. Auxiliary codes for fault prognosis of Tennessee Eastman process using a hybrid model (CPL1.0). SoftwareX, 10(??):Article 100309, July/December 2019. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018301924.

Drago:2021:CWP

[DKL^{+}21] Marco Drago, Sergey Klimenko, Claudia Lazzaro, Edoardo Milotti, Guenakh Mitselmakher, Valentin Necula, Brendan O’Brian, Giovanni Andrea Prodi, Francesco Salemi, Marek Szczepanczyk, Shubhanshu Tiwari, Vaibhav Tiwari, Gayathri V, Gabriele Vedovato, and Igor Yakushin. coherent WaveBurst, a pipeline for unmodeled gravitational-wave data analysis. SoftwareX, 14 (??):??, June 2021. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021000236.

Degenne:2016:OSP

[DL16] P. Degenne and D. Lo Seen. Ocelet: Simulating processes of landscape changes using interaction graphs. SoftwareX, 5(??):84–88, ???? 2016. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711016300103.

Doucet:2018:WIR

[DLH18] M. Doucet, R. M. Ferraz Leal, and T. C. Hobson. Web interface for reflectivity fitting. SoftwareX, 7(??):287–293, January/ June 2018. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018300657.

delosSantos-Villalobos:2017:PCV

[dlSVBLdA^{+}17] Sergio de los Santos-Villalobos, Claudio Bravo-Linares, Roberto Meigikos dos Anjos, Renan Cardoso, Max Gibbs, Andrew Swales, Lionel Mabit, and Gerd Dercon. The CSSIAR v.1.00 software: a new tool based on SIAR to assess soil redistribution using Compound Specific Stable Isotopes. SoftwareX, 6(??):7–12, ???? 2017. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711016300425.

delaTorre:2024:GGE

[dlTAJCÁBG24] Juan Carlos de la Torre, José M. Aragón-Jurado, Abdón Crespo-Álvarez, and Guillermo Bárcena-González. GAGI: Game engine for Artificial General Intelligence experimentation. SoftwareX, 26 (??):??, May 2024. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/

science/article/pii/
S2352711024000360.

Dhakal:2019:MBM

[DM19] Suresh Dhakal and Mohamed A. Moustafa. MC-BAM: Moment-curvature analysis for beams with advanced materials. SoftwareX, 9(??):175–182, January/June 2019. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018301432.

Decan:2020:SPL

[DM20] Alexandre Decan and Tom Mens. Sismic —

- a Python library for statechart execution and testing. SoftwareX, 12 (??):Article 100590, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020303034.

DeLaHoz-M:2022:GRS

[DMFG22] Javier De La Hoz-M, Susana Mendes, and María José Fernandez-Gómez. GeoWeightedModel: an R-Shiny package for Geographically Weighted Models. SoftwareX, 20 (??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022001686.

Oliveira:2021:IIA

[dMOH21] Artur André Almeida de Macedo Oliveira and Roberto Hirata. INACITY — INvestigate and Analyze a CITY. SoftwareX, 15(??):??, July 2021. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021000911.

Vitturi:2024:EPO

[dMVBTN24] Mattia de’ Michieli Vitturi, Andrea Bevilacqua, Alessandro Tadini, and Augusto Neri. ELICIPY 1.0: a Python online tool for expert elicitation. SoftwareX, 25(??): ??, February 2024. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711024000128.

Dmitrieff:2017:CMC

[DN17] Serge Dmitrieff and François

                                        Nédélec. ConfocalGN: a

minimalistic confocal image generator. SoftwareX, 6(??):237–242, ???? 2017. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711017300444.

Duncombe:2022:SOI

[DPD22] Todd A. Duncombe, Aaron Ponti, and Petra S. Dit-

trich.  SpectraSorter:

Ocean Insight spectrometer software application for high-throughput fullspectrum UV-VIS analysis and triggered sorting. SoftwareX, 19(??):??, July 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271102200098X.

Demidov:2021:UPP

[DPF^{+}21] Alexander G. Demidov, B. Lakshitha A. Perera, Michael E. Fortunato, Sibo Lin, and Coray M. Colina. Update 1.1 to “pysimm: a Python package for simulation of molecular systems”, (PII: S2352711016300395). SoftwareX, 15(??):??, July 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021000777. See

[FC17].

dePuiseau:2023:SPF

[dPPD^{+}23] Constantin Waubert de Puiseau, Jannik Peters, Christian Dörpelkus, Hasan Tercan, and Tobias Meisen.

schlably:  a Python

framework for deep reinforcement learning based scheduling experiments. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S2352711023000791.

Reis:2020:DSN

[dQRRBdSM20] Ana Waldila de Queiroz Ramiro Reis, Rodrigo Bird Burgos, and Margareth da Silva Magalhães. Development of a software for the numerical modeling of reinforced concrete sections — AlfaMCV. SoftwareX, 11 (??):Article 100444, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711018302061.

Drachal:2023:DRP

[Dra23] Krzysztof Drachal. “dynmix”: an R package for the estimation of dynamic finite mixtures. SoftwareX, 22 (??):??, May 2023. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023000845.

daSilvaGuerra:2020:PPT

[dRB^{+}20] João Victor da Silva Guerra, Helder Veras Ribeiro Filho, Leandro Oliveira Bortot, Rodrigo Vargas Honorato, José Geraldo de Carvalho Pereira, and Paulo Sérgio Lopes de Oliveira. ParKVFinder: a thread-level parallel approach in biomolecular cavity detection. SoftwareX, 12(??):Article 100606, July/December 2020. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020303198.

Dorado-Rojas:2023:MMP

[DRFB^{+}23] Sergio A. Dorado-Rojas, Fernando Fachini, Tetiana Bogodorova, Giuseppe Laera, Marcelo de Castro Fernandes, and Luigi

Vanfretti. ModelicaGridData:

Massive power system simulation data generation and labeling tool using Modelica and Python. SoftwareX, 21(??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022001765.

Dahling:2020:OAG

[DRM20] Stefan Dähling, Lukas Razik, and Antonello Monti. OWL2Go: Autogeneration of Go data models for OWL ontologies with integrated serialization and deserialization functionality. SoftwareX, 12(??):Article 100571, July/December 2020. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020302843.

Drvoderic:2021:CDC

[DRPS21] Matthias Drvoderic, Matthias Rettl, Martin Pletz, and Clara Schuecker. CrackDect: Detecting crack densities in images of fiberreinforced polymers. SoftwareX, 16(??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021001205.

Dosta:2020:MOS

[DS20] Maksym Dosta and Vasyl Skorych. MUSEN: an open-source framework for GPU-accelerated DEM simulations. SoftwareX, 12 (??):Article 100618, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711020303319.

Daole:2023:OXF

[DSB^{+}23] Mattia Daole, Alessio Schiavo, José Luis Corcuera Bárcena, Pietro Ducange, Francesco Marcelloni, and Alessandro Renda. OpenFL-XAI: Federated learning of explainable artificial intelligence models in Python. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/

science/article/pii/
S2352711023002017.

dosSantos:2017:ACC

[dSBSD17] Jhonatha R. dos Santos, Luiz F. N. Barreta, Maria E. Sbampato, and Marcelo G. Destro. ASAS: Computational code for Analysis and Simulation of Atomic Spectra. SoftwareX, 6 (??):198–202, ???? 2017. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711017300298.

deSousa:2023:GRP

[dSBSvE23] Kauê de Sousa, David Brown, Jonathan Steinke, and Jacob van Etten. gosset: an R package for analysis and synthesis of ranking data in agricultural experimentation. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023000985.

Drozin:2020:KCS

[DSI^{+}20] Dmitry Drozin, Sergey Sozykin, Natalia Ivanova, Tatiana Olenchikova, Tatyana Krupnova, Natalia Krupina, and Viacheslav Avdin. Kinetic calculation: Software tool for determining the kinetic parameters of the thermal decomposition process using the Vyazovkin method. SoftwareX, 11(??):Article 100359, January/June 2020. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711019301931.

Desai:2022:AEU

[DSJ^{+}22] Deshana Desai, Etai Shuchatowitz, Zhongshi Jiang, Teseo Schneider, and Daniele Panozzo. ACORNS: an easy-to-use code generator for gradients and Hessians. SoftwareX, 17(??): ??, January 2022. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021001540.

deSouza:2022:CCI

[dSM22] Otávio Augusto Peter de Souza and Letícia Fleck Fadel Miguel. CIOA: Circle-Inspired Optimization Algorithm, an algorithm for engineering optimization. SoftwareX, 19 (??):??, July 2022. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022001157.

DeSensi:2017:MHL

[DTD17] Daniele De Sensi, Massimo Torquati, and Marco Danelutto. Mammut: Highlevel management of system knobs and sensors. SoftwareX, 6(??):148–149, ???? 2017. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711017300225.

Drummond:2019:SRP

[DTDd19] Blair R. Drummond, Christian J. G. Tessier, Mathieu F. Dextraze, and Corrie J. B. daCosta. scbursts: an R package for analysis and sorting of single-channel bursts. SoftwareX, 10(??):Article 100285, July/December 2019. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019300810.

Ducoin:2023:GTI

[Duc23]  Jean-Grégoire  Ducoin.

Galclaim: a tool to identify host galaxy of astrophysical transient sources. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023002789.

DAcci:2023:MIU

[DV23] Luca S. D’Acci and Michele Voto. Morphogenesis of Isobenefit urbanism: Isobenefit-cities simulator.

SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023001048.

DiMauro:2022:WWE

[DVSD22] Anna Di Mauro, Salvatore Venticinque, Giovanni Francesco Santonastaso, and Armando Di

                                        Nardo.  WEUSEDTO —

Water End USE Dataset and TOols: an open water end use consumption dataset and data analytics tools. SoftwareX, 20 (??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022001327.

Das:2020:RID

[DW20] Subasish Das and L. D. White. RuralSpeedSafetyX: Interactive decision support tool to improve safety. SoftwareX, 11 (??):Article 100493, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711019303553.

Deniziak:2021:DLC

[DW21] Stanisław Deniziak and Mariusz Wiśniewski. DECOLib: a library of components for DECOmposition of discrete functions. SoftwareX, 16(??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021001023.

Dubey:2022:FXM

[DWO^{+}22] Anshu Dubey, Klaus Weide, Jared O’Neal, Akash Dhruv, Sean Couch, J. Austin Harris, Tom Klosterman, Rajeev Jain, Johann Rudi, Bronson Messer, Michael Pajkos, Jared Carlson, Ran Chu, Mohamed Wahib, Saurabh Chawdhary, Paul M. Ricker, Dongwook Lee, Katie Antypas, Katherine M. Riley, Christopher Daley, Murali Ganapathy, Francis X. Timmes, Dean M. Townsley, Marcos Vanella, John Bachan, Paul M. Rich, Shravan Kumar, Eirik Endeve, W. Raphael Hix, Anthony Mezzacappa, and Thomas Papatheodore. Flash-X: a multiphysics simulation software instrument. SoftwareX, 19(??):??, July 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022001030.

DeChavez:2021:OPM

[DyH21] Danjo De Chavez and Jun

ya Hasegawa. OpenMechanochem:

a Python module for mechanochemical simulations. SoftwareX, 16(??): ??, December 2021. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021001436.

Dhaene:2023:HWB

[DZH^{+}23] Zeno Dhaene, Nina Zizakić, Shaoguang Huang, Xian Li, and Aleksandra Pizurica. HSIToolbox: a web-based application for the classification of hyperspectral images. SoftwareX, 22 (??):??, May 2023. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023000365.

Dong:2022:DNT

[DZZ^{+}22] Li Dong, Yufan Zhang, Lingling Zhao, Ting Zheng, Weidong Wang, Jianfu Li, Diankun Gong, Tiejun Liu, and Dezhong Yao. DRT: a new toolbox for the Standard EEG Data Structure in large-scale EEG applications. SoftwareX, 17(??):??, January 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021001709.

Elhassan:2022:DME

[EAFJ^{+}22] Ammar Elhassan, Mohammad Al-Fawa’reh, Mousa Tayseer Jafar, Mohammad Ababneh, and Shifaa Tayseer Jafar. DFT-MF: Enhanced deepfake detection using mouth movement and transfer learning. SoftwareX, 19(??):??, July 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022000759.

Escudero-Arnanz:2023:DPP

[EAMSR^{+}23] Óscar Escudero-Arnanz, Antonio G. Marques, Cristina Soguero-Ruiz, Inmaculada Mora-Jiménez, and Gregorio Robles.

dtwParallel: a Python

package to efficiently compute dynamic time warping between time series. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023000602.

Erdmann:2024:KTR

[EB24] Stella Erdmann and Rebecca Betensky. KMstability: R tools to report the stability and precision of Kaplan–Meier estimates as well as measures of follow-up in time-to-event studies. SoftwareX, 26 (??):??, May 2024. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S2352711024000219.

ElOulaFrihi:2021:TSM

[EBGCT21] Zahrate El Oula Frihi, Julian Barreiro-Gomez, Salah Eddine Choutri, and Hamidou Tembine. Toolbox to simulate and mitigate COVID-19 propagation. SoftwareX, 14 (??):??, June 2021. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021000182.

Escribano:2024:PPP

[EBL24] Nicolas Escribano, José Manuel Bielsa, and Francisco

                                        Lahuerta. pymetamodels:

a Python package for metamodeling and design automation. SoftwareX, 26 (??):??, May 2024. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711024001067.

Ellul:2022:TCE

[ED22] Joshua Ellul and Carl James

                                        Debono. TinkercadNetConnector:

Connecting emulated IoT devices to the outside world. SoftwareX, 20 (??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022001364.

Esfahani:2023:DOS

[EDZ^{+}23] Mahdad Jafarzadeh Esfahani, Amir Hossein Daraie, Paul Zerr, Frederik D. Weber, and Martin Dresler.

Dreamento: an open-

source dream engineering toolbox for sleep EEG wearables. SoftwareX, 24 (??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023002911.

Evans:2023:HPA

[EFTS23] George S. Evans, Thomas F.

Fuerst, Chase N. Taylor, and Masashi Shimada. Hydrogen Permeation and Absorption Tool (HyPAT v2.0): an update to facilitate analyses of data from hydrogen absorption experiments. SoftwareX, 23 (??):??, July 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023001267.

El-Habr:2019:RPG

[EHGPT19] Camille El-Habr, Xavier Garcia, Pujana Paliyawan, and Ruck Thawonmas. Runner: a 2D platform game for physical health promotion. SoftwareX, 10 (??):Article 100329, July/ December 2019. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S235271101930024X.

El-khoury:2016:LCG

[Ek16] Jad El-khoury. Lyo code generator: a model-based code generator for the development of OSLCcompliant tool interfaces. SoftwareX, 5(??):183–189, ???? 2016. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711016300267.

Egert:2023:RRI

[EK23] Janine Egert and Clemens Kreutz. Rcall: an R interface for MAT-LAB. SoftwareX, 21(??): ??, February 2023. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022001947.

Emmanouil:2017:UAH

[EL17] Stergios Emmanouil and Andreas Langousis. UPStream: Automated hydraulic design of pressurized water distribution networks. SoftwareX, 6(??):243–247, ???? 2017. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711017300432.

Elkady:2020:ESE

[EL20] Ahmed Elkady and Dimitrios G. Lignos. EaRL — software for earthquake risk, loss and lifecycle analysis. SoftwareX, 12 (??):Article 100607, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020303204.

Elkady:2022:FOS

[Elk22] Ahmed Elkady. FM-2D — open-source platform for the 2-dimensional numerical modeling and seismic analysis of buildings. SoftwareX, 17(??):??, January 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021001679.

Elsner:2023:TPP

[Els23] Jean Elsner. Taming the Panda with Python: a powerful duo for seamless robotics programming and integration. SoftwareX, 24 (??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023002285.

Eschle:2020:ZSP

[ENCS20] Jonas Eschle, Albert Puig Navarro, Rafael Silva Coutinho, and Nicola Serra. zfit: Scalable Pythonic fitting. SoftwareX, 11(??):Article 100508, January/June 2020. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019303851.

Ertl:2022:POS

[EPF^{+}22] Matthias Ertl, Carlo Prelz, Daniel C. Fitze, Gerda Wyssen, and Fred W.

Mast. PlatformCommander

— an open source software for an easy integration of motion platforms in research laboratories. SoftwareX, 17(??):??, January 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271102100176X.

Elia:2023:PPL

[EPF^{+}23] Donatello Elia, Cosimo Palazzo, Sandro Fiore, Alessandro D’Anca, Andrea Mariello, and Giovanni Aloisio. PyOphidia: a Python library for high performance data analytics at scale. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023002340.

Englert:2023:AAP

[ESE^{+}23] Robert Englert, Manfred Schedlowski, Harald Engler, Winfried Rief, Christian Büchel, Ulrike Bingel, and Tamas Spisak. ALIIAS: Anonymization/ Pseudonymization with LimeSurvey integration and II-factor Authentication for Scientific research. SoftwareX, 24 (??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023002182.

Evans:2023:HGH

[EWF^{+}23] George S. Evans, Joseph M.

Watkins, Thomas F. Fuerst, Chase N. Taylor, and Masashi Shimada. HyPAT: a GUI for highthroughput gas-driven hydrogen permeation data analysis. SoftwareX, 21 (??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022002023.

Evans:2023:CAA

[EWT^{+}23] George S. Evans, Joseph M.

Watkins, Chase N. Taylor, Jagoda Urban-Klaehn, and Chuting T. Tsai. CDB-AP: an application for coincidence Doppler broadening spectroscopy analysis. SoftwareX, 23 (??):??, July 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023001711. See

corrigendum [EWT^{+}24].

Evans:2024:CCA

[EWT^{+}24] George S. Evans, Joseph M. Watkins, Chase N. Taylor, Jagoda Urban-Klaehn, and Chuting T. Tsai. Corrigendum to “CDB-AP: an application for coincidence Doppler broadening spectroscopy analysis” [SoftwareX 23 (2023) 101475]. SoftwareX, 25 (??):??, February 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023003059. See

[EWT^{+}23].

Fox:2020:PPR

[FAM^{+}20] Nathan Fox, Tom August, Francesca Mancini, Katherine E. Parks, Felix Eigenbrod, James M. Bullock, Louis Sutter, and Laura J. Graham.

“photosearcher” package

in R: an accessible and reproducible method for harvesting large datasets from Flickr. SoftwareX, 12 (??):Article 100624, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271102030337X.

Fernandez-Alvarez:2022:TTW

[FAPANG22] José C. Fernández-Alvarez, Albenis Pérez-Alarcón, Raquel Nieto, and Luis Gimeno. TROVA: TRansport Of water VApor. SoftwareX, 20(??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022001467.

Folgado:2022:TTS

[FBA^{+}22] Duarte Folgado, Marília Barandas, Margarida Antunes, Maria Lua Nunes, Hui Liu, Yale Hartmann, Tanja Schultz, and Hugo Gamboa. TSSEARCH: Time Series Subsequence Search Library. SoftwareX, 18 (??):??, June 2022. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022000425.

Fortunato:2017:PPP

[FC17] Michael E. Fortunato and Coray M. Colina. pysimm: a Python package for simulation of molecular systems. SoftwareX, 6 (??):1–6, ???? 2017. CODEN ???? ISSN 2352-7110. URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S2352711016300395.

See updates [DFC18, DPF^{+}21].

Fernandez-Candel:2024:ULU

[FCMTV24] Carlos Javier Fernández-Candel, Paula Muñoz, Javier Troya, and Antonio Vallecillo. UTypes: a library for uncertain datatypes in Python. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711024000475.

Fachini:2023:OOI

[FdCBV23] Fernando Fachini, Marcelo de Castro, Tetiana Bogodorova, and Luigi Vanfretti. OpenIMDML: Open Instance Multi-Domain Motor Library utilizing the Modelica modeling language. SoftwareX, 24 (??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S235271102300287X.

Fraszczak:2024:NCP

[FF24] Damian Fraszczak and Edyta Fraszczak. NetCenLib: a comprehensive Python library for network centrality analysis and evaluation. SoftwareX, 26 (??):??, May 2024. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024000700.

Fagundes:2020:CSM

[FG20] Daniel Leal Fagundes and Jair T. Goulart. ContHeart: Software for monitoring isolated cardiomyocyte shortening. SoftwareX, 12(??):Article 100547, July/December 2020. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019303826.

Fan:2023:PVN

[FH23] Hangchao Fan and Lina Hu. Pressure vessel nozzle local stress prediction software based on ABAQUS-machine learning. SoftwareX, 24(??): ??, December 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023002467.

Funk:2017:HNV

[FHA17] Sean P. Funk, Danny Hnatyshin, and Daniel S. Alessi. HYDROSCAPE: a new versatile software program for evaluating contaminant transport in groundwater. SoftwareX, 6(??): 255–260, ???? 2017. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S235271101730050X.

Fisher:2021:DTI

[FHB^{+}21] Ryan P. Fisher, Gary Hemming, Marie-Anne Bizouard, Duncan A. Brown, Peter F. Couvares, Florent Robinet, and Didier Verkindt. DQSEGDB: a timeinterval database for storing gravitational wave observatory metadata. SoftwareX, 14(??):??, June 2021. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021000224.

Farmer:2022:MTP

[FJ22] Jenny Farmer and Donald J. Jacobs. MATLAB tool for probability density assessment and nonparametric estimation. SoftwareX, 18(??):??, June 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022000231.

Floch:2023:MTB

[FK23] Kristóf Floch and Attila Kossa. Motion Tracker Beta: a GUI based opensource motion tracking application. SoftwareX, 23 (??):??, July 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023001206.

Fraccaroli:2022:SDT

[FLR22] Michele Fraccaroli, Evelina Lamma, and Fabrizio Riguzzi. Symbolic DNN-Tuner: a Python and ProbLog-based system for optimizing Deep Neural Networks hyperparameters. SoftwareX, 17(??): ??, January 2022. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021001825.

Fitzpatrick:2024:DPP

[FM24a] Paxton C. Fitzpatrick and Jeremy R. Manning. Davos: a Python package “smuggler” for constructing lightweight reproducible notebooks. SoftwareX, 25(??):??, February 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023003102.

Freeborn:2024:UDR

[FM24b] Todd J. Freeborn and Ja-

cob A. Mota. Ultrasound
DICOM Renamer: a MAT-

LAB graphical user interface for workflow improvement for DICOM ultrasound renaming. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711024001146.

Fernandez-Montes:2023:STE

[FMFCEG^{+}23] Alejandro Fernández-Montes, Damián Fernández-Cerero, Felipe Escalera-González, Agnieszka Jakóbik, Belén Bermejo, and Carlos Juiz.

                                        SimilarityTS: Toolkit for

the evaluation of similarity for multivariate time series. SoftwareX, 24 (??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023002236.

Filho:2022:IIPa

[FNX22] João Filho, Luiz Nunes, and José Xavier. iCorrVision-3D: an integrated Pythonbased open-source Digital Image Correlation Software for in-plane and outof-plane measurements (Part 2). SoftwareX, 19 (??):??, July 2022. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S235271102200084X.

Ferguson:2018:PPM

[FPBM18] Joseph C. Ferguson, Francesco Panerai, Arnaud Borner, and Nagi N. Mansour.

PuMA: the Porous Microstructure Analysis software. SoftwareX, 7(??):81– 87, January/June 2018. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018300281. See

update [FST^{+}21].

Flint:2022:BNS

[FPI^{+}22] Thomas F. Flint, Gowthaman Parivendhan, Alojz Ivankovic, Michael C. Smith, and Philip Cardiff. beamWeldFoam: Numerical simulation of high energy density fusion and vapourisationinducing processes. SoftwareX, 18(??):??, June 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022000504.

Forouli:2022:APP

[FPN^{+}22] Aikaterini Forouli, Anastasios Pagonis, Alexandros Nikas, Konstantinos Koasidis, Georgios Xexakis, Themistoklis Koutsellis, Christos Petkidis, and Haris Doukas. AUGMECON-Py: a Python framework for multi-objective linear optimisation under uncertainty. SoftwareX, 20 (??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S2352711022001388.

Fraszczak:2022:NNE

[Fra22a] Damian Fraszczak. NEFBDAA — .NET Environment for Building Dynamic Angular Applications. SoftwareX, 19(??):??, July 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022001017.

Fraszczak:2022:RRP

[Fra22b] Damian Fraszczak. RPaSDT — Rumor Propagation and Source Detection Toolkit. SoftwareX, 17 (??):??, January 2022. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022000085.

Fiorio:2021:IPP

[FRdN21] Luan Vinícius Fiorio, Chrystian Lenon Remes, and Yales Rômulo de No-

                                        vaes. impulseest: a

Python package for nonparametric impulse response estimation with input–output data. SoftwareX, 15(??):??, July 2021. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021000832.

Flint:2024:VLL

[FRE^{+}24] Thomas F. Flint, Joseph D.

Robson, Parsa Esmati, Nicolò Grilli, Gowthaman Parivendhan, and Philip Cardiff. Version 2.0 —

LaserbeamFoam:  Laser

ray-tracing and thermally induced state transition simulation toolkit. SoftwareX, 25(??):??, February 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023003084.

Frost:2024:CMI

[FRNW^{+}24] Emilie Frost, Malin Radtke, Marvin Nebel-Wenner, Frauke Oest, and Sanja

Stark.  cosima-mango:

Investigating Multi-Agent System robustness through integrated communication simulation. SoftwareX, 26 (??):??, May 2024. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024000384.

Flint:2023:LLR

[FRPC23] Thomas F. Flint, Joseph D.

Robson, Gowthaman Parivendhan, and Philip Cardiff.

laserbeamFoam:  Laser

ray-tracing and thermally induced state transition simulation toolkit. SoftwareX, 21(??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022002175.

Fruhstorfer:2021:PTD

[Fru21] Jens Fruhstorfer. ParSD — tool to design and analyze particle size distributions. SoftwareX, 15(??):??, July 2021. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021000790.

Flint:2019:HHE

[FS19] T. F. Flint and M. C. Smith. HEDSATS: High energy density semi-analytical thermal solutions. SoftwareX, 10(??):Article 100243, July/December 2019. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711019300688.

Filippidis:2023:MAA

[FS23] Achilleas Filippidis and Adam J. Sadowski. Modern analysis of axisymmetric shells with AQUINAS: a MATLAB finite element toolbox. SoftwareX, 23 (??):??, July 2023. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023001309.

Farook:2023:DLF

[FSAD23a] Taseef Hasan Farook, Farhan Hasin Saad, Saif Ahmed, and James Dud-

ley. Dental Loop FLT:

Facial landmark tracking. SoftwareX, 24(??): ??, December 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023002273.

Farook:2023:DLS

[FSAD23b] Taseef Hasan Farook, Farhan Hasin Saad, Saif Ahmed, and James Dud-

ley. Dental loop SnP:

Speech and phonetic pattern recognition. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271102300300X.

Fragkoulis:2016:PQS

[FSL16] Marios Fragkoulis, Diomidis Spinellis, and Panos Louridas. PiCO QL: a software library for runtime interactive queries on program data. SoftwareX, 5 (??):127–133, ???? 2016. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711016300176.

Ferreira-Saraiva:2023:PPP

[FSMCFP23] Bruno D. Ferreira-Saraiva, João P. Matos-Carvalho, Nuno Fachada, and Manuel Pita. ParShift: a Python package to study order and differentiation in group conversations. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023002509.

Ferguson:2021:UPP

[FST^{+}21] Joseph C. Ferguson, Federico Semeraro, John M. Thornton, Francesco Panerai, Arnaud Borner, and Nagi N. Mansour. Update 3.0 to “PuMA: the Porous Microstructure Analysis software”, (PII:S2352711018300281). SoftwareX, 15(??):??, July 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271102100090X. See

[FPBM18].

Ferenc:2020:DWF

[FVA^{+}20] Rudolf Ferenc, Tamás Viszkok, Tamás Aladics, Judit Jász, and Péter Hegedűs. Deep-water framework: the Swiss army knife of humans working with machine learning models. SoftwareX, 12(??):Article 100551, July/December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019303772.

Futia:2020:SSM

[FVD20] Giuseppe Futia, Antonio Vetrò, and Juan Carlos De Martin. SeMi: a semantic modeling machine to build knowledge graphs with graph neural networks. SoftwareX, 12(??):Article 100516, July/December 2020. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019302626.

Fox:2023:VRP

[FVVL23] Nathan Fox, Derek Van Berkel, Ramiro Serrano Verge, and Mark Lindquist.

vGameReviews: an R pack-

age for harnessing video game reviews for scientific research. SoftwareX, 23 (??):??, July 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271102300119X.

Fajardo:2021:ALW

[FWB^{+}21] Edgar Fajardo, Frank Wuerthwein, Brian Bockelman, Miron Livny, Greg Thain, James Alexander Clark, Peter Couvares, and Josh Willis. Adapting LIGO workflows to run in the Open Science Grid. SoftwareX, 14(??):??, June 2021. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021000248.

Feng:2023:LTL

[FWJ^{+}23] Xiaoyue Feng, Dashan Wei, Tianzhe Jiao, Chaopeng Guo, Dongqi Wang, and Jie Song. LiPTool: a tool for learning-based autonomous index placement in databases. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S235271102300239X.

Garouani:2022:AAE

[GABH22] Moncef Garouani, Adeel Ahmad, Mourad Bouneffa, and Mohamed Hamlich. AMLBID: an autoexplained Automated Machine Learning tool for Big Industrial Data. SoftwareX, 17(??):??, January 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021001631.

Garcia:2023:GSI

[GACS^{+}23] Iván García, Javier A. Albusac, José J. Castro-Schez, Vanesa Herrera, and David Vallejo. GVIDI: a system to improve the safety of groups of hikers based on real-time monitoring and dataset generation. SoftwareX, 24 (??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023002820.

Guillen:2023:BBA

[GAE23] Maria D. Guillen, Juan Aparicio, and Victor J.

España.  boostingDEA:

a boosting approach to Data Envelopment Analysis in R. SoftwareX, 24 (??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023002455.

Gagolewski:2021:GFR

[Gag21] Marek Gagolewski. genieclust:

Fast and robust hierarchical clustering. SoftwareX, 15(??):??, July 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021000649.

Gagolewski:2022:FBC

[Gag22] Marek Gagolewski. A framework for benchmarking clustering algorithms. SoftwareX, 20(??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022001881.

Genovese:2024:SSW

[GAM^{+}24] Maurilio Genovese, Agostino Arcasensa, Silvia Morbelli, Matteo Lenge, Carmen Barba, Laura Mirandola, Maria Eugenia Caligiuri, Massimo Caulo, Ferruccio Panzica, Francesco Cardinale, Camilla Rossi-Espagnet, Domenico Tortora, and Anna Elisabetta Vaudano. SWANe: Standardized workflow for advanced neuroimaging in epilepsy. SoftwareX, 26 (??):??, May 2024. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711024000748.

Gaza:2024:COS

[GB24] Haifa Gaza and Jaewook

                                        Byun. Chronoweb: an

open-source platform for analyzing temporal information diffusion on the web. SoftwareX, 26(??): ??, May 2024. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711024001092.

Gundlach:2023:PMS

[GBJ23] Jannek Gundlach, Maximilian Behnke, and Christian Jordan. PROVER-M: a simple model to project the disposal of fine sediments. SoftwareX, 23 (??):??, July 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023001036.

Godwin:2024:IDG

[GBW^{+}24] Ryan C. Godwin, Ayesha S.

Bryant, Brant M. Wagener, Timothy J. Ness, Jennifer J. DeBerry, LaShun L. Horn, Shanna H. Graves, Ashley C. Archer, and

Ryan L. Melvin. IRB-draft-generator:

a generative AI tool to streamline the creation of institutional review board applications. SoftwareX, 25(??):??, February 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023002972.

Gonzalez-Cely:2023:RPB

[GCBDDBF23] Aura Ximena Gonzalez-Cely, Cristian Felipe Blanco-Diaz, Camilo A. R. Diaz, and Teodiano Freire Bastos-Filho. Roborueda: Python-based GUI to control a wheelchair and monitor user posture. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S2352711023002510.

Gavidia-Calderon:2020:IJF

[GCC20] Carlos Gavidia-Calderon and César Beltrán Castañon. Isula: a Java framework for ant colony algorithms. SoftwareX, 11 (??):Article 100400, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711019300639.

Garcia-Cadena:2021:PLT

[GCdJAURO21] Carlos A. García-Cadena, Abraham de J. Aguilar-Uribe, and Luis F. Rojas-

                                        Ochoa. PhotonSTR-18:

a LabVIEW toolbox for photon correlation spectroscopy. SoftwareX, 13 (??):Article 100640, January 2021. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711020303538.

Gusain:2022:SME

[GCP22] Digvijay Gusain, Miloš Cvetković, and Peter Palensky. Simplifying multi-energy system cosimulations using energysim. SoftwareX, 18(??):??, June 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/

science/article/pii/
S2352711022000255.

Gong:2023:MMF

[GCW^{+}23] Qian Gong, Jieyang Chen, Ben Whitney, Xin Liang, Viktor Reshniak, Tania Banerjee, Jaemoon Lee, Anand Rangarajan, Lipeng Wan, Nicolas Vidal, Qing Liu, Ana Gainaru, Norbert Podhorszki, Richard Archibald, Sanjay Ranka, and Scott Klasky. MGARD: a multigrid framework for high-performance, errorcontrolled data compression and refactoring. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023002868.

Garcia:2016:OOO

[GdCF16] Álvaro López García, Enol Fernández del Castillo, and Pablo Orviz Fernández. ooi: OpenStack OCCI interface. SoftwareX, 5(??): 1–5, ???? 2016. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711016000029.

Galuzio:2020:MMO

[GdVSdSCM20] Paulo Paneque Galuzio, Emerson Hochsteiner de Vasconcelos Segundo, Leandro dos Santos Coelho, and Viviana Cocco Mariani. MOBOpt — multiobjective Bayesian optimization. SoftwareX, 12 (??):Article 100520, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711020300911.

Garate:2022:AMF

[GEG22] Gorka Garate, Julian Estevez, and Manuel Graña. ADITU: a mesh-free formulation for the solution of Helmholtz equation in bounded and unbounded domains. SoftwareX, 19 (??):??, July 2022. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022000735.

Graichen:2019:SPT

[GEH19] Uwe Graichen, Roland Eichardt, and Jens Haueisen.

                                        SpharaPy:  a Python

toolbox for spatial harmonic analysis of nonuniformly sampled data. SoftwareX, 10(??):Article 100289, July/December 2019. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711019301670.

Georgeon:2017:LAP

[Geo17] Olivier L. Georgeon. Little AI: Playing a constructivist robot. SoftwareX, 6 (??):155–160, ???? 2017. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711017300213.

Gamallo-Fernandez:2024:VPL

[GFRMVL24] Pedro Gamallo-Fernandez, Efrén Rama-Maneiro, Juan C. Vidal, and Manuel Lama. VERONA: a Python library for benchmarking deep learning in business process monitoring. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024001055.

Gunaratne:2021:NAB

[GG21] Chathika Gunaratne and Ivan Garibay. NL4Py: Agent-based modeling in Python with parallelizable NetLogo workspaces. SoftwareX, 16(??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021001035.

Gacitua:2023:LTS

[GGI23] Daniel Gacitúa and Roberto González-Ibáñez. LETICIA: a toolbox for search task deployment, execution, and user interaction logging on interactive information retrieval studies. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023001899.

Giroux:2021:TPP

[Gir21] Bernard Giroux. ttcrpy: a Python package for traveltime computation and raytracing. SoftwareX, 16 (??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021001217.

Giulioni:2019:GGA

[Giu19] Gianfranco Giulioni. GABRIELE: the General Agent Based Repast Implemented Extensible Laboratory for Economies. SoftwareX, 9(??):255–259, January/ June 2019. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711018300761.

Gutierrez:2022:GEA

[GJS^{+}22] Juan D. Gutiérrez, Antonio R. Jiménez, Fernando Seco, Fernando J. Álvarez, Teodoro Aguilera, Joaquín Torres-Sospedra, and Fran

                                        Melchor. GetSensorData:

an extensible Androidbased application for multi-sensor data registration. SoftwareX, 19 (??):??, July 2022. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022001121.

Gharari:2023:EPP

[GKK^{+}23] Shervan Gharari, Kasra Keshavarz, Wouter J. M. Knoben, Gouqiang Tang, and Martyn P. Clark. EASYMORE: a Python package to streamline the remapping of variables for Earth System models. SoftwareX, 24(??): ??, December 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023002431.

Gartner:2020:EWL

[GKM20] Jan Wilhelm Gärtner, Andreas Kronenburg, and Tobias Martin. Efficient WENO library for Open-FOAM. SoftwareX, 12 (??):Article 100611, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020303241.

Guerin:2015:MPE

[GL15] Greg R. Guerin and Andrew J. Lowe. Mapping phylogenetic endemism in R using georeferenced branch extents. SoftwareX, 3–4(??):13–21, December 2015. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711015000126.

Garcia-Lopez:2023:JTS

[GLCBHTHO23] Adrian García-López, Oscar Chávez-Bosquez, José Hernández-Torruco, and Betania Hernández-Ocaña. JMetaBFOP: a tool for solving global optimization problems. SoftwareX, 23 (??):??, July 2023. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023001486.

Garcia:2022:RIS

[GLGG22] Roberto García, Juan-Miguel López-Gil, and Rosa Gil. Rhizomer: Interactive semantic knowledge graphs exploration. SoftwareX, 20(??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022001534.

Gershanik:2022:PPS

[GLR22] Tzvi Gershanik, Itay Levin, and Daniel Rittel. 2BarG — a program to process split Hopkinson (Kolsky) bar test results. SoftwareX, 18(??):??, June 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022000644.

G:2022:GMG

[GLS^{+}22] Renjith G., Sonia Laudanna, Aji S., Corrado Aaron Visaggio, and Vinod P. GANG-MAM: GAN based enGine for Modifying Android Malware. SoftwareX, 18(??):??, June 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022000036.

Godwin:2024:TED

[GM24] Ryan C. Godwin and Ryan L. Melvin. Toward efficient data science: a comprehensive MLOps template for collaborative code development and automation. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024000943.

Gerber:2018:DRP

[GMF18] F. Gerber, K. Mösinger, and R. Furrer. dotCall64: an R package providing an efficient interface to compiled C, C++, and Fortran code supporting long vectors. SoftwareX, 7(??): 217–221, January/June 2018. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711018300785.

Gerst:2020:MSI

[GMF20] Ruman Gerst, Anna Medyukhina, and Marc Thilo Figge. MISA++: a standardized interface for automated bioimage analysis. SoftwareX, 11 (??):Article 100405, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711019302808.

Garnier:2019:ISR

[GMGG19] Guillaume F. G. Garnier, Clare A. Manderson, Saveen Giri, and Gil Garnier. Identi-Cyte: Simple red blood cell identification software. SoftwareX, 9(??):223–229, January/June 2019. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711018301377.

Gaete-Morales:2021:DPF

[GMKRS21] Carlos Gaete-Morales, Martin Kittel, Alexander Roth, and Wolf-Peter Schill. DIETERpy: a Python framework for the Dispatch and Investment Evaluation Tool with Endogenous Renewables. SoftwareX, 15(??):??, July 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021000947.

Gutierrez:2018:BAO

[GMNG^{+}18] Ronald R. Gutierrez, Jose A. Mallma, Francisco Núñez-González, Oscar Link, and Jorge D. Abad. Bedforms-ATM, an open source software to analyze the scale-based hierarchies and dimensionality of natural bed forms. SoftwareX, 7(??):184–189, January/June 2018. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018300451.

Goona:2023:DOS

[GNPP23] Nithin Kumar Goona, Shraddha M. Naik, Saidi Reddy Parne, and Anand Paul. DssPyLib: an open-source Python FEM software to solve Poisson equation in 2-D using distributed source scheme. SoftwareX, 21(??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023000043.

Gloster:2019:CCF

[GÓ19] Andrew Gloster and Lennon Ó Náraigh. cuSten — CUDA finite difference and stencil library. SoftwareX, 10(??):Article 100337, July/December 2019. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711019300561.

Goke:2021:AJJ

[Gök21] Leonard Göke. AnyMOD.jl: a Julia package for creating energy system models. SoftwareX, 16(??): ??, December 2021. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021001382.

Gorski:2024:SJP

[Gór24] Tomasz Górski. SmarTS: a Java package for smart contract test suite generation and execution. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711024000694.

Giuffre:2022:NIP

[GP22] Andrea Giuffre’ and Matteo Pini. NiceProp: an interactive Python-based educational tool for nonideal compressible fluid dynamics. SoftwareX, 17(??):??, January 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021001527.

Gomes:2023:PPT

[GPEM23] Eduardo Gomes, Lucas Pereira, Augusto Esteves, and Hugo Morais. PyECOM: a Python tool for analyzing and simulating Energy Communities. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023002765.

Ghimire:2021:CCR

[GPK^{+}21] Bhoj Raj Ghimire, Rishi Ram Parajuli, Bipin Khatiwada, Shobha Poudel, Kusum Sharma, and Bhogendra Mishra. Covira: a COVID-19 risk assessment, visualization and communication tool. SoftwareX, 16(??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021001400.

Godoy:2020:AAI

[GPW^{+}20] William F. Godoy, Norbert Podhorszki, Ruonan Wang, Chuck Atkins, Greg Eisenhauer, Junmin Gu, Philip Davis, Jong Choi, Kai Germaschewski, Kevin Huck, Axel Huebl, Mark Kim, James Kress, Tahsin Kurc, Qing Liu, Jeremy Logan, Kshitij Mehta, George Ostrouchov, Manish Parashar, Franz Poeschel, David Pugmire, Eric Suchyta, Keichi Takahashi, Nick Thompson, Seiji Tsutsumi, Lipeng Wan, Matthew Wolf, Kesheng Wu, and Scott Klasky. ADIOS 2: the Adaptable Input Output System. A framework for high-performance data management. SoftwareX, 12(??):Article 100561, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019302560.

Gladstein:2018:SPF

[GQCP^{+}18] Ariella L. Gladstein, Consuelo D. Quinto-Cortés, Julian L. Pistorius, David Christy, Logan Gantner, and Blake L. Joyce. Sim-Prily: a Python framework to simplify highthroughput genomic simulations. SoftwareX, 7(??): 335–340, January/June 2018. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018301213.

Ganguly:2023:QNS

[GRD^{+}23] Siddhartha Ganguly, Nakul Randad, Rihan Aaron D’Silva, Mukesh Raj S., and Debasish Chatterjee. QuITO: Numerical software for constrained nonlinear optimal control problems. SoftwareX, 24 (??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023002534.

Gonzalez-Ruiz:2023:FSP

[GRF23]  Vicente  González-Ruiz

and Jose-Jesus Fernández.

FlowDenoising: Structure-

preserving denoising in 3D electron microscopy (3DEM). SoftwareX, 23 (??):??, July 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023001097.

Gerum:2019:CPP

[GRW^{+}19] Richard C. Gerum, Sebastian Richter, Alexander Winterl, Christoph Mark, Ben Fabry, Céline Le Bohec, and Daniel P. Zitter-

bart. CameraTransform:

a Python package for perspective corrections and image mapping. SoftwareX, 10(??):Article 100333, July/December 2019. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711019302018.

Gregory:2021:SSW

[GSF21] Elizabeth D. Gregory, William C. Schneck, and Erik L. Frankforter. swSim: Solid wave simulation. SoftwareX, 14(??):??, June 2021. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021000431.

Garcia:2017:FSM

[GSP^{+}17] Dorleta Garcia, Sonia Sánchez, Raúl Prellezo, Agurtzane Urtizberea, and Marga Andrés. FLBEIA: a simulation model to conduct bio-economic evaluation of fisheries management strategies. SoftwareX, 6(??):135–140, ???? 2017. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711017300171.

Goldsteen:2023:APT

[GSS^{+}23] Abigail Goldsteen, Ola Saadi, Ron Shmelkin, Shlomit Shachor, and Natalia Razinkov. AI privacy toolkit. SoftwareX, 22(??):??, May 2023. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/

science/article/pii/
S2352711023000481.

Grilli:2021:NPS

[GTC21] Nicolò Grilli, Edmund Tarleton, and Alan C. F.

Cocks. Neper2CAE and

PyCiGen: Scripts to generate polycrystals and interface elements in Abaqus. SoftwareX, 13(??):Article 100651, January 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020303642.

Gross:2021:MTR

[GTG21] Felix Groß, Nick Träger, and Joachim Gräfe. MIEP — a time-resolved X-ray image evaluation program. SoftwareX, 15(??):??, July 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021000509.

Gurses-Tran:2023:PPL

[GTOM23] Gonca Gürses-Tran, Florian Oppermann, and Antonello Monti. ProLoaF: Probabilistic load forecasting for power systems. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023001838.

Gupta:2020:RNV

[GV20] Udit Gupta and Dionisios G. Vlachos. Reaction

Network Viewer (ReNView):

an open-source framework for reaction path visualization of chemical reaction systems. SoftwareX, 11 (??):Article 100442, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019302432.

Gomez:2019:CMC

[GVAO19] Francisco J. Gómez, Luigi Vanfretti, Miguel Aguilera, and Svein H. Olsen. CIM-2-mod: a CIM to Modelica mapping and model-2-model transformation engine. SoftwareX, 9(??):161–167, January/ June 2019. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018300554.

Gaugl:2023:GOS

[GWBF23] Robert Gaugl, Sonja Wogrin, Udo Bachhiesl, and Lukas Frauenlob. GridTool: an open-source tool to convert electricity grid data. SoftwareX, 21(??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/

science/article/pii/
S2352711023000109.

Gajowniczek:2022:HHO

[GWGB22] Krzysztof Gajowniczek, Jialin Wu, Soumyajit Gupta, and Chandrajit Bajaj. HOFS: Higher order mutual information approximation for feature selection in R. SoftwareX, 19(??):??, July 2022. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022000930.

Govender:2016:BDM

[GWK16] Nicolin Govender, Daniel N.

Wilke, and Schalk Kok.

Blaze-DEMGPU: Modular

high performance DEM framework for the GPU architecture. SoftwareX, 5(??):51–61, ???? 2016. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271101630005X.

German:2023:ESM

[GY23] Péter German and Dewen Yushu. Enabling scientific machine learning in MOOSE using Libtorch. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023001851.

Guo:2023:OPP

[GYWG23] Junjun Guo, Aijun Ye, Xiaowei Wang, and Zhong-

                                        guo Guan. OpenSeesPyView:

Python programming-based visualization and postprocessing tool for OpenSeesPy. SoftwareX, 21(??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022001960.

Gajowniczek:2021:IRPa

[GZ21a] Krzysztof Gajowniczek and Tomasz Zabkowski. ImbTreeAUC: an R package for building classification trees using the area under the ROC curve (AUC) on imbalanced datasets. SoftwareX, 15(??):??, July 2021. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021000807.

Gajowniczek:2021:IRPb

[GZ21b] Krzysztof Gajowniczek and Tomasz Zabkowski.

                                        ImbTreeEntropy:  an

R package for building entropy-based classification trees on imbalanced datasets. SoftwareX, 16 (??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/

science/article/pii/
S2352711021001242.

Gallo:2022:MAT

[GZA^{+}22] Vincent Gallo, Xin Zhou, Diane Abdallah, Manuela Angioi, Emma Redding, Beth Ackroyd, Chiara Galvan, and Elisabetta Ver-

sace. MotionPerfection:

an agile tool for the visualisation, analysis, annotation, and record of motor practice. SoftwareX, 20(??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022001650.

Galizia:2019:JGM

[GZR^{+}19] Antonella Galizia, Gabriele Zereik, Luca Roverelli, Emanuele Danovaro, Andrea Clematis, and Daniele D’Agostino. Json-GUI — a module for the dynamic generation of formbased web interfaces. SoftwareX, 9(??):28–34, January/June 2019. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018300505.

Galdyn:2023:RGG

[GZS23] Filip Gałdyn, Radosław Zajdel, and Krzysztof Sośnica. RINEXAV: GNSS global network selection open-source software based on qualitative analysis of RINEX files. SoftwareX, 22(??):??, May 2023. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023000687.

Gomez:2022:PPB

[GZW^{+}22] Alberto Gomez, Veronika A. Zimmer, Gavin Wheeler, Nicolas Toussaint, Shujie Deng, Robert Wright, Emily Skelton, Jackie Matthew, Bernhard Kainz, Jo Hajnal, and Julia Schnabel. PRETUS: a plug-in based platform for realtime ultrasound imaging research. SoftwareX, 17 (??):??, January 2022. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021001849.

Hahne:2019:POD

[HA19] Christopher Hahne and

                                        Amar Aggoun. PlenoptiSign:

an optical design tool for plenoptic imaging. SoftwareX, 10(??):Article 100259, July/December 2019. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711019300159.

Huang:2023:HGS

[HA23] Ming Huang and Cameron F.

                                        Abrams. HTPolyNet: a

general system generator for all-atom molecular simulations of amorphous crosslinked polymers. SoftwareX, 21(??): ??, February 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022002217.

Hacihabiboglu:2023:SPP

[Hac23] Hüseyin Hacihabiboglu.

sphstat: a Python package for inferential statistics on vectorial data on the unit sphere. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023002108.

Hankin:2020:IPR

[Han20] Robin K. S. Hankin. Intro-

ducing the permutations

R package. SoftwareX, 11 (??):Article 100453, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019302055.

Hasanzadeh:2018:IIA

[Has18] Kamyar Hasanzadeh. IASM:

Individualized activity space modeler. SoftwareX, 7(??):138–142, January/ June 2018. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018300542.

Herring:2020:BPL

[HBA^{+}20] Patrick Herring, Chirranjeevi Balaji Gopal, Muratahan Aykol, Joseph H. Montoya, Abraham Anapolsky, Peter M. Attia, William Gent, Jens S. Hummelshøj, Linda Hung, Ha-Kyung Kwon, Patrick Moore, Daniel Schweigert, Kristen A. Severson, Santosh Suram, Zi Yang, Richard D. Braatz, and Brian D. Storey. BEEP: a Python library for battery evaluation and early prediction. SoftwareX, 11 (??):Article 100506, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020300492.

Horvath:2016:DEL

[HBS16] Balázs Zsigmond Horváth, Bence Blaske, and Anita Szabó. Duenna — an experimental language teaching application. SoftwareX, 5(??):163–170, ???? 2016. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711016300231.

Hadjidoukas:2020:TST

[HBS^{+}20] P. E. Hadjidoukas, A. Bartezzaghi, F. Scheidegger, R. Istrate, C. Bekas, and A. C. I. Malossi. torcpy: Supporting task parallelism in Python. SoftwareX, 12(??):Article 100517, July/December 2020. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020300091.

Hackel:2024:CGP

[HCD24] Leonard Hackel, Kai Norman Clasen, and Begüm Demir. ConfigILM: a general purpose configurable library for combining image and language models for visual question answering. SoftwareX, 26 (??):??, May 2024. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271102400102X.

Hertel:2020:SRH

[HCS^{+}20] Lars Hertel, Julian Collado, Peter Sadowski, Jordan Ott, and Pierre Baldi. Sherpa: Robust hyperparameter optimization for machine learning. SoftwareX, 12(??):Article 100591, July/December 2020. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S2352711020303046.

Heller:2021:STS

[HDA21] William T. Heller, Mathieu Doucet, and Richard K. Archibald. Sas-temper: Software for fitting smallangle scattering data that provides automated reproducibility characterization. SoftwareX, 16 (??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S235271102100128X.

Helske:2022:EBG

[Hel22] Jouni Helske. Efficient Bayesian generalized linear models with timevarying coefficients: the walker package in R. SoftwareX, 18(??):??, June 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S235271102200022X.

Hammad:2019:RRE

[HESH19] Issam Hammad, Kamal El-Sankary, and Holly Hornibrook. RETSManager: Real-estate database builder and synchronizer. SoftwareX, 10(??):Article 100351, July/December 2019. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/

science/article/pii/
S235271101930250X.

Haidar:2021:PPP

[HFS^{+}21] Ali Haidar, Matthew Field, Jonathan Sykes, Martin Carolan, and Lois Holloway. PSPSO: a package for parameters selection using particle swarm optimization. SoftwareX, 15(??):??, July 2021. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021000510.

Harwood:2020:JOS

[HGG20] Adrian R. G. Harwood, James Gill, and Simeon Gill. JBlockCreator: an open source, pattern drafting framework to facilitate the automated manufacture of made-to-measure clothing. SoftwareX, 11 (??):Article 100365, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018302528.

Haustein:2017:DEM

[HGS17] Martin Haustein, Anton Gladkyy, and Rüdiger Schwarze. Discrete element modeling of deformable particles in YADE. SoftwareX, 6(??):107–117, ???? 2017. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711017300134.

Hassan:2018:SSP

[HGWM18] Rakib Hassan, Michael Gurnis, Simon E. Williams, and R. Dietmar Müller. SPGM: a Scalable PaleoGeomorphology Model. SoftwareX, 7(??):263–272, January/June 2018. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711018301341.

Handler:2021:PAM

[HH21] David C. L. Handler and Paul A. Haynes.

                                        PeptideMind — apply-

ing machine learning algorithms to assess replicate quality in shotgun proteomic data. SoftwareX, 13 (??):Article 100644, January 2021. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711020303575.

Haidri:2022:PPP

[HHB^{+}22a] Salman Haidri, Yaksh J. Haranwala, Vania Bogorny, Chiara Renso, Vinicius Prado da Fonseca, and Amilcar Soares. PTRAIL — a Python package for parallel trajectory data preprocessing. SoftwareX, 19 (??):??, July 2022. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022001066.

Heller:2022:DDR

[HHB^{+}22b] William T. Heller, John Hetrick, Jean Bilheux, Jose M. Borreguero Calvo, Wei-Ren Chen, Lisa DeBeer-Schmitt, Changwoo Do, Mathieu Doucet, Michael R. Fitzsimmons, William F. Godoy, Garrett E. Granroth, Steven Hahn, Lilin He, Fahima Islam, Jiao Lin, Kenneth C. Littrell, Marshall McDonnell, Jesse McGaha, Peter F. Peterson, Sai Venkatesh Pingali, Shuo Qian, Andrei T. Savici, Yingrui Shang, Christopher B. Stanley, Volker S. Urban, Ross E. Whitfield, Chen Zhang, Wenduo Zhou, Jay Jay Billings, Matthew J. Cuneo, Ricardo M. Ferraz Leal, Tianhao Wang, and Bin Wu. drtsans: the data reduction toolkit for small-angle neutron scattering at Oak Ridge National Laboratory. SoftwareX, 19(??):??, July 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022000681.

Haas:2022:RDE

[HHH22] Christoph Haas, Dörthe Holthusen, and Rainer

                                        Horn. The Rheological-Data
                                        Extraction Application:

a time-saving tool for the extraction of measured rheological data from a specific rheometerinherent software. SoftwareX, 18(??):??, June 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022000450.

Hansen:2020:CSV

[HHN20] Marcus H. Hansen, Peter Hokland, and Charlotte G. Nyvold. CNAplot — Software for visual inspection of chromosomal copy number alteration in cancer using juxtaposed sequencing read depth ratios and variant allele frequencies. SoftwareX, 11 (??):Article 100503, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711020300522.

Haro:2023:BMS

[HJRBM23] Sara Haro, Jonathan Jimenez-Reina, Ricardo Bermejo, and Liam Mor-

                                        rison.  BioIntertidal

Mapper software: a satellite approach for NDVIbased intertidal habitat mapping. SoftwareX, 24 (??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023002169.

Hafner:2018:SBR

[HKC^{+}18] Sasha D. Hafner, Konrad Koch, Hélène Carrere, Sergi Astals, Sören Weinrich, and Charlotte Rennuit. Software for biogas research: Tools for measurement and prediction of methane production. SoftwareX, 7(??): 205–210, January/June 2018. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018300840.

Hojati:2020:MPA

[HKF^{+}20] Sahar Hojati, Rahele Kafieh, Parisa Fardafshari, Masoud Aghsaei Fard, and Hatef Fouladi. A MAT-LAB package for automatic extraction of flow index in OCT-A images by intelligent vessel manipulation. SoftwareX, 12 (??):Article 100510, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020300364.

Horton:2019:GEA

[HKM^{+}19] Brandon K. Horton, Rajiv K. Kalia, Erick Moen, Aiichiro Nakano, Ken ichi Nomura, Michael Qian, Priya Vashishta, and Anders Hafreager. Game-Engine-Assisted Research platform for Scientific computing (GEARS) in virtual reality. SoftwareX, 9(??):112–116, January/ June 2019. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711018300633.

Harmouche:2022:XOS

[HKS^{+}22] Ahmed Harmouche, Ferenc Kövér, Sándor Szukits, Tamás Dóczi, Péter Bogner, and Arnold Tóth. XReport: an online structured reporting platform for radiologists. SoftwareX, 17(??):??, January 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022000115.

Habib:2024:LSB

[HKZ24] Mahir Habib, Muhammad Ashad Kabir, and Lihong Zheng. LEI2JSON: Schema-based validation and conversion of livestock event information. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023003254.

Holmi:2022:WMD

[HL22] Joonas Tapani Holmi and Harri Lipsanen. WITio: a MATLAB data evaluation toolbox to script broader insights into big data from WITec microscopes. SoftwareX, 18(??):??, June 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271102200019X.

Hamida:2022:OPF

[HLG22] Zachary Hamida, Blanche Laurent, and James-A. Goulet. OpenIPDM: a probabilistic framework for estimating the deterioration and effect of interventions on bridges. SoftwareX, 18(??):??, June 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022000565.

Haselsteiner:2019:VSC

[HLP^{+}19] Andreas F. Haselsteiner, Jannik Lehmkuhl, Tobias Pape, Kai-Lukas Windmeier, and Klaus-Dieter Thoben. ViroCon: a software to compute multivariate extremes using the environmental contour method. SoftwareX, 9(??):95–101, January/June 2019. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711018301420.

Herritt:2021:FFI

[HLR^{+}21] Matthew T. Herritt, Jacob C. Long, Mike D. Roybal, David C. Moller, Todd C. Mockler, Duke Pauli, and Alison L. Thompson. FLIP: FLuorescence Imaging Pipeline for field-based chlorophyll fluorescence images. SoftwareX, 14(??):??, June 2021. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021000303.

Huang:2016:UBI

[HLW^{+}16] Sheng-Cheng Huang, Sara Lee, Allen Wang, Scott B. Cantor, Clement Sun, Kaili Fan, Gregory P. Reece, Min Soon Kim, and Mia K. Markey. UT Biomedical Informatics Lab (BMIL) probability wheel. SoftwareX, 5(??): 203–210, ???? 2016. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711016300292.

Hebert:2022:NPP

[HM22] Raphaël Hebert and Emese Meglécz. NSDPY: a Python package to download DNA sequences from NCBI. SoftwareX, 18(??):??, June 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271102200036X.

Hafner:2023:SST

[HM23] Sasha D. Hafner and Kristina Mjöfors. STM: a software tool for predicting daily temperature in stored animal slurry. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271102300050X.

Hnaini:2024:SWB

[HMC^{+}24] Hiba Hnaini, Raúl Mazo, Joël Champeau, Paola Vallejo, and Jose Galindo. E-SCORE: a web-based tool for security requirements engineering. SoftwareX, 26 (??):??, May 2024. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271102400075X.

Hanwell:2015:VTV

[HMCA15] Marcus D. Hanwell, Kenneth M. Martin, Aashish Chaudhary, and Lisa S. Avila. The Visualization Toolkit (VTK): Rewriting the rendering code for modern graphics cards. SoftwareX, 1–2(??):3–8, September 2015. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711015000035.

Hoffmann:2023:SUF

[HMK23] Thomas J. Hoffmann, Christine Miaskowski, and Kord M. Kober. ShinyGAStool: a user-friendly tool for candidate gene association studies. SoftwareX, 21(??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022001923.

Heredia:2024:OPG

[HMY24] Cristobal Heredia, Sebastian Moreno, and Wilfredo Yushimito. ODMeans: an R package for global and local cluster detection for origin-destination GPS data. SoftwareX, 26 (??):??, May 2024. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711024001031.

Hernandez-Olivan:2023:MPL

[HOB23] Carlos Hernandez-Olivan and Jose R. Beltran. Musicaiz: a Python library for symbolic music generation, analysis and visualization. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023000614.

Harwood:2018:LMC

[HOM^{+}18] Adrian R. G. Harwood, Joseph O’Connor, Jonathan Sanchez Muñoz, Marta Camps Santasmasas, and Alistair J. Revell. LUMA: a many-core, fluid-structure interaction solver based on the Lattice-Boltzmann method. SoftwareX, 7 (??):88–94, January/June 2018. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018300219.

Hart:2020:MSM

[HR20] Kenneth A. Hart and Julian J. Rimoli. MicroStructPy: a statistical microstructure mesh generator in Python. SoftwareX, 12(??):Article 100595, July/December 2020. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020303083.

Hoy:2021:PCA

[HR21] Charlie Hoy and Vivien Raymond. PESummary: the code agnostic Parameter Estimation Summary page builder. SoftwareX, 15 (??):??, July 2021. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021000856.

Hsieh:2020:PRP

[HRC20] Nan-Hung Hsieh, Brad Reisfeld, and Weihsueh A. Chiu. pksensi: an R package to apply global sensitivity analysis in physiologically based kinetic modeling. SoftwareX, 12 (??):Article 100609, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711020303228.

Heinrich:2020:CVF

[HS20] Martin Heinrich and Rüdiger Schwarze. 3D-coupling of volume-of-fluid and Lagrangian particle tracking for spray atomization simulation in Open-FOAM. SoftwareX, 11 (??):Article 100483, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711020300303.

Heldens:2022:LPP

[HSD^{+}22] Stijn Heldens, Alessio Sclocco, Henk Dreuning, Ben van Werkhoven, Pieter Hijma, Jason Maassen, and Rob V. van Nieuw-

poort.  litstudy: a

Python package for literature reviews. SoftwareX, 20(??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271102200125X.

Haghgoo:2022:ESP

[HSMF22] Maliheh Haghgoo, Ilya Sychev, Antonello Monti, and Frank H. P. Fitzek. ENTIRETY — sEmanNTIc pRovisioning and govErning ioT devices in smart energY domain. SoftwareX, 18(??):??, June 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022000589.

Henelius:2018:MOS

[HT18] Andreas Henelius and Jari Torniainen. MI-DAS: Open-source framework for distributed online analysis of data streams. SoftwareX, 7(??):156–161, January/June 2018. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018300578.

Heder:2019:CBS

[HTB19] Mihály Héder, Szabolcs Tenczer, and Andrea Biancini. Collaboration between SAML federations and OpenStack clouds. SoftwareX, 9(??):44–48, January/June 2019. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711018302905.

Hrovatin:2022:PPP

[HTV22] Niki Hrovatin, Aleksandar Tosić, and Jernej Vicic. PPWSim: Privacy preserving wireless sensor network simulator. SoftwareX, 18(??):??, June 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022000516.

Ho:2023:CFR

[HTWT23] George To Sum Ho, Yung Po Tsang, Qin Wu, and Valerie Tang. ck-FARM: an R package to discover big data associations for business intelligence. SoftwareX, 22 (??):??, May 2023. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023000377.

Hansen:2016:MCC

[HVB16] Thomas Mejer Hansen, Le Thanh Vu, and Torben Bach. MPSLIB: a C++ class for sequential simulation of multiple-point statistical models. SoftwareX, 5(??):121–126, ???? 2016. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711016300164.

Herrera:2023:RIF

[HVCS^{+}23] Vanesa Herrera, David Vallejo, José J. Castro-Schez, Dorothy N. Monekosso, Ana de los Reyes, Carlos Glez-Morcillo, and Javier Albusac. Rehab-Immersive: a framework to support the development of virtual reality applications in upper limb rehabilitation. SoftwareX, 23 (??):??, July 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023001085.

Hai:2024:POQ

[HVH24] Vu Tuan Hai, Nguyen Tan Viet, and Le Bin Ho. < qo|op>: a quantum object optimizer. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024000979.

Haselsteiner:2022:UVS

[HWST22] Andreas F. Haselsteiner, Kai-Lukas Windmeier, Lena Ströer, and Klaus-Dieter Thoben. Update 2.0 to “ViroCon: a software to compute multivariate extremes using the environmental contour method”. SoftwareX, 20 (??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022001613.

Holm:2023:JDS

[HWT^{+}23] Stefan Holm, Marc Werder, Oliver Thees, Renato Lemm, and Janine Schweier. JuWaPfl: a decision support tool to estimate times and costs of processes related to young-forest maintenance. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023002777.

Huang:2022:IOD

[HX22] DongBo Huang and Xing Xu. IGAOD: an online design framework for interactive genetic algorithms. SoftwareX, 19(??):??, July 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022001236.

Hu:2024:GGG

[HX24] Hongyi Hu and Zhonghua

                                        Xu. Global-GMDs: the

global map of geomagnetic disturbances. SoftwareX, 25(??):??, February 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023003151.

Hussain:2023:OSF

[HYR^{+}23] Suhail S. M. Hussain, Chen Yaohao, Muhammad M. Roomi, Daisuke Mashima, and Ee-Chien Chang. An open-source framework for publishing/ subscribing IEC 61850 R-GOOSE and R-SV. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023001115.

Imgrund:2017:RLV

[IA17] M. Imgrund and A. Arth.

Rambrain — a library for virtually extending physical memory. SoftwareX, 6 (??):172–178, ???? 2017. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711017300274.

Ingley:2015:AOS

[IAW^{+}15] Spencer J. Ingley, Mohammad Rahmani Asl, Chengde Wu, Rongfeng Cui, Mahmoud Gadelhak, Wen Li, Ji Zhang, Jon Simpson, Chelsea Hash, Trisha Butkowski, Thor Veen, Jerald B. Johnson, Wei Yan, and Gil G. Rosenthal. anyFish 2.0: an open-source software platform to generate and share animated fish models to study behavior. SoftwareX, 3–4(??):6–12, December 2015. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711015000114.

Ibrahim:2024:COS

[ICS^{+}24] Ibrahim Anwar Ibrahim, Tanveer Choudhury, James Sargeant, Rakibuzzaman Shah, Md. Jahangir Hossain, and Syed Islam. CEREI: an open-source tool for cost-effective renewable energy investments. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024000797.

Iyaloo:2021:APM

[IDE^{+}21] Diana P. Iyaloo, Pascal Degenne, Khouaildi Bin Elahee, Danny Lo Seen, Ambicadutt Bheecarry, and Annelise Tran. ALBOMAURICE: a predictive model for mapping Aedes albopictus mosquito populations in Mauritius. SoftwareX, 13 (??):Article 100638, January 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020303514.

Ito:2023:SPF

[IHT23] Yuma Ito, Masanori Hirose, and Makio Toku-

naga.  Slitflow: a

Python framework for single-molecule dynamics and localization analysis. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023001589.

Iro:2023:VVB

[IIH^{+}23] Michael Iro, Dieter Ingerle, Sven Hampel, Ursula Fittschen, Vishal Dhamgaye, Oliver Fox, and Christina Streli. voxTrace: a voxel-based Monte-Carlo ray-tracing code for the simulation of X-ray fluorescence spectra. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023001772.

Izquierdo:2018:EMC

[IIS18] Luis R. Izquierdo, Segismundo S. Izquierdo, and William H. Sandholm. EvoDyn-3s: a Mathematica computable document to analyze evolutionary dynamics in 3-strategy games. SoftwareX, 7(??): 226–233, January/June 2018. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711018300864.

Ishii:2022:BBO

[IKYY22] Akimitsu Ishii, Ryunosuke Kamijyo, Akinori Yamanaka, and Akiyasu Yamamoto. BOXVIA: Bayesian optimization executable and visualizable application. SoftwareX, 18 (??):??, June 2022. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022000243.

Icenhour:2024:MEM

[ILP^{+}24] Casey T. Icenhour, Alexander D. Lindsay, Cody J. Permann, Richard C. Martineau, David L. Green, and Steven C. Shannon. The MOOSE electromagnetics module. SoftwareX, 25(??):??, February 2024. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023003175.

Nomura:2020:RSR

[iNKN^{+}20] Ken ichi Nomura, Rajiv K. Kalia, Aiichiro Nakano, Pankaj Rajak, and Priya Vashishta. RXMD: a scalable reactive molecular dynamics simulator for optimized time-to-solution. SoftwareX, 11 (??):Article 100389, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019300342.

Ikaheimonen:2023:NTB

[ITL^{+}23] Arsi Ikäheimonen, Ana M.

Triana, Nguyen Luong, Amirmohammad Ziaei, Jarno Rantaharju, Richard Darst, and Talayeh Aledavood. Niimpy: a toolbox for behavioral data analysis. SoftwareX, 23 (??):??, July 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023001681.

Imani:2023:RWT

[IVSV23] Saba Imani, Majid Vahed, Shreya Satodia, and Mohammad Vahed.

T-Ridership: a web tool

for reprogramming public transportation fleets to minimize COVID-19 transmission. SoftwareX, 22(??):??, May 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023000468.

Ibhadode:2021:ITO

[IZBT21] Osezua Ibhadode, Zhidong Zhang, Ali Bonakdar, and Ehsan Toyserkani. IbIPP for topology optimization — an image-based initialization and postprocessing code written in MATLAB. SoftwareX, 14 (??):??, June 2021. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021000467.

Jimenez:2020:SMT

[JAC20] Cristian Jimenez, Iván Amaya, and Rodrigo Cor-

                                        rea. SpinUpFlowDescriptor:

a MATLAB toolbox for ferrofluids materials under moderate and high amplitude and frequency of magnetic rotating fields in a spin-up geometry. SoftwareX, 12(??):Article 100567, July/December 2020. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711020302636.

Jimenez:2021:SSE

[JAC21] Cristian Jiménez, Ivan Amaya, and Rodrigo Correa. SCSV IPS: a software for the estimation of the shear coefficient of spin viscosity during the movement of a ferrofluid under the effect of an external rotating magnetic field. SoftwareX, 14(??):??, June 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021000133.

Jimenez:2023:QFT

[JAMP23] Miguel González Jiménez, Héctor Aguilera, Luis Moreno Merino, and Amalia Romero Prados. Q-Facies: a tool for the quantitative interpretation of groundwater hydrochemical facies. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023001462.

Janssen:2024:RFE

[Jan24] Rik D. T. Janssen. Ricgraph:

a flexible and extensible graph to explore research in context from various systems. SoftwareX, 26 (??):??, May 2024. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024001079.

Jung:2023:WFE

[JC23] Geunseong Jung and Jaehyuk Cha. A WebExtension framework for experimentation and evaluation of webpage segmentation methods. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023001978.

Jo:2024:STN

[JCFM24] Wonse Jo, Go-Eum Cha, Dan Foti, and Byung-

                                        Cheol Min. SMART-TeleLoad:

a new graphic user interface to generate affective loads for teleoperation. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711024001286.

Jozic:2022:DSW

[JFJM22] Kresimir Jozić, Nikolina Frid, Alan Jović, and Zeljka Mihajlović. DICOM SIVR: a web architecture and platform for seamless DICOM image and volume rendering. SoftwareX, 18(??):??, June 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022000498.

Jose-Garcia:2023:CMB

[JGGF23] Adán José-García and Wilfrido Gómez-Flores. CVIK: a Matlab-based cluster validity index toolbox for automatic data clustering. SoftwareX, 22 (??):??, May 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023000559.

Johnston:2019:SMP

[JHAMF19] Josiah Johnston, Rodrigo Henriquez-Auba, Benjamín Maluenda, and Matthias Fripp. Switch 2.0: a modern platform for planning highrenewable power systems. SoftwareX, 10(??):Article 100251, July/December 2019. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018301547.

Jakimow:2023:EBI

[JJT^{+}23] Benjamin Jakimow, Andreas Janz, Fabian Thiel, Akpona Okujeni, Patrick Hostert, and Sebastian van der Linden. EnMAP-Box: Imaging spectroscopy in QGIS. SoftwareX, 23 (??):??, July 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023002030.

Jalowiecki:2023:PPL

[JLP23] Konrad Jałowiecki, Paulina Lewandowska, and Lukasz Pawela. PyQBench: a Python library for benchmarking gate-based quantum computers. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023002546.

Joksas:2020:BPT

[JM20] Dovydas Joksas and Adnan Mehonic. badcrossbar: a Python tool for computing and plotting currents and voltages in passive crossbar arrays. SoftwareX, 12(??):Article 100617, July/December 2020. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711020303307.

Jordan:2022:BHE

[JMT22] Matthias Jordan, Markus Millinger, and Daniela

                                        Thrän.  Benopt-Heat:

an economic optimization model to identify robust bioenergy technologies for the German heat transition. SoftwareX, 18 (??):??, June 2022. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022000334.

Jalowiecki:2023:OEI

[JP23] Konrad Jałowiecki and Lukasz Pawela. Omnisolver: an extensible interface to Ising spin-glass and QUBO solvers. SoftwareX, 24 (??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023002558. See

corrigendum [JP24].

Jalowiecki:2024:COE

[JP24] K. Jałowiecki and L. Pawela.

Corrigendum to “Omnisolver: an extensible interface to Ising spinglass and QUBO solvers” [SoftwareX 24 (2023) 101559]. SoftwareX, 25 (??):??, February 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024000013. See

[JP23].

Jamal:2023:MDA

[JRP23] Shaji M. Jamal, Neelakantan T. R., and Meyyappan Palaniappan. MesoGen — development of an Android application for the Mesoscopic modeling of concrete. SoftwareX, 21 (??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022002011.

Johnson:2019:ACH

[JS19] Gabriel Johnson and Sandra Spiroff. Automating the calculation of the Hilbert–Kunz multiplicity and F -signature. SoftwareX, 9(??):35–38, January/June 2019. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018301833.

Janke:2020:PMP

[JSB20] T. Janke, R. Schwarze, and K. Bauer. Part2Track: a MATLAB package for double frame and time resolved Particle Tracking Velocimetry. SoftwareX, 11(??):Article 100413, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019303395.

Jericevich:2022:COS

[JSG22] Ivan Jericevich, Dharmesh Sing, and Tim Gebbie. CoinTossX: an opensource low-latency highthroughput matching engine. SoftwareX, 19(??): ??, July 2022. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022000875.

Jung:2021:AGB

[Jun21] Sang-Kyu Jung. AniLength: GUI-based automatic worm length measurement software using image processing and deep neural network. SoftwareX, 15 (??):??, July 2021. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271102100100X.

Jurkiewicz:2022:FMF

[Jur22] Piotr Jurkiewicz. flow-models:

a framework for analysis and modeling of IP network flows. SoftwareX, 17(??):??, January 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021001680.

Jurkiewicz:2023:FME

[Jur23a] Piotr Jurkiewicz. Flow-models 2.0: Elephant flows modeling and detection with machine learning. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023002029.

Jurkiewicz:2023:TRR

[Jur23b] Piotr Jurkiewicz. TopoHub:

a repository of reference Gabriel graph and realworld topologies for networking research. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023002364.

Jarmatz:2022:MEO

[JWJ^{+}22] Piet Jarmatz, Helene Wittenberg, Vahid Jafari, Amartya Das Sharma, Felix Maurer, Niklas Wittmer, and Philipp

                                        Neumann. MaMiCo 2.0:

an enhanced open-source framework for high-performance molecular-continuum flow simulation. SoftwareX, 20 (??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022001698.

Jivani:2022:GES

[JZGW22] Devyani Jivani, Jaroslaw Zola, Baskar Ganapathysubramanian, and Olga Wodo. GraSPI: Extensible software for the graphbased quantification of morphology in organic electronics. SoftwareX, 17(??):??, January 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021001898.

Kammer:2021:USB

[KAK21] David S. Kammer, Gabriele Albertini, and Chun-Yu Ke. UGUCA: a spectralboundary-integral method for modeling fracture and friction. SoftwareX, 15 (??):??, July 2021. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021000959.

Kim:2024:MBK

[KALS24] Dong-Hyun Kim, Seokho Ahn, Euijong Lee, and Young-Duk Seo. Morphemebased Korean text cohesion analyzer. SoftwareX, 26(??):??, May 2024. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271102400030X.

Kurrant:2021:MIA

[KAO^{+}21] Douglas Kurrant, Nasim Abdollahi, Muhammad Omer, Pedram Mojabi, Elise Fear, and Joe LoVetri. MWSegEval — an image analysis toolbox for microwave breast images. SoftwareX, 15(??):??, July 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021000674.

Kafieh:2019:AMM

[KAR^{+}19] Rahele Kafieh, Zahra Amini, Hossein Rabbani, Bahareh Kaviani Baghbaderani, Bahareh Salafian, Fatemeh Mazaheri, and Marzieh Mokhtari. Automatic multifaceted Matlab package for analysis of ocular images (AM-PAO). SoftwareX, 10 (??):Article 100339, July/ December 2019. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711019301852.

Karakoc:2021:RRB

[Kar21] Alp Karakoç. RegionTPMS — region based triply periodic minimal surfaces (TPMS) for 3-D printed multiphase bone scaffolds with exact porosity values. SoftwareX, 16(??): ??, December 2021. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021001175.

Kumar:2023:GGP

[KAS23] Manish Kumar, Roberto Alessi, and Enrico Salvati. GPFniCS: a generalised phase field method to model fracture. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S235271102300290X.

Kocher:2019:ESD

[KBB19] Uwe Köcher, Marius Paul Bruchhäuser, and Markus Bause. Efficient and scalable data structures and algorithms for goaloriented adaptivity of space-time FEM codes. SoftwareX, 10(??):Article 100239, July/December 2019. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018302978.

Kapsecker:2024:PMD

[KCNJ24] Maximilian Kapsecker, Nikita Charushnikov, Leon Nissen, and Stephan M. Jonas. PeakSwift: Mobile detection of r-peaks in single lead electrocardiograms. SoftwareX, 25 (??):??, February 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023003047.

Krupalija:2024:ERC

[KCP^{+}24] Ehlimana Krupalija, Emir Cogo, Damir Pozderac, Samir Omanović, Almir Karabegović, Razija Turčinhodžić Mulahasanović, and Ing-

mar Bešić. ETF-RI-CEG-Advanced:

a graphical desktop tool for black-box testing by using cause-effect graphs. SoftwareX, 25(??):??, February 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023003217.

Karathanasopoulos:2020:LDM

[KDHG20] N. Karathanasopoulos, F. Dos Reis, P. Hadjidoukas, and J. F. Ganghof-

                                        fer.  LatticeMech: a

discrete mechanics code to compute the effective static properties of 2D metamaterial structures. SoftwareX, 11 (??):Article 100446, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S235271101930370X.

Kafando:2023:SSN

[KDRT23] Rodrique Kafando, Rémy Decoupes, Mathieu Roche, and Maguelonne Teisseire. SNEToolkit: Spatial named entities disambiguation toolkit. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023001760.

Kiplimo:2024:DTS

[KEwMO24] Cedric Kiplimo, Collins Emasi Epege, Ciira wa Maina,

                                        and Billy Okal. DSAIL-TreeVision:

a software tool for extracting tree biophysical parameters from stereoscopic images. SoftwareX, 26 (??):??, May 2024. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024000323.

Khalil:2017:RRP

[KF17] Salim Khalil and Mohamed Fakir. RCrawler: an R package for parallel web crawling and scraping. SoftwareX, 6 (??):94–97, ???? 2017. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711017300110.

Kogon:2022:STM

[KF22] Rémi Kogon and David Faux. 3TM: Software for the 3-Tau Model. SoftwareX, 17(??):??, January 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022000048.

Khatami:2023:CCP

[KF23] Siamak Khatami and Christopher Frantz. Copatrec: a correlation pattern recognizer Python package for nonlinear relations. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023001528.

Khodke:2024:TLG

[KFA^{+}24] Sumedh Khodke, Evan Fox, Nafisa Anjum, Dineshkumar Katta, Amritha Subburayan, Sanjay Aravind Loganathan Ravichandran, MD Ahasan Habib, and Nazmus Sakib. TeamWellX: Leveraging genetic algorithms to optimize team health and well-being toward sustainable game development and strategy. SoftwareX, 25(??):??, February 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024000062.

Khuat:2023:HBP

[KG23] Thanh Tung Khuat and

Bogdan Gabrys. hyperbox-brain:

a Python toolbox for hyperbox-based machine learning algorithms. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023001218.

Kaim:2024:MLB

[KG24] Adrian Kaim and Katarzyna Gwó’zd’z. myDLTS: Lab-VIEW based software for deep level transient spectroscopy using impedance analyser. SoftwareX, 26 (??):??, May 2024. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024000505.

Krummenacher:2024:PHE

[KGF^{+}24] Marco Krummenacher, Moritz Gubler, Jonas A. Finkler, Hannes Huber, Martin Sommer-Jörgensen, and Stefan Goedecker. Performing highly efficient Minima Hopping structure predictions using the Atomic Simulation Environment (ASE). SoftwareX, 25(??):??, February 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024000037.

Krzemien:2020:JPF

[KGK^{+}20] W. Krzemien, A. Gajos, K. Kacprzak, K. Rakoczy, and G. Korcyl. J-PET Framework: Software platform for PET tomography data reconstruction and analysis. SoftwareX, 11 (??):Article 100487, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020300509.

Karathanasopoulos:2019:TOS

[KH19a] Nikolaos Karathanasopoulos and Panagiotis Had-

jidoukas.  TendonMech:

an open source high performance code to compute the mechanical behavior of tendon fascicles. SoftwareX, 9(??):324– 327, January/June 2019. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711018302929.

Kasprowski:2019:EVE

[KH19b] Paweł Kasprowski and Katarzyna Harezlak. ET-CAL — a versatile and extendable library for eye tracker calibration. SoftwareX, 8(??):71–76, ???? 2019. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711017300742.

Kawan:2021:RRE

[KHG21] Christoph Kawan, Sigurdur Freyr Hafstein, and Peter Giesl. ResEntSG: Restoration entropy estimation for dynamical systems via Riemannian metric optimization. SoftwareX, 15(??):??, July 2021. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021000741.

Kannisto:2022:SPM

[KHH^{+}22] Petri Kannisto, Ville Heikkilä, Otto Hylli, Mehdi Attar, Sami Repo, and Kari Systä. SimCES platform for modular simulation: Featuring platform independence, container ecosystem, and development toolkit. SoftwareX, 19(??):??, July 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022001133.

Krotka:2023:NRP

[KHJ^{+}23] Pavla Krotka, Katharina Hees, Peter Jacko, Dominic Magirr, Martin Posch, and Marta Bofill Roig. NCC: an R-package for analysis and simulation of platform trials with non-concurrent controls. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023001334.

Kayser:2022:OCP

[KHM^{+}22] Hendrik Kayser, Tobias Herzke, Paul Maanen, Max Zimmermann, Giso Grimm, and Volker Hohmann. Open community platform for hearing aid algorithm research: open Master Hearing Aid (open-MHA). SoftwareX, 17(??): ??, January 2022. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021001813.

Kisil:2020:MCL

[Kis20] Vladimir V. Kisil. MoebInv: C ++ libraries for manipulations in non-Euclidean geometry. SoftwareX, 11(??):Article 100385, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711019302523.

Kim:2023:SVC

[KJKK23] Dongha Kim, Yeongbin Jo, Taekyung Kim, and Hokeun Kim. SST v1.0.0 with C API: Pluggable security solution for the Internet of Things. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023000869.

Karczmarczyk:2021:OOO

[KJW21] Artur Karczmarczyk, Jarosław Jankowski, and Jarosław Watróbski. OONIS — Object-Oriented Network Infection Simulator. SoftwareX, 14(??):??, June 2021. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021000200.

Kaliszewski:2022:IWA

[KK22] Ignacy Kaliszewski and Olga Karelkina. IDOL: a web application for mixed integer linear multiobjective optimization. SoftwareX, 19(??):??, July 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022000723.

Kim:2024:AOS

[KK24] Jinseok Kim and Jenna Kim. ANDez: an opensource tool for author name disambiguation using machine learning. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024000906.

Karim:2021:DIS

[KKAUA21] Mohammad Bozlul Karim, Shigehiko Kanaya, and Md. Altaf-Ul-Amin. DPClusSBO: an integrated software for clustering of simple and bipartite graphs. SoftwareX, 16(??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271102100114X.

Kammouh:2022:MOS

[KKN^{+}22] Omar Kammouh, M. W.

A. (Maurits) Kok, Maria Nogal, Ruud Binnekamp, and A. R. M. (Rogier) Wolfert. MitC: Opensource software for construction project control and delay mitigation. SoftwareX, 18(??):??, June 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022000267.

Kien:2022:CSE

[KL22] Pham Ngoc Kien and Sang-Mook Lee. 1DCSEMQWE: 1D controlled source electromagnetic method in geophysics using quadrature with extrapolation. SoftwareX, 19(??):??, July 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022000838.

Klesk:2024:FEC

[Kle24] Przemysław Klesk. FastRealBoostBins: an ensemble classifier for fast predictions implemented in Python via

                                        numba.jit and numba.cuda.

SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711024000153.

Kouadio:2023:WML

[KLL23] Kouao Laurent Kouadio, Jianxin Liu, and Rong Liu. watex: machine learning research in water exploration. SoftwareX, 22 (??):??, May 2023. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/

science/article/pii/
S2352711023000638.

Kasprzhitskii:2021:XSC

[KLY21] Anton Kasprzhitskii, Georgy Lazorenko, and Victor Yavna. XMHFL: Software for calculating excited and ionized states of molecules by X-ray. SoftwareX, 13(??):Article 100647, January 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020303605.

Krasnopolsky:2021:XLS

[KM21] Boris Krasnopolsky and Alexey Medvedev. XAMG: a library for solving linear systems with multiple right-hand side vectors. SoftwareX, 14(??):??, June 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021000406.

Konur:2020:KSS

[KMIG20] Savas Konur, Laurentiu Mierla, Florentin Ipate, and Marian Gheorghe. kP-Workbench: a software suit for membrane systems. SoftwareX, 11 (??):Article 100407, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/

science/article/pii/
S2352711019302584.

Krishnamoorthy:2021:EPL

[KMK^{+}21] Aravind Krishnamoorthy, Ankit Mishra, Deepak Kamal, Sungwook Hong, Ken ichi Nomura, Subodh Tiwari, Aiichiro Nakano, Rajiv Kalia, Rampi Ramprasad, and Priya Vashishta. EZFF: Python library for multi-objective parameterization and uncertainty quantification of interatomic forcefields for molecular dynamics. SoftwareX, 13(??):Article 100663, January 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271102100008X.

Koot:2023:PVP

[KMLP^{+}23] Paul Koot, Miguel Angel Mendoza-Lugo, Dominik Paprotny, Oswaldo Morales-Nápoles, Elisa Ragno, and Daniël T. H. Worm. PyBanshee version (1.0): a Python implementation of the MATLAB toolbox BANSHEE for Non-Parametric Bayesian Networks with updated features. SoftwareX, 21 (??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022001972.

Kontopoulos:2023:TTR

[KMT23] Ioannis Kontopoulos, Antonios Makris, and Konstantinos Tserpes. TraClets: a trajectory representation and classification library. SoftwareX, 21(??): ??, February 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271102300002X.

Korontanis:2024:EML

[KMT24] Ioannis Korontanis, Antonios Makris, and Konstantinos Tserpes. EdgeCloud mon: a lightweight monitoring stack for K3s clusters. SoftwareX, 26(??): ??, May 2024. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024000463.

Kapsecker:2023:FLS

[KNW^{+}23] Maximilian Kapsecker, Daniel N. Nugraha, Christoph Weinhuber, Nicholas Lane, and Stephan M. Jonas. Federated learning with Swift: an extension of flower and performance evaluation. SoftwareX, 24 (??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023002297.

Kim:2021:ERP

[KOC21] Donghoh Kim, Hee-Seok Oh, and Guebin Choi. EPT: an R package for ensemble patch transform. SoftwareX, 14(??):??, June 2021. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021000492.

Korcyl:2021:NPS

[Kor21] Piotr Korcyl. Numerical package for solving the JIMWLK evolution equation in C++. SoftwareX, 16(??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021001473.

Kosztyan:2022:MMB

[Kos22] Zsolt T. Kosztyán. MFPP: Matrix-based flexible project planning. SoftwareX, 17 (??):??, January 2022. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022000012.

Kania:2024:JMM

[KOZ^{+}24] Adam Kania, Weronika Ormaniec, Dmytro Zhylko, Leszek Grzanka, Diana Piotrowska, and Andrzej

                                        Siódmok.  Joseph the

MoUSE — mouse ultrasonic sound explorer. SoftwareX, 25(??):??, February 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023003023.

Konkov:2019:QSS

[KP19] Vladimir Konkov and Roberto Peverati. QMC-SW: a simple workflow for quantum Monte Carlo calculations in chemistry. SoftwareX, 9(??): 7–14, January/June 2019. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018302218.

Koetje:2020:CIS

[KP20] Kara M. Koetje and Margaret L. Palmsten. Coastal imaging station design toolbox. SoftwareX, 11 (??):Article 100377, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019300354.

Krug:2023:MRP

[KP23] Rainer M. Krug and Owen L. Petchey. Microxanox: an R package for simulating an aquatic MICR obial ecosystem that can occupy OXic or ANOXic states. SoftwareX, 24 (??):??, December 2023.

CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023002352.

Kostenko:2020:PXR

[KPC^{+}20] A. Kostenko, W. J. Palenstijn, S. B. Coban, A. A. Hendriksen, R. van Liere, and K. J. Batenburg. Prototyping Xray tomographic reconstruction pipelines with FleXbox. SoftwareX, 11 (??):Article 100364, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711019301669.

Kurbucz:2024:LPL

[KPJ24] Marcell T. Kurbucz, Péter Pósfay, and Antal Jakovác. LLT: an R package for linear law-based feature space transformation. SoftwareX, 25(??):??, February 2024. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023003199.

Krasanakis:2023:JGN

[KPK23] Emmanouil Krasanakis, Symeon Papadopoulos, and Ioannis Kompatsiaris. JGNN: Graph Neural Networks on native Java. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023001553.

Krasanakis:2022:PPP

[KPKS22] Emmanouil Krasanakis, Symeon Papadopoulos, Ioannis Kompatsiaris, and Andreas L. Symeonidis. pygrank: a Python package for graph node ranking. SoftwareX, 20(??): ??, December 2022. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022001455.

Kiviluoma:2022:STF

[KPM^{+}22] Juha Kiviluoma, Fabiano Pallonetto, Manuel Marin, Pekka T. Savolainen, Antti Soininen, Per Vennström, Erkka Rinne, Jiangyi Huang, Iasonas Kouveliotis-Lysikatos, Maren Ihlemann, Erik Delarue, Ciara O’Dwyer, Terence O’Donnel, Mikael Amelin, Lennart Söder, and Joseph

Dillon. Spine Toolbox: a

flexible open-source workflow management system with scenario and data management. SoftwareX, 17(??):??, January 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021001886.

Kazantsev:2018:TSP

[KPN^{+}18] Daniil Kazantsev, Valery Pickalov, Srikanth Nagella, Edoardo Pasca, and Philip J. Withers. TomoPhantom, a software package to generate 2D– 4D analytical phantoms for CT image reconstruction algorithm benchmarks. SoftwareX, 7(??): 150–155, January/June 2018. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711018300335.

Kiesewetter:2016:XES

[KPOD16] Simon Kiesewetter, Rodney Polkinghorne, Bogdan Opanchuk, and Peter D. Drummond. xSPDE: Extensible software for stochastic equations. SoftwareX, 5(??):6–11, ???? 2016. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711016000030.

Kritikos:2023:SGT

[KPS23] Apostolos Kritikos, Prodromos Polychroniadis, and Ioannis Stamelos.

                                        Source-o-grapher: a tool

towards the investigation of software resilience in Open Source Software projects. SoftwareX, 22 (??):??, May 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271102300033X.

Kintsakis:2017:HSD

[KPSM17] Athanassios M. Kintsakis, Fotis E. Psomopoulos, Andreas L. Symeonidis, and Pericles A. Mitkas. Hermes: Seamless delivery of containerized bioinformatics workflows in hybrid cloud (HTC) environments. SoftwareX, 6 (??):209–216, ???? 2017. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711017300304.

Kazantsev:2019:CRT

[KPTW19] Daniil Kazantsev, Edoardo Pasca, Martin J. Turner, and Philip J. Withers.

CCPi-Regularisation toolkit

for computed tomographic image reconstruction with proximal splitting algorithms. SoftwareX, 9(??): 317–323, January/June 2019. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018301912.

Kramer:2020:GSC

[Krä20] Michel Krämer. GeoRocket:

a scalable and cloud-based data store for big geospatial files. SoftwareX, 11 (??):Article 100409, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711018302036.

Khan:2023:AAM

[KRA23] Abdur Rehman Khan, Umer Rashid, and Naveed Ahmed. AMED: Aggregated multimedia exploratory and discovery search software. SoftwareX, 21(??): ??, February 2023. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023000080.

Keller:2020:SSO

[KRB^{+}20] Johannes Keller, Volker Rath, Johanna Bruckmann, Darius Mottaghy, Christoph Clauser, Andreas Wolf, Ralf Seidler, H. Martin Bücker, and Norbert Klitzsch.

                                        SHEMAT-Suite: an open-

source code for simulating flow, heat and species transport in porous media. SoftwareX, 12(??):Article 100533, July/December 2020. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711020301357.

Krivenko:2022:LPQ

[Kri22] Igor Krivenko. libcommute/

                                        pycommute: a quantum

operator algebra domainspecific language and exact diagonalization toolkit. SoftwareX, 17(??):??, January 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021001722.

Karam:2021:BPS

[KS21] Mokbel Karam and Tony

Saad. BuckinghamPy: a

Python software for dimensional analysis. SoftwareX, 16(??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021001291.

Kang:2023:TPT

[KS23] Yong-Bin Kang and Timos

Sellis. TopicTracker: a

platform for topic trajectory identification and visualisation. SoftwareX, 22 (??):??, May 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023000262.

Kusanovic:2023:SVO

[KSA23] Danilo S. Kusanovic, Elnaz Seylabi, and Domniki

Asimaki. Seismo-VLAB:

an open-source finite element software for seismic meso-scale simulations. SoftwareX, 21(??): ??, February 2023. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022002187.

Kraljic:2018:GCA

[KSFG18] K. Kraljić, L. Strüngmann, E. Fimmel, and M. Gumbel. Genetic Code Analysis Toolkit: a novel tool to explore the coding properties of the genetic code and DNA sequences. SoftwareX, 7 (??):12–14, January/June 2018. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711017300572.

Krishnamani:2019:MCE

[KSP19] Venkatramanan Krishnamani, Mark A. Stamnes, and Robert C. Piper. MALTA: a calculator for estimating the coverage with shRNA, CRISPR, and cDNA libraries. SoftwareX, 9(??):154–160, January/June 2019. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711018301304.

Karam:2020:PPS

[KSS20] Mokbel Karam, James C. Sutherland, and Tony

                                        Saad.  PyModPDE: a

Python software for modified equation analysis. SoftwareX, 12(??):Article 100541, July/December 2020. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020300224.

Kizielewicz:2023:PUL

[KSS23] Bartłomiej Kizielewicz, Andrii Shekhovtsov, and Wojciech Sałabun. pymcdm — the universal library for solving multi-criteria decision-making problems. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271102300064X.

Kulikova:2020:STI

[Kul20] Sofya Kulikova. StimVis:

a tool for interactive computation of the TMSinduced effects over tractography data. SoftwareX, 12(??):Article 100594, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020303071.

Kursa:2021:PHP

[Kur21] Miron B. Kursa. Praznik:

High performance informationbased feature selection. SoftwareX, 16(??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021001138.

Kurbucz:2023:HHD

[Kur23] Marcell T. Kurbucz. hdData360r: a high-dimensional panel data compiler for governance, trade, and competitiveness indicators of World Bank Group platforms. SoftwareX, 21 (??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022002151.

Kulas:2023:ACE

[KW23] Daniel Kulas and Michal R.

                                        Wrobel. AffecTube —

Chrome extension for YouTube video affective annotations. SoftwareX, 23(??):??, July 2023. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023002005.

Kwary:2019:CPI

[Kwa19] Deny A. Kwary. A corpus platform of Indonesian academic language. SoftwareX, 9(??):102–106, January/June 2019. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/

science/article/pii/
S2352711018302413.

Kazantsev:2022:HPS

[KWB22] Daniil Kazantsev, Nicola Wadeson, and Mark Basham. High performance Savu software for fast 3D modelbased iterative reconstruction of large data at Diamond Light Source. SoftwareX, 19(??):??, July 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022000978.

Karzhaubayev:2022:EPS

[KWZ22] Kairzhan Karzhaubayev, Lian-Ping Wang, and Dauren Zhakebayev. An efficient parallel spectral code for 3D periodic flow simulations. SoftwareX, 20 (??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022001625.

Koutsellis:2023:CWB

[KXK^{+}23] Themistoklis Koutsellis, Georgios Xexakis, Konstantinos Koasidis, Natasha Frilingou, Anastasios Karamaneas, Alexandros Nikas, and Haris Doukas. Incognitive: a web-based Python application for fuzzy cognitive map design, simulation, and uncertainty analysis based on the Monte Carlo method. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023002091.

Karaduta:2018:SNT

[KZ18] Oleg Karaduta and Loutfouz Zaman. Shk-9: a new tool in approach of glycoprotein annotation. SoftwareX, 7(??):302–303, January/June 2018. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S235271101830178X.

Li:2023:GAS

[LAA23] Xiaotong Li, Karel Aámek, and Wes Armour. GPU accelerated singular value thresholding. SoftwareX, 23(??):??, July 2023. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023001966.

Lekinwala:2022:PPL

[LB22] Nirav L. Lekinwala and Mani Bhushan. pyGNMF: a Python library for implementation of generalised non-negative matrix factorisation method. SoftwareX, 20(??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022001753.

Lins:2020:XXR

[LBG20] Sergio A. Barcellos Lins, Boris Bremmers, and Giovanni E. Gigante. XIS-MuS — X-ray fluorescence imaging software for multiple samples. SoftwareX, 12 (??):Article 100621, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020303344.

Linsel:2020:GKD

[LBH^{+}20] Adrian Linsel, Kristian Bär, Joshua Haas, Jens Hornung, Matthias D. Greb, and Matthias Hinderer. GeoReVi: a knowledge discovery and data management tool for subsurface characterization. SoftwareX, 12(??):Article 100597, July/December 2020. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020303101.

Luciano:2021:ORP

[LBL^{+}21] Giorgio Luciano, Serena Berretta, Kristian Hovde Liland, Gavin J. Donley, and Simon A. Rogers. oreo: an R package for large amplitude oscillatory analysis. SoftwareX, 15 (??):??, July 2021. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S235271102100087X.

Lara:2021:PJP

[LBT^{+}21] José Daniel Lara, Clayton Barrows, Daniel Thom, Dheepak Krishnamurthy, and Duncan Callaway.

                                        PowerSystems.jl — a

power system data management package for large scale modeling. SoftwareX, 15(??):??, July 2021. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021000765.

Lay:2022:TUA

[LCGA22] Lam M. Lay, Kai-Cheng Chuang, Will Giles, and Justus Adamson. TARDIS: an updated artificial intelligence model to predict linear accelerator machine parameters at treatment delivery. SoftwareX, 19 (??):??, July 2022. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022000929.

Lelievre:2018:FSM

[LCMD^{+}18] Peter G. Lelièvre, Angela E. Carter-McAuslan, Michael W. Dunham, Drew J. Jones, Mariella Nalepa, Chelsea L. Squires, Cassandra J. Tycholiz, Marc A. Vallée, and Colin G. Farquharson. FacetModeller: Software for manual creation, manipulation and analysis of 3D surface-based models. SoftwareX, 7(??):41– 46, January/June 2018. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018300165.

Legleiter:2024:ROS

[LD24] Carl J. Legleiter and Michael Dille. A Robot Operating System (ROS) package for mapping flow fields in rivers via Particle Image Velocimetry (PIV). SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024000827.

Loraamm:2020:PTV

[LDAL20] Rebecca Loraamm, Joni Downs, James Anderson, and David S. Lamb. PyST-Prism: Tools for voxelbased space-time prisms. SoftwareX, 12(??):Article 100499, July/December 2020. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019303309.

Lindner:2020:VPP

[LDM20] Guilherme Lindner, Yann Devaux, and Sanja Miskovic.

                                        VortexFitting: a post-

processing fluid mechanics tool for vortex identification. SoftwareX, 12 (??):Article 100604, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711020303174.

Ladroit:2020:EOS

[LEFSO20] Yoann Ladroit, Pablo C. Escobar-Flores, Alexandre C. G. Schimel, and Richard L. O’Driscoll. ESP3: an open-source software for the quantitative processing of hydroacoustic data. SoftwareX, 12(??):Article 100581, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711020302946.

Lin:2022:XAW

[LEK22] Sibo Lin, Mohamed Elanany, and Motaz Khawaji. XTBDFT: Automated workflow for conformer searching of minima and transition states powered by extended tight binding and density functional theory. SoftwareX, 20(??): ??, December 2022. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022001601.

Lin:2015:MRP

[LF15] Yan-Xia Lin and Mark James

Fielding. MaskDensity14:

an R package for the density approximant of a univariate based on noise multiplied data. SoftwareX, 3–4(??):32–36, December 2015. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711015000163.

Lopez-Fernandez:2024:BNP

[LFGVGDBD24] Aurelio López-Fernández, Francisco A. Gómez-Vela, Jorge Gonzalez-Dominguez, and Parameshachari Bidare-Divakarachari.

bioScience:  a new

Python science library for high-performance computing bioinformatics analytics. SoftwareX, 26(??): ??, May 2024. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024000372.

Liu:2024:GAE

[LFK^{+}24] Lingbo Liu, Xiaokang Fu, Tobias Kötter, Kevin Sturm, Carsten Haubold, Weihe Wendy Guan, Shuming Bao, and Fahui Wang.

Geospatial analytics extension for KNIME. SoftwareX, 25(??):??, February 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023003230.

Li:2023:OSS

[LFP^{+}23] Jianning Li, André Ferreira, Behrus Puladi, Victor Alves, Michael Kamp, Moon Kim, Felix Nensa, Jens Kleesiek, Seyed-Ahmad Ahmadi, and Jan Egger. Open-source skull reconstruction with MONAI. SoftwareX, 23 (??):??, July 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023001280.

Lopez:2020:NHP

[LFR^{+}20] Jose Manuel López, Daniel Feldmann, Markus Rampp, Alberto Vela-Martín, Liang Shi, and Marc Avila. nsCouette — a highperformance code for direct numerical simulations of turbulent Taylor– Couette flow. SoftwareX, 11(??):Article 100395, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019302298.

Lyden:2021:PPM

[LFT21] Andrew Lyden, Graeme Flett, and Paul G. Tuohy. PyLESA: a Python modelling tool for planninglevel Local, integrated, and smart Energy Systems Analysis. SoftwareX, 14 (??):??, June 2021. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021000443.

Lindsay:2023:MNS

[LGG^{+}23] Alexander Lindsay, Guillaume Giudicelli, Peter German, John Peterson, Yaqi Wang, Ramiro Freile, David Andrs, Paolo Balestra, Mauricio Tano, Rui Hu, Ling Zou, Derek Gaston, Cody Permann, and Sebastian Schunert. MOOSE Navier–Stokes module. SoftwareX, 23(??): ??, July 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023001991.

Lindsay:2022:MEM

[LGP^{+}22] Alexander D. Lindsay, Derek R. Gaston, Cody J. Permann, Jason M. Miller, David Andrš, Andrew E. Slaughter, Fande Kong, Joshua Hansel, Robert W. Carlsen, Casey Icenhour, Logan Harbour, Guillaume L. Giudicelli, Roy H. Stogner, Peter German, Jacob Badger, Sudipta Biswas, Leora Chapuis, Christopher Green, Jason Hales, Tianchen Hu, Wen Jiang, Yeon Sang Jung, Christopher Matthews, Yinbin Miao, April Novak, John W. Peterson, Zachary M. Prince, Andrea Rovinelli, Sebastian Schunert, Daniel Schwen, Benjamin W. Spencer, Swetha Veeraraghavan, Antonio Recuero, Dewen Yushu, Yaqi Wang, Andy Wilkins, and Christopher

Wong. 2.0-MOOSE: En-

abling massively parallel multiphysics simulation. SoftwareX, 20(??): ??, December 2022. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022001200.

Lacombe:2022:KNP

[LH22] Francis Lacombe and Jean-Pierre Hickey. Krypton: Nonlinear Parabolized Stability Equation solver for transonic flow in curvilinear coordinates. SoftwareX, 20(??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022001248.

Lim:2023:RPR

[LH23] Youngbin Lim and Sangyul Ha. RufGen: a plugin for rough surface generation in Abaqus/CAE. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023000766.

Lange:2024:TIM

[LHB24] Ferdinand Lange, Tessa Habich, and Sascha Beutel. Think inside the (mk)box — a tool for creating and deploying application-specific Linux images utilized in a SiLA software development process. SoftwareX, 26(??): ??, May 2024. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024000761.

Lee:2018:DDA

[LHCK18] Deokjae Lee, S. Hwang, S. Choi, and B. Kahng. Decremental dynamic algorithm to trace mutually connected clusters. SoftwareX, 7(??):273–280, January/June 2018. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018301717.

Liao:2024:AOI

[LHL^{+}24] Yan Liao, Jialin Hua, Liangqing Luo, Weiying Ping, Xuewen Lu, and Yuansheng Zhong. APRCOIE: an open information extraction system for Chinese. SoftwareX, 26 (??):??, May 2024. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024000207.

Lu:2023:GSG

[LHY^{+}23] Binbin Lu, Yigong Hu, Dongyang Yang, Yong Liu, Liuqi Liao, Zuoyao Yin, Tianyang Xia, Zheyi Dong, Paul Harris, Chris Brunsdon, Lex Comber, and Guanpeng Dong. GWmodelS: a software for geographically weighted models. SoftwareX, 21 (??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022002096.

Liu:2023:MBM

[LHZT23] Guoyang Liu, Janet H. Hsiao, Weidong Zhou, and

Lan Tian. MartMi-BCI:

a Matlab-based real-time motor imagery braincomputer interface platform. SoftwareX, 22(??): ??, May 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023000675.

Liu:2023:RMP

[Liu23]  Yang Liu.  rwl read:

a MATLAB program to identify and correct formatting errors in tree-ring measurement files. SoftwareX, 21(??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022002035.

Liu:2024:EER

[Liu24]  Jun Liu. EBSDPolygonizer:

Enabling realistic microstructural modelling. SoftwareX, 25(??):??, February 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023003187.

Leeming:2020:HEM

[LIZ^{+}20] Michael G. Leeming, Andrew P. Isaac, Luke Zappia, Richard A. J. O’Hair, William A. Donald, and Bernard J. Pope. HiTIME: an efficient model-selection approach for the detection of unknown drug metabolites in LC-MS data. SoftwareX, 12(??):Article 100559, July/December 2020. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711019303796.

Lass:2020:MSP

[LJML20] Jakob Lass, Henrik Jacobsen, Daniel G. Mazzone, and Kim Lefmann. MJOLNIR: a software package for multiplexing neutron spectrometers. SoftwareX, 12(??):Article 100600, July/December 2020. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711020303137.

LeHoux:2021:OOS

[LK21] James Le Houx and Denis Kramer. OpenImpala: OPEN source IMage based PArallisable Linear Algebra solver. SoftwareX, 15 (??):??, July 2021. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021000662.

Lee:2023:VRV

[LKK23] Dasol Lee, Seokgyu Kwon, and Minkyung Kim. ViBA Rad: Visualization and basic analysis tools for radiative cooling. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/

science/article/pii/
S2352711023002583.

Labijak-Kowalska:2024:RJL

[LKK24] Anna Labijak-Kowalska and Miłosz Kadziński. robustDEA: a Java library for robust efficiency analysis. SoftwareX, 26(??): ??, May 2024. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024001201.

Lejeune:2020:FTF

[LKSS20] Emma Lejeune, Alex Khang, Jacob Sansom, and Michael S. Sacks. FM-Track: a fiducial marker tracking software for studying cell mechanics in a three-dimensional environment. SoftwareX, 11 (??):Article 100417, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019303474.

Loiseau:2020:FNG

[LLK^{+}20] Julien Loiseau, Hyun Lim, Mark Alexander Kaltenborn, Oleg Korobkin, Christopher M. Mauney, Irina Sagert, Wesley P. Even, and Benjamin K. Bergen. FleC-SPH: the next generation FleCSIble parallel computational infrastructure for smoothed particle hydrodynamics. SoftwareX, 12 (??):Article 100602, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711020303150.

Luciano:2020:TRP

[LLM^{+}20] G. Luciano, K. H. Liland, M. Marsotto, R. Svoboda, and S. Berretta. takos: an R package for thermal analysis calculations. SoftwareX, 12(??):Article 100637, July/December 2020. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711020303502.

Li:2024:PCF

[LLP^{+}24] Chunfang Li, Junli Lu, Yuchen Pei, Yushi Shen, Yuhang Hu, Yalv Fan, Yuanzhi Tian, Xiaoyu Linghu, Kun Wang, Zhuoqi Shi, and Jiangnan Sun.

                                        PyCinemetrics: Compu-

tational film studies tool based on deep learning and PySide2. SoftwareX, 26 (??):??, May 2024. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711024000578.

Lacroix:2019:MOG

[LMB^{+}19] Pierre Lacroix, Frédéric Moser, Antonio Benvenuti, Thomas Piller, David Jensen, Inga Petersen, Marion Planque, and Nicolas Ray. MapX: an open geospatial platform to manage, analyze and visualize data on natural resources and the environment. SoftwareX, 9 (??):77–84, January/June 2019. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018300177.

Lysko:2023:CCC

[ŁMF^{+}23] Andrzej Łysko, Witold Maćków, Paweł Forczmański, Paweł Terefenko, Andrzej Giza, Jakub Śledziowski, Grzegorz Stepień, and Arkadiusz Tomczak. CCMORPH — Coastal Cliffs Morphology Analysis Toolbox. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023000821.

Langer:2022:GRL

[LMM22] Tristan Langer, Richard Meyes, and Tobias Meisen.

Gideon Replay: a library

to replay interactions in web-applications. SoftwareX, 17(??):??, January 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S2352711021001862.

Leontaris:2018:ANM

[LMN18] Georgios Leontaris and Oswaldo Morales-Nápoles.

                                        ANDURIL — a MAT-

LAB toolbox for ANalysis and Decisions with UnceRtaInty: Learning from expert judgments. SoftwareX, 7(??):313–317, January/June 2018. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711018300608.

Lucchese:2020:REI

[LMN^{+}20] Claudio Lucchese, Cristina Ioana Muntean, Franco Maria Nardini, Raffaele Perego, and Salvatore Trani. RankEval: Evaluation and investigation of ranking models. SoftwareX, 12(??):Article 100614, July/December 2020. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711020303277.

Linner:2016:LBS

[LMS^{+}16] Elisabeth Schold Linnér, Max Morén, Karl-Oskar Smed, Johan Nysjö, and

                                        Robin Strand. LatticeLibrary
                                        and BccFccRaycaster:

Software for processing and viewing 3D data on optimal sampling lattices. SoftwareX, 5(??):

12–15, ???? 2016. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711016000042.

Loizides:2015:IVP

[LNS15] C. Loizides, J. Nagle, and P. Steinberg. Improved version of the PHOBOS Glauber Monte Carlo. SoftwareX, 1–2(??):9–12, September 2015. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711015000047.

Longo:2023:MVT

[LOM^{+}23] Giacomo Longo, Alessandro Orlich, Stefano Musante, Alessio Merlo, and Enrico Russo. MaCySTe: a virtual testbed for maritime cybersecurity. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271102300122X.

Ledee:2023:EOP

[LPG^{+}23] François Lédée, Pierryves Padey, Kyriaki Goulouti, Sébastien Lasvaux, and Didier Beloin-Saint-Pierre. EcoDynElec: Open Python package to create historical profiles of environmental impacts from regional electricity mixes. SoftwareX, 23 (??):??, July 2023. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023001814.

Lim:2024:VVB

[LPL^{+}24] Min Young Lim, Seong Hee Park, Soo-Hyun Lee, Jung Won Yoon, Pyo Min Hong, Hwajung Yoo, Kon-Woo Kwon, Jongwook Jeong, and Youn Kyu Lee. VELCRO: a visualbased programming tool for effortless deep learning model construction. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S235271102400027X.

Lonare:2021:ERP

[LPR21] Gunratan Lonare, Bharat Patil, and Nilesh Raut. edgar: an R package for the U.S. SEC EDGAR retrieval and parsing of corporate filings. SoftwareX, 16(??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021001369.

Li:2024:PPV

[LPS^{+}24] Chunfang Li, Yuchen Pei, Yushi Shen, Junli Lu, Yalv Fan, Xiaoyu Linghu, Yuanzhi Tian, and Kun

Wang.  PyVisVue3D3:

Python visualization from hierarchy tree to call graph. SoftwareX, 26 (??):??, May 2024. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024000608.

Luo:2022:JJB

[LQM22] Mimi Luo, Jiayu Qin, and Gang Mei. juSPH: a Juliabased open-source package of parallel Smoothed Particle Hydrodynamics (SPH) for dam break problems. SoftwareX, 19(??): ??, July 2022. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022000954.

Li:2023:PTP

[LRL23] Ye Li, Hongxiang Ren, and Haijiang Li. PyVT: a toolkit for preprocessing and analysis of vessel spatio-temporal trajectories. SoftwareX, 21 (??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023000122.

Laura:2018:APL

[LRPD18] Jason Laura, Kelvin Rodriguez, Adam C. Paquette, and Evin Dunn. AutoCNet: a Python library for sparse multi-image correspondence identification for planetary data. SoftwareX, 7(??):37–40, January/June 2018. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S235271101830013X.

Liland:2016:FRP

[LS16] Kristian Hovde Liland and

                                        Lars Snipen. fixedTimeEvents:

an R package for the distribution of distances between discrete events in fixed time. SoftwareX, 5 (??):216–226, ???? 2016. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711016300310.

Loverdos:2023:IGD

[LS23] Dimitrios Loverdos and Vasilis Sarhosis. Image2DEM: a geometrical digital twin generator for the detailed structural analysis of existing masonry infrastructure stock. SoftwareX, 22 (??):??, May 2023. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023000195.

Li:2020:OOS

[LSB^{+}20] Ting Li, Lawrence V.

Stanislawski, Tyler Brockmeyer, Shaowen Wang, and Ethan Shavers. OpenCLC: an open-source software tool for similarity assessment of linear hydrographic features. SoftwareX, 11(??):Article 100401, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018302747.

Li:2023:GGW

[LSF^{+}23] Hui Li, Bixia Sun, Yuan Fang, Mengyao Zhang, and Haitao Liu. GWGPM: Generator of wallpaper group patterns for molecular visualization. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023001668.

Leles:2019:NAS

[LSMG19] M. C. R. Leles, J. P. H.

Sansão, L. A. Mozelli, and H. N. Guimarães. A new algorithm in singular spectrum analysis framework: The Overlap-SSA (ov-SSA). SoftwareX, 8 (??):26–32, ???? 2019. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711017300596.

Lehtola:2018:RDL

[LSOM18] Susi Lehtola, Conrad Steigemann, Micael J. T. Oliveira, and Miguel A. L. Marques. Recent developments in libxc — a comprehensive library of functionals for density functional theory. SoftwareX, 7(??):1–5, January/June 2018. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711017300602.

Lee:2022:HSS

[LSPG22] Geon Hee Lee, Ali Sadollah, Sang Ho Park, and Zong Woo Geem. HS-Solver: Spreadsheet based harmony search algorithm solver for various optimization problems. SoftwareX, 20(??): ??, December 2022. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022001807.

Liu:2016:TTS

[LSSK16] Chenhui Liu, Anuj Sharma, Edward Smaglik, and Sirisha Kothuri. TraSER: a traffic signal event-based recorder. SoftwareX, 5 (??):150–155, ???? 2016. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/

science/article/pii/
S2352711016300218.

Lourenco:2024:PFL

[LVCF24] Bruno Lourenço, Cátia Vaz, Miguel E. Coimbra, and Alexandre P. Francisco. phyloDB: a framework for large-scale phylogenetic analysis of sequence based typing data. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024000396.

Lee:2021:ATS

[LVK21] Kong Aik Lee, Ville Vestman, and Tomi Kinnunen. ASVtorch toolkit: Speaker verification with deep neural networks. SoftwareX, 14(??):??, June 2021. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271102100042X.

La:2022:BPR

[LVT^{+}22] Viet-Phuong La, Quan-Hoang Vuong, Trung Tran, Minh-Hoang Nguyen, Manh-Tung Ho, and Manh-Toan Ho. The bayesvl package: an R package for implementing and visualizing Bayesian statistics. SoftwareX, 20(??): ??, December 2022. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022001637.

Li:2023:MEL

[LXJ^{+}23] Linmin Li, Weisen Xu, Bowen Jiang, Xiaojun Li, and Zuchao Zhu. A multiscale Eulerian–Lagrangian cavitating flow solver in OpenFOAM. SoftwareX, 21(??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022002229.

Liang:2023:SQQ

[LXK23] Kang Liang, Zhang Xiukai, and Oleg Krakhmalev.

                                        SLSL-QPSO:  Quantum-

behaved particle swarm optimization with shortlived swarm layers. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023002327.

Liu:2018:PPB

[LYX^{+}18] Benyuan Liu, Bin Yang, Canhua Xu, Junying Xia, Meng Dai, Zhenyu Ji, Fusheng You, Xiuzhen Dong, Xuetao Shi, and Feng Fu. pyEIT: a Python based framework for Electrical Impedance Tomography. SoftwareX, 7(??): 304–308, January/June 2018. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018301407.

Liu:2022:EUF

[LYZ^{+}22] Shun Liu, Junjie Yang, Xianxian Zeng, Haiying Song, Jian Cen, and Weichao Xu. An efficient and user-friendly software tool for ordered multi-class receiver operating characteristic analysis based on Python. SoftwareX, 19 (??):??, July 2022. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022001078.

Laarne:2021:ENL

[LZN21] Petri Laarne, Martha A.

Zaidan, and Tuomo Nieminen. ennemi: Non-linear correlation detection with mutual information. SoftwareX, 14(??):??, June 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021000315.

Montoya-Araque:2022:OSA

[MAAOZMAM22] Exneyder A. Montoya-Araque, A. J. Aparicio-Ortube, David G. Zapata-Medina, and Luis G. Arboleda-Monsalve. An open-source application software to determine the preconsolidation pressure of soils in incremental loading oedometer testing: pySigmaP. SoftwareX, 17(??):??, January 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022000097.

Macleod:2021:GPP

[MAC^{+}21] Duncan M. Macleod, Joseph S. Areeda, Scott B. Coughlin, Thomas J. Massinger, and Alexander L. Urban. GWpy: a Python package for gravitational-wave astrophysics. SoftwareX, 13 (??):Article 100657, January 2021. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021000029.

Manzano:2024:TPP

[MAG24] Martí Manzano, Claudia Ayala, and Cristina Gómez. TrustML: a Python package for computing the trustworthiness of ML models. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711024001110.

Miranda:2023:AMB

[MAGM^{+}23] Juan Carlos Miranda, Jaume Arnó, Jordi Gené-Mola, Spyros Fountas, and Eduard Gregorio.

AKFruitYield: Modular

benchmarking and video analysis software for Azure Kinect cameras for fruit size and fruit yield estimation in apple orchards. SoftwareX, 24 (??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023002443.

Maincon:2022:EJE

[Mai22] Philippe Mainçon. EspyInsideFunction.jl — extracting intermediate results from numerical functions. SoftwareX, 19 (??):??, July 2022. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022001194.

Mustapha:2023:DMW

[MAM^{+}23] Aatila Mustapha, Kadem Abdellah, Lachgar Mohamed, Lamhaddab Khalid, Hrimech Hamid, and Kartit Ali. DLDiagnosis: a mobile and web application for diseases classification using Deep Learning. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271102300184X.

McCann:2022:COS

[MAS^{+}22] Maile P. McCann, Dylan L. Anderson, Christopher R. Sherwood, Brittany Bruder, A. Spicer Bak, and Katherine L.

                                        Brodie. CoastalImageLib:

an open-source Python package for creating common coastal image products. SoftwareX, 20(??): ??, December 2022. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022001339.

Montoya-Araque:2018:PAS

[MASB18] Exneyder A. Montoya-Araque and Ludger O. Suarez-Burgoa. pyBIMstab: Application software for 2D slope stability analysis of block-in-matrix and homogeneous materials. SoftwareX, 7(??):383– 387, January/June 2018. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711018301638.

Milton:2021:OSC

[MB21] Matthew Milton and Andrea Benigni. ORTiS solver codegen: C++ code generation tools for high performance, FPGA-based, real-time simulation of power electronic systems. SoftwareX, 13(??):Article 100660, January 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021000054.

Mattia:2023:PFF

[MB23] Gabriele Proietti Mattia and Roberto Beraldi. P2PFaaS: a framework for FaaS peer-to-peer scheduling and load balancing in Fog and Edge computing. SoftwareX, 21 (??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022002084.

Marcolini:2022:HPL

[MBA^{+}22] Alessia Marcolini, Nicole Bussola, Ernesto Arbitrio, Mohamed Amgad, Giuseppe Jurman, and Cesare Furlanello. histolab: a Python library for reproducible Digital Pathology preprocessing with automated testing. SoftwareX, 20(??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022001558.

Marelli:2022:SFC

[MBC22] Davide Marelli, Simone Bianco, and Gianluigi Ciocca. SfM Flow: a comprehensive toolset for the evaluation of 3D reconstruction pipelines. SoftwareX, 17(??):??, January 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021001692.

Medvet:2020:VSS

[MBDS20] Eric Medvet, Alberto Bartoli, Andrea De Lorenzo, and Stefano Seriani. 2D-VSR-Sim: a simulation tool for the optimization of 2-D voxel-based soft robots. SoftwareX, 12(??):Article 100573, July/December 2020. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711020302867.

Mendes:2020:PIG

[MBF20] Luís Mendes, Alexandre Bernardino, and Rui M. L.

                                        Ferreira. piv-image-generator:

an image generating software package for planar PIV and optical flow benchmarking. SoftwareX, 12(??):Article 100537, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711020300339.

Muraru:2020:GTB

[MBI20] Sebastian Muraru, Jorge S. Burns, and Mariana Ionita. GOPY: a tool for building 2D graphene-based computational models. SoftwareX, 12(??):Article 100586, July/December 2020. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020302995.

Manrique-Bedoya:2022:CCA

[MBM22] Santiago Manrique-Bedoya and Marcelo Marucho. CACPPAF, a COMSOL application to characterize polyelectrolyte properties of actin filaments. SoftwareX, 20(??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022001777.

Michelon:2019:SAM

[MBU^{+}19] Gabriela Karoline Michelon, Claudio Leones Bazzi, Shrinivasa Upadhyaya, Eduardo Godoy de Souza, Paulo Sérgio Graziano Magalhães, Ligia Francielle Borges, Kelyn Schenatto, Ricardo Sobjak, Alan Gavioli, and Nelson Miguel Betzek. Software AgDataBox-Map to precision agriculture management. SoftwareX, 10 (??):Article 100320, July/ December 2019. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S2352711019300275.

Muravyev:2022:TOS

[MBY22] Kirill Muravyev, Andrey Bokovoy, and Konstantin

                                        Yakovlev. tx2 fcnn node:

an open-source ROS compatible tool for monocular depth reconstruction. SoftwareX, 17(??):??, January 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021001837.

Melenka:2018:BCB

[MC18] Garrett W. Melenka and Jason P. Carey. Braid CAM: Braided composite analytical model. SoftwareX, 7(??):23–27, January/June 2018. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711017300705.

Mitura:2024:MIC

[MCBM24] Jakub Mitura, Beata E. Chrapko, and Oliwia Bachanek-Mitura. MedVoxelHD: Improved CUDA-accelerated morphological Hausdorff distances in medical image analysis. SoftwareX, 26(??):??, May 2024. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711024001158.

MachadodeCampos:2020:UUF

[MCC20] Brunno Machado de Campos, Raphael Fernandes Casseb, and Fernando Cendes. UF^{2}C — userfriendly functional connectivity: a neuroimaging toolbox for fMRI processing and analyses. SoftwareX, 11(??):Article 100434, January/June 2020. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019301372.

Martin:2019:QPM

[MCGK19] R. D. Martin, Q. Cai, T. Garrow, and C. Kapahi. QExpy: a Python-3 module to support undergraduate physics laboratories. SoftwareX, 10(??):Article 100273, July/December 2019. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271101930144X.

Mittas:2023:CWB

[MCK^{+}23] Nikolaos Mittas, Fani Chatzopoulou, Efstratios Karagiannidis, Dimitrios Chatzidimitriou, Georgios Sianos, Lefteris Angelis, and Ioannis S. Vizirianakis. CRISSPAC: a webbased platform for predicting the SYNTAX Score and severity of coronary artery disease. SoftwareX, 21(??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023000067.

Molner:2024:SPB

[MCPG24] Antonio Molner, Francisco Carrillo-Perez, and Alberto Guillén. SnapperML: a Python-based framework to improve machine learning operations. SoftwareX, 26(??):??, May 2024. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711024000190.

Morovati:2019:NVF

[MD19] Vahid Morovati and Roozbeh Dargazany. NET v1.0: a framework to simulate permanent damage in elastomers under quasi-static deformations. SoftwareX, 10(??):Article 100229, July/December 2019. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711018300359.

Melchert:2021:PPT

[MD21] Oliver Melchert and Ayhan Demircan. pyGLLE: a Python toolkit for solving the generalized Lugiato– Lefever equation. SoftwareX, 15(??):??, July 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271102100073X.

Melchert:2022:GPG

[MD22] Oliver Melchert and Ayhan Demircan. GNLStools.py: a generalized nonlinear Schrödinger Python module implementing different models of input pulse quantum noise. SoftwareX, 20(??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022001509.

Medina:2022:WSP

[MDG22] Haritz Medina, Oscar Díaz, and Xabier Garmendia. WACline: a software product line to harness heterogeneity in Web annotation. SoftwareX, 18(??):??, June 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022000619.

McCaskey:2018:LHI

[MDL^{+}18] A. J. McCaskey, E. F.

Dumitrescu, D. Liakh, M. Chen, W. Feng, and T. S. Humble. A language and hardware independent approach to quantum–classical computing. SoftwareX, 7(??): 245–254, January/June 2018. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711018300700.

Malard:2022:EDP

[MDR^{+}22] Florian Malard, Laura Danner, Emilie Rouzies, Jesse G. Meyer, Ewen Lescop, and Stéphanie Olivier-Van Stichelen. EpyNN: Educational Python for Neural Networks. SoftwareX, 19(??):??, July 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022000905.

Moorthy:2017:CSP

[ME17] Arun S. Moorthy and Hermann J. Eberl. compuGUT: an in silico platform for simulating intestinal fermentation. SoftwareX, 6 (??):231–236, ???? 2017. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711017300201.

Martinez-Franco:2020:PDS

[MFCSÁM20] Juan Martínez-Franco, Edgar Céspedes-Sabogal, and David Álvarez-Martínez.

                                        PackageCargo:  a de-

cision support tool for the container loading problem with stability. SoftwareX, 12(??):Article 100601, July/December 2020. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020303149.

Mariggio:2020:ELI

[MFCV20] Gregorio Mariggiò, Sebastiano Fichera, Mauro Corrado, and Giulio Ventura. EQP — a 2D/3D library for integration of polynomials times step function. SoftwareX, 12(??):Article 100636, July/December 2020. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020303496.

Miedzinski:2023:ZLT

[MFJ23] Rafał Miedziński and Izabela Fuks-Janczarek. Z-lambda: a terminalbased software for simulation of heat transfer during Z-scan experiments in transparent solids. SoftwareX, 24 (??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023002315.

Meng:2021:MPP

[MFQ^{+}21] Siqin Meng, Zhendong Fu, Jianfei Qin, Xiaobai Ma, Yuqing Li, Lijie Hao, Yuntao Liu, Kai Sun, and Dongfeng Chen.

magcoilcalc: a Python

package for modeling and optimization of axisymmetric magnet coils generating uniform magnetic field for noble gas spinpolarizers. SoftwareX, 16 (??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021001060.

Mariotti:2022:WVB

[MG22] Gianfranco Mariotti and Roberto Giorgi. WebRISC-V: a 32/64-bit RISC-V pipeline simulation tool. SoftwareX, 18(??):??, June 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S235271102200070X.

Mijac:2023:RSF

[MGCS23] Marko Mijac, Antonio Garcia-Cabot, and Vjeran Strahonja. REFRAME — a software framework for managing reactive dependencies in object-oriented applications. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023002674.

Mardan:2023:PPP

[MGFO23] Amir Mardan, Bernard Giroux, and Gabriel Fabien-Ouellet. PyFWI: a Python package for full-waveform inversion and reservoir monitoring. SoftwareX, 22 (??):??, May 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023000808.

Martinez-Galicia:2024:CPM

[MGGHG^{+}24] David Martínez-Galicia, Alejandro Guerra-Hernández, Francisco Grimaldo, Nicandro Cruz-Ramírez, and Xavier Limón. ClassNoise: an R package for modeling, generating, and validating data with class noise. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024001237.

Meijer:2020:LTD

[MGK^{+}20] C. Meijer, M. W. Grootes, Z. Koma, Y. Dzigan, R. Gonçalves, B. Andela, G. van den Oord, E. Ranguelova, N. Renaud, and W. D. Kissling.

Laserchicken — a tool

for distributed feature calculation from massive Li-DAR point cloud datasets. SoftwareX, 12(??):Article 100626, July/December 2020. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020303393.

Miranda:2022:ADS

[MGMAG22] Juan Carlos Miranda, Jordi Gené-Mola, Jaume Arnó, and Eduard Grego-

                                        rio. AKFruitData: a dual

software application for Azure Kinect cameras to acquire and extract informative data in yield tests performed in fruit orchard environments. SoftwareX, 20(??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022001492.

Mateos:2022:LSP

[MHTZ22] Cristian Mateos, Matías Hirsch, Juan Manuel Toloza, and Alejandro

                                        Zunino. LiveDewStream:

a stream processing platform for running in-lab distributed deep learning inferences on smartphone clusters at the edge. SoftwareX, 20(??): ??, December 2022. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022001868.

Michail:2021:JOS

[Mic21] Dimitrios Michail. JHeaps: an open-source library of priority queues. SoftwareX, 16(??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021001370.

Montoya:2021:TTD

[MIHS21] Omar Montoya, Octavio Icasio-Hernández, and Joaquín Salas. TreeTool: a tool for detecting trees and estimating their DBH using forest point clouds. SoftwareX, 16(??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021001485.

Mulansky:2016:PPP

[MK16] Mario Mulansky and Thomas Kreuz. PySpike — a Python library for analyzing spike train synchrony. SoftwareX, 5 (??):178–182, ???? 2016. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711016300255.

Mann:2018:MMI

[MKB^{+}18] Martin Mann, Hans-Peter Kahle, Matthias Beck, Bela Johannes Bender, Heinrich Spiecker, and Rolf Backofen. MICA: Multiple interval-based curve alignment. SoftwareX, 7(??):53–58, January/June 2018. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711018300190.

Meyer:2023:IRS

[MKMK23] Elias Laurin Meyer, Constantin Kumaus, Michal Majka, and Franz Koenig. An interactive R-Shiny app for quickly visualizing a tidy, long dataset with multiple dimensions with an application in clinical trial simulations for platform trials. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023000432.

Miikki:2021:OSC

[MKR^{+}21] Kim Miikki, Alp Karakoç, Mahdi Rafiee, Duck Weon Lee, Jaana Vapaavuori, Jennifer Tersteegen, Laura Lemetti, and Jouni Paltakari. An open-source camera system for experimental measurements. SoftwareX, 14(??):??, June 2021. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021000339.

Mishra:2023:ETL

[MKS^{+}23] Renu Mishra, Inderpreet Kaur, Santosh Sahu, Sandeep Saxena, Nitima Malsa, and Mamta Narwaria. Establishing three layer architecture to improve interoperability in Medicare using smart and strategic API led integration. SoftwareX, 22 (??):??, May 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023000729.

Merkoureas:2023:SWA

[MKT^{+}23] Ilias Merkoureas, Antonia Kaouni, Georgia Theodoropoulou, Alexandros Bousdekis, Athanasios Voulodimos, and Georgios Miaoulis. Smyrida: a web application for process mining and interactive visualization. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023000237.

Manurung:2022:YRD

[MKU22] Auralius Manurung, Lisa Kristiana, and Nur Uddin. YADPF: a reusable deterministic dynamic programming implementation in MATLAB. SoftwareX, 17(??):??, January 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022000152.

Macha:2022:RPR

[MKWS22] Dawid Macha, Michał Kozielski, Lukasz Wróbel, and Marek Sikora. RuleXAI — a package for rulebased explanations of machine learning model. SoftwareX, 20(??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022001273.

Mrena:2024:TTD

[MKZ24] Michal Mrena, Miroslav Kvassay, and Elena Zaitseva. TeDDy: Templated decision diagram library. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711024000864.

Mukha:2018:EPP

[ML18] Timofey Mukha and Mattias Liefvendahl. Eddylicious: a Python package for turbulent inflow generation. SoftwareX, 7(??): 112–114, January/June 2018. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711018300487.

Martin:2020:OSM

[ML20] Andrea Quero Martín and Andrés Díaz Lantada. An open source medical passport based on an Android mobile application and near-field communication. SoftwareX, 11 (??):Article 100492, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019302407.

Mara:2022:EFN

[MLD22] Alexandru Mara, Jefrey Lijffijt, and Tijl De Bie. EvalNE: a framework for network embedding evaluation. SoftwareX, 17(??): ??, January 2022. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022000139.

Moreno-Lumbreras:2023:BFE

[MLGBR23] David Moreno-Lumbreras, Jesus M. Gonzalez-Barahona, and Gregorio Robles. BabiaXR: Facilitating experiments about XR data visualization. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023002832.

Masek:2018:SMT

[MLTF^{+}18] Martin Masek, Chiou Peng Lam, Cameron Tranthim-Fryer, Bas Jansen, and Kevin Baptist. Sleep monitor: a tool for monitoring and categorical scoring of lying position using 3D camera data. SoftwareX, 7(??):341–346, January/June 2018. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018301079.

Milani:2016:GVI

[MM16] G. Milani and F. Milani. GURU v2.0: an interactive Graphical User interface to fit rheometer curves in Han’s model for rubber vulcanization. SoftwareX, 5(??):62–66, ???? 2016. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711016300061.

McLean:2021:OOS

[MMCKK21] Thomas McLean, Christian Málaga-Chuquitaype, Nicos Kalapodis, and Georgios Kampas. OpenArch: an open-source package for determining the minimumthickness of arches under seismic loads. SoftwareX, 15(??):??, July 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021000686.

Maage:2024:SMT

[MMD24] Ingrid Måge, Josipa Matić, and Katinka Riiser Dankel. SEC2MWD: a MATLAB toolbox for derivation of molecular weight distributions from size exclusion chromatography. SoftwareX, 25(??):??, February 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023002984.

Micieli:2019:NTP

[MMG19] Davide Micieli, Triestino Minniti, and Giuseppe Gorini. NeuTomPy toolbox, a Python package for tomographic data processing and reconstruction. SoftwareX, 9(??):260–264, January/June 2019. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018302103.

Moore:2019:EHO

[MMH19] Jacob L. Moore, Nathaniel R.

Morgan, and Mark F. Horstemeyer. ELEMENTS: a high-order finite element library in C++. SoftwareX, 10(??):Article 100257, July/December 2019. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/

science/article/pii/
S235271101930113X.

Morra:2020:SCG

[MML20] Lia Morra, Francesco Manigrasso, and Fabrizio Lamberti. SoccER: Computer graphics meets sports analytics for soccer event recognition. SoftwareX, 12(??):Article 100612, July/December 2020. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020303253.

Montgomery:2023:COS

[MML23] David Montgomery, Federico Municchi, and Karin Leiderman. clotFoam: an open-source framework to simulate blood clot formation under arterial flow. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023001796.

Maniscalco:2020:AAS

[MMS20] Umberto Maniscalco, Antonio Messina, and Pietro Storniolo. ASS4HR — an artificial somatosensory system for a humanoid robot. The ROS package. SoftwareX, 11 (??):Article 100501, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019300536.

Miretti:2021:DDD

[MMS21] Federico Miretti, Daniela Misul, and Ezio Spessa. DynaProg: Deterministic Dynamic Programming solver for finite horizon multi-stage decision problems. SoftwareX, 14(??): ??, June 2021. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021000352.

Meng:2024:RSM

[MMT24] Han Meng, Gang Mei, and John C. Tipper. RealModel: a system for modeling and visualizing sedimentary rocks. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024001225.

Moshiri:2020:TMS

[Mos20] N. Moshiri. TreeSwift: a massively scalable Python tree package. SoftwareX, 11(??):Article 100436, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019300767.

Marcon:2017:PZO

[MP17] Yann Marcon and Au-

                                        tun Purser. PAPARA(ZZ)I:

an open-source software interface for annotating photographs of the deepsea. SoftwareX, 6(??): 63–68, ???? 2017. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711017300067.

Montanari:2018:OCC

[MP18] Mattia Montanari and Nik Petrinic. OpenGJK for C, C# and Matlab: Reliable solutions to distance queries between convex bodies in three-dimensional space. SoftwareX, 7(??):352–355, January/June 2018. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711018300591.

Medeiros:2020:MSP

[MP20] Marcelo S. Medeiros and Evandro Parente. MicroFEA 1.0 — a software package for Finite Element Analysis of functionally graded materials. SoftwareX, 11 (??):Article 100481, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/

science/article/pii/
S2352711019302481.

Mathur:2019:NSP

[MPAK19] Parul Mathur, Mauricio D. Perez, Robin Augustine, and Dhanesh G. Kurup. NDECOAX: a software package for nondestructive evaluation of stratified dielectric media. SoftwareX, 9(??):187– 192, January/June 2019. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018302280.

Melchor-Placencia:2021:OPC

[MPMC21] Carlos Melchor-Placencia and Christian Málaga-Chuquitaype. OpenMoist: a Python code for transient moisture transfer analysis. SoftwareX, 15 (??):??, July 2021. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021000571.

Michels:2024:CCM

[MPX^{+}24] Alexander C. Michels, Anand Padmanabhan, Zimo Xiao, Mit Kotak, Furqan Baig, and Shaowen

Wang. CyberGIS-Compute:

Middleware for democratizing scalable geocomputation. SoftwareX, 26 (??):??, May 2024. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711024000621.

Moina-Rivera:2023:VQM

[MRGAGP23]  Wilmer  Moina-Rivera,

Juan Gutiérrez-Aguado, and Miguel Garcia-Pineda. Video quality metrics toolkit: an open source software to assess video quality. SoftwareX, 23 (??):??, July 2023. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023001231.

Melchert:2019:OAS

[MRMD19] O. Melchert, B. Roth, U. Morgner, and A. Demircan. OptFROG — Analytic signal spectrograms with optimized time-frequency resolution. SoftwareX, 10(??):Article 100275, July/December 2019. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711019301323.

Mahmoudi:2023:OSW

[MSDP23] Amin Mahmoudi, Mahsa Sadeghi, Xiaopeng Deng, and Pengcheng Pan. OPA Solver: a web-based software for Ordinal Priority Approach in multiple criteria decision analysis using JavaScript. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271102300242X.

Maurer:2024:PPT

[MSK24] Valentin J. Maurer, Marc Siggel, and Jan Kosinski. PyTME (Python Template Matching Engine): a fast, flexible, and multi-purpose template matching library for cryogenic electron microscopy data. SoftwareX, 25(??):??, February 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024000074.

Marrero:2023:DTG

[MSLH23] Alejandro Marrero, Eduardo Segredo, Coromoto León, and Emma Hart. DIGNEA: a tool to generate diverse and discriminatory instance suites for optimisation domains. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023000511.

Marticorena-Sanchez:2024:UDA

[MSLNSMOG24] Raúl Marticorena-Sánchez, Carlos López-Nozal, Ana Serrano-Mamolar, and Alicia Olivares-Gil. UBUMonitor: Desktop application for visual e-learning student clustering with Moodle. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711024000980.

Mercier:2023:BPE

[MSS^{+}23] Arthur Mercier, Nikita Smolin, Oliver Sihlovec, Stefanos Koffas, and Stjepan Picek. Backdoor Pony: Evaluating backdoor attacks and defenses in different domains. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023000833.

Menvouta:2023:DPP

[MSV23] Emmanuel Jordy Menvouta, Sven Serneels, and Tim Verdonck. direpack: a Python 3 package for state-of-the-art statistical dimensionality reduction methods. SoftwareX, 21 (??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S235271102200200X.

Mateo:2019:STF

[MT19] Carlos Mateo and Juan Antonio Talavera. Short-Time Fourier Transform with the Window Size Fixed in the Frequency Domain (STFT-FD): Implementation. SoftwareX, 8(??):5–8, ???? 2019. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711017300638.

Millinger:2022:MCG

[MTJ^{+}22] Markus Millinger, Philip Tafarte, Matthias Jordan, Frazer Musonda, Katrina Chan, Kathleen Meisel, and Danial Esmaeili Aliabadi. A model for costand greenhouse gas optimal material and energy allocation of biomass and hydrogen. SoftwareX, 20 (??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022001820.

Meng:2018:MPP

[MTPHH18] Siqin Meng, Rasmus Toft-Petersen, Lijie Hao, and Klaus Habicht. multiflexxlib: a Python package for data reduction and visualization for the cold-neutron multi energy wide angle analyzer MultiFLEXX. SoftwareX, 7(??):309–312, January/June 2018. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018301055.

McCubbine:2018:GPC

[MTS^{+}18] Jack McCubbine, Fabio Caratori Tontini, Vaughan Stagpoole, Euan Smith, and Grant O’Brien. Gsolve, a Python computer program with a graphical user interface to transform relative gravity survey measurements to absolute gravity values and gravity anomalies. SoftwareX, 7(??):129–137, January/ June 2018. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711018300566.

M:2024:FLL

[MTS^{+}24] Veena K. M., Venkat Tummala, Yuva Sahith Varma Sangaraju, Marreddy Sai Vineel Reddy, Preetham Kumar, Veena Mayya, Uma Kulkarni, Sulatha Bhandary, and Shailaja S. FFA-Lens: Lesion detection tool for chronic ocular diseases in fluorescein angiography images. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711024000177.

Muller:2023:PSC

[Mül23] Matthias Müller. PyblioNet — software for the creation, visualization and analysis of bibliometric networks. SoftwareX, 24 (??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023002613.

Mentaschi:2019:AAE

[MVBF19] Lorenzo Mentaschi, Michalis Vousdoukas, Giovanni Besio, and Luc Feyen. alpha-BetaLab: Automatic estimation of subscale transparencies for the Unresolved Obstacl’s Source Term in ocean wave modelling. SoftwareX, 9(??): 1–6, January/June 2019. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018301456.

Maertens:2024:DEC

[MVG^{+}24] Rien Maertens, Maarten Van Neyghem, Maxiem Geldhof, Charlotte Van Petegem, Niko Strijbol, Peter Dawyndt, and Bart Mesuere. Discovering and exploring cases of educational source code plagiarism with Dolos. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024001262.

Mirz:2019:DDP

[MVRM19] Markus Mirz, Steffen Vogel, Georg Reinke, and Antonello Monti. DPsim — a dynamic phasor realtime simulator for power systems. SoftwareX, 10 (??):Article 100253, July/ December 2019. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711018302760.

Ma:2023:TTF

[MW23] Jie Ma and Qianlian

                                        Wang. TransRNM — a

traffic flow and demand prediction tool for transportation network with ridesharing services. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023001644.

Merzky:2015:SSA

[MWJ15] Andre Merzky, Ole Weidner, and Shantenu Jha. SAGA: a standardized access layer to heterogeneous Distributed Computing Infrastructure. SoftwareX, 1–2(??):1–2, September 2015. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711015000023.

Mauro:2020:KTC

[MWM20] Yihong Z. Mauro, Collin J.

Wilkinson, and John C.

Mauro. KineticPy: a

tool to calculate long-time kinetics in energy landscapes with broken ergodicity. SoftwareX, 11 (??):Article 100393, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019303516.

Motoyama:2022:MLM

[MYKT22] Yuichi Motoyama, Kazuyoshi Yoshimi, Takeo Kato, and

Synge Todo. MateriApps
LIVE! and MateriApps

Installer: Environment for starting and scaling up materials science simulations. SoftwareX, 20 (??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022001285.

Ma:2023:PPT

[MZL^{+}23] Jinfeng Ma, Hua Zheng, Ruonan Li, Kaifeng Rao, Yanzheng Yang, and

Weifeng Li. PyVecContour:

a Python toolkit for vectorized isosurface mapping. SoftwareX, 21(??): ??, February 2023. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023000134.

Mertel:2021:HGA

[MZSH21] Adam Mertel, David Zbíral, Zdenek Stachon, and Hana Horínková. Historical geocoding assistant. SoftwareX, 14(??): ??, June 2021. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021000273.

Naito:2017:BBT

[Nai17] O. Naito. A browserbased tool for conversion between Fortran NAMELIST and XML/ HTML. SoftwareX, 6 (??):19–24, ???? 2017. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S235271101630036X.

Nasser:2020:PMT

[Nas20] Mohamed M. S. Nasser.

                                        PlgCirMap: a MATLAB

toolbox for computing conformal mappings from polygonal multiply connected domains onto circular domains. SoftwareX, 11 (??):Article 100464, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/

science/article/pii/
S2352711019303073.

Nobre:2019:NMA

[NBCC19] Ricardo Nobre, João Bispo, Tiago Carvalho, and João M. P. Cardoso. Nonio — modular automatic compiler phase selection and ordering specialization framework for modern compilers. SoftwareX, 10 (??):Article 100238, July/ December 2019. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018301614.

Nejahi:2019:GGO

[NBM^{+}19] Younes Nejahi, Mohammad Soroush Barhaghi, Jason Mick, Brock Jackman, Kamel Rushaidat, Yuanzhe Li, Loren Schwiebert, and Jeffrey Potoff. GOMC: GPU Optimized Monte Carlo for the simulation of phase equilibria and physical properties of complex fluids. SoftwareX, 9(??):20– 27, January/June 2019. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018301171. See

update [NBS^{+}21].

Nejahi:2021:UGG

[NBS^{+}21] Younes Nejahi, Mohammad Soroush Barhaghi, Gregory Schwing, Loren Schwiebert, and Jeffrey Potoff. Update 2.70 to “GOMC: GPU Optimized Monte Carlo for the simulation of phase equilibria and physical properties of complex fluids”. SoftwareX, 13(??):Article 100627, January 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271102030340X. See

[NBM^{+}19].

Niembro-Cecena:2024:SME

[NCGLRR24] José A. Niembro-Ceceña, Roberto A. Gómez-Loenzo, and Juvenal Rodríguez-

Reséndiz. SoftCtrlDC-M:

Embedded control software for brushed direct current motors. SoftwareX, 25(??):??, February 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024000141.

Nordstrand:2022:EST

[ND22] Johan Nordstrand and Joydeep Dutta. ELC: Software and tutorial for finite-element modeling of electrochemical desalination. SoftwareX, 20(??): ??, December 2022. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022001522.

Neverov:2017:XGA

[Nev17] V. S. Neverov. XaNSoNS:

GPU-accelerated simulator of diffraction patterns of nanoparticles. SoftwareX, 6(??):54–62, ???? 2017. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711017300043.

Noleto-Filho:2021:NFO

[NFASC21] Eurico Mesquita Noleto-Filho, Ronaldo Angelini, Jeroen Steenbeek, and Adriana Rosa Carvalho. New, flexible and opensource fisheries self-reporting

app: the Shiny4SelfReport.

SoftwareX, 16(??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021001254.

Nasinski:2023:HIC

[NG23] Maciej Nasiński and Krzysztof Gajowniczek. Handling an inconsistently coded categorical variable in a longitudinal dataset with cat2cat. SoftwareX, 24 (??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/

science/article/pii/
S2352711023002212.

Neic:2020:AIB

[NGK^{+}20] Aurel Neic, Matthias A. F. Gsell, Elias Karabelas, Anton J. Prassl, and Gernot Plank. Automating image-based mesh generation and manipulation tasks in cardiac modeling workflows using Meshtool. SoftwareX, 11 (??):Article 100454, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271101930295X.

Nieto:2024:PSC

[NGOT24] Maria T. Alvela Nieto, Hannes Gelbhardt, Jan-Hendrik Ohlendorf, and Klaus-Dieter Thoben. ProQualAI: a software to compute qualitative and quantitative product features in process industries using deep learning image processing algorithms. SoftwareX, 25 (??):??, February 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024000025.

Najafabadi:2023:APP

[NHR23] Mohsen Yoosefzadeh Najafabadi, Ali Heidari, and Istvan Rajcan. AllInOne

Pre-processing: a com-

prehensive preprocessing framework in plant field phenotyping. SoftwareX, 23(??):??, July 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023001607.

Nakajima:2023:DSS

[NIP^{+}23] Evandro Alves Nakajima, Antonio Augusto Ignacio, Suzan Kelly Borges Piovesan, Helton Jose Alves, and Carlos Eduardo

Borba.  DRMSimulator:

Software for simulation of the Dry Reforming of Methane (DRM) process. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023001875.

Nisius:2020:BCC

[Nis20] Richard Nisius. BLUE:

Combining correlated estimates of physics observables within ROOT using the Best Linear Unbiased Estimate method. SoftwareX, 11(??):Article 100468, January/June 2020. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020300169.

Nikezic:2016:CPP

[NIY16] D. Nikezic, M. Ivanovic, and K. N. Yu. A computer program TRACK P for studying proton tracks in PADC detectors. SoftwareX, 5(??):67–73, ???? 2016. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711016300073.

Nguyen:2018:AGB

[NK18] Duc T. Nguyen and Blair

                                        Kaneshiro. AudExpCreator:

a GUI-based Matlab tool for designing and creating auditory experiments with the Psychophysi’s Toolbox. SoftwareX, 7(??): 328–334, January/June 2018. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711018301985.

Nebiu:2020:SIF

[NK20] Dhurata Nebiu and Hiqmet Kamberaj. Symbolic

                                        Information Flow Measurement

(SIFM): a software for measurement of information flow using symbolic analysis. SoftwareX, 11 (??):Article 100470, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S235271101930305X.

Narajewski:2021:TRP

[NKHZ21] Michał Narajewski, Jens Kley-Holsteg, and Florian Ziel. tsrobprep — an R package for robust preprocessing of time series data. SoftwareX, 16 (??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021001084.

Niskanen:2023:CPB

[NL23] Matti Niskanen and Timo Lähivaara. COMPOSTI: a Python-based program for seismic trans-dimensional inversion. SoftwareX, 21 (??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022002163.

Ninni:2020:MSM

[NM20] Davide Ninni and Miguel A.

Mendez. MODULO: a software for Multiscale Proper Orthogonal Decomposition of data. SoftwareX, 12 (??):Article 100622, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020303356.

Navarrete:2018:BWB

[NMLM18] Claudio Bustos Navarrete, María Gabriela Morales Malverde, Pedro Salcedo Lagos, and Alejandro Díaz Mujica. Buhos: a webbased systematic literature review management software. SoftwareX, 7(??): 360–372, January/June 2018. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018300293.

Najafi:2023:HAB

[NMT23] Amin Najafi and Masoud Masih-Tehrani. Hybrid adaptive battery parameter estimation approach for equivalent circuit model toolbox. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023002303.

Nardini:2023:ESG

[NNDS23] Giovanni Nardini, Alessandro Noferi, Pietro Ducange, and Giovanni Stea. Exploiting Simu5G for generating datasets for training and testing AI models for 5G/6G network applications. SoftwareX, 21(??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271102300016X.

Nar:2019:RCF

[NOÖÇ19] Fatih Nar, Osman Erman Okman, Atilla Özgür, and Müjdat Çetin. RmSAT-CFAR: Fast and accurate target detection in radar images. SoftwareX, 8(??):39–42, ???? 2019. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271101730047X.

Nabil:2016:PIF

[NR16] Mahdi Nabil and Alexan-

der S. Rattner. interThermalPhaseChangeFoam

— a framework for twophase flow simulations with thermally driven phase change. SoftwareX, 5(??):211–215, ???? 2016. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711016300309.

Nowakowski:2023:FIS

[NR23] Piotr Nowakowski and Przemysław Rokita. FieldView: an interactive software tool for exploration of three-dimensional vector fields. SoftwareX, 22 (??):??, May 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023001024.

Nino-Ruiz:2023:ADA

[NRC23] Elías D. Nino-Ruiz and Randy Consuegra. AMLCS-DA: a data assimilation package in Python for Atmospheric General Circulation Models. SoftwareX, 22(??):??, May 2023. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023000705.

Nabhani:2022:TOO

[NS22] Abbas Nabhani and Hanne K. Sjølie. TreeSim: an object-oriented individual tree simulator and 3D visualization tool in Python. SoftwareX, 20 (??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S235271102200139X.

Nordstrom:2016:GCA

[NSLD16] Johan Nordström, Johan Schött, Inka L. M. Locht, and Igor Di Marco. A GPU code for analytic continuation through a sampling method. SoftwareX, 5(??):171–177, ???? 2016. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711016300243.

Nieto:2021:BAA

[NSO21] Marcos Nieto, Orti Senderos, and Oihana Otaegui. Boosting AI applications: Labeling format for complex datasets. SoftwareX, 13(??):Article 100653, January 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020303666.

Nagaraju:2019:PSR

[NSS^{+}19] Vidhyashree Nagaraju, Venkateswaran Shekar, Joshua Steakelum, Melanie Luperon, Ying Shi, and Lance Fiondella. Practical software reliability engineering with the Software Failure and Reliability Assessment Tool (SFRAT). SoftwareX, 10(??):Article 100357, July/December 2019. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019302134.

Ozcevik:2023:MSM

[ÖA23] Yusuf Özçevik and Os-

man Altay. MetricHunter:

a software metric dataset generator utilizing Source-Monitor upon public GitHub repositories. SoftwareX, 23 (??):??, July 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S2352711023001954.

Olufsen:2020:OST

[OAF20] Sindre Nordmark Olufsen, Marius Endre Andersen, and Egil Fagerholt. μDIC: an open-source toolkit for digital image correlation. SoftwareX, 11 (??):Article 100391, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711019301967.

Orozco-Arroyave:2019:N

[OAVCVB^{+}19] Juan Rafael Orozco-Arroyave, Juan Camilo Vásquez-Correa, Jesús Francisco Vargas-Bonilla, R. Arora, N. Dehak, P. S. Nidadavolu, H. Christensen, F. Rudzicz, M. Yancheva, H. Chinaei, A. Vann, N. Vogler, T. Bocklet, M. Cernak, J. Hannink, and Elmar Nöth. NeuroSpeech. SoftwareX, 8 (??):69–70, ???? 2019. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711017300341.

Overberg:2023:EOS

[OBWM23] Florian A. Overberg, Philipp C. Böttcher, Dirk Witthaut, and Simon Morgenthaler. Emipy: an opensource Python-based tool to analyze industrial emissions in Europe. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023001541.

Ortner:2020:MFP

[OC20] Michael Ortner and Lucas Gabriel Coliado Ban-

deira.  Magpylib: a

free Python package for magnetic field computation. SoftwareX, 11 (??):Article 100466, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020300170.

Oh:2022:DRS

[OCK22] Hee-Seok Oh, Guebin Choi, and Donghoh Kim. Decomp2d: an R solution for image decomposition. SoftwareX, 19(??):??, July 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022001108.

Oldford:2023:FVA

[OCK^{+}23] Greig Oldford, David Cote, Dan G. Kehler, Gabrielle R. Riefesel, and Yolanda F. Wiersma.

FIPEX v10.4: an ArcGIS

Desktop Add-in for assessing impacts of fish passage barriers and longitudinal connectivity of rivers. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023001656.

ODriscoll:2021:SSC

[O’D21] Benjamin O’Driscoll. SCRAMBLE: Sweep Comparison Research Application for Multiple Back-Gated FieLd Effect measurements of graphene field effect transistors. SoftwareX, 15 (??):??, July 2021. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021000819.

Oikonomou:2022:CNO

[ODE^{+}22] George Oikonomou, Simon Duquennoy, Atis Elsts, Joakim Eriksson, Yasuyuki Tanaka, and Nicolas Tsiftes. The Contiki-NG open source operating system for next generation IoT devices. SoftwareX, 18(??):??, June 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022000620.

Ortin:2021:CPA

[OE21] Francisco Ortin and Javier Escalada. Cnerator: a Python application for the controlled stochastic generation of standard C source code. SoftwareX, 15 (??):??, July 2021. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271102100056X.

Ortin:2022:SPL

[OGPSQ22] Francisco Ortin, Miguel Garcia, Baltasar Garcia Perez-Schofield, and Jose Quiroga. The StaDyn programming language. SoftwareX, 20(??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022001297.

Ohnishi:2021:GGC

[Ohn21] Seiki Ohnishi. Gxsview:

Geometry and cross section viewer for calculating radiation transport. SoftwareX, 14(??):??, June 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021000261.

Oishi:2016:BOS

[OHO16] A. Christopher Oishi, David A. Hawthorne, and

Ram Oren. Baseliner:

an open-source, interactive tool for processing sap flux data from thermal dissipation probes. SoftwareX, 5 (??):134–138, ???? 2016. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711016300188.

Oosterbeek:2022:STA

[OJ22] Reece N. Oosterbeek and Jonathan R. T. Jeffers.

                                        StrutSurf: a tool for

analysis of strut morphology and surface roughness in additively manufactured lattices. SoftwareX, 18(??):??, June 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022000383.

Oh:2023:MVO

[OKJ23] Hongseok Oh, Hyunsoo Kim, and Hyerin Jo. MeaSSUre: I–V: Open software for transistor characterization using sourcemeter units. SoftwareX, 21(??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023000146.

Orella:2021:HTA

[OLRLB21] Michael Julian Orella, McLain Evan Leonard, Yuriy Román-Leshkov, and Fikile Richard Brushett. High-throughput analysis of contact angle goniometry data using DropPy.

SoftwareX, 14(??):??, June 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021000108.

Ocampo-Pineda:2022:TJE

[OPPZ22] Mario Ocampo-Pineda, Roberto Posenato, and Francesca Zerbato. TimeAwareBPMN-js: an editor and temporal verification tool for Time-Aware BPMN processes. SoftwareX, 17(??): ??, January 2022. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021001734.

Ocaya:2016:CLP

[OT16] R. O. Ocaya and J. J.

Terblans. C-language package for standalone embedded atom method molecular dynamics simulations of fcc structures. SoftwareX, 5(??): 101–106, ???? 2016. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711016300139.

Owoyele:2022:MPM

[OTdMP22] Babajide Owoyele, James Trujillo, Gerard de Melo,

and Wim Pouw. Masked-Piper:

Masking personal identities in visual recordings while preserving multimodal information. SoftwareX, 20(??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022001546.

Osca:2024:SSO

[OV24] Javier Osca and Jiri Vala. SOQCS: a stochastic optical quantum circuit simulator. SoftwareX, 25 (??):??, February 2024. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023002996.

Omidi:2021:PPM

[OYW^{+}21] Parsa Omidi, Lawrence C. M. Yip, Hui Wang, Mamadou Diop, and Jeffrey J. L. Carson. PhaseWare: Phase map retrieval for fringe projection profilometry and off-axis digital holographic interferometry. SoftwareX, 13(??): Article 100652, January 2021. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711020303654.

Ozbasaran:2017:SVS

[Ozb17] Hakan Ozbasaran. solveTruss v1.0: Static, global buckling and frequency analysis of 2D and 3D trusses with Mathematica. SoftwareX, 6 (??):128–134, ???? 2017. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271101730016X.

Ozkurt:2023:PUF

[Ozk23] Nalan Ozkurt. PrepAnnECG:

a user friendly MATLAB ECG preprocessing and annotation GUI for health professionals. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023002856.

Pachouly:2022:STC

[PAK22] Jalaj Pachouly, Swati Ahirrao, and Ketan Kotecha. SDPTool: a tool for creating datasets and software defect predictions. SoftwareX, 18(??):??, June 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022000358.

Park:2022:PPC

[Par22] Jin Seok Park. pyBaram:

Parallel compressible flow solver in high-performance Python for teaching and research. SoftwareX, 20 (??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S235271102200190X.

Pasca:2022:PPP

[PARS22] Dag Pasquale Pasca, Angelo Aloisio, Marco Martino Rosso, and Stefanos Sotiropoulos. PyOMA and PyOMA GUI: a Python module and software for Operational Modal Analysis. SoftwareX, 20(??): ??, December 2022. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022001340.

Powers:2021:MOS

[PBL^{+}21] Connor Powers, Lindsay Bassman, Thomas M. Linker, Ken ichi Nomura, Sahil Gulania, Rajiv K. Kalia, Aiichiro Nakano, and Priya Vashishta. MIS-TIQS: an open-source software for performing quantum dynamics simulations on quantum computers. SoftwareX, 14(??):??, June 2021. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021000418.

Pathak:2022:CMU

[PBM22] Chintan Pathak, David Beck, and Don MacKenzie. ChargEVal — a multi-user framework for simulating and analysing charging station deployment scenarios using agent-based modelling. SoftwareX, 20 (??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022001728.

Pawlowicz:2024:HTC

[PBS^{+}24] Mateusz Pawlowicz, Bartosz Balis, Adam Szymkiewicz, Jirka Šimunek, Anna Gumuła-Kawecka, and Beata Jaworska-Szulc. HMSE: a tool for coupling MOD-FLOW and HYDRUS-1D computer programs. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024000517.

Pancino:2022:GKB

[PBSB22] Niccolò Pancino, Pietro Bongini, Franco Scarselli, and Monica Bianchini. GNNkeras: a Keras-based library for Graph Neural Networks and homogeneous and heterogeneous graph processing. SoftwareX, 18(??):??, June 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022000486.

Perret:2019:HHG

[PCC^{+}19] B. Perret, G. Chierchia, J. Cousty, S. J. F.

Guimarães, Y. Kenmochi, and L. Najman. Higra: Hierarchical graph analysis. SoftwareX, 10 (??):Article 100335, July/ December 2019. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S235271101930247X.

Peixoto:2023:CGM

[PCdJSdFR^{+}23] João Paulo Just Peixoto, Daniel G. Costa, Washington de J. S. da Franca Rocha, Paulo Portugal, and Francisco Vasques. CityZones: a geospatial multi-tier software tool to compute urban risk zones. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S235271102300105X.

Pereira:2022:PTP

[PCL22] Fábio Pereira, Paul Crocker, and Valderi R. Q. Leithardt. PADRES: Tool for PrivAcy, Data REgulation and Security. SoftwareX, 17(??):??, January 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021001515.

Paniego:2024:BMO

[PCPC24] Sergio Paniego, Roberto Calvo-Palomino, and JoséMaría

Cañas. Behavior metrics:

an open-source assessment tool for autonomous driving tasks. SoftwareX, 26 (??):??, May 2024. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024000736.

Peng:2016:TRP

[PDH16] Jun Peng, ZhiBao Dong, and FengQing Han. tgcd: an R package for analyzing thermoluminescence glow curves. SoftwareX, 5 (??):107–111, ???? 2016. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711016300140.

Pizarro:2022:VVS

[PDM22] Alonso Pizarro, Silvano F.

Dal Sasso, and Salvatore Manfreda. VISION: VIdeo StabilisatION using automatic features selection for image velocimetry analysis in rivers. SoftwareX, 19 (??):??, July 2022. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022001054.

Pintor:2022:SSE

[PDS^{+}22] Maura Pintor, Luca Demetrio, Angelo Sotgiu, Marco Melis, Ambra Demontis, and Battista Biggio. secml: Secure and explainable machine learning in Python. SoftwareX, 18(??):??, June 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022000656.

Predein:2017:CSP

[PDTG17] Peter A. Predein, Anna A. Dobrynina, Tsyren A. Tubanov, and Eugeny I. German. CodaNorm: a software package for the body-wave attenuation calculation by the codanormalization method. SoftwareX, 6(??):25–29, ???? 2017. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711016300413.

Petukhova:2022:TPP

[PF22] Alina Petukhova and Nuno Fachada. TextCL: a Python package for NLP preprocessing tasks. SoftwareX, 19(??):??, July 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022000802.

Petrovic:2023:KTT

[PF23] Filip Petrovic and Jirí Filipovic. Kernel Tuning Toolkit. SoftwareX, 22 (??):??, May 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271102300081X.

Pankhurst:2018:ETD

[PFC^{+}18] M. J. Pankhurst, R. Fowler, L. Courtois, S. Nonni, F. Zuddas, R. C. Atwood, G. R. Davis, and P. D. Lee. Enabling threedimensional densitometric measurements using laboratory source X-ray micro-computed tomography. SoftwareX, 7(??):115– 121, January/June 2018. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018300360.

Patterson:2019:IOS

[PFJM19] Zachary Patterson, Kyle Fitzsimmons, Stewart Jackson, and Takeshi Mukai.

Itinerum:  the open

smartphone travel survey platform. SoftwareX, 10 (??):Article 100230, July/ December 2019. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018300980.

Palatella:2018:ENA

[PG18] Luigi Palatella and Fabio Grasso. The EKF-AUS-NL algorithm implemented without the linear tangent model and in presence of parametric model error. SoftwareX, 7 (??):28–33, January/June 2018. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018300025.

Permann:2020:MEM

[PGA^{+}20] Cody J. Permann, Derek R. Gaston, David Andrš, Robert W. Carlsen, Fande Kong, Alexander D. Lindsay, Jason M. Miller, John W. Peterson, Andrew E. Slaughter, Roy H. Stogner, and Richard C. Martineau. MOOSE: Enabling massively parallel multiphysics simulation. SoftwareX, 11 (??):Article 100430, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019302973.

Pato:2017:GNC

[PI17] Miguel Pato and Fabio Iocco. galkin: a new compilation of Milky Way rotation curve data. SoftwareX, 6(??):48–53, ???? 2017. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711016300437.

Plietzsch:2022:PJE

[PKA^{+}22] Anton Plietzsch, Raphael Kogler, Sabine Auer, Julia Merino, Asier Gil de Muro, Jan Liße, Christina Vogel, and Frank Hellmann.

PowerDynamics.jl — an

experimentally validated open-source package for the dynamical analysis of power grids. SoftwareX, 17(??):??, January 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021001345.

Penkov:2020:XRC

[PKKQ20] Oleksiy V. Penkov, Igor A.

Kopylets, Mahdi Khadem, and Tianzuo Qin. X-Ray Calc: a software for the simulation of X-ray reflectivity. SoftwareX, 12 (??):Article 100528, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019303681.

Palaseanu-Lovejoy:2023:IOS

[PL23] Monica Palaseanu-Lovejoy.

iBluff: an open-source R package for geomorphic analysis of coastal bluffs/cliffs. SoftwareX, 21(??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S2352711023000213.

Petitjean:2024:MPE

[PLC^{+}24] Quentin Petitjean, Silène Lartigue, Mélina Cointe, Nicolas Ris, and Vincent Calcagno. MoveR: an R package for easy processing and analysis of animal video-tracking data. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711024000451.

Platas-Lopez:2023:DRP

[PLGHG^{+}23] Alejandro Platas-López, Alejandro Guerra-Hernández, Francisco Grimaldo, Nicandro Cruz-Ramírez, Efrén Mezura-Montes, and Marcela Quiroz-Castellanos. dplbnDE: an R package for discriminative parameter learning of Bayesian Networks by Differential Evolution. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023001383.

Petkovic:2023:CDT

[PLK^{+}23] Matej Petković, Jurica Levatić, Dragi Kocev, Martin Breskvar, and Saso Dzeroski. CLUSplus: a decision tree-based framework for predicting structured outputs. SoftwareX, 24 (??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023002224.

Pho:2023:RSO

[PM23] Patrick Pho and Alexander V. Mantzaris. reg-sgc: an open-source software for regularized Simple Graph Convolution. SoftwareX, 21(??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022002114.

Perez-Mercado:2023:CCA

[PMBM^{+}23] Rubén Pérez-Mercado, Antonio Balderas, Andrés Muñoz, Juan Francisco Cabrera, Manuel Palomo-Duarte, and Juan Manuel

Dodero.  ChatbotSQL:

Conversational agent to support relational database query language learning. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023000420.

Pachon:2022:SIP

[PMFÁM22] Juan Carlos Pachón, Juan Martínez-Franco, and David Álvarez-Martínez. SIC: an intelligent packing system with industry-grade features. SoftwareX, 20 (??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022001595.

Polanco-Martinez:2023:NRP

[PMLM23] Josué M. Polanco-Martínez and José L. López-Martínez.

                                        NonParRolCor: an R pack-

age for estimating rolling correlation for two regular time series. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023000493.

Pereira:2016:PFV

[PMM16] G. Pereira, M. McGugan, and L. P. Mikkelsen.

                                        FBG SiMul V1.0: Fibre

Bragg grating signal simulation tool for finite element method models. SoftwareX, 5(??):156–162, ???? 2016. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S235271101630022X.

Pallonetto:2019:SSC

[PMMF19] Fabiano Pallonetto, Eleni Mangina, Federico Milano, and Donal P. Finn. SimApi, a smartgrid cosimulation software platform for benchmarking building control algorithms. SoftwareX, 9(??): 271–281, January/June 2019. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018300463.

Paprotny:2020:BMT

[PMNWR20] Dominik Paprotny, Oswaldo Morales-Nápoles, Daniël T. H. Worm, and Elisa Ragno. BANSHEE — a MATLAB toolbox for non-parametric Bayesian networks. SoftwareX, 12 (??):Article 100588, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020303010.

Pina-Martins:2016:NMS

[PMP16] F. Pina-Martins and O. S.

Paulo. NCBI Mass Sequence Downloader — large dataset downloading made easy. SoftwareX, 5(??):74–79, ???? 2016. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711016300085.

Palaniappan:2023:ZEI

[PMS23] Arunkumar Palaniappan, Ramaswamy Muthiah, and Murugesh Tiruchi Sundaram. ZigBee enabled IoT based intelligent lane control system for autonomous agricultural electric vehicle application. SoftwareX, 23 (??):??, July 2023. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S235271102300208X.

Prince:2024:MOM

[PMY^{+}24] Zachary M. Prince, Lynn Munday, Dewen Yushu, Max Nezdyur, and Murthy Guddati. MOOSE optimization module: Physicsconstrained optimization. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711024001250.

Picone:2021:WGP

[PMZ21] Marco Picone, Marco Mamei, and Franco Zambonelli. WLDT: a general purpose library to build IoT digital twins. SoftwareX, 13(??):Article 100661, January 2021. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021000066.

Pujol-Nadal:2023:ORT

[PNC23] Ramón Pujol-Nadal and Gabriel Cardona. OTSunWebApp: a ray tracing web application for the analysis of concentrating solarthermal and photovoltaic solar cells. SoftwareX, 23 (??):??, July 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023001450.

Pilia:2021:EOS

[PNL^{+}21] Nicolas Pilia, Claudia Nagel, Gustavo Lenis, Silvia Becker, Olaf Dössel, and Axel Loewe. ECGdeli: an open source ECG delineation toolbox for MAT-LAB. SoftwareX, 13(??): Article 100639, January 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020303526.

Pistone:2019:NFA

[PO19] L. Pistone and M. Onorato. nlchains: a fast and accurate time integration of 1-D nonlinear chains on GPUs. SoftwareX, 10 (??):Article 100255, July/ December 2019. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019300585.

Portioli:2023:DOB

[Por23] Francesco P. A. Porti-

oli. DynABlock 2D: an

optimization-based MAT-LAB application for rocking dynamics, nonlinear static and limit analysis of masonry block structures. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023001504.

Posenato:2022:CTJ

[Pos22] Roberto Posenato. CSTNU Tool: a Java library for checking temporal networks. SoftwareX, 17(??): ??, January 2022. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021001564.

Pop:2019:NNA

[PP19] Cristian Pop and Alexandru Popa. NewsCompare — a novel application for detecting news influence in a country. SoftwareX, 10(??):Article 100305, July/December 2019. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S235271101930175X.

Praschl:2021:IFI

[PPBZ21] Christoph Praschl, Andreas Pointner, David Baumgartner, and Gerald Adam Zwettler. Imaging framework: an interoperable and extendable connector for imagerelated Java frameworks. SoftwareX, 16(??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021001357.

Popova:2017:OSS

[PPRE17] Olga Popova, Boris Popov, Dmitry Romanov, and Marina Evseeva. Optimel: Software for selecting the optimal method. SoftwareX, 6(??):225–230, ???? 2017. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711017300316.

Perinelli:2019:NPI

[PR19] Alessio Perinelli and Leonardo

Ricci. NetOnZeroDXC: a

package for the identification of networks out of multivariate time series via zero-delay crosscorrelation. SoftwareX, 10 (??):Article 100316, July/ December 2019. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019301335.

Pagano:2023:SSE

[PRC^{+}23] Francesco Pagano, Andrea Romdhana, Davide Caputo, Luca Verderame, and Alessio Merlo. SEBASTiAn: a static and extensible black-box application security testing tool for iOS and Android applications. SoftwareX, 23 (??):??, July 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023001449.

Pham:2019:FHM

[PRSS19] Hung Tien Pham, Wolfram Rühaak, Valerian Schuster, and Ingo Sass. Fully hydro-mechanical coupled plug-in (SUB+) in FEFLOW for analysis of land subsidence due to groundwater extraction. SoftwareX, 9 (??):15–19, January/June 2019. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018301286.

Petschke:2018:DLS

[PS18a] Danny Petschke and Torsten E. M. Staab. DLT-PulseGenerator: a library for the simulation of lifetime spectra based on detector-output pulses. SoftwareX, 7(??):122–128, January/June 2018. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018300530.

Petschke:2018:UVDa

[PS18b] Danny Petschke and Torsten E. M. Staab. Update (v1.1) to DLTPulseGenerator: a library for the simulation of lifetime spectra based on detector-output pulses. SoftwareX, 7(??): 171–173, January/June 2018. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018300694.

Petschke:2018:UVDb

[PS18c] Danny Petschke and Torsten E. M. Staab. Update (v1.2) to DLTPulseGenerator: a library for the simulation of lifetime spectra based on detector-output pulses. SoftwareX, 7(??): 259–262, January/June 2018. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018301092.

Petschke:2019:DSA

[PS19a] Danny Petschke and Torsten E. M. Staab. DDRS4PALS: a software for the acquisition and simulation of lifetime spectra using the DRS4 evaluation board. SoftwareX, 10(??):Article 100261, July/December 2019. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S2352711019300676.

Petschke:2019:UVD

[PS19b] Danny Petschke and Torsten E. M. Staab. Update (v1.3) to DLTPulseGenerator: a library for the simulation of lifetime spectra based on detector-output pulses. SoftwareX, 9(??): 183–186, January/June 2019. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S235271101930038X.

Plewe:2019:PTC

[PS19c] Kaden Plewe and Amanda D. Smith. PECT: a tool for computing the temporal and spatial variation of externalities related to power generation in the United States. SoftwareX, 9 (??):61–67, January/June 2019. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711018300475.

Pagano:2022:EEM

[PS22] Davide Pagano and Lorenzo Sostero. EmiR: Evolutionary minimization for R. SoftwareX, 18(??):??, June 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022000590.

Panayiotou:2022:CEU

[PTS22] Konstantinos Panayiotou, Emmanouil Tsardoulias, and Andreas L. Symeonidis. Commlib: an easyto-use communication library for Cyber-Physical Systems. SoftwareX, 19 (??):??, July 2022. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022001091.

Panagiotidis:2020:HPH

[PTT20]  Kyriakos  Panagiotidis,

Naoum Tsolakis, and Alexander Tsigkas. HC-4-PM: a heterarchical communication framework for project management. SoftwareX, 12(??):Article 100557, July/December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019303243.

Price:2022:HEH

[PZ22] James Price and Marianne

Zeyringer. highRES-Europe:

the high spatial and temporal Resolution Electricity System model for Europe. SoftwareX, 17(??): ??, January 2022. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022000164.

Podvesovskii:2021:DMC

[PZKK21] Aleksandr Podvesovskii, Alena Zakharova, Dmitriy Korostelyov, and Alek-

                                        sandr Kuzin. DecisionMaster:

a multi-criteria decision support system with ability to combine different decision rules. SoftwareX, 16(??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021001114.

Qin:2023:SDP

[QTL23] Xianan Qin, Zhiwen Tu,

                                        and Hao Liu. SBTAnalyzer:

a data processing software package for singlebacterium tracking microscopy at material surfaces. SoftwareX, 22(??): ??, May 2023. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023000730.

Ragni:2017:MCM

[Rag17] Matteo Ragni. Mr.CAS — a minimalistic (pure) Ruby CAS for fast prototyping and code generation. SoftwareX, 6 (??):124–127, ???? 2017. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711017300146.

Robinet:2020:OTC

[RAL^{+}20] Florent Robinet, Nicolas Arnaud, Nicolas Leroy, Andrew Lundgren, Duncan Macleod, and Jessica McIver. Omicron: a tool to characterize transient noise in gravitational-wave detectors. SoftwareX, 12 (??):Article 100620, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020303332.

Ramos:2022:OOS

[RAL22] Andrés Ramos, Erik F. Alvarez, and Sara Lumbreras. OpenTEPES: Opensource Transmission and Generation Expansion Planning. SoftwareX, 18(??): ??, June 2022. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271102200053X.

Raskovalov:2019:ASN

[Ras19a] Anton A. Raskovalov. az-TotMD: Software for nonconstant force field molecular dynamics. SoftwareX, 10(??):Article 100233, July/ December 2019. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018301389.

Rodriguez:2019:FMP

[RAS19b] Miguel A. Rodriguez, Christoph M. Augustin, and Shawn C. Shadden. FEniCS mechanics: a package for continuum mechanics simulations. SoftwareX, 9(??): 107–111, January/June 2019. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711018300979.

Rasmussen:2020:PLR

[Ras20] Leandro Lima Rasmussen. UnBlocks^{gen}: a Python library for 3D rock mass generation and analysis. SoftwareX, 12(??):Article 100577, July/December 2020. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711020302909.

Raston:2021:HPC

[Ras21] Paul L. Raston. HeNDS: a program for calculating average Helium NanoDroplet Sizes. SoftwareX, 14(??):??, June 2021. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021000480.

Robinson:2017:PBM

[RB17] Martin Robinson and Maria Bruna. Particlebased and meshless methods with Aboria. SoftwareX, 6(??):165–171, ???? 2017. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711017300250.

Ramos:2024:SKB

[RBQ^{+}24] Leo Ramos, Mike Bermeo, Andrés Quelal, Anthony Crespo, Gabriela Cajamarca, and Francklin Rivas-Echeverría. SustainMeter: a knowledge-based system for assessing sustainability of SMEs in the fashion industry. SoftwareX, 25(??):??, February 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023003072.

Ramasawmy:2020:EMT

[RCT20] Danny R. Ramasawmy, Ben T. Cox, and Bradley E.

Treeby. ElasticMatrix:

a MATLAB toolbox for anisotropic elastic wave propagation in layered media. SoftwareX, 11 (??):Article 100397, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019303048.

Ramos-Carreno:2023:DDC

[RCT23] Carlos Ramos-Carreño and José L. Torrecilla. dcor: Distance correlation and energy statistics in Python. SoftwareX, 22 (??):??, May 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023000225.

Rodenberg:2021:FPC

[RDH^{+}21] Benjamin Rodenberg, Ishaan Desai, Richard Hertrich, Alexander Jaust, and Benjamin Uekermann.

FEniCS-preCICE: Cou-

pling FEniCS to other simulation software. SoftwareX, 16(??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021001072.

Rettenmaier:2019:LBA

[RDO^{+}19] Daniel Rettenmaier, Daniel Deising, Yun Ouedraogo, Erion Gjonaj, Herbert De Gersem, Dieter Bothe, Cameron Tropea, and Holger Marschall. Load balanced 2D and 3D adaptive mesh refinement in Open-FOAM. SoftwareX, 10 (??):Article 100317, July/ December 2019. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018301699.

Reinartz:2022:PPP

[RE22] Christopher Reinartz and Thomas T. Enevoldsen. pyTEP: a Python package for interactive simulations of the Tennessee Eastman process. SoftwareX, 18(??):??, June 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022000449.

Reis:2017:FFB

[REFB17] Marcelo S. Reis, Gustavo Estrela, Carlos Eduardo Ferreira, and Junior Barrera. featsel: a framework for benchmarking of feature selection algorithms and cost functions. SoftwareX, 6 (??):185–192, ???? 2017. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711017300286.

Reininghaus:2023:CCL

[Rei23] Maximilian Reininghaus.

cnpy++: a C++17 library for reading and writing .npy/.npz files. SoftwareX, 21(??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023000201.

Raptis:2023:CAA

[REK^{+}23] Emmanuel K. Raptis, Konstantinos Englezos, Orfeas Kypris, Marios Krestenitis, Athanasios Ch. Kapoutsis, Konstantinos Ioannidis, Stefanos Vrochidis, and Elias B. Kosmatopoulos. CoFly: an automated, AI-based open-source platform for UAV precision agriculture applications. SoftwareX, 23(??):??, July 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023001103.

Robles-Enciso:2023:SPE

[RERES23] Alberto Robles-Enciso, Ricardo Robles-Enciso, and Antonio F. Skarmeta.

Sim-PowerCS: an extensi-

ble and simplified opensource energy simulator. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023001632.

Rettl:2024:CCC

[RFP^{+}24] Matthias Rettl, Siegfried Martin Frankl, Martin Pletz, Markus Tauscher, and Clara Schuecker. ConForce: Computation of configurational forces for FEM results. SoftwareX, 26 (??):??, May 2024. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271102400089X.

Rosa-Gallardo:2023:NRR

[RGdlTQ^{+}23] Daniel J. Rosa-Gallardo, Juan Carlos de la Torre, Fernando M. Quintana, Juan P. Dominguez-Morales, and Fernando Perez-Peña. NESIM-RT: a real-time distributed spiking neural network simulator. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023000456.

Robertson:2022:RPP

[RGP^{+}22] Hayden Robertson, Isaac J.

Gresham, Stuart W. Prescott, Grant B. Webber, Erica J. Wanless, and Andrew Nelson. refellips: a Python package for the analysis of variable angle spectroscopic ellipsometry data. SoftwareX, 20 (??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022001431.

Radke:2023:SOS

[RHR^{+}23] Karl Ludger Radke, Janina Hußmann, Lena Röwer, Dirk Voit, Jens Frahm, Gerald Antoch, Dirk Klee, Frank Pillekamp, and Hans-Jörg Wittsack.

                                        ``shortCardiac'' — An

open-source framework for fast and standardized assessment of cardiac function. SoftwareX, 23(??): ??, July 2023. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023001498.

Richard:2019:IAN

[Ric19] Celine Richard. Introducing article numbering to SoftwareX. SoftwareX, 10 (??):Article 100374, July/ December 2019. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711019303759.

Redondo-Iglesias:2023:DDA

[RIHVP23] Eduardo Redondo-Iglesias, Marwan Hassini, Pascal Venet, and Serge Pelissier. DATTES: Data analysis tools for tests on energy storage. SoftwareX, 24 (??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023002807.

Roszkowska:2024:IPI

[RJDKJ24] Ewa Roszkowska, Bartłomiej Jefmański, Andrzej Dudek, and Marta Kusterka-Jefmańska. IFMCDM: an R package for intuitionistic fuzzy multi-criteria decision making methods. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271102400092X.

Renaud:2020:IMS

[RJH^{+}20] Nicolas Renaud, Yong Jung, Vasant Honavar, Cunliang Geng, Alexandre M. J. J. Bonvin, and Li C. Xue. iScore: an MPI supported software for ranking protein-protein docking models based on a random walk graph kernel and support vector machines. SoftwareX, 11 (??):Article 100462, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019303061.

Renders:2023:IOS

[RJN^{+}23] Jens Renders, Ben Jeurissen, Anh-Tuan Nguyen, Jan De Beenhouwer, and Jan Sijbers. ImWIP: Opensource image warping toolbox with adjoints and derivatives. SoftwareX, 24 (??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S2352711023002200.

Rose:2019:PSB

[RK19] Michael E. Rose and John R. Kitchin. pybliometrics: Scriptable bibliometrics using a Python interface to Scopus. SoftwareX, 10 (??):Article 100263, July/ December 2019. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711019300573.

Ruszczak:2023:OOT

[RKA^{+}23] Bogdan Ruszczak, Krzysztof Kotowski, Jacek Andrzejewski, Christoph Haskamp, and Jakub Nalepa. OXI: an online tool for visualization and annotation of satellite time series data. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023001723.

Randrianasoa:2021:ABE

[RKDP21] Jimmy Francky Randrianasoa, Camille Kurtz, Éric Desjardin, and Nicolas Passat. AGAT: Building and evaluating binary partition trees for image segmentation. SoftwareX, 16(??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/

science/article/pii/
S2352711021001308.

Rudolph-Lilith:2019:CMT

[RL19] Michelle Rudolph-Lilith.

ChessY: a Mathematica toolbox for the generation, visualization and analysis of positional chess graphs. SoftwareX, 9 (??):39–43, January/June 2019. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018301687.

Rodriguez:2021:SSO

[RLF^{+}21] Néstor Rodríguez, David López, Alberto Fernández, Salvador García, and Francisco Herrera. SOUL: Scala Oversampling and Undersampling Library for imbalance classification. SoftwareX, 15(??):??, July 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021000868.

Ruff:2023:PCV

[RLJS23] Zachary J. Ruff, Damon B. Lesmeister, Julianna M. A. Jenkins, and Christopher M. Sul-

livan. PNW-Cnet v4: Au-

tomated species identification for passive acoustic monitoring. SoftwareX, 23 (??):??, July 2023. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023001693.

Rahman:2018:QCT

[RLK18] Md Mushfiqur Rahman, Yu Lei, and Georgios Kalantzis. QALMA: a computational toolkit for the analysis of quality protocols for medical linear accelerators in radiation therapy. SoftwareX, 7(??): 101–106, January/June 2018. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711018300311.

Ribeiro:2024:SBM

[RLM^{+}24] Laura Cozzi Ribeiro, Matheus Pereira Libório, Hasheem Mannan, Sandro Laudares, Petr Iakovlevich Ekel, Douglas Alexandre Gomes Vieira, and Cristiane Neri Nobre. Software for building and measuring the quality of composite indicators using ordered weighted averaging: So-called S-CI-OWA. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711024000311.

Ribeiro:2021:TMS

[RLN21] Jorge Ribeiro, Pedro Lima, and Francisco Nunes.

                                        Trial Monitor:  Scaf-

folding personalised Web dashboards for Human-Computer Interaction field trials. SoftwareX, 16 (??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271102100145X.

Righolt:2018:SCP

[RMM18] Christiaan H. Righolt, Barret A. Monchka, and Salaheddin M. Mahmud. From source code to publication: Code Diary, an automatic documentation parser for SAS. SoftwareX, 7(??):222–225, January/June 2018. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018300669.

Raj:2023:FPT

[RMM23] Surbhi Raj, Jimson Mathew, and Arijit Mondal. FDT: a Python toolkit for fake image and video detection. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023000912.

Rossi:2021:NTD

[RMMG21] Riccardo Rossi, Andrea Murari, Luca Martellucci, and Pasquale Gaudio.

NetCausality: a time-

delayed neural network tool for causality detection and analysis. SoftwareX, 15(??):??, July 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021000893.

Reyes:2024:SPL

[RMV^{+}24] Axel Reyes, Marcelo Mendoza, Camila Vera, Francesca Lucchini, Jan Dimter, Felipe Gutiérrez, Naim Bro, Hans Lobel, and Ariel

Reyes. SpatialCluster:

a Python library for urban clustering. SoftwareX, 26 (??):??, May 2024. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024001109.

Ruben:2021:MST

[RNG^{+}21] Lucas Dominguez Ruben, Kensuke Naito, Ronald Roger Gutierrez, Ricardo Szupiany, and Jorge Darwin Abad. Meander Statistics Toolbox (MStaT): a toolbox for geometry characterization of bends in large meandering channels. SoftwareX, 14(??): ??, June 2021. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021000194.

Rahman:2021:FFE

[RNK21] Md Motiur Rahman, Tahmina Tasnim Nahar, and Dookie Kim. FeView: Finite element model (FEM) visualization and post-processing tool for OpenSees. SoftwareX, 15 (??):??, July 2021. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021000789. See

corrigendum [RNK22].

Rahman:2022:CFF

[RNK22] Md Motiur Rahman, Tahmina Tasnim Nahar, and Dookie Kim. Corrigendum to “FeView: Finite element model (FEM) visualization and post-processing tool for OpenSees” [SoftwareX 15 (2021) 100751]. SoftwareX, 18(??):??, June 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022000401. See

[RNK21].

Ranathunga:2017:MPP

[RNR17] D. Ranathunga, H. Nguyen, and M. Roughan. MGtoolkit: a Python package for implementing metagraphs. SoftwareX, 6(??):85–90, ???? 2017. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711017300080.

Roald:2023:MLC

[Roa23] Marie Roald. MatCoupLy: Learning coupled matrix factorizations with Python. SoftwareX, 21 (??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022002102.

Roomi:2022:OIM

[ROMH22] Muhammad M. Roomi, Wen Shei Ong, Daisuke Mashima, and Suhail S. M.

                                        Hussain. OpenPLC61850:

an IEC 61850 MMS compatible open source PLC for smart grid research. SoftwareX, 17(??):??, January 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S235271102100162X.

Roughan:2019:PSS

[Rou19] Matthew Roughan. Practically surreal: Surreal arithmetic in Julia. SoftwareX, 9(??):293–298, January/June 2019. CO-DEN ???? ISSN 2352-

                                        7110.  URL https://
                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711018302152.

Rodrigues:2020:NDA

[RPV^{+}20] Marco S. Rodrigues, Rui M. S. Pereira, Mikhail I. Vasilevskiy, Joel Borges, and Filipe Vaz. NANOP-TICS: In-depth analysis of NANomaterials for OPTI-Cal localized surface plasmon resonance Sensing. SoftwareX, 12(??):Article 100522, July/December 2020. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271102030042X.

Rubensson:2022:CTM

[RRKA22] Emanuel H. Rubensson, Elias Rudberg, Anastasia Kruchinina, and Anton G. Artemov. The Chunks and Tasks Matrix Library. SoftwareX, 19(??):??, July 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022000991.

Rohlfs:2018:WTD

[RRS18] Wilko Rohlfs, Manuel Rietz, and Benoit Scheid. WaveMaker: the threedimensional wave simulation tool for falling liquid films. SoftwareX, 7(??): 211–216, January/June 2018. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271101830075X.

Rudberg:2018:EOS

[RRSK18] Elias Rudberg, Emanuel H. Rubensson, Paweł Sałek, and Anastasia Kruchinina. Ergo: an open-source program for linear-scaling electronic structure calculations. SoftwareX, 7(??): 107–111, January/June 2018. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711018300402.

Raskovalov:2022:AMD

[RS22] Anton Raskovalov and Platon Surkov. azTotMD 2.0: Molecular dynamics with the radiative thermostat and temperaturedependent force field (CUDA version). SoftwareX, 17 (??):??, January 2022. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022000127.

Razeghi:2020:CIM

[RSLL^{+}20] Orod Razeghi, José Alonso Solís-Lemus, Angela W. C. Lee, Rashed Karim, Cesare Corrado, Caroline H. Roney, Adelaide de Vecchi, and Steven A. Niederer. CemrgApp: an interactive medical imaging application with image processing, computer vision, and machine learning toolkits for cardiovascular research. SoftwareX, 12 (??):Article 100570, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020302831.

Roy:2020:POS

[RSMW20] M. J. Roy, N. Stoyanov, R. J. Moat, and P. J. Withers. pyCM: an open-source computational framework for residual stress analysis employing the contour method. SoftwareX, 11 (??):Article 100458, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019303668.

Rongen:2020:UAA

[RtHLMN20] Guus Rongen, Cornelis Marcel Pieter ’t Hart, Georgios Leontaris, and Oswaldo Morales-Nápoles. Update (1.2) to ANDURIL and ANDURYL: Performance improvements and a graphical user interface. SoftwareX, 12(??):Article 100497, July/December 2020. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020301837.

Ruby:2016:SFO

[Rub16] Michael Ruby. SpectraFox: a free open-source data management and analysis tool for scanning probe microscopy and spectroscopy. SoftwareX, 5(??):25–30, ???? 2016. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711016300012.

Ravasi:2020:PLO

[RV20] Matteo Ravasi and Ivan Vasconcelos. PyLops — a linear-operator Python library for scalable algebra and optimization. SoftwareX, 11(??):Article 100361, January/June 2020. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711019301086.

Razakh:2021:PPI

[RWJ^{+}21] Taufeq Mohammed Razakh, Beibei Wang, Shane Jackson, Rajiv K. Kalia, Aiichiro Nakano, Ken ichi Nomura, and Priya Vashishta. PND: Physicsinformed neural-network software for molecular dynamics applications. SoftwareX, 15(??):??, July 2021. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/

science/article/pii/
S2352711021000972.

Ren:2023:MMS

[RXZ^{+}23] Weijie Ren, Wenjia Xie, Ye Zhang, Hang Yu, and Zhengyu Tian. MSAT: Matrix stability analysis tool for shock-capturing schemes. SoftwareX, 24 (??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023002625.

Soderstrom:2017:SOC

[SA17] Ken Soderstrom and Ali Alalawi. Software for objective comparison of vocal acoustic features over weeks of audio recording:

KLFromRecordingDays. Soft-

wareX, 6(??):267–270, ???? 2017. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711017300523.

Siddharth:2020:RPI

[SA20] R. Siddharth and G. Aghila.

RandPro — a practical implementation of random projection-based feature extraction for high dimensional multivariate data analysis in R. SoftwareX, 12(??):Article 100629, July/December 2020. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S2352711020303423.

Silva:2018:HFP

[SAA18] D. J. Silva, J. S. Amaral, and V. S. Amaral. Heatrapy: a flexible Python framework for computing dynamic heat transfer processes involving caloric effects in 1.5D systems. SoftwareX, 7(??): 373–382, January/June 2018. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711018301298.

Schwarz:2023:OSE

[SAA^{+}23a] Johannes Schwarz, Kristjan Axelsson, Daniel Anheuer, Martin Richter, Johanna Adam, Martin Heinrich, and Rüdiger Schwarze. An OpenFOAM solver for the extended Navier–Stokes equations. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023000742.

Sola:2023:ATS

[SAA^{+}23b] Fernando Sola, Daniel Ayala, Rafael Ayala, Inma Hernández, Carlos R. Rivero, and David Ruiz. AYNEXT — tools for streamlining the evaluation of link prediction techniques. SoftwareX, 23 (??):??, July 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271102300170X.

Sleem:2020:PPB

[SABEh20] Ahmed Sleem, Mohamed Abdel-Baset, and Ibrahim El-henawy. PyIVNS: a Python based tool for interval-valued neutrosophic operations and normalization. SoftwareX, 12 (??):Article 100632, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020303459.

Sani:2021:RTD

[SAC^{+}21] Laura Sani, Michele Amoretti, Stefano Cagnoni, Monica Mordonini, and Riccardo Pecori. ReSS: a tool for discovering relevant sets in complex systems. SoftwareX, 14(??):??, June 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021000388.

Stanziola:2023:JWO

[SACT23] Antonio Stanziola, Simon R. Arridge, Ben T. Cox, and Bradley E. Treeby. j-Wave: an opensource differentiable wave simulator. SoftwareX, 22 (??):??, May 2023. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023000341.

Sandoval:2022:IJI

[SAD22] Steven Sandoval, Hasan Alshammari, and Mamta Dalal. ISA.jl: Instantaneous spectral analysis in Julia. SoftwareX, 20 (??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022001571.

Soto:2023:RRP

[SADU23] Jonatan Soto, Johan Aparicio, Aquiles Darghan, and Milan Oldrich Ur-

                                        ban.  RankspeQ: an

R package platform for genotype characterization and performance-ranking based on MultispeQ measurements. SoftwareX, 24 (??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023002406.

Salles:2016:BPB

[Sal16] T. Salles. Badlands: a parallel basin and landscape dynamics model. SoftwareX, 5(??):190–194, ???? 2016. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711016300279.

Studiawan:2022:DFT

[SAS^{+}22] Hudan Studiawan, Tohari Ahmad, Bagus J. Santoso, Ary M. Shiddiqi, and Baskoro A. Pratomo.

DroneTimeline: Foren-

sic timeline analysis for drones. SoftwareX, 20 (??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271102200173X.

Stocco:2021:ASG

[SB21] Davide Stocco and Enrico Bertolazzi. Acme: a small 3D geometry library. SoftwareX, 16 (??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021001266.

Soni:2022:DDS

[SB22] Yash Soni and Terry Brett.

Data driven simulations of infectious diseases: Exploring facial recognition approach in predicting the infection severity. SoftwareX, 18(??):??, June 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S2352711022000474.

Shanmugam:2017:EAI

[SBCK17] Janaki Shanmugam, Konstantin B. Borisenko, Yu-Jen Chou, and Angus I.

                                        Kirkland. eRDF Analyser:

an interactive GUI for electron reduced density function analysis. SoftwareX, 6 (??):179–184, ???? 2017. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711017300249.

Stephens:2023:SSM

[SBJ^{+}23] Victoria B. Stephens, Joshua Bedwell, Alex J. Josephson, Keturah Oldham, and David O. Lignell. SootLib: a soot model library for combustion simulation. SoftwareX, 22 (??):??, May 2023. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023000717.

Smyl:2019:OMP

[SBL19] Danny Smyl, Sven Bossuyt, and Dong Liu. Open-QSEI: a MATLAB package for quasi static elasticity imaging. SoftwareX, 9 (??):73–76, January/June 2019. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/

science/article/pii/
S2352711018300256.

Syrmos:2022:MMS

[SBT22] Evangelos Syrmos, Dimitrios Bechtsis, and Naoum Tsolakis. MIROR: a middleware software tool for interfacing mobile industrial robots with optimization routing algorithms. SoftwareX, 17(??):??, January 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021001552.

Stanghellini:2017:SCS

[SC17] Giuseppe Stanghellini and Gabriela Carrara. Segy-change: the Swiss Army knife for the SEG-Y files. SoftwareX, 6(??):36–41, ???? 2017. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711017300031.

Speranza:2019:RNO

[SC19] Giorgio Speranza and Roberto Canteri. RxpsG: a new open project for photoelectron and electron spectroscopy data processing. SoftwareX, 10 (??):Article 100282, July/ December 2019. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019300378.

Salmon:2023:PAM

[SC23] Fabien Salmon and Ludovic Chatellier. pyMeshFOAM: Automated meshing for CFD and fluid-structure simulations. SoftwareX, 23 (??):??, July 2023. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023001279.

Soler:2024:QIQ

[SCD^{+}24] David Soler, Iván Cillero, Carlos Dafonte, Manuel Fernández-Veiga, Ana Fernández Vilas, and Francisco J.

                                        Nóvoa. QKDNetSim+: Im-

provement of the quantum network simulator for NS-3. SoftwareX, 26 (??):??, May 2024. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711024000566.

Silva:2020:DRD

[SCG^{+}20] Vítor Silva, Vinícius Campos, Thaylon Guedes, José Camata, Daniel de Oliveira, Alvaro L. G. A. Coutinho, Patrick Valduriez, and Marta Mattoso. DfAnalyzer: Runtime dataflow analysis tool for computational science and engineering applications. SoftwareX, 12 (??):Article 100592, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020303058.

Skinner:2023:DDP

[SCK^{+}23] Grace Skinner, Rob Cooke, Junghyuk Keum, Andy Purvis, Cristina Raw, Ben A. Woodcock, and Joseph Millard. Dynameta: a dynamic platform for ecological meta-analyses in R Shiny. SoftwareX, 23 (??):??, July 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023001358.

Steenbeek:2018:ENA

[SCPC18] Jeroen Steenbeek, Xavier Corrales, Mark Platts, and Marta Coll. Ecosampler: a new approach to assessing parameter uncertainty in Ecopath with Ecosim. SoftwareX, 7(??):198–204, January/June 2018. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018300803.

Silva-Cancino:2024:AGU

[SCSB24] Nathalia Silva-Cancino, Fernando Salazar, and Ernest Bladé. ACROPOLIS: a graphical user interface for classification of risk for off-stream reservoirs using machine learning. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711024000281.

Stankowski:2023:IPS

[SD23] Jakub Stankowski and Adrian Dziembowski. IV-PSNR: Software for immersive video objective quality evaluation. SoftwareX, 24 (??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023002881.

Soler-Dominguez:2019:WTT

[SDCA19] José L. Soler-Domínguez, Manuel Contero, and Mariano Alcañiz. Workflow and tools to track and visualize behavioural data from a virtual reality environment using a lightweight GIS. SoftwareX, 10(??):Article 100269, July/December 2019. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711018300931.

Strijbol:2024:BES

[SDG^{+}24] Niko Strijbol, Robbe De Proft, Klaas Goethals, Bart Mesuere, Peter Dawyndt, and Christophe Scholliers. Blink: an educational software debugger for Scratch. SoftwareX, 25(??):??, February 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023003138.

Skorych:2020:DOS

[SDH20] Vasyl Skorych, Maksym Dosta, and Stefan Heinrich. Dyssol — an opensource flowsheet simulation framework for particulate materials. SoftwareX, 12(??):Article 100572, July/December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020302855.

Sebastianelli:2021:ADB

[SDL21] Alessandro Sebastianelli, Maria Pia Del Rosso, and Silvia Liberata Ullo. Automatic dataset builder for machine learning applications to satellite imagery. SoftwareX, 15(??):??, July 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021000728.

Semper:2020:ERF

[SDP^{+}20] S. Semper, M. Döbereiner, S. Pawar, M. Landmann, and G. Del Galdo. eadf: Representation of farfield antenna responses in Python. SoftwareX, 12 (??):Article 100583, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S235271102030296X.

Shams:2023:GPR

[SE23] Mosayeb Shams and Ahmed H.

                                        Elsheikh. Gym-preCICE:

Reinforcement learning environments for active flow control. SoftwareX, 23 (??):??, July 2023. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023001425.

Sosa-Espadas:2023:IIR

[SEdCCCRG23] Cristian E. Sosa-Espadas, Mauricio G. Orozco del Castillo, Nora Cuevas-Cuevas, and Juan A. Recio-Garcia. IREX: Iterative Refinement and Explanation of classification models for tabular datasets. SoftwareX, 23 (??):??, July 2023. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023001164.

Sadollah:2016:WCA

[SEL^{+}16] Ali Sadollah, Hadi Eskandar, Ho Min Lee, Do Guen Yoo, and Joong Hoon Kim. Water cycle algorithm: a detailed standard code. SoftwareX, 5 (??):31–36, ???? 2016. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711016300024.

Sandoval-Espino:2023:ISP

[SEZLMS^{+}23] Jorge Arturo Sandoval-Espino, Alvaro Zamudio-Lara, José Antonio Marbán-Salgado, J Jesús Escobedo-Alatorre, Omar Palillero-Sandoval, and J. Guadalupe Velásquez Aguilar. Image set preparation: a platform to prepare a myoelectric signal to train a CNN. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023002054.

Schumacher:2016:AMT

[SF16] Florian Schumacher and Wolfgang Friederich. ASKI: a modular toolbox for scattering-integral-based seismic full waveform inversion and sensitivity analysis utilizing external forward codes. SoftwareX, 5(??):243–251, ???? 2016. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711016300346.

Sanders:2019:DLA

[SFF^{+}19] Jeremiah W. Sanders, Justin R. Fletcher, Steven J. Frank, Ho-Ling Liu, Jason M. Johnson, Zijian Zhou, Henry Szu-Meng Chen, Aradhana M. Venkatesan, Rajat J. Kudchadker, Mark D. Pagel, and Jingfei Ma. Deep learning application engine (DLAE): Development and integration of deep learning algorithms in medical imaging. SoftwareX, 10(??):Article 100347, July/December 2019. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711019302535.

Schneck:2021:PMA

[SFG21] William C. Schneck, Erik L. Frankforter, and Elizabeth D. Gregory. PanNDE: a modular architecture for high-performance NDE simulation. SoftwareX, 15 (??):??, July 2021. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021000923.

Shimojo:2019:QOS

[SFK^{+}19] Fuyuki Shimojo, Shogo Fukushima, Hiroyuki Kumazoe, Masaaki Misawa, Satoshi Ohmura, Pankaj Rajak, Kohei Shimamura, Lindsay Bassman, Subodh Tiwari, Rajiv K. Kalia, Aiichiro Nakano, and Priya Vashishta. QXMD: an opensource program for nonadiabatic quantum molecular dynamics. SoftwareX, 10 (??):Article 100307, July/ December 2019. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019300512.

Sisniega:2024:FOS

[SG24] Jaime Céspedes Sisniega and Álvaro López García. Frouros: an open-source Python library for drift detection in machine learning systems. SoftwareX, 26 (??):??, May 2024. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024001043.

Shkurti:2016:PPB

[SGA^{+}16] Ardita Shkurti, Ramon Goni, Pau Andrio, Elena Breitmoser, Iain Bethune, Modesto Orozco, and Charles A. Laughton. pyPcazip: a PCA-based toolkit for compression and analysis of molecular simulation data. SoftwareX, 5(??):37–43, ???? 2016. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S2352711016300036.

Santos:2022:MUM

[SGB^{+}22] Javier E. Santos, Alex Gigliotti, Abhishek Bihani, Christopher Landry, Marc A. Hesse, Michael J. Pyrcz, and Maša Prodanović. MPLBM-T: Multiphase LBM library for permeable media analysis. SoftwareX, 18(??):??, June 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022000668.

Sartori:2018:DTL

[SGBH18] Alberto Sartori, Nicola Giuliani, Mauro Bardelloni, and Luca Heltai. deal2lkit: a toolkit library for high performance programming in deal.II. SoftwareX, 7(??):318–327, January/June 2018. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711018302048.

Spillner:2022:TRS

[SGDC22] Josef Spillner, Panagiotis Gkikopoulos, Pamela Delgado, and Christine Choirat. Towards reproducible software studies with MAO and Renku. SoftwareX, 17(??):??, January 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021001783.

Sgarrella:2023:SSD

[SGP23] Joe Sgarrella, Farshad Ghanbari, and Christian

Peco. I-STL2MOOSE: From

STL data to integrated volumetrical meshes for MOOSE. SoftwareX, 21 (??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022001911.

Steppa:2019:HPH

[SH19] Constantin Steppa and Tim L. Holch. HexagDLy — processing hexagonally sampled data with CNNs in PyTorch. SoftwareX, 9(??):193–198, January/ June 2019. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018302723.

Steffen:2023:OSS

[SH23] Lennart Steffen and Reinhard Hinkelmann. hms^{++}: Open-source shallow water flow model with focus on investigating computational performance. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S2352711023000936.

Shaaban:2023:VSE

[Sha23] Mostafa Shaaban. Viability of the social-ecological agroecosystem (ViSA). SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023000560.

Sheu:2019:ASF

[She19] G. Y. Sheu. aXBRL: Search of fraudulent XBRL instance documents with an Android app. SoftwareX, 9(??):308–316, January/ June 2019. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711018302334.

Shadab:2023:PGS

[SHH23] Mohammad Afzal Shadab, Eric Hiatt, and Marc Andre Hesse. PKgui: GUI software for Polubarinova– Kochina’s solutions of steady unconfined groundwater flow. SoftwareX, 24 (??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023002698.

Stippinger:2023:BUH

[SHK^{+}23] Marcell Stippinger, Dávid Hanák, Marcell T. Kurbucz, Gergely Hanczár, Olivér M. Törteli, and Zoltán Somogyvári. BiometricBlender: Ultra-high dimensional, multi-class synthetic data generator to imitate biometric feature space. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023000626.

Shoshany:2024:CTP

[Sho24]  Barak Shoshany.  A

C++17 thread pool for high-performance scientific computing. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271102400058X.

Sipper:2023:EKE

[SHTE23] Moshe Sipper, Tomer Halperin, Itai Tzruia, and Achiya Elyasaf. EC-KitY: Evolutionary computation tool kit in Python with seamless machine learning integration. SoftwareX, 22 (??):??, May 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023000778.

Sclocco:2020:ART

[SHvW20] A. Sclocco, S. Heldens, and B. van Werkhoven. AMBER: a real-time pipeline for the detection of single pulse astronomical transients. SoftwareX, 12(??):Article 100549, July/December 2020. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711018302826.

Salazar:2024:SSD

[SIC24] Fernando Salazar, Joaquín Irazábal, and André Conde. SOLDIER: SOLution for Dam behavior Interpretation and safety Evaluation with boosted Regression trees. SoftwareX, 25(??):??, February 2024. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023002947.

Silvestre:2020:OOT

[Sil20] Daniel Silvestre. OPTool — an optimization toolbox for iterative algorithms. SoftwareX, 11 (??):Article 100371, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711019302547.

Shabani:2022:HOI

[SK22] Amirhosein Shabani and Mahdi Kioumarsi. Hyperomet: an OpenSees interface for nonlinear analysis of unreinforced masonry buildings. SoftwareX, 20(??): ??, December 2022. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022001480.

Sharma:2022:DRP

[SKD22] Parichit Sharma, Hasan Kurban, and Mehmet Dalkilic. DCEM: an R package for clustering big data via data-centric modification of Expectation Maximization. SoftwareX, 17(??):??, January 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021001771.

Shigarov:2019:TSP

[SKM19] A. Shigarov, V. Khristyuk, and A. Mikhailov. TabbyXL: Software platform for rule-based spreadsheet data extraction and transformation. SoftwareX, 10 (??):Article 100270, July/ December 2019. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018302966.

Shekhovtsov:2023:VPU

[SKS23] Andrii Shekhovtsov, Bartłomiej Kizielewicz, and Wojciech Sałabun. Version [1.1] — [pymcdm — the universal library for solving multicriteria decision-making problems]. SoftwareX, 24 (??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023002157.

Scheffel:2018:SES

[SL18] Jan Scheffel and Kristoffer Lindvall. SIR — an efficient solver for systems of equations. SoftwareX, 7 (??):59–62, January/June 2018. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711018300062.

Sinuco-Leon:2020:OLM

[SL20a] German A. Sinuco-Leon. OPENMMF: a library for multimode driven quantum systems. SoftwareX, 12 (??):Article 100603, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711020303162.

Stansbury:2020:PAF

[SL20b] Conrad Stansbury and Alessandra Lanzara. PyARPES: an analysis framework for multimodal angle-resolved photoemission spectroscopies. SoftwareX, 11 (??):Article 100472, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019301633.

Stephens:2021:ODT

[SL21] Victoria B. Stephens and David O. Lignell. Onedimensional turbulence (ODT): Computationally efficient modeling and simulation of turbulent flows. SoftwareX, 13(??):Article 100641, January 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271102030354X.

Scoggins:2020:MMT

[SLBC^{+}20] James B. Scoggins, Vincent Leroy, Georgios Bellas-Chatzigeorgis, Bruno Dias, and Thierry E. Magin. Mutation^{++}: multicomponent thermodynamic and transport properties for ionized gases in C++. SoftwareX, 12(??):Article 100575, July/December 2020. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020302880.

Schwertfeger:2023:IBC

[SLL23] Benjamin Thomas Schwertfeger, Gerrit Lohmann, and Henrik Lipskoch. Introduction of the BiasAdjustCXX command-line tool for the application of fast and efficient bias corrections in climatic research. SoftwareX, 22 (??):??, May 2023. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023000754.

Szatkowski:2023:LGB

[SLM23] Mateusz Szatkowski, Przemysław Litwin, and Jan Masajada. LBSA: a GUIbased Matlab software for the advanced laser beam shaping with spatial light modulators. SoftwareX, 22 (??):??, May 2023. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023000389.

Schwindt:2020:RA

[SLPR20] Sebastian Schwindt, Kenneth Larrieu, Gregory B. Pasternack, and Geoff Rabone. River Architect. SoftwareX, 11 (??):Article 100438, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/

science/article/pii/
S235271101930281X.

Sundararaman:2017:JSJ

[SLWS^{+}17] Ravishankar Sundararaman, Kendra Letchworth-Weaver, Kathleen A. Schwarz, Deniz Gunceler, Yalcin Ozhabes, and T. A. Arias. JDFTx: Software for joint density-functional theory. SoftwareX, 6 (??):271–277, ???? 2017. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711017300559.

Stocchi:2019:FWT

[SM19] Marco Stocchi and Michele Marchesi. Fast wavelet transform assisted predictors of streaming time series. SoftwareX, 8(??): 1–4, ???? 2019. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711017300481.

Shen:2023:VAO

[SMD23] Yang Shen, Robert H.

Moore, and Ankit Deo. Visualizing Abaqus output database in ParaView: a universal converter in Python and C++. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/

science/article/pii/
S2352711023000274.

Szydelko:2024:EVM

[SMD^{+}24] Błażej Szydełko, Dawid Mieloch, Adrian Dziembowski, Jakub Stankowski, Dominika Klóska, Jun Young Jeong, and Gwangsoon Lee. ECPC — versatile multicamera system calibration framework for immersive video applications. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024000402.

Song:2023:BTQ

[SMKL23] Jaejoon Song, Satoshi Morita, Ying-Wei Kuo, and J. Jack Lee. BayesESS: a tool for quantifying the impact of parametric priors in Bayesian analysis. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023000547.

Soto:2022:WGO

[SMR22] Andrés Soto, Héctor Mora, and Jaime A. Riascos.

Web Generator: an open-

source software for synthetic web-based user interface dataset generation. SoftwareX, 17(??): ??, January 2022. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022000073.

Sheikh-Mohammad-Zadeh:2024:SPG

[SMZSW24] Abbas Sheikh-Mohammad-Zadeh, Nicolas Saunier, and E. O. D. Waygood. STUDIO: a Python graphical tool for analyzing street user observations from video data. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024001134.

Spitzenberger:2020:PDV

[SNHS20] Andy Spitzenberger, Sebastian Neumann, Martin Heinrich, and Rüdiger Schwarze. Particle detection in VOF-simulations with OpenFOAM. SoftwareX, 11(??):Article 100382, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271101930264X.

Saz-Navarro:2024:CAN

[SNLFGVRB24] Dulcenombre M. Saz-Navarro, Aurelio López-Fernández, Francisco A. Gómez-Vela, and Domingo S. Rodriguez-Baena. CyEnGNet-App: a new Cytoscape app for the reconstruction of large co-expression networks using an ensemble approach. SoftwareX, 25(??):??, February 2024. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711024000050.

Schlueter:2021:GSM

[SNW^{+}21] Martin Schlueter, Mehdi Neshat, Mohamed Wahib, Masaharu Munetomo, and Markus Wagner. GTOPX space mission benchmarks. SoftwareX, 14(??):??, June 2021. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S235271102100011X.

Siala:2021:PPP

[SO21] Kais Siala and Leonhard Odersky. pyGRETA, pyCLARA, pyPRIMA: a pre-processing suite to generate flexible model regions for energy system models. SoftwareX, 16 (??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021001333.

Solvang:2023:TRP

[SO23a] Hiroko Kato Solvang and Mineaki Ohishi. trec: an R package for trend estimation and classification to support integrated ecosystem assessment of the marine ecosystem and environmental factors. SoftwareX, 21 (??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023000055.

Stepien:2023:TPT

[SO23b] Igor Stepień and Mariusz Oszust. TIQA-PSI: Toolbox for perceptual Image Quality Assessment of Pan-Sharpened Images. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023001905.

Sejdic:2019:SCC

[SOS19] Ervin Sejdić, Irena Orović, and Srdjan Stanković. A software companion for compressively sensed timefrequency processing of sparse nonstationary signals. SoftwareX, 8(??): 9–10, ???? 2019. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711017300365.

Shahdi:2019:SWS

[SP19] Arya Shahdi and Ekarit Panacharoensawad. SP-Wax: Solid–liquid equilibrium thermodynamic modeling software for paraffinic systems. SoftwareX, 9(??):145–153, January/June 2019. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711018302395.

Stringari:2022:PLC

[SP22] Caio Eadi Stringari and Hannah Power. Picoastal: a low-cost coastal video monitoring system. SoftwareX, 18(??):??, June 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022000541.

Slaughter:2023:MST

[SPG^{+}23] Andrew E. Slaughter, Zachary M. Prince, Peter German, Ian Halvic, Wen Jiang, Benjamin W. Spencer, Somayajulu L. N. Dhulipala, and Derek R.

                                        Gaston. MOOSE Stochastic

Tools: a module for performing parallel, memoryefficient in situ stochastic simulations. SoftwareX, 22 (??):??, May 2023. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023000419.

Soto-Perdomo:2023:ODG

[SPMGA^{+}23] Juan Soto-Perdomo, Juan Morales-Guerra, Juan David Arango, Sebastian Montoya Villada, Pedro Torres, and Erick Reyes-Vera.

OptiGUI DataCollector:

a graphical user interface for automating the data collecting process in optical and photonics labs. SoftwareX, 24 (??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023002170.

Subbiah:2017:PCR

[SR17] M. Subbiah and V. Ra-

jeswaran. proportion:

a comprehensive R package for inference on single binomial proportion and Bayesian computations. SoftwareX, 6(??): 30–35, ???? 2017. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711017300018.

Saad:2019:PPS

[SR19] Tony Saad and Giovanna Ruai. PyMaxEnt: a Python software for maximum entropy moment reconstruction. SoftwareX, 10 (??):Article 100353, July/ December 2019. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019302456.

Siripanich:2020:DMM

[SR20] Amarin Siripanich and Taha Hossein Rashidi. Dymium: a modular microsimulation modelling framework for integrated urban modelling. SoftwareX, 12(??):Article 100555, July/December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S235271102030008X.

Stojanovic:2023:MLS

[SRK23] Mario Stojanović, Francesco Romanò, and Hendrik C. Kuhlmann. MaranStable: a linear stability solver for multiphase flows in canonical geometries. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023001012.

Souza:2017:IMT

[SRML17] Roberto Souza, Letícia Rittner, Rubens Machado, and Roberto Lotufo. iamxt: Max-tree toolbox for image processing and analysis. SoftwareX, 6(??): 69–80, ???? 2017. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711017300079.

Sundararaghavan:2017:MIB

[SS17] Veera Sundararaghavan and Siddhartha Srivas-

tava. MicroFract: an

image based code for microstructural crack path prediction. SoftwareX, 6 (??):91–93, ???? 2017. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711017300092.

Salis:2019:RFS

[SS19] Vitalis Salis and Diomidis Spinellis. RepoFS: File system view of Git repositories. SoftwareX, 9(??): 288–292, January/June 2019. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018300712.

Spencer:2023:AGG

[SSC23] Max Spencer, Shohreh Sheiati, and Xiao Chen. AQUADA GUI: a graphical user interface for automated quantification of damages in composite structures under fatigue loading using computer vision and thermography. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023000882.

Setnikar:2022:RSR

[SSM22] Grégory Setnikar, Julien Samson, and Marie-Aude

                                        Méasson. Raman Selection
                                        Rules Calculator:  a

simplified selection rules calculator for Raman spectroscopy experiment. SoftwareX, 19(??):??, July 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022000942.

Serrao:2021:OTG

[SSP21] Prince Henry Serrao, Stefan Sandfeld, and Aruna Prakash. OptiMic: a tool to generate optimized polycrystalline microstructures for materials simulations. SoftwareX, 15 (??):??, July 2021. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021000534.

Scott:2016:SFP

[SSSH16] Erin Scott, Natalia Serpetti, Jeroen Steenbeek, and Johanna Jacomina Heymans. A Stepwise Fitting Procedure for automated fitting of Ecopath with Ecosim models. SoftwareX, 5(??):16–24, ???? 2016. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/

science/article/pii/
S2352711016000054.

Seiffarth:2024:OMO

[SSW^{+}24] Johannes Seiffarth, Tim Scherr, Bastian Wollenhaupt, Oliver Neumann, Hanno Scharr, Dietrich Kohlheyer, Ralf Mikut, and Katharina

Nöh. ObiWan-Microbi:

OMERO-based integrated workflow for annotating microbes in the cloud. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024000098.

Szeremeta:2021:GME

[ST21] Lukasz Szeremeta and Dominik Tomaszuk. Generating molecular entities as structured data. SoftwareX, 15(??):??, July 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021000698.

Stockdale:2018:KCL

[STC^{+}18] Gabriel Stockdale, Simone Tiberti, Daniela Camilletti, Gessica Sferrazza Papa, Ahmad Basshofi Habieb, Elisa Bertolesi, Gabriele Milani, and Siro Casolo. Kinematic collapse load calculator: Circular arches. SoftwareX, 7(??): 174–179, January/June 2018. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S235271101830061X.

Segura-Tinoco:2023:AAA

[STC23] Andrés Segura-Tinoco and Iván Cantador. ARGAEL: ARGument Annotation and Evaluation tooL. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023001061.

Sakib:2021:SSS

[STH^{+}21] Nazmus Sakib, Shiyu Tian, Md Munirul Haque, Rumi Ahmed Khan, and Sheikh Iqbal Ahamed. SepINav (Sepsis ICU Navigator): a data-driven software tool for sepsis monitoring and intervention using Bayesian Online Change Point Detection. SoftwareX, 14(??):??, June 2021. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021000340.

Sanz:2023:MML

[SU23] Victorino Sanz and Alfonso Urquia. MSGLib: a Modelica library for message passing communication. SoftwareX, 23 (??):??, July 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023001346.

Schwarz:2021:PPF

[SUM21] Sebastian Schwarz, Sebastian Alexander Uerlich, and Antonello Monti.

pycity scheduling — a

Python framework for the development and assessment of optimisationbased power scheduling algorithms for multi-energy systems in city districts. SoftwareX, 16(??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021001230.

Suriano:2020:SSS

[Sur20] Domenico Suriano. SentinAir system software: a flexible tool for data acquisition from heterogeneous sensors and devices. SoftwareX, 12(??):Article 100589, July/December 2020. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020303022.

Solmaz:2023:ATC

[SV23] Serkan Solmaz and Tom Van Gerven. Acrossim: a toolkit for cross-platform integration of CFD simulation data in computer graphics. SoftwareX, 24 (??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023002819.

Strijbol:2023:TET

[SVM^{+}23] Niko Strijbol, Charlotte Van Petegem, Rien Maertens, Boris Sels, Christophe Scholliers, Peter Dawyndt, and Bart Mesuere. TESTed — an educational testing framework with languageagnostic test suites for programming exercises. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023001000.

Schuster:2023:CDP

[SvZvdA23] Daniel Schuster, Sebastiaan J. van Zelst, and Wil M. P. van der Aalst. Cortado: a dedicated process mining tool for interactive process discovery. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023000699.

Sharaf:2020:BRP

[SWCP20] Taysseer Sharaf, Theren Williams, Abdallah Chehade, and Keshav Pokhrel. BLNN: an R package for training neural networks using Bayesian inference. SoftwareX, 11(??):Article 100432, January/June 2020. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271101930322X.

Seinstra:2015:E

[SWK15] Frank Seinstra, David Wallom, and Kate Keahey. Editorial. SoftwareX, 1–2(??):i, September 2015. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711015000072.

Schoeder:2019:EHP

[SWK19] S. Schoeder, W. A.

Wall, and M. Kronbichler. ExWave: a high performance discontinuous Galerkin solver for the acoustic wave equation. SoftwareX, 9(??):49– 54, January/June 2019. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018302024.

Semaan:2020:SSC

[SY20] Richard Semaan and Vikas Yadav. SCOUT: Signal Correction and Uncertainty Quantification Toolbox in MATLAB. SoftwareX, 11(??):Article 100474, January/June 2020. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711019303711.

Tudisco:2017:TLD

[TACH17] Erika Tudisco, Edward Andò, Rémi Cailletaud, and Stephen A. Hall. TomoWarp2: a local digital volume correlation code. SoftwareX, 6(??): 261–266, ???? 2017. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711017300511.

Tagusari:2023:SRP

[Tag23] Junta Tagusari. shinyHugePlot: an R package for displaying the features of data with a huge sample size. SoftwareX, 24 (??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023002790.

Takekawa:2022:FPC

[Tak22] Takashi Takekawa. Fast parallel calculation of modified Bessel function of the second kind and its derivatives. SoftwareX, 17(??):??, January 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/

science/article/pii/
S2352711021001655.

Taqi:2016:VFP

[Taq16] Ali H. Taqi. A visual Fortran 90 program for the two-particle or two-hole excitations of nuclei: the PPRPA program. SoftwareX, 5(??):44–50, ???? 2016. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711016300048.

Talamantes-Becerra:2021:OTF

[TBCG21] Berenice Talamantes-Becerra, Jason Carling, and Arthur Georges. omicR: a tool to facilitate BLASTn alignments for sequence data. SoftwareX, 14(??):??, June 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021000479.

Thorp:2023:VPF

[TBD23] Kelly R. Thorp, Josh Brekel, and Kendall C. De-Jonge. Version 1.2.0 — pyfao56: FAO-56 evapotranspiration in Python. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023002145.

Terezol:2024:OPP

[TBO24] Morgane Térézol, Anaïs Baudot, and Ozan Ozisik. ODAMNet: a Python package to identify molecular relationships between chemicals and rare diseases using overlap, active module and random walk approaches. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024000724.

Tsiotas:2020:VVE

[TC20] Dimitrios Tsiotas and Avraam Charakopoulos. VisExpA: Visibility expansion algorithm in the topology of complex networks. SoftwareX, 11 (??):Article 100379, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019302237.

Thuptimdang:2023:MMC

[TCCK23] Wanwara Thuptimdang, Patjanaporn Chalacheva, Thomas D. Coates, and Michael C. K. Khoo. McDAPS: a multi-channel physiological signals display and analysis system for clinical researchers. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023001784.

Trujillo:2022:GCC

[TCH^{+}22] Leonardo Trujillo, Jose Manuel Muñoz Contreras, Daniel E. Hernandez, Mauro Castelli, and Juan J. Tapia. GSGP-CUDA — a CUDA framework for Geometric Semantic Genetic Programming. SoftwareX, 18(??):??, June 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022000607.

Taranti:2022:CHD

[TCPC22]  Pier-Giovanni  Taranti,

Carlos Alberto Nunes Cosenza, Leonardo Antonio Monteiro Pessôa, and Rodrigo Abrunhosa Col-

lazo.  coppeCosenzaR:

a hierarchical decision model. SoftwareX, 17(??): ??, January 2022. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021001539.

Triana-Camacho:2023:PGS

[TCQOGM23] Daniel A. Triana-Camacho, Jorge H. Quintero-Orozco, and Enrique García-Macías.

Pypiezo-GO: a software

tool for processing electromechanical measurements of piezoelectric reduced graphene oxidecement composites. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023001474.

Thakur:2019:VNV

[TDG19] Abhishek Thakur, Arnav Dhamija, and Tejeshwar Reddy G. VECTORS — VidEo Communication Through Opportunistic Relays and Scalable video coding. SoftwareX, 9 (??):55–60, January/June 2019. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711018300918.

Temiz:2023:DDL

[Tem23] Hakan Temiz. DeepSR: a deep learning tool for image super resolution. SoftwareX, 21(??): ??, February 2023. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022001790.

Thrussell:2023:EPL

[TF23] Jasper Thrussell and Jim Michael Ferguson. ExactPack: a Python library of exact analytic solutions. SoftwareX, 24 (??):??, December 2023.

CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271102300256X.

Thaler:2024:JMS

[TFCZ24] Stephan Thaler, Paul Fuchs, Ana Cukarska, and Julija Zavadlav. JaxSGMC: Modular stochastic gradient MCMC in JAX. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024000931.

Tsai:2019:UIA

[TGS^{+}19] Hsieh-Fu Tsai, Joanna Gajda, Tyler F. W. Sloan, Andrei Rares, and Amy Q. Shen. Usiigaci: Instance-aware cell tracking in stain-free phase contrast microscopy enabled by machine learning. SoftwareX, 9(??):230– 237, January/June 2019. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018301882.

Tesse:2023:GMP

[THG^{+}23] Robin Tesse, Cédric Hernalsteens, Eustache Gnacadja, Nicolas Pauly, Eliott Ramoisiaux, and Marion Vanwelde. Georges: a modular Python library for seamless beam dynamics simulations and optimization. SoftwareX, 24 (??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023002753.

Hart:2019:UAM

[tHLMN19] Cornelis Marcel Pieter ’t Hart, Georgios Leontaris, and Oswaldo Morales-Nápoles. Update (1.1) to ANDURIL — a MAT-LAB toolbox for ANalysis and Decisions with UnceRtaInty: Learning from expert judgments: ANDURYL. SoftwareX, 10(??):Article 100295, July/December 2019. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711019302419.

Thorp:2022:PFE

[Tho22] Kelly R. Thorp. pyfao56: FAO-56 evapotranspiration in Python. SoftwareX, 19(??):??, July 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022001261.

Trinh:2024:MFQ

[THSW24] Thuat T. Trinh, Morten Hammer, Vishist Sharma, and Øivind Wilhelmsen. Mie-FH: a quantum corrected pair potential in the LAMMPS simulation package for hydrogen mixtures. SoftwareX, 26 (??):??, May 2024. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024000876.

Torre:2020:MIP

[THT20] Iván G. Torre, Richard J.

Heck, and A. M. Tarquis. MULTIFRAC: an ImageJ plugin for multiscale characterization of 2D and 3D stack images. SoftwareX, 12(??):Article 100574, July/December 2020. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020302879.

Ternstrom:2018:FNS

[TJS18] Sten Ternström, Dennis Johansson, and Andreas Selamtzis. FonaDyn — - a system for real-time analysis of the electroglottogram, over the voice range. SoftwareX, 7 (??):74–80, January/June 2018. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271101830030X.

Ternstrom:2019:UFS

[TJS19] Sten Ternström, Dennis Johansson, and Andreas Selamtzis. Update 2.0 to FonaDyn — a system for real-time analysis of the electroglottogram, over the voice range. SoftwareX, 10(??):Article 100343, July/December 2019. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019302559.

Tan:2023:NAD

[TK23] Ziya Tan and Mehmet Karaköse. A new approach for drone tracking with drone using Proximal Policy Optimization based distributed deep reinforcement learning. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023001930.

Tranter:2019:PSE

[TKLG19] T. G. Tranter, M. D. R. Kok, M. Lam, and J. T. Gostick. pytrax: a simple and efficient random walk implementation for calculating the directional tortuosity of images. SoftwareX, 10(??):Article 100277, July/December 2019. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019302286.

Tian:2019:AAP

[TLDM19] Liang Tian, Lin Li, Jun Ding, and Normand Mousseau. ART data analyzer: Automating parallelized computations to study the evolution of materials. SoftwareX, 9(??):238– 243, January/June 2019. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018302851.

Tseng:2023:XER

[TMGW23] Mattan Tseng, Jack Miller, Dante Goldner, and Eliot Winer. XRAE: Extended Reality Acuity Examination. SoftwareX, 23(??): ??, July 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023001255.

Thieu:2023:MOS

[TOPC23] Nguyen Van Thieu, Diego Oliva, and Marco Pérez-

Cisneros. MetaCluster:

an open-source Python library for metaheuristicbased clustering problems. SoftwareX, 24(??): ??, December 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023002935.

Truster:2018:DDE

[Tru18] Timothy J. Truster. DEIP, discontinuous element insertion Program–Mesh generation for interfacial finite element modeling. SoftwareX, 7(??):162–170, January/June 2018. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711018300797.

Tartarini:2020:PPP

[TS20] Federico Tartarini and Ste-

                                        fano Schiavon. pythermalcomfort:

a Python package for thermal comfort research. SoftwareX, 12(??):Article 100578, July/December 2020. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711020302910.

Thrun:2021:FCA

[TS21] Michael C. Thrun and Quirin Stier. Fundamental clustering algorithms suite. SoftwareX, 13(??): Article 100642, January 2021. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711020303551.

Tetzlaff:2022:MSA

[TS22] Laurens M. Tetzlaff and Gero Szepannek. mlr3shinystate-of-the-art machine learning made easy. SoftwareX, 20(??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022001649.

Tartarini:2020:CTC

[TSCH20] Federico Tartarini, Stefano Schiavon, Toby Cheung, and Tyler Hoyt. CBE

Thermal Comfort Tool:

Online tool for thermal comfort calculations and visualizations. SoftwareX, 12(??):Article 100563, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020302454.

Tsapetis:2023:UVU

[TSG^{+}23] Dimitrios Tsapetis, Michael D.

Shields, Dimitris G. Giovanis, Audrey Olivier, Lukas Novak, Promit Chakroborty, Himanshu Sharma, Mohit Chauhan, Katiana Kontolati, Lohit Vandanapu, Dimitrios Loukrezis, and Michael Gardner. UQpy v4.1: Uncertainty quantification with Python. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023002571.

Terefenko:2024:CVN

[TŚG^{+}24] Paweł Terefenko, Jakub Śledziowski, Andrzej Giza, Kamran Tanwari, Witold Maćków, and Andrzej

                                        Łysko. CCMORPH v2.0 —

new coastal cliffs morphology analysis and visualization toolbox. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711024000499.

Tenorio-Suarez:2022:MOS

[TSGOC^{+}22] Mauricio Ivan Tenorio-Suárez, Arturo Gómez-Ortega, Horacio Canales, Saul Piedra, and James Pérez-Barrera. MaSMaker: an open-source, portable software to create and integrate maze-like surfaces into arbitrary geometries. SoftwareX, 19(??):??, July 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022001212.

Tsoulos:2019:GPT

[TSMT19] Ioannis G. Tsoulos, Vasileios Stavrou, Nikolaos E. Mastorakis, and Dimitrios

                                        Tsalikakis. GenConstraint:

a programming tool for constraint optimization problems. SoftwareX, 10 (??):Article 100355, July/ December 2019. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019302250.

Tervydis:2024:CTO

[TSNRM24] Paulius Tervydis, Linas Svilainis, Zilvinas Nakutis, and Alberto Rodríguez-

Martínez. CloudEdgeAssetOptimizer:

Tool to optimize the Cloud–Edge computing network resources at given requirements of processing delay, battery capacity and cost. SoftwareX, 26(??):??, May 2024. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024000852.

Tsypin:2017:CRD

[TT17] Lev M. Tsypin and Aaron P.

Turkewitz. The coregulation Data Harvester: Automating gene annotation starting from a transcriptome database. SoftwareX, 6(??):161–164, ???? 2017. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711017300237.

Toth:2023:VBP

[TTHZ23] Róbert Tóth, Bálint Tóth, Miklós Hoffmann, and

Marianna Zichar. viskillz-blender

— a Python package to generate assets of Mental Cutting Test exercises using Blender. SoftwareX, 22 (??):??, May 2023. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023000249.

Tsoulos:2019:NTB

[TTT19] Ioannis G. Tsoulos, Alexandros Tzallas, and Dimitris Tsalikakis. NNC: a tool based on grammatical evolution for data classification and differential equation solving. SoftwareX, 10(??):Article 100297, July/December 2019. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711019301128.

Tsang:2022:BOL

[TWL22] Yung Po Tsang, Chun Ho Wu, and Carman Ka Man

                                        Lee. BlockTrainHK: an

online learning game for experiencing blockchain concepts. SoftwareX, 19 (??):??, July 2022. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022001005.

Urbikain:2020:MCM

[UdL20] G. Urbikain and L. N. López de Lacalle. MoniThor: a complete monitoring tool for machining data acquisition based on FPGA programming. SoftwareX, 11 (??):Article 100387, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018300645.

Ulloa:2016:CSE

[Ull16] Roberto Ulloa. CulSim:

a simulator of emergence and resilience of cultural diversity. SoftwareX, 5 (??):144–149, ???? 2016. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711016300206.

Usseglio-Viretta:2022:MOS

[UVPB^{+}22] F. L. E. Usseglio-Viretta, P. Patel, E. Bernhardt, A. Mistry, P. P. Mukherjee, J. Allen, S. J. Cooper, J. Laurencin, and K. Smith. MATBOX: an Open-source Microstructure Analysis Toolbox for microstructure generation, segmentation, characterization, visualization, correlation, and meshing. SoftwareX, 17(??): ??, January 2022. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021001618.

Vanfretti:2016:RME

[VBA^{+}16] Luigi Vanfretti, Maxime Baudette, Achour Amazouz, Tetiana Bogodorova, Tin Rabuzin, Jan Lavenius, and Francisco José Goméz-López. RaPId: a modular and extensible toolbox for parameter estimation of Modelica and FMI compliant models. SoftwareX, 5 (??):139–143, ???? 2016. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S235271101630019X.

Vidmar:2018:QPP

[VC18] R. Vidmar and N. Creati. QCOBJ: a Python package to handle quantityaware configuration files. SoftwareX, 7(??):347–351, January/June 2018. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711018302383.

Vallejo:2024:EWB

[VCA^{+}24] Paola Vallejo, Daniel Correa, Juan Carlos Arbeláez, Marta S. Tabares, Santiago Ruiz-Arenas, Elizabeth Rendon-Velez, David Ríos-Zapata, and Joan Alvarado. EcoMind: Webbased waste labeling tool. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024000554.

Vishart:2021:PPB

[VCLS21] Jonas Lynge Vishart, Jaime Castillo-León, and Winnie E. Svendsen. pyEIA: a Python-based framework for data analysis of electrochemical methods for immunoassays. SoftwareX, 15(??):??, July 2021. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021000637.

Valenza:2020:WMA

[VDCL20] Andrea Valenza, Luca Demetrio, Gabriele Costa, and Giovanni Lagorio.

WAF-A-MoLE: an adver-

sarial tool for assessing ML-based WAFs. SoftwareX, 11(??):Article 100367, January/June 2020. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019302997.

vandenOord:2020:PID

[vdOJP^{+}20] Gijs van den Oord, Fredrik Jansson, Inti Pelupessy, Maria Chertova, Johanna H. Grönqvist, Pier Siebesma, and Daan Crommelin. A Python interface to the Dutch Atmospheric Large-Eddy Simulation.

SoftwareX, 12(??):Article 100608, July/December 2020. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711020303216.

Veltri:2019:NFH

[VDP^{+}19] Luca Veltri, Luca Davoli, Riccardo Pecori, Armando Vannucci, and Francesco Zanichelli. NEMO: a flexible and highly scalable network EMulatOr. SoftwareX, 10(??):Article 100248, July/December 2019. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711019300135.

vonDomaros:2018:PMC

[vDPI^{+}18] Michael von Domaros, Eva Perlt, Johannes Ingenmey, Gwydyon Marchelli, and Barbara Kirchner. Peacemaker 2: Making clusters talk about binary mixtures and neat liquids. SoftwareX, 7(??):356–359, January/June 2018. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711018301845.

VanBerkel:2023:PRP

[VEF^{+}23] Derek Van Berkel, Thomas Estabrook, Nathan Fox, Rahul Agrawal Bejarano, Lisa Maillard, Devin Gill, Erica Akemi Goto, and Maria Carmen Lemos. PPGISr: an R package for Public Participatory GIS. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023000857.

Velasquez:2023:TAB

[Vel23] Juan D. Velasquez. TechMiner:

Analysis of bibliographic datasets using Python. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271102300153X.

Villoria:2018:WBA

[VEM^{+}18] Nelson B. Villoria, Joshua Elliott, Christoph Müller, Jaewoo Shin, Lan Zhao, and Carol Song. Webbased access, aggregation, and visualization of future climate projections with emphasis on agricultural assessments. SoftwareX, 7 (??):15–22, January/June 2018. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711017300626.

Verma:2022:PBT

[Ver22] Ashok Kumar Verma. A Python-based tool for constructing observables from the DSN’s closed-loop archival tracking data files. SoftwareX, 19(??):??, July 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022001145.

Vieira:2023:PIP

[VFFS23] Gabriel S. Vieira, Afonso U. Fonseca, Julio C. Ferreira, and Fabrizzio Soares.

                                        ProtectLeaf: an insect

predation analyzer for agricultural crop monitoring. SoftwareX, 24 (??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023002339.

Vanfretti:2018:APS

[VJA^{+}18] L. Vanfretti, G. M. Jónsdóttir, M. S. Almas, E. Rebello, S. R. Firouzi, and M. Baudette. Audur — a platform for synchrophasor-based power system wide-area control system implementation. SoftwareX, 7(??): 294–301, January/June 2018. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711018301730.

vanKessel:2020:NMC

[vKH20] L. van Kessel and C. W. Hagen. Nebula: Monte Carlo simulator of electronmatter interaction. SoftwareX, 12(??):Article 100605, July/December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020303186.

vanKuppevelt:2020:MAD

[vKMH^{+}20] D. van Kuppevelt, C. Meijer, F. Huber, A. van der Ploeg, S. Georgievska, and V. T. van Hees. Mcfly: Automated deep learning on time series. SoftwareX, 12(??):Article 100548, July/December 2020. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019300202.

Vargas-Martinez:2024:MDM

[VMRVF^{+}24] Manuel Vargas-Martínez, Nelson Rangel-Valdez, Eduardo Fernández, Claudia Gómez-Santillán, Gilberto Rivera, and Fausto Balderas.

MOSA/D-O and MOSAD/D-O-II:

Performance analysis of decomposition-based algorithms in many objective problems. SoftwareX, 25 (??):??, February 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023003060.

Petegem:2023:DLC

[VMS^{+}23] Charlotte Van Petegem, Rien Maertens, Niko Strijbol, Jorg Van Renterghem, Felix Van der Jeugt, Bram De Wever, Peter Dawyndt, and Bart Mesuere. Dodona: Learn to code with a virtual coteacher that supports active learning. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023002741.

Vila-Perez:2022:EGD

[VPVNP22] Jordi Vila-Pérez, R. Loek Van Heyningen, Ngoc-Cuong Nguyen, and Jaume Peraire. Exasim: Generating discontinuous Galerkin codes for numerical solutions of partial differential equations on graphics processors. SoftwareX, 20(??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022001303.

Vanfretti:2016:IPS

[VRBM16] L. Vanfretti, T. Rabuzin, M. Baudette, and M. Murad. iTesla Power Systems Library (iPSL): a Modelica library for phasor timedomain simulations. SoftwareX, 5(??):80–83, ???? 2016. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/

science/article/pii/
S2352711016300097. See

update [BCR^{+}18].

Vencels:2019:ELO

[VRG19] Juris Vencels, Peter Råback, and Vadims Geza. EOF-Library: Open-source Elmer FEM and OpenFOAM coupler for electromagnetics and fluid dynamics. SoftwareX, 9(??):68–72, January/June 2019. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018302164.

Varvia:2023:MRP

[VRP23] Petri Varvia, Janne Räty, and Petteri Packalen. mgpr: an R package for multivariate Gaussian process regression. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023002595.

Villada:2023:AMB

[VRVAC23] Sebastian Montoya Villada, Erick Reyes-Vera, and Mauricio Arias-Correa. AnIMAGE: a MATLABbased tool for generating microstrip antennas with complex shapes. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S235271102300198X.

Vidoni:2019:RRP

[VV19] Melina Vidoni and Aldo Vecchietti. “rsppfp”: an R package for the shortest path problem with forbidden paths. SoftwareX, 9(??):265–270, January/ June 2019. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711018300955.

Donckt:2022:TFT

[VVDV22] Jonas Van Der Donckt, Jeroen Van Der Donckt, Emiel Deprost, and Sofie Van Hoecke. tsflex: Flexible time series processing and feature extraction. SoftwareX, 17(??): ??, January 2022. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021001904.

Wallerberger:2023:SIO

[WBH^{+}23] Markus Wallerberger, Samuel Badr, Shintaro Hoshino, Sebastian Huber, Fumiya Kakizawa, Takashi Koretsune, Yuki Nagai, Kosuke Nogaki, Takuya Nomoto, Hitoshi Mori, Junya Otsuki, Soshun Ozaki, Thomas Plaikner, Rihito Sakurai, Constanze Vogel, Niklas Witt, Kazuyoshi Yoshimi, and Hiroshi Shinaoka. sparse-ir: Optimal compression and sparse sampling of manybody propagators. SoftwareX, 21(??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022001844.

Watrobski:2023:VPM

[WBNS23] Jarosław Watróbski, Aleksandra Baczkiewicz, Kesra Nermend, and Wojciech Sałabun. Version [1.2] —

[pyrepo-mcda — reference

objects based MCDA software package]. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023002716.

Wacha:2021:PPE

[WBS21] András Wacha and Tamás Beke-Somfai. PmlBeta: a PyMOL extension for building β-amino acid insertions and β-peptide sequences. SoftwareX, 13 (??):Article 100654, January 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020303678.

Watrobski:2022:PMR

[WBS22a] Jarosław Watróbski, Aleksandra Baczkiewicz, and Wojciech Sałabun. pyrepo-mcda — reference objects based MCDA software package. SoftwareX, 19(??):??, July 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022000711.

Watrobski:2022:VPM

[WBS22b] Jarosław Watróbski, Aleksandra Baczkiewicz, and Wojciech Sałabun. Ver-

                                        sion [1.1] — [pyrepo-mcda

— reference objects based MCDA software package]. SoftwareX, 19(??):??, July 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022001170.

Wu:2023:SST

[WC23] Lingyan Wu and Gang Cai. SNNTool: a software tool for sampling neural networks algorithms implementation. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023002704.

Weisrock:2023:CGE

[WCW^{+}23] Antoine Weisrock, Victor Couty, Jean-François Witz, Lieven Thorrez, and Pauline Lecomte-Grosbras. CRAPPY goes embedded: Including lowcost hardware in experimental setups. SoftwareX, 22(??):??, May 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023000444.

Werys:2020:TMT

[WDZ^{+}20] Konrad Werys, Iulius Dragonu, Qiang Zhang, Iulia Popescu, Evan Hann, Henrike Puchta, Agata Kubik, Dogan Polat, Cody Wu, Niall O. Moon, Ahmet Barutcu, Vanessa M. Ferreira, and Stefan K.

Piechnik. Total Mapping
Toolbox (TOMATO): an

open source library for cardiac magnetic resonance parametric mapping. SoftwareX, 11 (??):Article 100369, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711019302225.

Weber:2017:UTO

[Web17] Tobias Weber. Update 1.5 to “Takin: an opensource software for experiment planning, visualisation, and data analysis”, (PII: S2352711016300152). SoftwareX, 6(??):141–147, ???? 2017. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711017300183. See

[WGB16].

Weber:2021:UTO

[Web21] Tobias Weber. Update 2.0 to “Takin: an opensource software for experiment planning, visualisation, and data analysis”, (PII: S2352711016300152). SoftwareX, 14(??):??, June 2021. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021000121. See

[WGB16].

Weber:2023:MPT

[Web23] Tobias Weber. Motion planning for triple-axis spectrometers. SoftwareX, 23(??):??, July 2023. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023001516.

Wette:2020:SPO

[Wet20] Karl Wette. SWIGLAL: Python and Octave interfaces to the LAL-Suite gravitational-wave data analysis libraries. SoftwareX, 12(??):Article 100634, July/December 2020. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020303472.

Wijaya:2023:NJD

[WFD23] Agung Danu Wijaya, Dedy Farhamsa, and Darmawati Darwis. netDFT: JAVA Density Functional Theory for solid. SoftwareX, 23 (??):??, July 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023001413.

Wittke:2023:EEO

[WFL^{+}23] Samantha Wittke, Anne Fouilloux, Petteri Lehti, Juuso Varho, Arttu Kivimäki, Maiju Karhu, Mika Karjalainen, Matti Vaaja, and Eetu Puttonen. EODIE — Earth Observation Data Information Extractor. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023001176.

Weber:2016:TOS

[WGB16] Tobias Weber, Robert Georgii, and Peter Böni. Takin: an open-source software for experiment planning, visualisation, and data analysis. SoftwareX, 5(??):112–120, ???? 2016. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711016300152. See

updates [Web17, Web21].

Willberg:2023:UMI

[WHG^{+}23] Christian Willberg, Jan-Timo Hesse, Marc Garbade, Martin Rädel, Falk Heinecke, Andreas Schuster, and Anna Pernatii. A user material interface for the Peridynamic Peridigm framework. SoftwareX, 21(??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023000183.

Willberg:2024:PPL

[WHP24] Christian Willberg, Jan-Timo Hesse, and Anna Pernatii. PeriLab — peridynamic laboratory. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711024000712.

Willis:2017:ONS

[Wil17] Ashley P. Willis. The

                                        Openpipeflow Navier–

Stokes solver. SoftwareX, 6(??):118–123, ???? 2017. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711017300158.

Wieckowski:2024:VHD

[WKCS24] Jakub Wieckowski, Bartłomiej Kizielewicz, Witold Chmielarz, and Wojciech Sałabun. Version [1.1] — [handling decision-making in intuitionistic fuzzy environment: PyIFDM package]. SoftwareX, 25(??): ??, February 2024. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023003163.

Wu:2020:CBS

[WKR^{+}20] Wensi Wu, Justyna Kosianka, Heather Reed, Christopher Stull, and Christopher Earls. CU-BENs: a structural modeling finite element library. SoftwareX, 11(??):Article 100485, January/June 2020. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711020300133.

Wieckowski:2022:PPL

[WKS22] Jakub Wieckowski, Bartłomiej Kizielewicz, and Wojciech Sałabun. pyFDM: a Python library for uncertainty decision analysis methods. SoftwareX, 20(??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/

science/article/pii/
S2352711022001893.

Wieckowski:2023:HDM

[WKS23] Jakub Wieckowski, Bartłomiej Kizielewicz, and Wojciech Sałabun. Handling decision-making in Intuitionistic Fuzzy environment: PyIFDM package. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023000407.

Wang:2016:OCS

[WLP16] Shaowen Wang, Yan Liu, and Anand Padmanabhan. Open cyberGIS software for geospatial research and education in the big data era. SoftwareX, 5(??):i, ???? 2016. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711015000138.

Weng:2023:AMC

[WLW23] Hao Weng, Zhengyang Liu, and Dayong Wang.

ALPHA MODEL: a C++ code

for modelling and analyzing hydrate-saturation-dependent permeability of hydrate-bearing porous media. SoftwareX, 23 (??):??, July 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/

science/article/pii/
S2352711023001292.

Weinhold:2021:PMT

[WM21a] Richard Weinhold and Robert Mieth. Power Market Tool (POMATO) for the analysis of zonal electricity markets. SoftwareX, 16(??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021001394.

Wilkinson:2021:EPM

[WM21b] Collin J. Wilkinson and John C. Mauro. Explorer.py: Mapping the energy landscapes of complex materials. SoftwareX, 14 (??):??, June 2021. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021000285.

Walczak:2024:KTF

[WMA24] Jakub Walczak, Marco Mancini, and Shahbaz Alvi. Kit4DL: Towards fast prototyping and experimentation in machine learning and deep learning. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024000785.

Wilkinson:2018:RPC

[WMM18] Collin J. Wilkinson, Yihong Z. Mauro, and John C. Mauro. RelaxPy: Python code for modeling of glass relaxation behavior. SoftwareX, 7(??):255– 258, January/June 2018. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018301146.

Wu:2023:DAP

[WMZH23] DaJiang Wu, Zhiyuan Mei, Yi Zhu, and Haozhong Hu. Development of an ABAQUSTM plug-in for predicting composite plates stiffness with inplane periodicity. SoftwareX, 21(??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022001996.

Weise:2020:PSU

[WPM^{+}20] Konstantin Weise, Lucas Poßner, Erik Müller, Richard Gast, and Thomas R. Knösche. Pygpc: a sensitivity and uncertainty analysis toolbox for Python. SoftwareX, 11 (??):Article 100450, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/

science/article/pii/
S2352711020300078.

Wijayasari:2023:MRB

[WRV23] Winda Wijayasari, Faizal Immaddudin Wira Rohmat, and Sparisoma Viridi. Malaccha: an R-based end-to-end Markov transition matrix extraction for land cover datasets. SoftwareX, 21(??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023000110.

Wieckowski:2024:MOS

[WS24a] Jakub Wieckowski and Wojciech Sałabun. MakeDecision: Online system for the graphical design of decisionmaking models in crisp and fuzzy environments. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024000293.

Wieckowski:2024:VPP

[WS24b] Jakub Wieckowski and Wojciech Sałabun. Version [1.1]-[pyFDM: a Python library for uncertainty decision analysis methods]. SoftwareX, 25(??): ??, February 2024. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S2352711023003035.

Wu:2023:VGB

[WSAC^{+}23] Rui Wu, Connor Scully-Allison, Chase Carthen, Andy Garcia, Roger Hoang, Christopher Lewis, Ronn Siedrik Quijada, Jessica Smith, Sergiu M. Dascalu, and Frederick C. Harris. vFirelib: a GPU-based fire simulation and visualization tool. SoftwareX, 23 (??):??, July 2023. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023001073.

Wagner:2022:FEL

[WSK22] Christoph W. Wagner, Sebastian Semper, and Jan Kirchhof. fastmat: Efficient linear transforms in Python. SoftwareX, 18 (??):??, June 2022. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022000218.

Wojnar:2024:RLR

[WSRC24] Maksymilian Wojnar, Szymon Szott, Krzysztof Rusek, and Wojciech

                                        Ciezobka. Reinforced-lib:

Rapid prototyping of reinforcement learning solutions. SoftwareX, 26 (??):??, May 2024. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024000773.

Wac:2023:CCN

[WSRMB23] Marceli Wac, Raul Santos-Rodriguez, Chris McWilliams, and Christopher Bourdeaux. CATS: Cloudnative time-series data annotation tool for intensive care. SoftwareX, 24 (??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023002893.

Wogrin:2022:LOS

[WTAG^{+}22] Sonja Wogrin, Diego Alejandro Tejada-Arango, Robert Gaugl, Thomas Klatzer, and Udo Bachhiesl. LEGO: the opensource Low-carbon Expansion Generation Optimization model. SoftwareX, 19 (??):??, July 2022. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022000887.

Wagner:2021:TEC

[WTZ^{+}21] Sebastian Wagner, Konstantin Thierbach, Thomas Zerjatke, Ingmar Glauche, Ingo Roeder, and Nico Scherf. TraCurate: Efficiently curating cell tracks. SoftwareX, 13(??):Article 100656, January 2021. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021000017.

Williams:2017:SBS

[WW17] Ian H. Williams and Philippe B. Wilson. SULISO: the Bath suite of vibrational characterization and isotope effect calculation software. SoftwareX, 6 (??):i, ???? 2017. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711016300358.

Wolgast:2023:ALA

[WWB^{+}23] Thomas Wolgast, Nils Wenninghoff, Stephan Balduin, Eric Veith, Bastian Fraune, Torben Woltjen, and Astrid Nieße. ANALYSE — Learning to attack cyber-physical energy systems with intelligent agents. SoftwareX, 23 (??):??, July 2023. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023001802.

Xu:2023:PPT

[XGW23] Jiachen Xu and Moritz Grosse-Wentrup. PyTES: a Python toolbox for closed-loop transcranial electrical stimulation. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023000997.

Xie:2022:RRP

[Xie22] Fangzhou Xie. rethnicity:

an R package for predicting ethnicity from names. SoftwareX, 17(??): ??, January 2022. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021001874.

Xu:2021:SSP

[XSC^{+}21] Qimen Xu, Abhiraj Sharma, Benjamin Comer, Hua Huang, Edmond Chow, Andrew J. Medford, John E. Pask, and Phanish Suryanarayana. SPARC: Simulation Package for Ab-initio Real-space Calculations. SoftwareX, 15(??):??, July 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021000546.

Xu:2020:MSM

[XSS20] Qimen Xu, Abhiraj Sharma, and Phanish Suryanarayana. M-SPARC: Matlab-simulation package for ab-initio realspace calculations. SoftwareX, 11(??):Article 100423, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711019303966.

Xu:2024:PUF

[XTD24] Yumeng Xu, Shubhanshu Tiwari, and Marco Drago. PycWB: a userfriendly, modular, and Python-based framework for gravitational wave unmodelled search. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711024000104.

Xu:2022:OPO

[XYC22] Hongyan Xu, Ayten Yigiter, and Jie Chen. onlineBcp: an R package for online change point detection using a Bayesian approach. SoftwareX, 17(??): ??, January 2022. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022000140.

Xiong:2024:GGP

[XZL24] Hongqiang Xiong, Zhiyu Zhang, and Jing Li. GPRlab: a ground penetrating radar data processing and analysis software based on MATLAB. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024000918.

Yaacov:2023:BBP

[Yaa23] Tom Yaacov. BPpy: Behavioral programming in Python. SoftwareX, 24 (??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023002522.

Yayan:2023:TMB

[YB23] Ugur Yayan and Cem Baglum. Tailored mutationbased software fault injection tool (IM-FIT). SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023001590.

Yang:2022:PTD

[YBVC^{+}22] Rui Yang, Nathan Blanc, Ariel Vardi-Chouchana, Noam Ouzana, and Guy Z. Ramon. PC-TAS: a design environment for phasechange and classical thermoacoustic systems. SoftwareX, 19(??):??, July 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022000899.

Yorulmaz:2019:DUF

[YC19] Onur Yorulmaz and A. Enis Cetin. Deconvolution using Fourier Transform phase, l_{1} and l_{2} balls, and filtered variation. SoftwareX, 8(??):11–17, ???? 2019. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711017300651.

Yurin:2020:PKB

[YD20] Aleksandr Yu. Yurin and Nikita O. Dorodnykh. Personal knowledge base designer: Software for expert systems prototyping. SoftwareX, 11(??):Article 100411, January/June 2020. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711019303334. See

update [YDN21].

Youssefi:2015:SRT

[YDMC15] S. Youssefi, S. Denei, F. Mastrogiovanni, and G. Cannata. Skinware 2.0: a real-time middleware for robot skin. SoftwareX, 3–4 (??):1–5, December 2015. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711015000102.

Yurin:2021:UPK

[YDN21] Aleksandr Yurievich Yurin, Nikita Olegovich Dorodnykh, and Olga Anatolievna Nikolaychuk. Update (4.2020.0303) to “Personal Knowledge Base Designer: Software for expert systems prototyping”, (PII: S2352711019303334). SoftwareX, 16(??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021001163. See

[YD20].

Yang:2024:VPS

[YFVL24] Xiaohao Yang, Nathan Fox, Derek Van Berkel, and Mark Lindquist. Viewscape: an R package for the spatial analysis of landscape perception and configurations in viewsheds of landscapes. SoftwareX, 26 (??):??, May 2024. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024000335.

Yetgin:2019:FES

[YG19] Ömer Emre Yetgin and Ömer Nezih Gerek. Feature extraction, selection and classification code for power line scene recognition. SoftwareX, 8(??): 43–47, ???? 2019. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711017300560.

Yang:2022:AMB

[YHC^{+}22] Tao Yang, Lizhu Hu, Anja Chen, Maroš Tunák, Shangyong Zhang, Deyou Yu, Xiaodong Tan, Michal Petru, Ivan Mašín, and Sundaramoorthy Palanisamy. AFDeter: a MATLABbased tool for simple and rapid determination of the structural parameters and the airflow-related properties of fibrous materials. SoftwareX, 20(??): ??, December 2022. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022001315.

Yildiz:2019:RTS

[YKC^{+}19]  Çagatay Yildiz, Baris

Kurt, Taha Yusuf Ceritli, Bülent Sankur, and Ali Taylan Cemgil. A real-time SIP network simulation and monitoring system. SoftwareX, 8 (??):21–25, ???? 2019. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711017300614.

Yadav:2019:PPB

[YKKD19] Vinay Yadav, Subhankar Karmakar, Pradip P. Kalbar, and A. K. Dikshit. PyTOPS: a Python based tool for TOPSIS. SoftwareX, 9(??):217–222, January/June 2019. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018302279.

Yi:2024:SXS

[YLM^{+}24] Di-Fan Yi, Qian Liu, Rui-Ting Ma, Bin-Long Wang, Hong-Bang Liu, Fei Xie, Huan-Bo Feng, Zu-Ke Feng, and Yang-Heng Zheng. Star-XP: a simulation framework for Polar-2/low energy Xray polarization detector. SoftwareX, 25(??): ??, February 2024. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023003229.

Yerima:2018:MMA

[YLS^{+}18] Suleiman Y. Yerima, Michael Loughlin, Sakir Sezer, John Moriarty, Mark Mc-Cann, Helen McAneney, Leeanne O’Hara, Mark A. Tully, Paul S. Ell, Robert Miller, and Geraldine Macdonald. MobiQ: a modular Android application for collecting social interaction, repeated survey, GPS and photographic data. SoftwareX, 7(??): 143–149, January/June 2018. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711018300670.

Yu:2023:REP

[YMA23] Jiaxin Yu, Tapan Mukerji, and Per Avseth.

                                        rockphypy: an exten-

sive Python library for rock physics modeling. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023002637.

Yahiaoui:2023:ITS

[YMCB23] Malik Yahiaoui, Denis Mazuyer, and Juliette Cayer-Barrioz. IssyOsborne: transIent Squeeze and Starvation in elastohydrodynamic — OScillating Bodies sOlving Reynolds oNe-dimensional Equation. SoftwareX, 23(??): ??, July 2023. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711023001188.

Yerlikaya-Ozkurt:2023:CPP

[YÖYB23] Fatma Yerlikaya-Özkurt, Ceyda Yazici, and Inci Batmaz. cmaRs: a powerful predictive data mining package in R. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023002492.

Yuan:2022:ISM

[YSM22] Yu Yuan, Gabriel Stockdale, and Gabriele Milani. Interactive software for mapping concentrated displacements in masonry arches. SoftwareX, 19 (??):??, July 2022. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022000966.

Yang:2022:SSR

[YYL^{+}22] Jin Yang, Yue Yin, Alexander K. Landauer, Selda Buyukozturk, Jing Zhang, Luke Summey, Alexander McGhee, Matt K. Fu, John O. Dabiri, and Christian Franck.

SerialTrack: ScalE and

rotation invariant augmented Lagrangian particle tracking. SoftwareX, 19 (??):??, July 2022. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022001224.

Zajc:2020:MTV

[Zaj20] Luka Cehovin Zajc. A modular toolkit for visual tracking performance evaluation. SoftwareX, 12 (??):Article 100623, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711020303368.

Zdybal:2020:PPS

[ZAPS20] Kamila Zdybał, Elizabeth Armstrong, Alessandro Parente, and James C. Sutherland. PCAfold: Python software to generate, analyze and improve PCA-derived lowdimensional manifolds. SoftwareX, 12(??):Article 100630, July/December 2020. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711020303435.

Zhang:2020:GPP

[ZC20] Yifei Zhang and Jia Cao. GSimPy: a Python package for measuring group similarity. SoftwareX, 12 (??):Article 100526, July/ December 2020. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711019303590.

Zermeno-Campos:2022:PCB

[ZCCGGV^{+}22] Nestor Alexander Zermeño-Campos, Daniel Cuevas-González, Juan Pablo García-Vázquez, Roberto López-Avitia, Miguel Enrique Bravo-Zanoguera, Marco A. Reyna, and Arnoldo Díaz-Ramírez. PÉEK: a cloud-based application for automatic electrocardiogram prediagnosis. SoftwareX, 19 (??):??, July 2022. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022000814.

Zine-El-Abidine:2023:OIM

[ZEADR23] Mouad Zine-El-Abidine, Helin Dutagaci, and David

Rousseau. Ordinalysis:

Interpretability of multidimensional ordinal data. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023000390.

Zekollari:2017:TMF

[Zek17] Harry Zekollari. TopoZeko:

a MATLAB function for 3-D and 4-D topographical visualization in geosciences. SoftwareX, 6 (??):278–284, ???? 2017. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711017300535.

Zhu:2019:IRL

[ZGZvB19] Bing Zhu, Zihan Gao, Junkai Zhao, and Seppe K. L. M. vanden Broucke. IRIC: an R library for binary imbalanced classification. SoftwareX, 10 (??):Article 100341, July/ December 2019. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711019301700.

Zambo:2018:M

[ZHP^{+}18] Samantha Zambo, Todd Holland, Nathaniel Plant, Kevin Duvieilh, Paul Elmore, Will Avera, Brian Bourgeois, A. Louise Perkins, and David Lalejini. MergeBathy (2015). SoftwareX, 7(??):180–183, January/June 2018. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711018300888.

Zhu:2015:GCP

[Zhu15] Ru Zhu. Grace: a crossplatform micromagnetic simulator on graphics processing units. SoftwareX, 3–4(??):22–26, December 2015. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S235271101500014X.

Zienert:2019:CTP

[Zie19] Tilo Zienert. cp-tools: a Python library for predicting heat capacity of crystalline substances. SoftwareX, 9(??):244–247, January/June 2019. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018301791.

Zhang:2023:VMS

[ZJKS23] Boqin Zhang, Xin Jing, Shashikant Kumar, and Phanish Suryanarayana. Version 2.0.0 — M-SPARC: Matlab-Simulation Package for Ab-initio Realspace Calculations. SoftwareX, 21(??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022002138.

Zhao:2024:MSI

[ZKJ^{+}24] Jun Zhao, Yu-Kwang Kim, Yu-Hyeong Jang, Jae Ho Chang, Sang Kyu Lee, and Hyoung-Moon Kim. MLEce: Statistical inference for asymptotically efficient closed-form estimators in R. SoftwareX, 26 (??):??, May 2024. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024000268.

Zhao:2023:CMP

[ZL23] Yang Zhao and Qing Liu. Causal ML: Python package for causal inference machine learning. SoftwareX, 21(??): ??, February 2023. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022002126.

Zou:2023:MMT

[ZLLT23] Jiaqi Zou, Zonghao Li, Xuanying Liu, and Haonan Tong. MSCPDPLab: a MAT-LAB toolbox for transfer learning based multisource cross-project defect prediction. SoftwareX, 21(??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022002047.

Zhu:2018:OPL

[ZMS18] Minjie Zhu, Frank McKenna, and Michael H. Scott. OpenSeesPy: Python library for the OpenSe’s finite element framework. SoftwareX, 7(??):6–11, January/June 2018. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711017300584.

Zhuang:2017:NVT

[ZNS17] Kelin Zhuang, Gerald R. North, and Mark J. Stevens. A NetCDF version of the two-dimensional energy balance model based on the full multigrid algorithm. SoftwareX, 6 (??):193–197, ???? 2017. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711017300262.

Zitzlsberger:2023:RRS

[ZPM23] Georg Zitzlsberger, Michal Podhoranyi, and Jan Martinovic. rsdtlib: Remote sensing with deeptemporal data library. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023000651.

Zupko:2019:FSE

[ZR19] R. Zupko and M. Rouleau.

ForestSim: Spatially explicit agent-based modeling of non-industrial forest owner policies. SoftwareX, 9(??):117–125, January/ June 2019. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711018302310.

Zambrano:2020:TDT

[ZRBCI20] V. Zambrano, R. Rodríguez-Barrachina, S. Calvo, and S. Izquierdo. TWINKLE: a digital-twin-building kernel for real-time computeraided engineering. SoftwareX, 11(??):Article 100419, January/June 2020. CO-DEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711019300664.

Zare-Rami:2019:MVP

[ZRK19] Keyvan Zare-Rami and Yong-Rak Kim. MIDAS-VT-Pre: Software to generate 2D finite element model of particle/ fiber embedded composites with cohesive zones. SoftwareX, 10(??):Article 100292, July/December 2019. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711018303054.

Zolotov:2021:PFO

[ZRK21] Oleg Zolotov, Yulia Romanovskaya, and Maria Knyazeva. pyFIRI — a free and open source Python software package of the non-auroral Earth’s lower ionosphere. SoftwareX, 16(??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021001461.

Zolotov:2022:VPF

[ZRK22] Oleg Zolotov, Yulia Romanovskaya, and Maria Knyazeva. Version 2.0pyFIRI — a free and open source Python software package of the nonauroral Earth’s lower ionosphere. SoftwareX, 20 (??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711022001819.

Zippo:2022:NEP

[ZS22] Raffaele Zippo and Giovanni Stea. Nancy: an efficient parallel Network Calculus library. SoftwareX, 19(??):??, July 2022. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S235271102200108X.

Zhang:2024:MCT

[ZS24] Lei Zhang and Markus Stricker. MatNexus: a comprehensive text mining and analysis suite for materials discovery. SoftwareX, 26(??):??, May 2024. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711024000256.

Zampieri:2021:ARA

[ZTC^{+}21] Matteo Zampieri, Andrea Toreti, Andrej Ceglar, Pierluca De Palma, Thomas Chatzopoulos, and Melania Michetti. Analysing the resilience of agricultural production systems with ResiPy, the Python production resilience estimation package. SoftwareX, 15(??):??, July 2021. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021000716.

Zubarev:2022:MNN

[ZVP22] Ivan Zubarev, Gavriela Vranou, and Lauri Parkkonen. MNEflow: Neural networks for EEG/ MEG decoding and interpretation. SoftwareX, 17(??):??, January 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711021001795.

Zhang:2022:PIP

[ZWCQ22] Yinsheng Zhang, Haiyan Wang, Yongbo Cheng, and Xiaolin Qin. pyCLAMs: an integrated Python toolkit for classifiability analysis. SoftwareX, 18(??):??, June 2022. CODEN ???? ISSN 2352-7110. URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S2352711022000188.

Zarski:2021:KTL

[ZWM21] Mateusz Zarski, Bartosz Wójcik, and Jarosław Adam Miszczak. KrakN: Transfer learning framework and dataset for infrastructure thin crack detection. SoftwareX, 16 (??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711021001503.

Zhang:2023:FRA

[ZY23] Yanhu Zhang and Lijuan Yan. Face recognition algorithm based on particle swarm optimization and image feature compensation. SoftwareX, 22 (??):??, May 2023. CO-DEN ???? ISSN 2352-7110. URL http://

www.sciencedirect.com/
science/article/pii/
S2352711023000018.